A Bibliography of Publications in The Computer Journal: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

22 October 2022 Version 1.99

Title word cross-reference

(2, 2) [LTC^{+}15]. (C_{s}, P_{t}) [HJP15]. (n, k) [ZHW19]. 1 [Dow15, TRY16]. 2 [DDG^{+}15, FGR17, KSA12, LL14, MBRM15, PG11, WWW16, XZLL18, ZX16]. 2^{2n+1} − 1 [BG15]. 2^{2n+2} − 1 [BG15]. 2^{2q} [AJ15]. 2^{n} [BG15]. 2^{n+1} − 1 [HS19]. 2^{n+k} [HS19]. 2^{n} − 1 [HS19]. 2^{q} ± 1 [AJ15]. 2^{q} ± 3 [AJ15]. 3 [DB13, GHFY18, GB10, jLbLzH18, LJ15, ZZLL18]. 5 [AJ15]. ^{∗} [NHC13]. ^{+} [YB16]. ^{2} [Cha10b, Hua14]. ^{3} [Cha10b, SC10]. ^{e} [NHC13]. _{cyclical} [YLLS16]. c [KRDH13]. C^{3}P [EFV15]. ℓ [ZTL15]. g [XZL17, ZLX^{+}19]. i^{∗} [SKK18]. K [EA17, LWPZ13, ABM12, ALH17, APW11, DLV10, Fan10, Fan11, GN19, GYDX12, KVX12, Kuo10, LLF17, MP18, WCL15, WWJ18]. L(2, 1) [Cal11a]. L(h, k) [Cal11b]. L(p, q) [ZQ13]. L_{p} [KV16]. m [Fan10]. μ [Jia14]. N [YC19, Fan10, Fan11, LLF17]. O(k) [DLV10]. p [BPK10, DD10b]. π [Cao10, HY11]. ±1 [HZW^{+}14]. q [CZCD18]. QR [ACG^{+}11]. S [LJ15]. t [Kor11, WCCL13].

-Adic [BPK10, DD10b]. -Ary [LLF17, CZCD18, Fan10, Fan11, Kor11, WCCL13]. -bit [KVX12]. -boxes [LJ15]. -Bubble-Sort [ZHW19]. -calculus [Cao10, HY11]. -Clustering [DLV10]. -Coteries [Kuo10]. -Cover [LWPZ13]. -Covered [ABM12]. -cube [Fan10, Fan11]. -Cubes [LLF17]. -D [DB13]. -Dimensional [Dow15]. -Diversity [ZTL15]. -Extra [WWW16, GHFY18]. -Free [HJP15].

-Good-Neighbor [XZL17, ZLX^{+}19]. -Labeling [Cal11a, ZQ13]. -Labelling [Cal11b]. -Layer [DDG^{+}15]. -Means [KRDH13]. -Medoids [EA17]. -Metric [TRY16]. -Moduli [AJ15]. -Multiple [LTC^{+}15]. -Nearest [GYDX12, WCL15]. -Neighborhood [KSA12]. -NMF [MP18]. -Pancycle-Connectivity [Fan10]. -policy [YC19]. -Security [Jia14]. -Time [DLV10]. -tree [NHC13]. -trees [XZLL18].

1 [BV15]. 128 [LJ18a, LYD^{+}18]. 1950s [Day11, Wet10]. 1960s [Day11].

2 [AAZ13]. 2-Party [BBKL19]. 2009 [GG10]. 2010 [Jar11]. 24th [GG10]. 2D [HIDFGPC15].

3 [NYT^{+}11]. 3D [AOS^{+}15]. 3D-RP [AOS^{+}15].

4 [YYO15].

5G [RASM17].

6 [XHC^{+}15]. 64/128 [LJ18a].

7 [AAZ13]. 754 [AAHTH10].

802.11 [OKA17, RHF^{+}15]. 802.11p [GH17].

978-0-262-02649-9 [Lar10]. 978-0-387-33333-5 [Gaz10]. 978-0-521-88038-1 [Maj10]. 978-1-4020-5529-4 [Jas10]. 978-1-4200-4757-8 [Joh10]. 978-981-277-171-1 [Lev10a].

= [AD16].

AADL [BCK^{+}11]. AAS [ALZ^{+}17]. abatement [HHH^{+}18]. ABE [QZZ18]. abelian [HWS^{+}19]. ABO [ZYY19]. ABO-LTFs [ZYY19]. Absolute [WXLL18].

Abstract [BDT10, MKW11]. Abstraction [WJ16, YWY10]. Accelerate [LGC19]. Accelerated [LR12]. Accelerating [VRD10, VO16]. Acceleration [PHM^{+}12]. Accelerator [WOLP15]. Accelerometer [CAV17]. Access [CLG^{+}19, CK15, HSMY12, HFP^{+}19, JCSZ13, KHC15, KHR^{+}19, LWDZ16, LPL15, LLLW17, RHF^{+}15, RASM17, XTH11, ZTBW11, ZVH11, ZDL^{+}17]. accessibility [NK19]. Accurate [PB14, RT12, YB16]. Achievable [Alm19]. Achieve [DBC18, Tan15]. Achieving [ABG^{+}12, BN14, KCZJ14]. ACO [TL19]. ACO-TS [TL19]. ACORN [ZFL18]. across [PSS10]. Action [WWHL12]. Active [CLM16, LJ15, OJSO14, WJ19]. Activities [BY14, NHMI13, WL18]. Activity [BY16, CLLH13, DSTC12, GRK13, KTC^{+}11, MGBD15, WXLL18, ZHL^{+}17]. Activity-Aware [DSTC12]. actor [SKS19]. Acyclic [BBB^{+}15, PGBFW14]. Ad [BBM10, BSK19, BAFF11, GH17, GGZC11, HC15, MK19, SJS12, SGG^{+}13, WCKH10, YWSH10, YDE11, YT11, ZYR^{+}13]. Adaptability [LSW10]. Adaptation [CMSML16, Kha11, RDB14b, TY14, ZSX10]. Adapted [GRK13, HM13]. Adaptive [AS11, ALA19, AAH10, ABCG11, BACD13, EFYS19, FFH17, FXV13, GN10, GTM15, HZW^{+}14, HT16, HLAZ15, HXLX18, HXLX22, ISH13, JDAS12, KXS^{+}10, KCZJ14, LCH16, LH11, LZL^{+}15, LL11b, MDY15, SVP13, WHYH12, WLI^{+}14, XYL^{+}11, ZLX^{+}15]. Adder [BG15]. Adder-Based [BG15]. Additive [Yas19]. Adic [BPK10, DD10b]. AdSelector [LS17]. Advanced [JSP13, KHC14]. Advances [Ano10, HXZ12]. Advent [Day11]. Adverse [NNF19, NK19]. adverse-tisements [NK19]. Advertisement [LS17]. Advertisements [NK19]. Advertising [LNBFPA13, WXZ^{+}12]. AES [BW16, SY15, VGA19, WJ19]. AES-Like [BW16, WJ19]. AET [HTC^{+}15]. Affiliation [XLM^{+}12, XGLM14, XZLW15]. Affiliation-Hiding [XLM^{+}12, XGLM14, XZLW15]. Affine [LYL^{+}18]. Against [BVS^{+}13, BL15b, BL16, CW12a, CMA14, GDCC16, HLJ^{+}15, LA12, LÖ10, SGH15, SLY^{+}16, WSA15, Che15a, HLLG18]. Age [ATA19]. Age-Group [ATA19]. Agent [AFGG11, BL11, CFMR14, GK17, KTTRJ10, LR14, Nic11, PXG^{+}17, RA14, Rog11, SNG^{+}10, TKB11, ZLG15, dFHP^{+}11, LBZ19]. Agent-Based [BL11, KTTRJ10, Nic11, TKB11]. Agents [BH10, Cor11, FT11, SZB15, SYH11, ZC10]. Aggregate [AGF15, WCD19]. Aggregating [MMH18]. Aggregation [EKOS19, SJ18a]. Aggregations [CTIAP12]. Agreeing [vdALM^{+}10]. Agreement [Chi16, MDS15, XLM^{+}12, XGLM14, XZLW15, dAEN^{+}18]. Ahead [JMB12, NH19]. Ahead-of-Time [JMB12]. AI [BT18]. Aided [Alm19, GMSV14, LNWZ19, MV19]. Air [XYL^{+}11]. Air-Cushion [XYL^{+}11]. al [LLSW16]. Alan [Lav12]. Algebra [DH12b, ZYF17]. Algebraic [WCXZ17]. Algebras [HTG12]. Algol [Gra12]. Algorithm [AK12, AUB11, BZS^{+}16, BKP11, Cai12, CW11, Che14, CFJ^{+}13, CGVP15, DA14, DLV10, DB13, FS18, GÁVRRL16, HM16, HLJ^{+}15, HQL17, IEBS19, KJ11, KR14, KV15b, LL11a, LL14, LR14, LYL^{+}18, LYC11, LV17, MSH^{+}11, MK19, OR12, RJ18, SC11, SZW^{+}18, SLL15, WZCC18, WS10, WLZ^{+}18, XHTH13, XYL^{+}11, ZWJ^{+}14, yZdZhZ18, ZSJ10, ZTTM18]. Algorithmic [BT18, ET19, Mur10b]. Algorithms [BBM10, BCG12, BMRS11, CC11, CHL14, CMSML16, DE10, GF17, HJK13, HK15, KRDH13, Kha16, KTTRJ18, LS14, Mar10a, MCT19, MBRM15, PB14, SJA17, STW^{+}18, SSK12, TKM11, Tah11, Tim11, WOV^{+}10, WCW^{+}18, YLW^{+}17, YDE11, YS15, ZBY^{+}10, ZW15, wZfG15, ZX16, ZDZ^{+}15a, Gon07]. Alignment [IA15, VRD10]. All-Against-All [LA12]. All-But-Many [CCL^{+}19]. Alliance [Ano10]. Allocation [BAFF11, CLH^{+}14, CMKJ10, CL16, Do11, FGS15, HGZ10, KV15a, KL10, KZY16, KCZJ14, LZL^{+}17, LS14, NNN^{+}14, OBA16, PZZ^{+}17, PCC^{+}16, RAJ15, ŞLV^{+}11, SKK^{+}12, SZB15, TL19, ZJH^{+}15, LJWL19, TXJ^{+}19]. Almost [GDCC16]. Almost-Tight [GDCC16]. Amazon [ÁHFE18]. Ambient [CvdT10, LLV10, PSS10, SS10a, SSY15, vDBvEW10]. Ami [SKK^{+}12]. Ammann [Maj10]. AMPS [GTM15]. Analogies [NLDH11]. Analysing [GSS14]. Analysis [ASCTFP16, Alh19, AHM15, BV15, BKFP19, BFCRH14, BCK^{+}11, CTIAP12, CW11, CWRZ18, CCY10, CCHL18, CAV17, DLM^{+}14, Das17, DBHC15, DHW10, Dim13, EDH^{+}18, FP19, Fra11, GRVD^{+}15, GH17, GMS^{+}12, HJL16, IAG^{+}14, LZZZ13, LSTC11, ML13, Mal10, Meg18, Meg19, MBRM15, Mur10a, NSA15, NP16, Pek12, PTOM18, RHF^{+}15, RMGT11, SY15, SSS^{+}12b, SKK18, Tah11, VKC15, WGS17, WHP^{+}13, XS11, YC19, ZH15, ZDM^{+}15, ZYY^{+}13, jZ18, ZL15, KAZ18, TZ11]. Analysis-Based [Meg18]. Analytic [ZW15]. Analytical [LDK11]. Analytics [AGF15, ALA19]. Analyzer [SSS^{+}12b]. Analyzing [DLL^{+}13, SS10b]. Anchor [NZ14]. and/or [YLA^{+}13]. ANEEC [PWY^{+}13]. Animal [KTC^{+}11]. Animation [SPS^{+}18]. Annealing [HGZ10]. Annotated [ATS15, Cal11b, SHH^{+}15]. Announcement [ALZ^{+}17]. Anomalies [ZYWW13]. Anomaly [GBBK11]. Anonymity [IDVGMP^{+}13]. Anonymization [MP18, SWLZ12]. Anonymous [Chi12, HH14, LSQX19, Ver17, Wan14, WYML16, YZJH12, ZJ14, ZMW16, ALZ^{+}17]. Answerin [CJYY17]. Answers [GN19]. Ant [WS10, ZDZ^{+}15b]. Ant-based [WS10].

Antcycle [HB11]. Anti [SK18a, TPG^{+}15]. Anti-Packet [TPG^{+}15]. Anti-Smishing [SK18a]. Antidimensional [TRY16]. API [QF19]. Appearance [SA11]. Appliance [CLLH13]. Appliance-Aware [CLLH13]. Application [BPK10, CCL^{+}19, CLRJ14, CBA18, CWZ19, CRGM14, DAOG14, Dim13, Küp15, MGBD15, NM19, RDB14b, TAC^{+}18, TEP^{+}16, WWXH18, WHP^{+}13, XHTH13, ZH15, ZM16]. Application-Level [CRGM14]. Application-Specific [DAOG14, TEP^{+}16]. Applications [AAA19, AAH10, Ang13, ABL^{+}18, Awa13, BDC11, BBP13, BFMT15, CXF^{+}15, CCC^{+}10, CRGM14, DG15a, ETR^{+}16, GB15, KHC14, KLL14, LWDZ16, LR14, MGM12, NP16, Nic11, PB12, PHM^{+}12, PL18, RAKJ17, RHH12, RR16, TS19, WOV^{+}10, XZA14, ZZQ^{+}19, ZYY19, RM08, Lev10a]. Applied [FGR17, GÁVRRL16, WN11]. Approach [AKA15, AZHASD14, AGP10, ALA19, BDT10, BS16, BFMT15, CFM17, CTD18, CJYY17, CQS13, Chi16, CW12b, CRGM14, DG15a, ELS11, FP18, FET17, GN19, HLAZ15, HS11, HY11, HZJS17, KHC15, KAS13, KZY16, LH11, LAP11, LfHmXjL11, LJ18b, LLV10, MDY15, MKN13, MMPB10, MEH19, MHMSGH16, OLL15, ÖKA11, PK18, PRG^{+}10, PTWB14, RHH12, RLVRGÁ15, SAPS19, ŞLV^{+}11, SH15, SK18a, TLRE11, TA16a, VMF^{+}14, VKC15, WGZW14, WXZ^{+}12, YMS^{+}15, YLLS16, ZW15, ZHY^{+}14, CPSK07, Gaz10]. Approaches [CQL10, EFYS19, GTB10, JHHC15, LCMC11, SZB19, TB11]. Approximate [Che14, IJY^{+}14, YHGL17]. Approximation [LS14, Mar10a, Gon07]. Approximations [TEP^{+}16]. Arbitrary [GDKP10]. Arbitration [Küp15]. Arbitrator [WSA15]. Arc [ZM19]. Arc-Connected [ZM19]. Archimedean [Ana10]. Architectural [Fra11]. Architecture [ACW13, AS11, AV16, ASG15, CXF^{+}15, CCCS11, CP16, CLL10, EOIH15, EDH^{+}18, GDKP10, GIB12, HMM11, yHRT^{+}12, IMS10, KJ11, KS18, LCX14, NPTZ16, PCC^{+}16, RMP10, SKKM15, TV12, VBMH10, WF10, WWZ^{+}17, ZLG15]. Architecture-Level [yHRT^{+}12]. Architecture/OS [CLL10]. Architectures [ABS12, GMS^{+}12, HYZ17, HMH18, OLF^{+}17, PHM^{+}12, ZZX16]. Archival [HZQ^{+}19]. ArchSORS [OLF^{+}17]. Area [AK12, BKPS10, CQS13, DDLM17, yHRT^{+}12, IAG^{+}14, KL14, KSPR15, LSCG10, PL16, SAKOK11, SJS12, XLXZ17, YZJH12]. Area-Feature [BKPS10]. Area-Thickness [DDLM17]. Argumentation [JHHC15]. Argumentation-Driven [JHHC15]. Arithmetic [HS19, LSTC11, Par15]. Arithmetic-Friendly [HS19]. Armed [LV17]. Arrangement [CQS13, MRPR15]. Array [PZ19]. Arrays [PC12]. Article [SMLM14]. Articles [CWWK14]. Artificial [GV16, Lev11b, LLV10, PW19, SS10a]. Ary [LLF17, CZCD18, Fan10, Fan11, Kor11, WCCL13]. Ascending [Mer13]. ASCENT [BCKM17]. Ashman [SLW^{+}17]. ASIC [NYT^{+}11]. ASICs [Jas10, Nur07]. Aspects [Nil10]. assembly [WCL^{+}11]. Assessing [ARR^{+}16]. Assessment [RMB15, THY^{+}18, TKB18]. Asset [Tim11]. Asset-Task [Tim11]. Assets [GTN10]. Assignment [GTN10, NG17, PTWB14, SJ18b, Tim11, WGL^{+}18, ZHL^{+}17]. Assignments [AÇPD11]. Assistance [EMB19]. Assisting [MGBD15]. Associate [Llo13]. Associated [CW11]. Association [GBA18, HK13]. Associative [ZSJ10]. Assortative [Meg16]. Assortativity [Meg19]. Assumption [CCL^{+}19, GWW^{+}13]. Assumptions [EKOS19, IAG^{+}14, PDNH15, RKBY15, YLL^{+}17]. Assurance [RBNB15]. Assured [LMA^{+}15, Tan15]. Asymmetric [OBA16, SPdGPM18, XLM^{+}12, XGLM14, XZLW15, ZZQ^{+}19]. Asynchronous [DGFGHZ13, Hie13, KW11, LAP11, ZLX^{+}15]. Atmospheric [LWKB15]. Atomic [DGFGHZ13, MPSP17]. Attack [CWZ19, DHT^{+}19, DSB15, GDCC16, HLJ^{+}15, HLAZ15, KH10, LLSW16, LJ19, PP17, ZFL18, TYL^{+}18]. Attackers [BL15b, BL16]. Attacking [YZJH12]. Attacks [BS16, BKBK14, CZ19, Che15a, CL18, CMA14, HLLG18, LJ18a, LYD^{+}18, LSG^{+}19, LÖ10, SY15, SP15, SH15, SGH15, TV12, YL17]. Attribute [CLG^{+}19, CD16, CHH^{+}19, GSW^{+}16, HSMY12, HSMY14, HBC^{+}19, LW16, WDCL18, WLH15b, WHLH16, ZZM17a, ZZM17b, ZDL^{+}17, Ver17]. Attribute-Based [CLG^{+}19, CD16, CHH^{+}19, GSW^{+}16, HSMY12, HSMY14, HBC^{+}19, LW16, WDCL18, WLH15b, ZZM17a, WHLH16, Ver17]. Auction [CZLY19, CLH^{+}14, FZCL18, KZY16, LZWY18, YHGL17, ZJH^{+}15]. Auction-Based [KZY16, LZWY18]. Auctions [FXV13, Vel10]. Audio [QF19]. Auditing [LLLW17]. Augmentation [WHSW15]. Augmentation-Based [WHSW15]. Augmented [BDL^{+}13, CCY10, CKH18, DLL^{+}13]. Authenticated [FVS17, XLM^{+}12, XGLM14, XZLW15]. Authenticating [OKG^{+}12]. Authentication [HLLC11, Jia17, KS18, MBC15, WZXL12, WZCC18, WT10]. Authority [XZLW15, ZDL^{+}17]. Authorization [LMGC17, SRD^{+}12, YKK18]. Authorized [GHY18, HTC^{+}15, LLSW16, Ma17]. Auto [TS19]. Auto-scaling [TS19]. Automata [BLS16, BE12, Che15b, Dan11, ISD15, Kap11, KV15b, PC12, YEFVJ15]. Automated [CXH14, DLM^{+}14, GLBS13, KTC^{+}11, NBN14, PBH^{+}13, Vel10]. Automatic [AFKT12, BKFP19, BPFK19, FAFD15, IDVGMP^{+}13, LfHmXjL11, MT11, PWY^{+}13, TA16a, YLLS16]. Automatically [NC16]. Automating [ET19, SMM^{+}19]. Automation [BT18, FWC13]. Automaton [GJ16, LAP11]. Autonomic [KAS13]. Autonomous [AKL^{+}19, DB15, HHV17, WYL^{+}13]. Availability [CWRZ18, LFHF14, TXJ^{+}19]. Available [ZDCZ18]. Average [KMNA^{+}16]. Avoidance [CRGM14, SM16]. AVX [GK16]. AVX2 [MKL18]. Aware [ACG^{+}11, AGP10, BZS^{+}16, Cha10a, CZL^{+}18, CLLH13, CP16, CK10, CL16, CMY17, Cor11, DJAJ15, DSTC12, Do11, GM11, GHMP18, HZWT15, JG15, KHC15, KHR^{+}19, KSPR15, LR14, LWS^{+}14, MSH^{+}11, NSRP15, RAKJ17, RASM17, RR16, RRCC^{+}15, SSY15, SCT18b, VO16, Wak17, WLW^{+}18, WCKH10, XZY^{+}10, XLX17, YDE11, YGLW15, ZZM17b, AV16, HMZ15, OB18, TYL^{+}18, WS10]. Awareness [RL11]. Axes [Whi12a]. Axial [VBVP14].

B [EDH^{+}18, KOA15, RT12]. B-spline [RT12]. B-Tree [EDH^{+}18]. Baccelli [Kon10, Pen10]. Back [Har10b]. Backbone [DE10]. Background [KS16]. Backoff [IAG^{+}14]. Backtracking [LZ19]. Backward [LCX14, XHTH13]. Bad [KCC15]. Baier [Lar10]. Balanced [KV19, XHC^{+}15]. Balancing [CMY17, LMMP16, RLTZ17, YWR^{+}14]. Ballots [CW12a]. Bandit [LV17, PANH10]. Bandwidth [CLLL17, LFHF14, LWZ^{+}18, SKK^{+}12, SLW^{+}17, WCKH10]. Bandwidth-Availability-Based [LFHF14]. Bandwidth-Based [WCKH10]. Bank [KV16]. Barrier [KSH^{+}14, NK19]. Base [ISST19]. Based [AOS^{+}15, AAA19, AFGG11, AV16, AZHASD14, AJ17, ALA19, ABS12, AÇPD11, ASS15, BL11, BWLA16, BBM17, BG15, BDC11, BÜ11, BGM^{+}13, BBKL19, BACD13, BP19, Cai12, CLL14, CCUA14, CBA18, CZLC14, CL17, CZCD18, CLG^{+}19, CLND19, CCC^{+}10, CZL^{+}18, CHDP17, Chi12, CD16, CHH^{+}19, CSS16, DE10, DH12a, DJAJ15, DHT^{+}19, DSBB19, DA18, DB13, EA17, ED09, ED10, EMB19, EMTSM18, Erg11, FEDHL16, FNP12, FFH17, FVS17, FS18, GWW^{+}13, GWWC15, GSW^{+}16, GLBS13, GK17, GDCC16, GYDX12, GY13, GJJ15, HSMY12, HSMY14, HLJ^{+}15, HPG^{+}15, HQL17, HBC^{+}19, HCZ^{+}19, HHL10, HZX15, HWS^{+}19, HLC10a, HZQ^{+}19, HuRH^{+}15, HHHC16, HH14, HP17, IEBS19, ISH13, IDVGMP^{+}13, JDAS12, JD12, JHHC15, JJO^{+}17, KS18, KHC15, KAS13, KTTRJ18, KTTRJ10, KZY16, KVX12, LMGC17, LMG^{+}18, LYY^{+}18a, LHYW12, LP14, LDLJ15, LTH^{+}15, LDZ16, LTZY16, LYPL17, LBZ19, LPL15, LSLW15]. Based [LS14, LY10, LL11b, LYL^{+}18, LZ19, LNWZ19, LCLL12, LWW13, LPD13, LFHF14, LGHD15, LW16, LCXZ16, LZWY18, LLV10, LV17, LGPRH14, LNBFPA13, LDB^{+}15, LLS17, LLF17, LLH18, Ma17, ML13, MBC15, MKN13, Meg18, Meg19, MGBD15, MK19, NSRP15, NNF19, Ni16, Nic11, NL19, ÖKA11, PABD10, PB12, PZ19, PiLCH11, PR11, PYM^{+}15, PDNH15, PYS18, PL18, Pop11, PP17, QF19, RHH12, RDZ^{+}16, RSW14, RAJ15, RRCC^{+}15, RJV13, dMRGAS18, SV15, SAKOK11, SJ18a, SBV19, SM16, SL10a, SH15, SLW^{+}17, SK18b, SJ18b, SZB19, SSK19, SKK18, SKS19, SZL15, SGH15, STBB14, TLRE11, Tan11, TPG^{+}15, TL19, TA16a, TA16b, TNWT14, TT12, TTH15, TV12, TKB11, UKW^{+}18, VBVP14, VGA15, VGA19, Wak17, Wan14, WS15, WZCC18, WT18, WDCL18, WLH15b, WCKH10, WT10, WMS^{+}12, WCW^{+}14, WHSW15, XLM^{+}14, XXW11, XGLM14]. Based [XZW^{+}17, YC11, YGFL15, YWR^{+}14, YYO15, YMWS11, YHS^{+}17, ZTBW11, ZWJ^{+}14, ZDM^{+}15, ZXZ^{+}11, ZCL^{+}12, ZCL13, ZMW16, ZZ17, ZZM17a, jZ18, ZCX^{+}16, ZVH11, ZVG16, ZYM18, ZDCZ18, ZYH^{+}19, ZSJ10, ZZZ14, ZHL15, ZDZ^{+}15b, BWR12, CWZ19, FM11, GH17, GN19, HY11, Hsu12, HHH^{+}18, IA15, KJ11, LSW10, MS14, NS16, RLVRGÁ15, SZW^{+}18, TYL^{+}18, WHLH16, WS10, WXZ^{+}12, WWJ18, XLXZ17, YWFQ18, Ver17]. Bases [Sta18]. Basis [BBP13, Bro10, FGR17, Mel13]. Bat [SZW^{+}18]. Bayesian [Cha11, GOR^{+}10, SF17]. BC [LSG^{+}19]. BCPL [Ric13]. BE [VRD10]. Bees [RLVRGÁ15, XYL^{+}11]. Before [SWLZ12]. Behavior [CLC^{+}19, CLJ^{+}22, HCZ^{+}19, SCKH18, TKB11]. Behavior-Obfuscation [CLC^{+}19, CLJ^{+}22]. Behavioral [Cao10, Cao14, GIP^{+}12a, GIP^{+}12b]. Behaviors [GAF^{+}15, LBD^{+}19, VB16]. Behaviour [WDW12]. Behaviours [RiCH10]. Belief [SBV19]. Benchmarking [Jar12, MSWI^{+}12, TU17]. Benchmarks [LPV10, WT12]. Bending [Xie11]. Better [HM16, JG15]. Between [JLS11, RSW14, SPdGPM18, ZC10, LCX14]. Betweenness [Che14]. Beyond [Roc12]. Bézier [GTS^{+}11]. BFT [CNV13]. BFT-TO [CNV13]. Bibliography [Cal11b]. Bidders [FXV13, Vel10]. Bidirectional [HC15]. Big [NPTZ16, NP16, XLX17]. Big-Data [NPTZ16]. BigFeel [FP19]. Bilinear [ASS15, IL15]. Binary [AÇPD11, FET17, LYC11, Mer13, RCS16, SSS16, SK18b, Sta18, TPV18, YTV16]. Binding [ARR^{+}16, CK10]. Bio [ABG^{+}12, VGA19]. Bio-Inspired [ABG^{+}12]. Bio-Key [VGA19]. Bioinspiration [XYL^{+}11]. Biological [Mit12, STW^{+}18]. BioMedical [AJ17]. Biometric [NGAuHQ16, YYK^{+}17]. BIP [AÇPD11]. Bipancycle [Fan10]. Bipancycle-Connectivity [Fan10]. Bipancyclicity [Fan11]. Bipartite [WHS^{+}16]. Bird [HQL17]. Birds [HQL17]. Birthmark [PiLCH11, YWFQ18]. Biswapped [CL17, LC14, XS11]. Bit [CHL14, GGZC11, JJO^{+}17, KTM19, OLL15, YLL^{+}12, KVX12]. Bit-Parallel [CHL14]. Bit-Vectors [OLL15]. Biterm [LZL^{+}19]. Bits [Sin12, YCL17]. BlackjackBench [DLM^{+}14]. Blind [BCPV11, LGPRH14, Tan11, YMWS11]. Block [GSS19, STW^{+}18, YCL17, ZX16]. Blockchain [BT18, PTOM18]. Blockchains [RM19]. Blogs [HY15]. BMC [YDHW18]. Board [EÇGK16]. Body [BY14, FFH17, KL14, KSPR15]. Body-Worn [BY14]. Bone [SBV19]. Book [Gaz10, Jas10, Joh10, Lar10, Lav12, Lev10a, Lev11a, Maj10, Mar10a, Uli11]. Boolean [AGR15, SZL15, ZZQ^{+}19]. Boosted [ÖKA11]. Boosting [FNP12, ZZM17b]. Bootstrapping [WWXH18]. both [SDN15]. Botnet [NSA15]. Bottom [BGM^{+}13]. Bottom-Up [BGM^{+}13]. Bound [RMR^{+}15a, WJ19]. Boundary [BKPS10]. Bounded [KLA^{+}15, PDNH15, QZZ18, ZYT13]. Bounding [CTIAP12, MPLDV13]. Bounds [ASCTFP16, GF17, Jia17, LJ15, PB14]. BOUQUET [MMH18]. Box [BW16, LYL^{+}18, RMP10]. Boxes [WJ19, LJ15]. BPEL [MK15, aSPW^{+}17]. Brain [VBVP14]. Branch [ZYY19]. Branches [YLC15]. Branching [GF13, WCW10]. Breakdowns [YC19]. Breaking [CLS15]. Bribery [CW12a]. Bridge [yZdZhZ18]. British [GG10]. Broadcast [FYF^{+}18, LMGC17, LMG^{+}18, MPSP17]. Broker [SBBB12]. Browser [QF19]. Browsing [YZJH12]. BT [WT12]. Bubble [WWW16, ZHW19, WW19]. Bubble-Sort [WWW16, WW19]. Bucket [CC19]. Buffering [NHC13]. Bug [ZJLC16]. Building [DFG10, HHCL10, JG15, Lav12, RMB11, SRD^{+}12]. Built [Gra12]. Bulk [BV15]. Burnt [SZL16]. Business [DBC18, LDB^{+}15, WDW12]. Butterfly [RMP^{+}16]. Bypassing [WZ17]. Bytecode [BDT10].

C [AD16, LCMC11, YWY10]. C# [PS17]. C-like [LCMC11]. C-Planarity [AD16]. C3ware [LPL14]. Cache [CP16, HLAZ15, Kha11, LGHD15, MDB^{+}18, SSS16, SY15, YC11, ZWJ^{+}14]. Cache-Sensitive [SSS16]. Caching [GRVD^{+}15, HGRV15, HLAZ15, YIUH14]. Calculation [NYT^{+}11]. Calculus [HY11, Mis14, Cao10]. Call [HLC10a]. Cambridge [Maj10, Har11]. Camera [FFH17]. Can [dRFMD^{+}17, YZLC15]. Cancellation [BBP13]. CAP [MEdJMGE^{+}19]. Capabilities [DBHC15, Lop15a]. Capability [DBC18, IA15, SDN15]. Capacitor [Mar10b]. Capacity [CP16, HZWT15, ZDCZ18]. Capsule [Kam10, Kam11a, Kam11b, Kam11c, Kam11d, Kam11e, Kam11f, Kam11g, Kam11h, Kam11i, Kam11j, Kam12a, Kam12b, Kam12c, Kam12d, Kam12e, Kam12f, Kam12g, Kam12h, Kam12i, Kam12j, Kam12k, Kam13]. Capture [BP10, ZDZ^{+}15a]. Capturing [CXH14]. Carbon [HMZ15, MSWI^{+}12, RATB^{+}13]. Carbon-aware [HMZ15]. Carlo [WL13]. Carry-Select [LSTC11]. Cartesian [SAK16, WZF18, YC14b]. Cascadable [BHAC10]. Cascade [Kot11, ZL19]. Cascaded [MZW^{+}18]. Case [Bla13, OS18, PRG^{+}10, RMGT11, SY15, YL17]. Cases [EFYS19, GB14, SHH^{+}15]. Cash [Tan11, YMWS11]. Casting [CW12a]. Categorization [CZC10, PWY^{+}13, Zam19]. Caterpillars [CFS14]. Causal [ALH17, YWDW12]. Causality [Win11]. Cayley [XZLL18, ZH19]. CBR [KAS13]. CBR-Based [KAS13]. CCA [BWLA16, CBJX19, CZLC14, GWW^{+}13, HWS^{+}19, LLPY19, LTZY16, LSLW15, PDNH15, ZYY19, ZY17]. CCA-Secure [BWLA16, CZLC14, GWW^{+}13, LTZY16]. CCA2 [LLSW16]. CCSA [NP16]. CDH [PDNH15]. CDNs [HZWT15]. Celebration [Har11]. Cell [CCUA14, VRD10]. Cells [LZL^{+}17, STW^{+}18]. Cellular [Dan11, LAP11, LZZZ13, ZYY^{+}13]. Center [IEBS19, JWCZ13, LFLJ18, PCC^{+}16, SLW^{+}17]. Centers [CCHL18, LGC19]. Centrality [Che14, LZ19, Meg19]. Centres [RATB^{+}13]. Centric [GRVD^{+}15, HGRV15, LGC19, ZVG16]. Centroid [GYDX12]. cepstral [CC11]. Cerebral [PKM18]. Certificate [BP19, CGE^{+}14, GWWC15, KSH^{+}14, LTH^{+}15, LDZ16, LLS17, WMS^{+}12]. Certificate-Based [GWWC15, LTH^{+}15, LDZ16, LLS17, WMS^{+}12]. Certificateless [GWWC15, HMS^{+}12, IL15, LSQZ17, LSQL18a, RSD19, SZS14, TCL15, WMS^{+}12, YY17, ZM18]. Certificates [HP17]. Certification [BF19, LDB^{+}15, Ver17]. Certifying [SW14]. CFD [CXF^{+}15]. Chain [WL13]. Chained [JC10]. Chains [VM14, YLW^{+}17]. Challenges [AFG^{+}17, Fra15, RJS^{+}17, WRSV12, ZJLC16]. Chan [MPP15]. Chance [SA11]. Change [BPK10, CDYC11, GK17, TSC^{+}17]. Changepoints [GOR^{+}10]. Channel [CLL14, KH10, SJ18b, TT12, Whi12a, YL17, ZYY^{+}13, ZJH^{+}15]. Channel-Recommendation [CLL14]. Channels [Cao10, Hie13, QZXR15]. Chaotic [Erg11, PC12]. Characteristics [SDN15, WT12]. Characterization [CXF^{+}15, DLM^{+}14]. Characterizations [Yan19]. Characterizing [TRY16]. Charging [LSY^{+}16]. Cheating [DD10a]. Checking [BK08, Das17, FYMY15, JLDJ19, KLA^{+}15, YL17, Lar10]. Checkpointing [RMGT11]. Checkpoints [BDL^{+}13]. Chess [Lev11b]. Children [PKM18]. Chinese [SY13]. Chip [AS11, ADML^{+}13, BHAC10, CCCS11, DAOG14, Jas10, JC10, Nur07]. Choices [HZJS17]. Chris [Mal10]. Christel [Lar10]. Churn [IK17]. CHURNs [RBNB15]. Cios [Gaz10]. Cipher [BW16, DM18, DG12, Hey17, LJ19, YCL17].

Ciphers [DJG^{+}15, Hey17, LJ16, ZH15]. Ciphertext [CHH^{+}19, JMG^{+}16, PDNH15]. Ciphertext-Policy [CHH^{+}19]. Ciphertexts [LLPY19]. Circuits [LAP11, WLZ^{+}15]. Circulant [GSRM17]. Circular [CHL14, IEBS19, LA12, LJA15, PZ19]. Circus [ZLCW14]. City [GTK^{+}19]. Civil [ET19]. Clairvoyant [SW14]. Class [BCG12, DTFT11, DTFT12, DCLN11, JZ13, KH18, NM19, SY15, TZ11]. Classes [DP16, DGV17, PS17]. Classification [CC19, CHH^{+}19, FET17, FGR17, HPG^{+}15, IK17, JYP^{+}15, JS15, LR10, PT13, SBV19, SPJA11, Yil12, ZCL^{+}12, ZSJ10, KAZ18]. Classifier [FXV13, GYDX12, JD12]. Classifiers [Tah11]. Clause [TA16a, TA16b]. Clause-Based [TA16a, TA16b]. Clauses [WJ16]. Clients [Chi16, LLPY19]. Clique [DP16]. Clique-Width [DP16]. Closing [Den12a, Kap11]. Cloud [AJA16, BBM17, BGD^{+}10, CFM17, CLL14, CLH^{+}14, CC14, CTD18, CCHL18, CL15, CL16, CMY17, DSBB19, DB15, EFV15, EV16, ETR^{+}16, FPY15, GA18, GB15, HSMY14, HLZ15, HuRH^{+}15, IJY^{+}14, KMSM15, KS18, KKMG15, KHC14, KCZJ14, Küp15, LCH16, LPL14, LLPY19, LfHmXjL11, LDLJ15, LWS^{+}14, LWZ^{+}18, LNBFPA13, MDS15, MGM12, NPTZ16, NP16, NNN^{+}14, PXG^{+}17, PCC^{+}16, RLTZ17, RMFM15, RR16, RJ18, RAJ15, SL14, TV15, URHK19, Wak17, WRSV12, XTH11, ZZQ^{+}19, ZVH11, ZVG16, ZDL^{+}17, ZDZ^{+}15b, dAEN^{+}18, HHH^{+}18, NP16]. Cloud-Based [DSBB19, KS18, LNBFPA13]. Cloud-Distributed [NPTZ16]. Clouds [AD11, Jay12, LLLW17, XLX17]. Cluster [BP19, EDH^{+}18, LWZ^{+}18, SM16, EDH^{+}18]. Cluster-Based [BP19, SM16]. Clustered [EB12, HYZ17, KS18, SJ18b]. Clustering [AGF15, AV16, CTD18, DE10, DLV10, EA17, ISST19, KRDH13, LH11, LLN^{+}15, PRJS11, PKM18, SSS12a, SVG^{+}15, ZDZ^{+}15b]. Clusters [HHV17, HZQ^{+}19, PXG^{+}17, WT12]. Co [HY15, LLZY15, MMAY19, SK18b, ZYY^{+}13, LHYW12]. Co-Channel [ZYY^{+}13]. Co-creation [HY15]. Co-independent [MMAY19]. Co-occurrence [SK18b]. Co-Occurring [LLZY15]. CO-SVC-MDC-Based [LHYW12]. Coalgebraic [CKP^{+}11]. Coalition [Ano10, LWW13]. Codd [HLL11]. Code [CCL^{+}13, DD10b, – DG13, GDKP10, LLDL17, QO17, SV15, SJ14, Sta18, Tah11, WCCL13, XHC^{+}15, ZZX16]. Code-Based [SV15]. Coded [ECL15, HZQ^{+}19, HXQ^{+}19, XHQX18]. Codes [FAFD15, KBN10, KS19, ZSL19]. Coding [JYL18, LR12, PBL14, Whi12a, WCXZ17]. Coefficient [Meg19]. Coefficient-Based [Meg19]. Coefficients [GB10]. Coercion [CW12a]. Coexistence [AZHASD14, PR11]. Cognitive [Alm19, DA18, KV16, MDN^{+}11, NB12, SJA17, Sak10, SJ18b, ZJHJ17, ZJHJ19, ZJH^{+}15]. Cohesive [FMRS17]. Collaborate [NM19]. Collaboration [HLZ^{+}17, RATB^{+}13]. Collaborative [KJR15, LPL14, MK11, ÖKA11, SNG^{+}10, STBB14, ZWC^{+}19, TYL^{+}18]. Collection [CCF11, CCC^{+}10, IDVGMP^{+}13, WBS15]. Collective [HQL17]. Collision [HHL10, SM16]. Collocative [MKW11]. Colony [ZDZ^{+}15b]. Color [BCPV11, BÜ11, FET17, KYU11, LL11b]. Color-Texture [FET17]. Colored [HWCZ16, jZ18]. Coloring [LL14]. Coloured [HJL16]. Colouring [HJP15]. Combinatorial [MMAY19]. Combined [Chi14, OJSO14, SSK12]. Combined-Semantics [Chi14]. Combining [ASCTFP16, HM13, HHCL10, JLDJ19, LZHS14, LBD^{+}19]. Comment [Ver17]. Comments [GG10, TCL15]. Commerce [LNBFPA13, AAA19].

Common [CZC10, LWC15, PiLCH11].

Communication

[ADBPLV13, AV16, Awa13, BP19, Das17, Dim13, HCZ^{+}19, KSPR15, Kon10, OKG^{+}12, QS15, RTE^{+}13, SZB15, VO16, WNNZ17]. Communication-Aware [VO16]. Communications [CL13, EMB19, GH17, HH17, LZZZ13, RSD19]. Communities [AAZ13, FMRS17, LH13, WCW^{+}14, YMS^{+}15, ZL15]. Community [CJYY17, DLL^{+}13, HBS^{+}19, Jun12, KCC15, LBD^{+}19, LLV10, RMB11, XLM^{+}14]. Community-Based [LLV10]. Community-Topic [DLL^{+}13]. Commutability [DPZ11]. Compact [BF19, EB12, LSQX19, ZMW16]. Compaction [Sin12]. Comparative [GÁVRRL16, KV16, MCT19, TKB18]. Comparing [HBDJ13, HMM11, MS11, SZL15]. Comparison [AHM15, CQL10, Do11, RCS16, STW^{+}18, XLXZ17]. Comparisons [JS15]. Compatibility [CFS13, CFS14, WDW12]. Competitive [ADBPLV13]. Compiler [JMB12, MKW11, PHB15, RCS16]. Compiling [LCMC11]. Complete [KM14, LMMP16, XZL17]. Complex [AHM15, Cro10, LBZ19, Meg16, Meg18, SY13, SSS^{+}12b, VGA15, WHYH12, WDW12, YZLC15]. Complexity [GdJ13, HHS^{+}15, HJP15, AUB11]. Compliant [YT16a]. Component [Bro10, HMM11, ML13, MV16, TLRE11]. Component-Based [ML13, TLRE11]. Component-Oriented [Bro10]. Components [EFY16, YEFVJ15]. Composing [TLRE11]. Composite [Elg15, NB17]. Composition [BZS^{+}16, NRZQ15, WXP^{+}10, ZSX10]. Compositional [HS11, YDHW18]. Compositionally [YEFVJ15]. Compositions [MK15, Mer13, YEFVJ15]. Comprehensive [RDB^{+}14a]. Compressed [JJO^{+}17]. Compression [BMG12, CC19, DB13, FNP12, GHMP18, KXS^{+}10, KBN10, LJF19, MMB13, Pop11, PH15, SY13, TS17, WGZW14]. Computable [Bla13]. Computation [Abd15, Aho12, ABL^{+}18, Bac12, Baj12, BBDF11, BE12, Buz12, Che14, Con12, Den12a, DW12, Den12b, Den12c, Fra12, FGS15, Fre12, Gel12, LLZY15, LHM^{+}15, Mit12, NSMS14, RR16, Ros12a, Ros12b, SH10, SCD15, WLH15a, Weg12, ZWC^{+}19]. Computational [Aho12, KV15a, MMAY19, NBN14, Nil10, TSK17, Tra12, WLHH18]. Computed [STW^{+}18]. Computer [Bra11, CZC10, GG10, Gra12, Ham12, HS11, KHC14, LL15, NLDH11, SM12, Trc10, BTHS12, GG10, Mal10, Mil10, Pen10]. Computers [FGG13, LPD13, Lav12]. Computing [ACW13, AKL^{+}19, AJA16, BFCRH14, BGD^{+}10, BGM^{+}11, BD16, CFM17, CL15, CCCS11, DB15, DN16, EFV15, ETR^{+}16, GA18, Gur15, HSMY14, HHCL10, HuRH^{+}15, IJY^{+}14, IJM14, JAAA^{+}17, Jar12, Jas10, JSP13, KMSM15, KHC14, KCZJ14, LHL16, MDS15, MHW10, MCT19, MGM12, NP16, Nur07, OS16, PB12, PSP14, PXG^{+}17, RMFM15, RAJ15, Ros12a, Ros12b, SMLM14, Wak17, XTH11, XZA14, YCL15, ZSX10, ZWC^{+}19, dAEN^{+}18, ÁHFE18]. computing-intensive [ÁHFE18]. CON [WGL^{+}18]. CON/SLK [WGL^{+}18]. Concentrations [LWKB15]. Concentric [PZ19]. Concept [CHDP17, DBHC15, DSZZ15, MS14, TMC15, ZDCZ18]. Concepts [PTP10]. Conceptual [SAPS19, SSS12a]. Conclusive [GdJ13]. Concurrency [YDHW18]. Concurrent [ER14, HLC10b]. Condition [LJC11, SAK16, XZL17]. Conditional [LK18, LLTY13, LSLW15, SZL16, XZLL18, ZLX^{+}19]. Conditionally [ZJ14]. Conditions [MK15]. Confidence [dMRGAS18]. Confidentiality [HLLC11].

Configurable [EFV15]. Configure [MT11]. Conflict [HFP^{+}19]. Conflicts [CZLY19]. Congestion [HCL15, LWDZ16, LR14]. Congestion-Aware [LR14]. Congruence [HJL10]. Conjunctive [Chi14]. Connected [HYZ17, ZM19]. Connection [CW12a]. Connections [SMLM14, ZL19]. Connectivity [Fan10, GHFY18, OKA17, WWW16, WZF18, XZLL18, ZHW19, ZH19, ZWFW15, WW19]. Conquer [VvdAMG17]. Consensus [BD14, YZLC15]. Consideration [Fre12]. considerations [LJWL19]. Considered [Fre12]. Considering [KCC15, SGG^{+}13]. Consistency [KLS18, MEdJMGE^{+}19]. Consistent [CL15, CFJ^{+}10]. Constant [AEHS15, KOTY17, LSQX19, ZMW16]. Constant-Round [KOTY17]. Constant-Size [AEHS15]. Constants [CW11]. Constellation [OJSO14]. Constrained [CLSV15, JMG^{+}16, KÖ14, KO15, LWC15, ZLYX10]. Constrained-Version [KÖ14, KO15]. Constraint [BBGM14, KLS18]. Constraints [BBGM14, KV16, KKPB14, QS15, SZB15, WWHL12, WS15, ZC10]. Constructed [ZH15]. Constructing [Dun11, KÖ14, KO15]. Construction [BWLA16, BPBRT16, CFJ^{+}10, EEK17, GWW^{+}13, GWWC15, KM14, KTA12, SMM^{+}19, WMS^{+}12]. Constructions [KOTY17, YLL^{+}17]. Constructive [CFJ^{+}13]. Constructs [TKM11]. Consumption [AG12, GGZC11, LNBFPA13, PHB15, RATB^{+}13]. Contact [WBS15]. Contagion [TNWT14]. Contagion-Based [TNWT14]. Contained [ZLL^{+}14]. Container [HHV17]. Containerized [TS19]. Contemporaries [Lav12]. Content [AAZ13, AGP10, AGM^{+}16, GRVD^{+}15, GLBS13, HGRV15, ÖKA11, PW12, PA15, PZPS15, PH15, SK18b, SMLM14, VBVP14, WZXL12, XLM^{+}14, ZXZ^{+}11].

Content-Based [SK18b, VBVP14, XLM^{+}14]. Content-Boosted [ÖKA11]. Content-Centric [GRVD^{+}15, HGRV15]. Content-Modelling [AAZ13]. Contention [CWCS14, PR11, ZTBW11]. Contention-Based [PR11, ZTBW11]. Context [Cha10a, CL16, DG15a, KHC15, KHR^{+}19, KS19, KBMA12, MHW10, PCLU12, RCTK18, RL11, SVP13, SSY15, Swa11, ZZM17b, ZTTM18]. Context-Adaptive [SVP13]. Context-Aware [Cha10a, CL16, KHC15, KHR^{+}19, SSY15, ZZM17b]. Context-Awareness [RL11]. Context-Driven [DG15a]. Context-Free [ZTTM18]. Contexts [SMLM14]. Contextual [WXZ^{+}12]. Continuity [PSS10]. Continuous [Dow15, EV16, NH19, Par15, Tra12, ZY17, ZYM18, ZYH^{+}19]. Continuous-Digit [Par15]. Contour [CLM16]. Contract [DGFGHZ13]. Contracting [JZ13]. Contracts [vdALM^{+}10]. Contrast [JDAS12]. Contrasting [LPP^{+}13]. Contribution [Mal10]. Contributors [Ma17]. Control [ATS15, CCUA14, Cha10b, Che15b, CHDP17, HSMY12, HBC^{+}19, HHS^{+}15, HLC10a, HCL15, HFP^{+}19, JCSZ13, KHC15, KHR^{+}19, LWW13, LLLW17, LKG10, SC10, WN11, XTH11, YDE11, YWFQ18, ZTBW11, ZLYX10, ZVH11, ZDL^{+}17, ZLG15]. Controllability [Cha10b, DH12a]. Controllable [WGL^{+}18, ZHL^{+}17]. Controlled [GTS^{+}11, WP17]. Controller [HXLX18, HXLX22]. Controllers [MT11]. Convergence [BE12, CLM16]. Conversion [GJ16]. Converter [BG15]. Convex [GF17, OJSO14, PL16]. Convivial [CvdT10]. Convolutional [jLbLzH18, TYL^{+}18]. Cooling [ZLG15]. Cooperation [NdMCdMM16, SGG^{+}13]. Cooperation-Oriented [NdMCdMM16]. Cooperative [AV16, DA18, EMB19, LHYW12, LE13, SJA17, WN11, ZYY^{+}13, ZLX^{+}15]. Coordinate [YKK18]. Coordinated [MEH19, TMOO11]. Coordination [KTTRJ10, RHF^{+}15, RFMJ10, SVP13]. CORDIC [AK12, KJ11]. Core [CXF^{+}15, EDH^{+}18, GMS^{+}12, PHM^{+}12, RTE^{+}13, XZY^{+}10, YGH^{+}14, YS15, CLL10]. Corona [QLZ18]. Corps [RMB11]. Correcting [ABS14, Yas19]. Correction [Yas19]. Correlation [LBD^{+}19, Meg19, XTH11, YCL17]. Correspondence [Mur10a]. Correspondences [WDW12]. Corrigendum [ED10, GIP^{+}12a, HXLX22, HSZS18, KO15]. Cost [ÁHFE18, DSBB19, HZWT15, IEBS19, LWS^{+}14, LGHD15, LV17, PP17, TXJ^{+}19]. Cost-Aware [HZWT15, LWS^{+}14]. Cost-driven [ÁHFE18]. Cost-Efficient [DSBB19, LGHD15]. Cost-Sensitive [LV17]. Costs [HJM12, MSWI^{+}12]. Coteries [Kuo10]. Could [Sab11]. Counter [BGM^{+}13, SPRR^{+}17]. Counter-Based [BGM^{+}13]. Countermeasures [PZPS15]. Counting [KR14, ST17]. Counts [DHW10, Mal10, MKL18]. Coupled [Erg11]. Covariance [AAH10]. Cover [ISD15, LWPZ13]. Coverage [CSS16, SP10, TU17, TMOO11, WZ17]. Coverage-Driven [SP10]. Covered [ABM12]. Covering [BBB^{+}15]. Covert [NSA15]. CPL [Ric13]. CPPC [RMGT11]. CPU [EDH^{+}18]. Crash [KSA12]. Crawlers [TU17]. CRC [Joh10]. CRC/Taylor [Joh10]. creation [HY15]. Creative [PCLU12]. Credible [ZW15]. CRESCENT [Elg15]. Crick [ZTTM18]. Crisis [GST15]. Criteria [PYS18, BHR10]. Criteria-Based [PYS18]. Critical [Cro10, RMB15, Sta18, TKB18, YWSH10, ZLCW14]. CRM [LHM^{+}15]. Cross [CCF11, DSB15, Erg11, KOA15, LHM^{+}15, MV16, OB18, PCLU12, YGFL15, ZLYX10].

Cross-Context [PCLU12]. Cross-Coupled [Erg11]. Cross-Layer [LHM^{+}15, OB18]. Cross-Network [CCF11]. Cross-Resolution [KOA15]. Cross-Site [DSB15]. Crossbar [JC10]. Crossed [CKH18]. Crowd [FZCL18, LZWY18, PL18, TKB11, WWJ18]. Crowd-outsourcing [WWJ18]. Crowd-Powered [PL18]. Crowdsourced [Alm19]. Crustal [NHMI13]. Cry [Day11]. Cryptanalysis [DG12, DJG^{+}15, LJF16, LSQL18a, LSQX19, LJ16, MV19, YCL17, YMWS11]. Cryptocurrency [RM19]. Cryptographic [RMP10, YS15]. Cryptography [LWL10, VGA19, YL17]. Cryptosystems [CLND19]. CS [LJ19]. CT [LJ18b]. cube [Fan10, Fan11]. Cubes [CKH18, CFJ^{+}13, LLF17]. Cubic [BK12a, BK14, RT12, XZL17]. Cues [HZAZ18]. Current [RJS^{+}17]. Curve [ABS12, MSTA17]. Curvelet [SK18b]. Curves [BWR12, GTS^{+}11, LL11a]. Cushion [XYL^{+}11]. Customer [HY15]. Customization [PCC^{+}16]. Cut [DA14]. Cuts [Yan19]. Cutting [CC19]. Cyber [OS18, WYL^{+}13]. Cyber-Physical [WYL^{+}13]. Cybersecurity [TK15]. Cycle [EFV15, LLF17, MKN13]. Cycle-Based [MKN13]. Cycle-sharing [EFV15]. Cycles [CL17]. Cyclic [YLLS16]. Cycling [ZLX^{+}15, HB11].

D [DB13, FGR17, GB10, Hua14, KV16, jLbLzH18, LJ15, MBRM15, PG11, ZZLL18, ZX16]. D-like [LJ15]. Daily [BY14]. Data [AJ17, AJBTT19, ABCG11, BPFK19, BdBG^{+}17, CZLY19, Cao14, CCF11, CCHL18, CLG^{+}19, CCC^{+}10, CPSK07, CHH^{+}19, DCA18, Dow15, EKOS19, ET19, EV16, ETR^{+}16, FYMY15, FPY15, GTK^{+}19, GRK13, GAFP^{+}14, HSMY14, HZHC11, HZQ^{+}19, JDAZN16, JYP^{+}15, JRC^{+}10, JWCZ13, KRDH13, Kha11, KSPR15, KLT^{+}15, Kot11, LLPY19, LJA13, LLZY15, LFLJ18, LCMC11, LHFF13, LM17, LGC19, MDY15, Ma17, MP18, MMH18, MP17, MDB^{+}18, Meg18, MDSF12, Mur10a, NTSA16, NH19, NPTZ16, NP16, NC16, ÖKA11, PB12, PSP14, PG11, PCC^{+}16, PZL12, Pyl19, QZZ18, RR16, RJ18, RATB^{+}13, RM08, SAPS19, SMM^{+}19, SJ18a, SLW^{+}17, SSS^{+}12b, SWLZ12, Tan15, WLH15a, WZCC18, WBS15, XLX17, YB16, ZVH11, ZH14, ZDL^{+}17, ZTL15, Gaz10, Lev10a]. Data-Driven [GRK13]. Data-Intensive [EV16, ETR^{+}16, RR16]. Data-Plane [MMH18]. Database [Cha10a, SC11]. Databases [AJ17, ABL^{+}18, GBA18, WP17]. Datacenters [RAKJ17]. Dataflow [WS15]. Dataflow-Based [WS15]. Datagram [HCL15]. Datasets [GRK13]. Date [WGL^{+}18]. DBM [CW11]. DCCP [RDB14b]. DCell [LFLJ18, WEFJ15]. DCF [YT11, HJL16]. DCOPs [KTTRJ18]. DCT [GB10]. DDoS [CZL^{+}18, SP15]. Deadline [CLSV15]. Deadline-Constrained [CLSV15]. Deadlock [Cha10b, Das17, YEFVJ15]. Deadlock-Free [YEFVJ15]. Dealing [VN16]. Decades [Koç10]. Decentralized [BCKM17, BDL^{+}13, CMKJ10, CD16, JRC^{+}10, MEH19, RFMJ10, SDW13, VMF^{+}14]. Decimal [KJ11]. Decision [ATA19, Ano10, DG15a, Lev10a, LV17, OS18, SS10b, Tah11, XXW11, Yil12, RM08]. Decision-Making [DG15a, SS10b]. Decisional [CCL^{+}19]. Decisions [JD12]. Decoder [PBL14]. Decomposed [VvdAMG17]. Decomposition [ACG^{+}11, FS18, KVX12, LR12, NS16]. Decoupling [HK13]. Decreasing [WM19]. Deduplicated [BBM17]. Deduplication [LDLJ15, MDY15, VB16]. Deduplication-Based [LDLJ15]. Deep [KK18, Mur10a, SBV19, TD12, VWR11, dLGCML14, KAZ18]. Defects [DD19].

Defence [Nic11]. Defense [CL18]. Defined [AFG^{+}17, Ano17, DP16, dRFMD^{+}17, GHMP18, JAAA^{+}17, LZL^{+}17, LLDL17, RJS^{+}17, RASM17, WB16]. Degradation [Kuo10]. Degree [AHM15, LPV10]. Degrees [HBDJ13]. Delay [AG12, AK12]. Delegated [MZHY15, TMC15]. Delegation [YAM^{+}15]. Deletion [TPG^{+}15]. Delimited [PK18]. Delivering [SPRR^{+}17]. Delivery [AGP10, ABCG11, LCH16]. Delta [BPK10]. Demand [CDYC11, CP16]. Demand-Aware [CP16]. Dendrograms [Bra10]. Deniable [HLLC11]. Denial [BKBK14, LÖ10, PP17]. Dense [ABH15, LZL^{+}17]. Density [ZH14]. Dependability [BCK^{+}11]. Dependable [Llo13]. Dependencies [BL15a, FSGS15, LLTY13, SKS19]. Dependency [DJAJ15]. Dependent [DB15, Dim13, PT13, WXLL18, ZHL^{+}17, GdM16, JJ18, WGL^{+}18]. depending [JJ18]. Deployed [ED09, ED10]. Deploying [Cor11, DCLN11, FT11]. Deployment [HL15, PZ18, SZB19, YHS^{+}17, ZWJ^{+}14]. Depth [AUB11]. Depth-optimal [AUB11]. Derived [HK15]. Deriving [EFYS19, Mel13, YEFVJ15]. Describing [DBC18]. Description [YT16a]. Design [AK12, BGM^{+}11, CvdT10, Cro10, DAOG14, DPZ11, EFV15, FWC13, Fra11, HS19, HLC10b, Jas10, JC10, LJA13, Lop12, MGBD15, Nur07, RTE^{+}13, SKKM15, YAQ12]. Designation [Che15a, LSQL18b]. Designing [BH10, OLF^{+}17, PSD15]. Desktop [LCH16]. Detect [LJ18b, NSA15, WOLP15]. Detecting [BSK19, BKBK14, CZ19, GN19, HCZ^{+}19, MS11, PiLCH11, SH15, TV12, WHS^{+}16, WCW^{+}14, YSC^{+}15, ZYWW13, ZL15]. Detection [ATS15, BPK10, BS16, BBK11, CDYC11, CZL^{+}18, CLC^{+}19, CLJ^{+}22, DA14, DSB15, DA18, DD19, DSZZ15, – DG13, EA17, GBBK11, GBBK14, GAFP^{+}14, HLJ^{+}15, HZAZ18, HL15, HBS^{+}19, LH13, LJA13, NSMS14, PCLU12, PP17, RCTK18, RiCH10, SDW13, SZW^{+}18, SP15, SSK19, WM19, WGS17, YB16, YWFQ18]. Detector [DSZZ15, GSAS12, RRDC^{+}18, dMRGAS18, KT18, TYL^{+}18]. Detectors [NPTZ16]. Deteriorating [WL18, WXLL18, LJWL19]. Determination [ŞLV^{+}11]. Determine [Meg18]. Determinism [HBDJ13]. Deterministic [KK18, KV15b, PS15, XLC19, EFYS19]. Determinization [BLS16]. Deuce [Wet10]. Developing [LWKB15, PRG^{+}10, TBBH18]. Development [GK17, KL14, LMA^{+}15, PW19, ZC10]. Device [LZZZ13]. Device-to-Device [LZZZ13]. Devices [GDKP10, GdM16, GMSV14, JMG^{+}16, LS17, MK13, PSS10, WLI^{+}14, WT10]. DFA [VWR11]. DHT [AOS^{+}15]. DHT-Based [AOS^{+}15]. Diagnosability [GHFY18, LFLJ18, SZL16, WWW16, ZLX^{+}19]. Diagrams [Mer13]. Dialogue [OS18]. Diameter [LPV10]. Diameters [CYTP18]. DICOM [LJ18b]. Dictionary [KS12, Pop11]. Dictionary-Based [Pop11]. Difference [XLXZ17]. Difference-Comparison-based [XLXZ17]. Differences [SR10, WXLL18]. Different [GÁVRRL16, ZL15]. Differential [BFF^{+}15, LJF16, LJ15]. Differentiated [GN10]. Diffie [Chi16, GWW^{+}13]. Diffusion [FLCT10, LBZ19, ZHL15]. Digit [Cha10a, Par15]. Digital [FGR17, MS12, MHMSGH16, Par15, Shi08, Joh10]. Digraphs [RH17, ZM19]. Dijkstra [Day11]. Dilation [RMR^{+}15a]. Dimension [GSRM17, NdMCdMM16]. Dimensional [ADML^{+}13, Dow15, LGC19]. Dimensions [NC16]. Dioxide [RATB^{+}13]. Direct [KM14, ASG15]. Directed [BBB^{+}15]. Direction [Hua14]. Directional [KLS18]. Directions [BKBK14, ZJLC16]. Directories [GVVL12]. Dirichlet [MZW^{+}18]. Disabled [PRG^{+}10]. Disabling [KKPB14]. Disaster [NRZQ15, SZB15, SZB19]. Discovering [DBHC15, Jun12]. Discovery [Gaz10, GVVL12, HHV17, LPP^{+}13, LLTY13, LBD^{+}19, LJA15, MS14, PCLU12, Suz13, VL13, VvdAMG17, WCKH10, YNP15, CPSK07]. Discrete [PG11, Xie11]. Discriminant [GY13]. Discriminating [GB10]. Discriminative [CBA18, ZDM^{+}15]. Discussant [GG10, Mal10, Mil10, Pen10]. Discussion [QS15]. Disjoint [ABH15, BKP11, BK12a, BK12b, BK14, HJK13, LC14, WS10]. Disjoint-Path [BKP11, BK12a, BK12b]. Disk [GFPC16]. Displaying [CWWK14]. Dispute [BT18]. Dissemination [LLDL17, SJ14, UKW^{+}18]. Dissortative [Meg16]. Distance [CBA18, CLW11, GRK13, GSS19, Hua14, HBS^{+}19, IDVGMP^{+}13, KMNA^{+}16, LL14, LPD13, MPLDV13, MCT19, SH10, SAK16, ZZ17]. Distance- [LL14]. Distance-Based [CBA18, IDVGMP^{+}13]. Distance-Bounding [MPLDV13]. Distinguisher [ZZ17]. Distinguishers [LJF19]. Distinguishing [EFYS19, HT15, HT16, HT17]. Distortion [Jia17]. Distributed [ACG^{+}11, AUB11, AZHASD14, BD14, BEG^{+}16, BCC^{+}19, BBGM14, BKBK14, CC14, CCHL18, CFJ^{+}10, DE10, DA14, DFG10, FZCL18, FP19, GTN10, GHXW16, GBA18, HGZ10, HT16, Hie16, HT17, HLC10a, KR14, KTTRJ10, LL15, LR14, LWW13, LZN^{+}16, MT11, MPH14, NPTZ16, PB12, PHM^{+}12, RHF^{+}15, RJV13, RMB11, SU18, SC11, WN11, XLX17, ZZX16]. Distributed-Memory [ZZX16]. Distribution [AHM15, AGM^{+}16, CGE^{+}14, CLH^{+}14, LCLL12, LWL^{+}17, Lop15b, TPG^{+}15, THP^{+}11, THP^{+}12]. Distributions [VM14]. DiVA [TEP^{+}16]. Dive [TEP^{+}16]. Divergence [FET17]. Diverse [CC14, HZJS17]. Diversity [ZTL15]. Divide [VvdAMG17]. Dividends [CW11]. Dividing [GF17]. Division [CW11]. DNA [Mar10b]. DNA/RNA [Mar10b]. DNS [HLAZ15]. DO [YLLS16]. DoA [PZ19]. DoA-Based [PZ19]. Document [GTL13, MHMSGH16, SY13]. Document-Enriched [GTL13]. Documents [ABS14, SSS12a, Thi11, WGZW14]. Does [NTSA16]. Domain [Bla13, JS15, PRJS11]. Domains [Bla13, TNWT14]. Domestic [RiCH10]. Domination [MMAY19]. Dot [SAK16]. Dot-Cartesian [SAK16]. Dot-Lexicographic [SAK16]. Double [CZLY19, ZJH^{+}15]. Down [BGM^{+}13]. Download [ZGC16]. DPPACS [RR16]. Drawings [BD16, DDLM17]. Drift [DSZZ15]. Driven [BFMT15, BS10a, DG15a, GRK13, GK17, JHHC15, LMA^{+}15, LLZY15, LHFF13, QS15, SP10, SNG^{+}10, ZSX10, dFHP^{+}11, dLGCML14, ÁHFE18]. Driver [ZDZ^{+}15a, CLLL17]. DROP [WWZ^{+}17]. Drug [NNF19]. DS [PS15]. DS-Methods [PS15]. DSC [LJ19]. DTKI [YCR16]. Dual [BWR12, HK13, HHS18, MCT19, NG17, PT13, PP17]. Dual-Form [HHS18]. Dual-Tree [MCT19]. Duality [FSGS15, ZSL19]. Due [WGL^{+}18, ZHL^{+}17]. Due-Window [ZHL^{+}17]. DuelMerge [MM17]. Durable [Elg15]. during [SVP13]. Duty [HB11, ZLX^{+}15]. Duty-cycling [HB11]. DV [LYPL17]. DV-Hop [LYPL17]. DWT [THY^{+}18]. Dynamic [ADML^{+}13, BCC^{+}19, BAFF11, BACD13, CMKJ10, CWS^{+}10, CMSML16, Dan11, GK17, HBS^{+}19, JCSZ13, Kap11, Kha11, KTTRJ10, KS16, LR10, LR12, jLbLzH18, LHM^{+}15, MSH^{+}11, NM19, RH17, SSS16, SSK12, SL10b, SLW^{+}17, SZB15, TS19, TV12, YDE11, YWFQ18, ZZZ14]. Dynamical [Nil10]. Dynamically [ASG15, Ort11, QO17]. Dynamics [YZJH12].

E-commerce [AAA19, LNBFPA13].

E-Voting [LGPRH14]. Each [YLL^{+}12]. EAR [DSTC12]. Early [Day11, HuRH^{+}15]. Earthquakes [NHMI13]. EC2 [ÁHFE18]. Economy [XYL^{+}11, BS10b, Uli11]. Ecosystems [LDB^{+}15]. Edge [BCH^{+}15, CQS13, DD19, Fan11, HWCZ16, JAAA^{+}17, JC10, LL14, TS17, WZF18, ZWC^{+}19]. Edge-Colored [HWCZ16]. Edge-Fault-Tolerant [Fan11]. Edges [HM17]. Edit [GSS19]. Editor [Jay12, Llo13, RA14, Suz13]. Editorial [Ang13, Awa13, BDMS13, BBMW13, Gel10, HXZ12, Jar12, JK12, LSQZ17, NP16, SS10a, ST17, WGS17, ZNQR15]. EDSAC [Bar11, Har11, Swa11]. Educational [AJ17]. Effect [SR10, Sta18, TXJ^{+}19]. Effective [BH10, CLS15, GN19, KRDH13, LLTY13, MS12, PK18, TPV18, WGZW14, XLC19]. Effects [WGL^{+}18, YLSL19]. Efficiency [Chi16, GTM15, HYZ17, JWCZ13, LZZZ13, MSWI^{+}12, SPdGPM18, SGG^{+}13, ZYY^{+}13]. Efficient [AS11, AGM^{+}16, BWLA16, BGD^{+}10, BBKL19, BACD13, CLS15, Cha10b, CTD18, Che14, CZCD18, CLG^{+}19, CCC^{+}10, CFJ^{+}10, CMY17, DA14, DSBB19, FP18, GWW^{+}13, GLL^{+}13, GBA18, GJJ15, HHL10, HZX15, HIDFGPC15, HL15, HLC10a, HWXD14, HLZ^{+}17, IJY^{+}14, JDAZN16, KLT^{+}15, LAP11, LJA13, LDLJ15, LSLW15, LWPZ13, LLN^{+}15, LGHD15, LSY^{+}16, LZZ^{+}17, LBIC14, MH11, MDSF12, MGZ18, NdMCdMM16, NHC13, OVGG14, OKA17, PSP14, PK18, PZZ^{+}17, PCC^{+}16, RSD19, RJ18, RM19, SJ14, SZS14, SJ18b, SHL^{+}15, TPV18, TT12, WL13, WLH15a, WT18, WDCL18, WSR11, YC14a, YDE11, YLA^{+}13, YS15, ZTBW11, ZYR^{+}13, ZX16, ZDL^{+}17, vDBvEW10, TCL15]. Efficiently [BdBG^{+}17, SLY^{+}16, WCL15]. Effort [AAA19]. EHRs [LLLW17]. EigenBots [EÇGK16]. Elderly [PRG^{+}10]. Electric [ALZ^{+}17, LSY^{+}16, WYL^{+}13]. Electricity [JG15]. Electronic [Tan11, TAC^{+}18, YMWS11]. Electrostatic [NYT^{+}11, YLSL19]. Elliptic [ABS12, MSTA17]. Embedded [CLL10, EFY16, HGZ10, HYZ17, JMB12, MSH^{+}11, MK11, PHB15, RH17, YGH^{+}14, YS15]. Embedding [DDL^{+}15, GY13, HLZ15, RMR^{+}15a]. Embeddings [LZL^{+}19, LLF17, RSW14]. Emergency [DFG10, FGG13, HLKL15, UKW^{+}18]. Emergent [Cro10]. Emerging [OS16]. Emissions [MSWI^{+}12, RATB^{+}13]. Empirical [DCLN11, JWCZ13, aSPW^{+}17, WCCL17]. EMS [ZTBW11]. EMS-MAC [ZTBW11]. Enabled [URHK19]. Enabling [JAAA^{+}17, KJR15, NH19]. Encoded [LWC15]. Encoding [FNP12, TJZF12, VBVP14, YDHW18]. Encodings [JJO^{+}17]. Encrypted [CHH^{+}19, DCA18, Lop12, ZVG16]. Encryption [BVS^{+}13, BWLA16, BWR12, Che15a, Chi12, CD16, GWWC15, GSW^{+}16, GDCC16, HLLG18, HWS^{+}19, HTC^{+}15, Jia14, LMGC17, LMG^{+}18, LLSW16, LLPY19, LTZY16, LSLW15, LSQL18b, LNWZ19, LW16, LYY^{+}18b, LLH18, MZHY15, NMS14, PDNH15, PYS18, RDZ^{+}16, SZS14, SGH15, TCL15, TMC15, TT12, WP17, WDCL18, WMS^{+}12, XY18, ZZQ^{+}19, ZYT13, ZWTM15, ZMW16, ZZM17a, ZY17, ZYM18, ZYH^{+}19, GZXA19, Wan14]. Encryptions [SLY^{+}16]. End [MK11, MHMSGH16]. End-to-End [MK11, MHMSGH16]. Endpoint [CC19]. Endpoint-Cutting [CC19]. Enduring [For12]. Energy [ACG^{+}11, AG12, AKL^{+}19, AV16, AGM^{+}16, ARVR15, BGD^{+}10, CLLH13, DA14, DSTC12, Do11, GM11, GHMP18, GLL^{+}13, GTM15, GGZC11, HYZ17, JBM^{+}19, JG15, JLS11, JWCZ13, KV15a, LZZZ13, LDLJ15, LWPZ13, LLN^{+}15, LZZ^{+}17, LBIC14, LSCG10, MSH^{+}11, MSWI^{+}12, OKA17, PHB15, PSP14, PK18, PZZ^{+}17, RATB^{+}13, SDN15, SPdGPM18, SJ18b, SHL^{+}15, TPV18, WYL^{+}13, WS10, Xie11, YDE11, ZTBW11, ZYY^{+}13, ZYR^{+}13, ZJHJ17, ZJHJ19, ZLYX10, ZNQR15]. Energy-Aware [ACG^{+}11, Do11, GM11, GHMP18, MSH^{+}11, AV16, WS10]. Energy-Constrained [ZLYX10]. Energy-Efficient [BGD^{+}10, DA14, GLL^{+}13, LWPZ13, LLN^{+}15, LBIC14, PK18, PZZ^{+}17, SHL^{+}15, YDE11, ZYR^{+}13]. Enforce [QS15]. Enforcement [Tan15]. Enforcing [WWHL12, ZVH11]. Engine [EB12]. Engineering [Awa13, BS10a, Bro10, Ham12, Jar11, JK12, LMA^{+}15, RLJ15, RMR15b, SL10a, TB10, dLGCML14]. Engineers [Har10a]. Engines [HWXD14, Lev11a, CMS10]. Enhance [CLLL17, DHT^{+}19, NB17, RL11]. Enhanced [DLL^{+}13, GHXW16, LQZ^{+}10, RHF^{+}15, SHL^{+}15, TV15]. Enhancement [DG15a, JDAS12, VB16]. Enhancements [GRVD^{+}15]. Enhancing [BDT10, ER14, IA15, WWZ^{+}17, YS15]. Enriched [GTL13]. enriching [PRJS11]. Ensemble [DSZZ15, IK17, SZW^{+}18]. Entailment [QS15]. Enterprise [HMZ15, HMH18, WRSV12, YHS^{+}17]. Enterprise-Ready [WRSV12]. Enterprises [KJR15]. Entities [CWWK14]. Entity [PWY^{+}13]. Entropy [GIP^{+}12a, GIP^{+}12b, YGFL15, ZDZ^{+}15b]. Entropy-Based [YGFL15]. enTTS [YL17]. Enumerating [YLW^{+}17]. Environment [CC14, CDYC11, CLLL17, CL15, FYF^{+}18, FT11, FP19, IJM14, KLT^{+}15, KZY16, LDLJ15, LJYL13, NNN^{+}14, PZ19, PZL12, TV15, XTH11, YHS^{+}17, ZSX10, ZZLL18]. Environments [ARVR15, BY14, DSTC12, GB15, HLZ15, HLKL15, JSP13, LfHmXjL11, RLTZ17, RAJ15, RiCH10, SZB15, SZB19, SSK19, WWJ18]. Ephemerizer [Tan15]. Epsilon [GJ16]. Epsilon-Free [GJ16]. Equality [CHH^{+}19, HTC^{+}15, LLSW16, MZHY15].

Equations [BFF^{+}15, GF13]. Equi [Ma17]. Equi-join [Ma17]. Equijoin [WP17]. Equilibrium [SPJA11]. Equivalence [Chi14, HJL10, LYL^{+}18, WDW12, ZL19]. Equivalences [Cao10]. Erasure [HZQ^{+}19, HXQ^{+}19]. Erasure-Coded [HZQ^{+}19, HXQ^{+}19]. Ergodic [Ana10]. Erratum [CLJ^{+}22, DTFT12, Ros12b, THP^{+}12]. Error [FLCT10, LJA13, Ni16, PB14, Yas19]. Error-Diffusion [FLCT10]. Errors [Cro10, LJA13, Yas19]. ESORICS [Ver17]. Essay [CXH14]. Establishment [HH17, YNN11, dAEN^{+}18, Mit19]. Estimating [GTB10, WCCL17]. Estimation [AAA19, ATA19, CMSML16, GIB12, IS13, KLL14, LPD13, Ni16, PK18, SBV19]. Estimators [Dow15]. EU [Zam19]. Evacuation [DFG10]. Evading [RCS16]. Evaluating [SZL15, ZLL^{+}14]. Evaluation [ALA19, AD11, BUB13, BBKL19, Bra11, ETR^{+}16, HBDJ13, ISH13, JMB12, KV16, LZHS14, LFLJ18, LYL^{+}18, MKN13, MDS15, MK13, SHR^{+}11, WT18, XLX17, XXW11, ZDCZ18, ZDZ^{+}15b]. Evaluations [ZM16]. Even [Fan11]. Event [ALH17, BL11, HL15, HZJS17, KH18, KW11, LHFF13, PZ19, PBH^{+}13, RSW14, WGS17, KOA15]. Event-B [KOA15]. Event-Based [RSW14]. Event-Driven [LHFF13]. Events [CWWK14, KHYC15, NNF19, SDW13, Win11, ZYF17]. Eventually [GSAS12]. Evolution [MT11, PC12, Weg12]. Evolutionary [BE12, FS18, KNHK12, SC11]. Evolved [Ric13]. Evolving [BJY11, SDW13, ZCL13]. Exact [CHL14, HLZ15, STW^{+}18]. Exactly [QLZ18]. Example [ED09, ED10]. Example-Based [ED09, ED10]. Exchange [DG15b, FVS17, WSA15, WT10, YLL^{+}17, vDBvEW10]. Exchanged [ZLX^{+}19]. Excited [Erg11]. Execution [CWS^{+}10, LLpC16, NHC13, QS15, Tim11, YWY10, YS15, ÁHFE18]. Exemplars [SLZ14]. Exercises [SPS^{+}18]. Expanding [BLS16]. Expansion [LTC^{+}15]. Expectation [CTD18]. Expected [KOTY17]. Experience [HXLX18, HXLX22, LCH16]. Experiment [TKM11, Tah11]. Experimental [WGS17]. Experiments [dRFMD^{+}17, RDB14b, RHG^{+}11, SLP11, SZL15]. Expert [YMS^{+}15]. Explicit [HP17, KLA^{+}15]. Explicit-State [KLA^{+}15]. Exploiting [RL11, SSS12a, VB16, BFF^{+}15]. Exploits [ZGC16]. Exploration [BGM^{+}11, HMH18, KLA^{+}15, MEH19, ZZLL18]. Explorator [KKBF12]. Explore [FT11]. Exploring [GIP^{+}12a, GIP^{+}12b, KLS18, YWDW12]. Exponential [AAHTH10]. Exposing [YSC^{+}15]. Exposure [BVS^{+}13]. Exposures [CZC10]. Expressing [ZV15]. Expression [GJ16, HBDJ13, MZW^{+}18]. Expressions [AGR15, KV15b, PB14, XLC19]. Expressiveness [BE12, WVGP11]. Extend [TMC15]. Extended [BCK^{+}11, BMG12, HZW^{+}14, KV15b, QLZ18, SH15, TS17]. Extending [FSMT19, dLGCML14]. Extension [OJSO14, SVS15]. Extensions [LWL10]. External [LHCN11]. Extra [WWW16, GHFY18]. Extractable [CZLC14]. Extracting [CWWK14]. Extraction [AFKT12, AHM15, BWLA16, CC11, GLBS13, NLDH11, PA15, PWY^{+}13]. Extractor [WLHH18]. Extreme [AAA19].

F5 [LLY^{+}12]. Fabric [DD19]. Face [CC11, CW12b, GB10]. Facebook [WCCL17]. Faces [HM17]. Facial [MZW^{+}18]. Facilitate [QO17]. Facilitating [KLA^{+}15, WSR11]. Factor [CLH^{+}14, CL17]. Factorization [HWS^{+}19, YAM^{+}15]. Factors [RMGT11]. Failure [CRGM14, GSAS12, dMRGAS18, WNNZ17, KT18]. Failures [Cro10, WLI^{+}14, XHQX18, YAQ12]. Fair [DG15b, PR11, PZZ^{+}17, SKK^{+}12, WSA15]. Fair-Exchange [DG15b]. Fairness [JBM^{+}19, SPdGPM18]. Fake [JLS11]. Families [HHL10, HLL11]. Family [CBJX19, DJG^{+}15, LYY^{+}18b, YCL17]. Far [dRFMD^{+}17]. Fare [IDVGMP^{+}13]. Farms [Do11, Mit10]. Fast [CLL14, CC19, GTN10, GK16, Kor11, KVX12, LH13, LK14, NYT^{+}11, VM14, XHC^{+}15, XHQX18, YTV16, YB16]. Faster [MKL18]. FastSpMM [OVGG14]. Fault [CSS16, Fan11, HZHC11, SP10, Sin12, SZL15, WZF18, WCD19, WM19, WLC^{+}19, YWR^{+}14, ZFL18, ZMSM13, ZX16, ZM19]. Fault-Based [SZL15]. Fault-Tolerance [WLC^{+}19]. Fault-Tolerant [WCD19, YWR^{+}14, ZX16]. Faults [GOR^{+}10, HWCZ16, LLF17]. Faulty [DA18, GHFY18, LYY^{+}16]. FD [dMRGAS18]. Feature [AHM15, ARR^{+}16, ATA19, BKPS10, CC11, CZL^{+}18, HPG^{+}15, JD12, JS15, MBBA16, NS16, NLDH11, PA15, ZYWW13]. Features [BS16, KYU11, LL11b, THY^{+}18, TA16b, YWDW12, ZCL^{+}12, ZTL15]. February [GG10]. FEC [TY14]. Federated [SBBB12]. Feedback [Hey17, LYPL17, PYM^{+}15, YLSL19, ZH15, ZL19]. Feedback-Based [PYM^{+}15]. FEIPS [DG15b]. Femtocell [ISST19]. Femur [SBV19]. Fewer [Cha10b, MM17]. FHE [WT18, WWXH18]. FHSD [SP15]. Fibonacci [KBN10]. Fictitious [SL10b]. Fields [PG11, YTV16]. File [GHXW16, HYZ17, LY10, WHP^{+}13, XXW11, ZGC16, ZSL19]. File-Sharing [LY10]. Files [PH15]. Filling [BWR12]. Filter [IK17, KV16]. Filtering [Cai12, HSMY14, HGRV15, KXS^{+}10, KVX12, ÖKA11]. Filtering-Based [Cai12]. Find [FSGS15]. Finding [KCC15]. Fine [KL10, ZDL^{+}17]. Fine-Grain [KL10]. Fine-Grained [ZDL^{+}17]. Finger [JHBA17]. Fingerprinting [QF19]. Fingerprints [YYK^{+}17]. Finite [EFY16, EFYS19, HT15, HWS^{+}19, KV15b, Ros14, Whi12a, YTV16]. Finite-State [EFYS19, Ros14]. First [BBDF11, Har11, Lav12, LSQZ17, LHFF13]. Fixed [JJO^{+}17, NK14]. Flash [KS19, LHCN11, MH11]. Flaw [SH15]. Flexible [ARR^{+}16, OBA16]. Floating [AAHTH10, PB14]. Floating-Point [AAHTH10, PB14]. Flock [HQL17]. Flood [DHT^{+}19]. floors [ISST19]. Flow [ATS15, HBC^{+}19, KL10, LZZZ13, SLW^{+}17, YWFQ18, ZQ13]. Flow-Level [LZZZ13]. flowshop [TXJ^{+}19]. Folded [CYTP18, YLC15]. Folksonomies [Jun12]. Foraging [XYL^{+}11]. Forbidden [DP16]. Force [YLSL19]. Forecasting [CZL^{+}18]. Forensics [MS12]. Forest [BCC^{+}19]. Form [HHS18, WVGP11]. Formal [DBHC15, LSTC11, SHH^{+}15, mAYL10, ZW15]. Formalism [Das17]. Formalization [LNWZ19, STW^{+}18]. Formalized [YCR16]. Format [BPFK19, SLP11]. Formation [DE10, HSZS17, HSZS18, LWW13]. Formats [HJL10]. Forming [FMRS17, HLZ^{+}17]. Forms [HM14, KMZ16]. Formulae [Dun11]. Formulation [PP17]. Formulations [Gur15]. Forums [HY15]. Forward [BVS^{+}13, LTH^{+}15, LCX14, NMS14, WLH15b, XHTH13]. Forward-Secure [BVS^{+}13, LTH^{+}15, NMS14, WLH15b]. Forwarding [CL18, MMH18]. Foundations [HB11, Trc10]. FOX [LJF16]. FPGA [CC19, LCMC11, YT16b]. FPGAs [Jas10, Nur07]. Fractal [HMM11]. Fractional [ZSL19]. Fractures [LJ18b]. Fragment [WVGP11]. Fragmentation [PSP14, SLW^{+}17]. Fragmentation-Based [SLW^{+}17]. Fragmentations [WWB17]. Fragments [WWB17]. Framework [AKA15, AAZ13, AHH13, BS10a, CZC10, Elg15, EFV15, FWC13, Fra11, GM11, GMS^{+}12, GBA18, yHRT^{+}12, Hsu12, HuRH^{+}15, HHHC16, KHR^{+}19, KKM^{+}15, KTA12, LV17, MGZ18, PZ18, PA15, PXG^{+}17, RMFM15, RMR15b, SRD^{+}12, SKK18, SPJA11, TAC^{+}18, URHK19, VvdAMG17, WN11, ZJ14, HHH^{+}18]. Frameworks [RMGT11]. Francis [Joh10]. Francois [Pen10]. Frank [Joh10]. Free [BPBRT16, GJ16, HJP15, IL15, LSQZ17, LSQL18a, MDY15, RDB14b, TTH15, YEFVJ15, YWSH10, YHGL17, YY17, ZM18, ZTTM18, HBS^{+}19]. Frequent [BBM17, LLZY15, MDSF12, ZCX^{+}16]. Freshness [MDB^{+}18, RBNB15]. Friendly [HS19, KCC15]. Frog [GÁVRRL16]. FSM [CSS16]. FSM-Based [CSS16]. FSMs [PS15]. Fuel [XYL^{+}11]. Full [PG11, KH18]. Fully [AFKT12, HLLC11, LMGC17, LSLW15, SGH15]. Function [AAHTH10, BBKL19, CQS13, GF17, GHY18, HLC10a, LJF19, NS16]. Functional [AKA15, BL15a, FSGS15, LLTY13, MBBA16, ZYT13, ZWTM15, SKS19]. Functions [BUB13, CCL^{+}19, SH10, SLY^{+}16, WCXZ17]. Fundamental [Bac12]. Fundamentals [Joh10, Shi08]. Funding [Zam19]. Fusion [ATA19, CZL^{+}18, JHBA17, Nic11, YYK^{+}17]. Fusions [Mis14]. Future [BKBK14, HJL10, JSP13, SSY15, WOV^{+}10, ZJLC16]. Fuzzy [AV16, ALA19, CW12b, GJ16, HXLX18, HXLX22, JDAS12, KRDH13, SVS15, WLHH18, mAYL10, jZ18, ZFZ12].

G [DTFT11, DTFT12]. G-networks [DTFT11, DTFT12]. GA [LH11]. Gait [CAV17, PKM18]. Galleries [BBM14]. Gambling [AGP10]. Game [EÇGK16, FM11, LWW13, Möl13, SF17, SKK18, SKS19, TNWT14]. Game-based [FM11]. Game-Theoretic [TNWT14]. Games [CMKJ10, CLRJ14]. Gap [HJP15]. Garbage [WLH15a]. Gates [BBKL19]. Gathering [HZHC11, KLT^{+}15, Meg18]. Gauge [NHMI13]. Gaussian [ABH15, KLL14]. GB [CHDP17]. GB-PMIPv6 [CHDP17]. GDL [KTTRJ18].

GDL-Based [KTTRJ18]. Gene [LH11]. General [Cha10b, HWXD14, KOTY17, Kuo10, LPL15, SZL15]. General-Purpose [HWXD14]. Generalization [Day11, GF17, Kot11, Pop11]. Generalized [LZHS14, LPL15, PC12, SH10, WLC^{+}19, ZHW19, ZH19]. Generalizing [PS15]. Generate [HM13]. Generated [XZLL18]. Generating [CQS13]. Generation [AGR15, FAFD15, FSMT19, ISD15, Kor11, LHCN11, LTC^{+}15, MKK15, PBH^{+}13, SP10, WCW10, ZZX16]. Generationwise [ELS11]. Generative [RG14]. Generator [Erg11, XLC19]. Generators [Ana10]. Generic [BWLA16, BBP13, Chi16, GWWC15, KCC10, SY15, TLRE11, YLL^{+}17]. Genetic [BZS^{+}16, DP13, DD10b, GA18, HM16, SC11, SKKM15, WLZ^{+}18, yZdZhZ18, ZH14]. Genome [DD10b]. Genuine [WCW^{+}14]. Geo [AJBTT19]. Geographic [SJS12]. Geolocation [FPY15]. Geometric [DDL^{+}15, HGZ10]. Geometry [BBM10, NB12, SA11, Kon10]. Geospatial [CWWK14, ZZLL18]. Geotagging [GST15]. GHZ [CCL^{+}13]. GHZ-State [CCL^{+}13]. GI [YC19]. GI/M/1/K [YC19]. GIFT [CWZ19]. GIS [TST^{+}11]. Given [CGVP15, LPV10]. Global [BBGM14, LHL16, ÖKA11, WHYH12, ZZM17b]. GML [WGZW14]. Go [dRFMD^{+}17, Sab11]. Going [Sab11]. Good [XZL17, ZLX^{+}19]. GORMANN [GV16]. GOST [LJF19]. Governance [RMFM15]. Government [ET19]. GPGPU [PBL14]. GPS [TPG^{+}15, ZWFW15]. GPS-Based [TPG^{+}15]. GPU [NPTZ16, NYT^{+}11, RT12, VO16]. GPUs [OVGG14, VGF11]. Gradient [LPD13]. Gradient-Based [LPD13]. Grain [KL10]. Grained [ZDL^{+}17]. Grammar [ZTTM18]. Grammars [Che15b, KMZ16, KKM19, ZTTM18]. Granularity [PXG^{+}17]. GRAPE [NYT^{+}11]. Graph [ATS15, CFM17, DP16, DE10, GH17, Gur15, HPG^{+}15, Meg18, MGZ18, WHS^{+}16, WWW16, YWFQ18]. Graph-based [GH17]. Graphics [LR12, NdMCdMM16]. Graphs [ABS13, BBB^{+}15, BHR10, BD16, CFS13, CFS14, CCY10, CQS13, GLK^{+}16, DDLM17, GSRM17, HM17, HWCZ16, HJP15, KMNA^{+}16, MMAY19, Meg16, QYZ19, SAK16, TRY16, WZF18, XZLL18, YC14b, ZQ13, ZHW19, ZH19, WW19]. Gravitational [HQL17]. Gravitationally [GV16]. Gray [GA18, SV15, WCCL13]. Gray-Code [WCCL13]. Green [CL13, JWCZ13, KV15a, LZL^{+}17, ZNQR15]. GreenOCR [LLN^{+}15]. Grid [KV15a, LPL15, SAKOK11, SJ18a, ZNQR15, ARVR15, IJM14, JBM^{+}19, KHC14, KCZJ14, LP14, SKK^{+}12, XLXZ17, YWSH10]. Grid-Based [LPL15, SAKOK11, SJ18a]. Grids [Cal11a, EMTSM18]. Grooming [RLVRGÁ15]. Group [ATA19, AEHS15, CLLL17, CZCD18, CHDP17, DT13, FVS17, HH17, KZY16, LWL^{+}17, ST17, SYH11, XLM^{+}12, XGLM14, XZLW15, Mit19, WLWL18]. Group-Based [CHDP17]. Grouped [HLZ^{+}17]. Grouping [OR12]. Groups [HWS^{+}19, WHS^{+}16, BLRT10]. Growing [Har10a, VBMH10]. Growth [SV15]. GSA [RJ18]. GSOS [GF13]. GSPNs [BBDF11]. GSW [WT18]. Guarantee [HL15, NH19]. Guaranteed [CSS16, LWZ^{+}18]. Guaranteeing [YWR^{+}14]. Guarding [BBM14, BCH^{+}15]. Guessing [Che15a]. Guest [BBMW13, NP16, Jay12, RA14, Suz13].

H.264 [MMB13]. Hacker [ZGC16]. Half [BBKL19]. Halftone [FLCT10]. Hamiltonian [LLF17, WL13, WEFJ15]. Hammer [Sta18]. Hamming [Sta18]. Handbook [Gon07, Mar10a]. Handelman [Tam18]. Handling [KW11, Kot11, mAYL10]. Handoff [LQZ^{+}10]. Handover [CHDP17]. Handwritten [Cha10a, GdM16]. Hankin [Mal10]. Hard [MSH^{+}11, ZWTM15, wZfG15]. Hard-to-Invert [ZWTM15]. Hardcover [Gaz10, Jas10, Joh10, Lar10, Lev10a, Maj10]. Hardness [APW11, BLRT10]. Hardware [DLM^{+}14, GAFP^{+}14, LCMC11, RMP10, WOLP15]. Harmful [Fre12]. Harmony [RKBY15]. Harrison [GG10]. Harvesting [ZJHJ17, ZJHJ19, ZGC16]. Hash [CZLC14, HHL10, HLC10a, LYY^{+}18a, LJF19, MSTA17, NS16, RMB11, ZZM17a]. Hashed [GWW^{+}13]. Hashing [LK14, THY^{+}18]. HCSP [WNNZ17]. HDH [PDNH15]. Headline [YGFL15]. Healing [THP^{+}11, THP^{+}12]. Health [BEG^{+}16, ZVG16]. Healthcare [BN14, HLKL15, MMPB10]. Heap [EEK17]. Heartbeat [IA15]. Heating [ZLG15]. Hellman [Chi16, GWW^{+}13]. Hess [HP17]. Hess-Like [HP17]. Heterogeneous [BHR10, DAOG14, GTL13, HWY11, KV19, LQZ^{+}10, MSWI^{+}12, PR11, QLZ18, SZB15, TMOO11, XHQX18, dFHP^{+}11]. Heuristic [BLRT10, EFYS19, HL15, KV19, TB11]. Heuristics [DDG^{+}15, KÖ14, KO15]. Hexagonal [YLW^{+}17]. HIBE [LSQX19]. Hidden [CLG^{+}19, XHTH13, YLL^{+}12, YT11, ZYT13, ZYY19]. Hiding [DCA18, JDAZN16, XLM^{+}12, XGLM14, XZLW15]. Hierarchical [BKP11, BK12a, BK12b, BK14, LSLW15, LJ18b, NMS14, PABD10, QYZ19, SSS12a, WYML16, ZCL13, ZMW16]. Hierarchy [Cao10, VN16, ZSL19]. High [ASG15, BGM^{+}11, CWRZ18, DN16, EB12, ECL15, GIB12, Jar12, LHFF13, LGHD15, MPSP17, MDSF12, NdMCdMM16, PW12]. High-Dimension [NdMCdMM16]. High-Performance [BGM^{+}11, DN16, ECL15, Jar12, LGHD15, EB12]. High-Priority [LHFF13]. High-Speed [ASG15, PW12, GIB12, MDSF12].

High-Throughput [MPSP17]. Higher [ZZ17]. Highly [AS11, BCC^{+}19, PBL14, ZX16, DT13]. Highway [GH17]. Hill [SJS12]. Hill-Area-Restricted [SJS12]. Hillston [BTHS12]. HISS [DT13]. Histogram [LL11b]. Histograms [ASCTFP16, MCT19]. Historical [CWWK14]. Histories [QS15]. Hit [MS14]. Hitch [CHDP17]. HMM [DA18]. HMM-Based [DA18]. Hoc [BAFF11, GH17, GGZC11, HC15, MK19, SJS12, SGG^{+}13, YWSH10, YDE11, YT11, BBM10, BSK19, WCKH10, ZYR^{+}13]. HOL [AAHTH10]. Hole [WZ17]. Holes [BSK19]. Hollow [IEBS19]. Home [MMPB10, OKA17]. Homogeneous [HWCZ16, AG12]. Homomorphic [GHY18, WT18, WCXZ17]. Honey [RLVRGÁ15]. Honeybee [XYL^{+}11]. Hop [LYPL17, YT11]. Horn [WJ16]. Horse [Sta18]. Hotness [DSBB19]. Hours [HSZS17, HSZS18]. HPC [WS15]. HSI [FSMT19]. Huang [LLSW16]. Hull [PL16]. Human [HHS^{+}15, IA15, Lev11b, SLZ14, WWHL12]. Humans [RBNB15, RB17]. Hybrid [ABCG11, CLL14, CJYY17, CP16, FET17, FYF^{+}18, GBBK14, GÁVRRL16, HH17, JYP^{+}15, KSH^{+}14, KV19, LWPZ13, LWYZ17, LSTC11, LGHD15, MEH19, MK19, NGAuHQ16, Ort11, ŞLV^{+}11, SSK19, TL19, WNNZ17, WT12, YC11, YC14b, YB16]. Hyper [CYTP18, KÖ14, KO15, YLC15]. Hyper-heuristics [KÖ14, KO15]. Hyper-Stars [CYTP18, YLC15]. Hyperbolic [AK12]. Hypercube [KSA12, WLC^{+}19]. Hypercubes [BKP11, BK12b, MRPR15, Yan19, ZLX^{+}19]. Hypergraphs [FSGS15].

I/O [AD11, DCLN11, GFPC16, LMR18, WHP^{+}13]. I/Os [XHC^{+}15]. IaaS [ETR^{+}16]. IB [CZLC14]. Iceberg [YC14a]. ID [LMGC17, TT12, TTH15, WT10, YLX^{+}11, ZCL13]. ID-Based [LMGC17, TT12, TTH15, WT10, ZCL13]. Ideas [PTP10]. Identification [BS16, CZCD18, CAV17, FLWL19, GBBK11, NPTZ16, TA16b, VGA15, YGFL15, YKK18, FFH17]. Identifying [CZ19, FXV13, PHB15]. Identity [ASS15, BWLA16, CZLC14, CLND19, Chi12, GDCC16, GJJ15, HZX15, LMG^{+}18, LTZY16, LSLW15, RDZ^{+}16, SGH15, Wan14, ZMW16, ZYM18, ZYH^{+}19]. Identity-Based [ASS15, BWLA16, CZLC14, CLND19, Chi12, GJJ15, HZX15, LMG^{+}18, LTZY16, LSLW15, RDZ^{+}16, SGH15, Wan14, ZMW16, ZYM18, ZYH^{+}19]. Ideological [WCCL17]. Idioms [ARR^{+}16]. IDS [GBBK14]. IEEE [AAHTH10, AZHASD14, HJL16, OKA17, RHF^{+}15, YT11]. IEEE-754 [AAHTH10]. IGM [CQL10]. IIDness [Cao14]. ILP [MS14]. ILP-based [MS14]. Image [AD11, BWR12, Cha10a, CLM16, CW12b, ED09, ED10, HNAS18, HZAZ18, IJM14, JDAS12, KYU11, LL11b, MBC15, MPP15, RDMRM12, SK18b, THY^{+}18, TS17, VBVP14, VGA19, WZXL12, YLL^{+}12, ZXZ^{+}11]. Images [BCPV11, FLCT10, FGR17, LJ18b, LLY^{+}12, SBV19, SY13]. Imbalanced [Kot11]. IMDS [Das17]. iMIG [LZL^{+}15]. Immune [DD10a]. Impact [Har10b, JWCZ13, RCTK18, dMRGAS18, YCL15]. impacting [RMGT11]. Impacts [LLpC16]. impairments [NK19]. Impatience [HJM12]. Implementation [AKL^{+}19, AAHTH10, BW16, CLS15, Fra11, GK16, Hie13, Hie16, KKMG15, LAP11, LYL^{+}18, LKG10, PSS10, PS17, RMP10, VBMH10, VGF11, YGH^{+}14]. Implementations [EKOS19, WT12]. Implementing [XZY^{+}10, vdALM^{+}10]. Implicit [HP17]. Importance [YL17]. Important [STBB14]. Impossible [LJF16]. Imprecision [LTL10]. Imprecision-Tolerant [LTL10]. Improve [KAS13, YZLC15]. Improved [AKL^{+}19, Chi12, KR14, KV15b, LJF16, LJF19, LYD^{+}18, LSG^{+}19, SP15, WCD19, WXZ^{+}12, KT18, Wan14]. Improvement [LYL^{+}18, LJ19, LLS17, NNN^{+}14]. Improving [Abd15, BBM17, BSK19, CLS15, Chi16, GFPC16, HXZ^{+}16, HYZ17, LSW10, MS14, RRDC^{+}18, SDN15, SWG13, SGG^{+}13]. Impulse [FET17]. Impulse-Noise [FET17]. IMS [LQZ^{+}10]. In-Kernel [GFPC16]. In-Memory [KTA12, HXQ^{+}19]. Incentive [LZWY18]. Including [SLL15]. Incomplete [HLL11, HT15, MRPR15, Mis14]. Inconsistencies [YSC^{+}15]. Incorporating [GK17]. Increasing [ELS11]. Incremental [BLS16, EFYS19, LM17, SP10]. Independent [CFJ^{+}13, HJK13, NZ14, PT13, Sin12, WW17, YC14b, YLC15, ZTL15, MMAY19]. Index [Cha11]. Indexed [AC14]. Indexes [KTA12, NHC13]. Indexing [LGC19]. Indices [CBA18]. Indirect [NB17]. Indistinguishability [FYMY15, NBN14]. Individualized [SCKH18]. Individuals [HTG12]. Indoor [ISST19, SCT18a]. Induced [DP16]. Induction [Yil12]. Infection [ZFZ12]. Infer [LH11]. Inference [ALA19, JDAS12, KKM^{+}15, LM17, Ort11, QO17, Rig14]. Inferring [KHYC15]. Infinite [DLL^{+}13, Whi12a]. Influence [CLM16, LBZ19, YCL15]. Information [ACB17, Ano10, Baj12, BP10, Cha11, GTB10, GTL13, HLJ^{+}15, HBC^{+}19, IF16, JRC^{+}10, KHR^{+}19, Lev11a, Mel13, Mur10a, Mur10b, Nic11, Roc12, RKBY15, RLJ15, SRD^{+}12, TK15, TD12, TSK17, TKB11, Uli11, YMS^{+}15, ZHL15, vDBvEW10, BS10b, CMS10]. Informative [LCXZ16]. Infrastructure [KJR15, TKB18, YAQ12]. Infrastructures [RMB15]. Inherently [KSA12]. Initialization [RDB^{+}14a]. Initiatives [GTK^{+}19]. Injection [GDKP10]. Inner [LMG^{+}18]. Innovation [Uli11, BS10b]. Innovative [SHR^{+}11]. Input [Kap11, RHH12]. Input/Output [Kap11]. Insecurity [HZX15, Wan14]. Insider [AJA16]. Insights [TBBH18]. Inspection [PW12, VWR11, XLXZ17]. Inspired [ABG^{+}12, DP13, FS18]. Instances [SW14, mAYL10]. Instantiation [DPZ11, LNWZ19, MGBD15]. Instantiations [LYY^{+}16]. Instruction [KL10]. Instructions [GK16, MKL18]. Instrumented [FT11]. Insulated [LDZ16, LLS17]. Integer [AÇPD11, CCY10, KTM19, RAJ15, WT18, ZQ13]. Integral [IEBS19, KKMG15, LJF19, PG11]. Integrated [CXF^{+}15, CL16, LfHmXjL11, RMR15b]. Integrating [AJ17, AJBTT19, TKB11]. Integration [DPZ11, FP19]. Integrity [FYMY15, MV16]. Intel [CXF^{+}15]. Intelligence [Lev11b, LLV10, PW19, PSS10, SS10a, vDBvEW10]. Intelligent [Alh19, LE13, MMPB10, TMOO11, VBBR16, WLW^{+}18]. Intensive [EV16, ETR^{+}16, RR16, ÁHFE18]. Inter [BY16, HS19, SSK12, SKS19]. Inter- [SSK12]. Inter-Activity [BY16]. Inter-actor [SKS19]. Inter-Modulo [HS19]. Inter-Subject [BY16]. Interaction [GLL^{+}13, KHYC15, SHR^{+}11]. Interactive [Fra11, LBD^{+}19, URHK19, LK18]. Interchange [SLP11]. Interconnection [BLRT10, FWC13, KMNA^{+}16, SAKOK11]. Interest [CZLY19, CQL10]. Interface [YGH^{+}14]. Interfaces [PRG^{+}10, KAZ18]. Interfacing [JYP^{+}15]. Interference [QZXR15, YDE11, ZYY^{+}13]. Interference-Aware [YDE11]. Interflow [QZXR15]. Interleaving [TY14]. International [Ano10]. Internet [CW12a, Cro10, DG15b, HZWT15, HLC10a, MDB^{+}18, NNF19, PZ18, SSK19]. Interorganizational [vdALM^{+}10]. Interplay [SPdGPM18]. Interpolation [FLCT10, RT12]. Interpretation [BDT10].

Interpretative [MKW11]. Interpreting [SVP13, TD12]. Interrogating [HLC10a]. Interrogating-Call [HLC10a]. Interval [Bla13]. Intra [SSK12]. Intra-Task [SSK12]. Introduction [AO08, Ano10, DW12, Jay12, Lev10b, Llo13, Pek12, RA14, SS10a, Suz13, Maj10]. Intrusion [CNV13, GBBK14, HLJ^{+}15, NSMS14, SZW^{+}18, SSK19]. Invariant [BÜ11, NS16]. Invariants [LWYZ17]. Inventing [Swa11]. Inversions [YTV16]. Invert [ZWTM15]. Inverted [KTA12]. Invertible [SLY^{+}16]. Inverting [DKB^{+}14]. Investigating [BY16]. Investigation [JWCZ13, ZHL^{+}17]. Invisibility [BN14]. Invited [BTHS12]. Involving [OLL15, RB17]. IoT [CLLH13, PZ19]. IP [ASG15, EB12, FEDHL16, OKA17, SP15, TJZF12, WB16]. IP-Connectivity [OKA17]. IPE [ZM16]. IPTV [CLL14]. IPv4 [NK14]. IPv6 [ECL15, LE13]. Irrational [Sta18]. Irredundant [ZLL^{+}14]. ISBN [Gaz10, Jas10, Joh10, Lar10, Lev10a, Maj10]. ISBN-13 [Gaz10, Jas10, Joh10, Lar10, Lev10a, Maj10]. ISGcloud [RMFM15]. Isolated [YS15]. ISP [ZWJ^{+}14]. Issue [Ano10, Ano17, Jay12, Llo13, Pek12, RA14, RLJ15, SS10a, Suz13, XZA14]. Issues [AFG^{+}17, AD11, FT11, LE13, PZPS15, Mit19]. Item [CZ19]. Items [CZ19, DJAJ15, SVG^{+}15]. Itemset [MDSF12]. Iterative [LCLL12, VGF11]. iTrust [CMSML16].

J [Gaz10]. Jammer [FLWL19]. Jane [BTHS12]. Japan [NHMI13]. Jari [Jas10]. Java [AFGG11, BDT10, JMB12, KW11, PiLCH11, RTE^{+}13, ZLCW14]. Java-Based [AFGG11]. Jeff [Maj10]. jInfer [KKM^{+}15]. Job [SDN15, WGL^{+}18, WLWL18]. Job-dependent [WGL^{+}18]. Jobs [LLpC16, WL18, JJ18, LJWL19]. join [Ma17]. Joint [FGS15, SA11, ZJHJ19, ZLYX10, ZJH^{+}15]. Joltik [LSG^{+}19]. Joltik-BC [LSG^{+}19]. Joost [Lar10]. Joost-Pieter [Lar10]. Journal [BTHS12, GG10, Mal10, Mil10, Pen10]. Jumping [KKM19]. Jungle [Roc12].

Kaaniche [Ver17]. KAD [CGE^{+}14]. Karhunen [BCPV11]. Katoen [Lar10]. KDM [CBJX19]. KEM [CZLC14]. Kernel [GFPC16, XZY^{+}10, XXW11, ZDM^{+}15]. Kernels [IEBS19]. Key [BN14, BVS^{+}13, Che15a, CLND19, CWZ19, Chi16, CMA14, ELS11, FVS17, GSW^{+}16, HLLG18, HH17, HWS^{+}19, HWY11, HTC^{+}15, Jia14, LLSW16, LDZ16, LTW10, LSQL18b, LCLL12, LWL^{+}17, LYY^{+}18b, LLS17, LLH18, MZHY15, MV19, PDNH15, SGH15, SLY^{+}16, TMC15, THP^{+}11, THP^{+}12, VGA19, WP17, WT10, WCXZ17, XLM^{+}12, XGLM14, XZLW15, YLL^{+}17, YL17, YNN11, ZCL13, ZY17, Mit19]. Key-Evolving [ZCL13]. Key-Insulated [LDZ16, LLS17]. Key-Policy [GSW^{+}16]. Keying [BCPV11]. Keys [ABL^{+}18, HLL11, LSQX19, ZMW16]. Keystroke [XTH11]. Keyword [Che15a, GN19, LSQL18b, WDCL18]. Khudra [CWZ19]. Kiasu [LSG^{+}19]. Kiasu-BC [LSG^{+}19]. Kind [WJ19, XZLL18]. Kinds [ZH19]. Klepto [XY18]. KNN [ZDM^{+}15]. Knowledge [DP13, Gaz10, JK12, KKBF12, LYY^{+}16, Möl13, ST17, SSS12a, WXZ^{+}12, YNP15, CPSK07]. KRAMER [STBB14]. Krzysztof [Gaz10]. Kullback [FET17]. Kurgan [Gaz10]. Kurtosis [YYO15]. KVM [LZL^{+}15].

Label [PT13]. Labeling [BKPS10, Cal11a, SAK16, ZQ13]. Labelling [Cal11b]. Ladder [Mar10b]. Laguerre [DKB^{+}14]. Landmark [WHSW15].

Language [jLbLzH18, ML13, Ort11, PH15, Sab11, SVS15, TKM11]. Languages [LCMC11, PSS10]. LANs [HM16]. Large [CL18, FFH17, KTTRJ18, LH13, LW16, LLDL17, LPV10, MDY15, WSR11, WT12, WCW^{+}14, ZHY^{+}14, ZWFW15]. Large-Scale [KTTRJ18, LH13, LLDL17, LPV10, MDY15, WSR11, WT12, WCW^{+}14, ZHY^{+}14, ZWFW15]. Laser [DB13]. Late [Day11, Wet10]. Latency [RASM17, YLLS16, ZMSM13]. Latency-Aware [RASM17]. Latency-Resistant [YLLS16]. Latent [KKBF12, LR10, XLM^{+}14]. Later [EEK17]. Lattice [LNWZ19]. Lattice-Based [LNWZ19]. Lattices [ACB17, LYY^{+}18a]. Laurent [Ver17]. Layer [DDG^{+}15, JYL18, LHM^{+}15, OB18, RDB14b, ZLYX10, HNAS18]. Layered [HFP^{+}19, IMS10, PZ18]. Layering [YYK^{+}17]. Layout [Gur15, ISST19, SSS16]. LBP [VBVP14]. Leakage [DCA18, HHS18, IL15, LTZY16, LSQZ17, SGH15, TTH15, ZYT13, ZWTM15, ZM16, ZZM17a, ZYY19, ZY17, ZYM18, ZYH^{+}19]. Leakage-Free [IL15, LSQZ17, TTH15]. Leakage-Resilient [HHS18, LTZY16, ZYT13, ZZM17a, ZY17, ZYM18, ZYH^{+}19]. Leaping [GÁVRRL16]. Learn [TA16a]. Learning [BY14, Cao14, CCUA14, CLRJ14, ISD15, JBM^{+}19, LV17, LKG10, RG14, SK18a, WGL^{+}18, WLZ^{+}18, Zam19, TXJ^{+}19]. Least [KTM19]. lecture [Kon10, BTHS12, GG10, Mal10, Mil10, Pen10]. Legacy [For12]. Leibler [FET17]. Length [LWC15, MPLDV13, PDNH15]. Less [CNV13]. Lessen [QZXR15]. Level [ATA19, CRGM14, GBBK14, yHRT^{+}12, JHBA17, LZZZ13, MDS15, MV16, TY14, ZYWW13, dAEN^{+}18, MZW^{+}18, TS19]. Levenberg [BMRS11]. Leveraging [GVVL12, SMLM14, URHK19, PBL14]. Lexicographic [SAK16]. Library [OVGG14, VBBR16, VBBR16]. Lie [HWS^{+}19]. Life [MKN13, RG14]. Life-Long [RG14]. Lifecycle [Tan15]. Lifetime [KAAE11]. Lightpath [PTWB14]. Lightweight [GMSV14]. Like [BW16, HP17, WJ19, LCMC11, LJ15, LJ16]. Limitations [MK15]. Limited [PK18, TXJ^{+}19]. Line [BÜ11, DDLM17, YMWS11, HHL10, Tan11]. line/Off [HHL10]. Linear [CCY10, CMA14, DP13, GF13, Gur15, HJK13, KH18, KXS^{+}10, LJ15, LJ16, MRPR15, RAJ15, YCL17, ZZZ14]. Linear-Time [HJK13]. Linearizability [WSY19]. Lines [FSMT19, PS17]. Linesman [Pyl19]. Lingual [Jun12]. Link [AKL^{+}19, CBA18, PCLU12, YAQ12, ZWJ^{+}14, ZHY^{+}14]. Linkable [YLA^{+}13]. Linked [AJ17, AJBTT19]. Links [ACB17]. Linux [XZY^{+}10]. Lior [Lev10a]. List [AEHS15, CGE^{+}14]. Literature [PCLU12, ZJLC16]. Live [LZL^{+}15]. Liveness [LJC11]. LMaFit [XZW^{+}17]. LMaFit-Seed [XZW^{+}17]. Load [CMY17, KV19, MK15, RLTZ17, YWR^{+}14]. Load-Balanced [KV19]. Load-Balancing [CMY17]. Local [BGM^{+}11, FET17, FMRS17, GYDX12, IAG^{+}14, LSCG10, ÖKA11, SK18b, VGA15, YWSH10, YZJH12]. Local-Minimum-Free [YWSH10]. Locality [GY13, XLX17]. Locality-Aware [XLX17]. Locality-Based [GY13]. Localization [FLWL19, HM16, HJS^{+}13, IEBS19, LYPL17, PZ18, PZ19, ZBY^{+}10]. Locating [ADBPLV13]. Location [JLS11, KTC^{+}11, LTL10, LZWY18, NH19, Ni16, OKT^{+}16, PSD15, RL11, WWJ18]. Location-based [WWJ18]. Location-Privacy [PSD15]. Locations [LWKB15]. Locomotion [WCL^{+}11]. Loève [BCPV11]. Log [YKK18]. Log-polar [YKK18]. Logging [BCKM17, FEDHL16]. Logic [HXLX18, HXLX22, JHHC15, KH18, PL18, Rig14, WLZ^{+}15]. Logic-Based [JHHC15]. Logical [Bro10, MMPB10]. LogicCrowd [PL18]. Logics [ACB17, CKP^{+}11, HY11]. Logo [SA11]. Logs [YCL15]. Loiss [DG12]. Long [Kha16, RG14]. Longer [YLC15]. Longest [LWC15]. Lookup [ASG15, EB12, ECL15]. Loop [BPBRT16]. Loop-Free [BPBRT16]. Loopless [WCW10]. Loss [GF17]. Lossless [MMB13]. Lossy [CCL^{+}19]. Low [AK12, AUB11, BF19, FEDHL16, IEBS19, RDB^{+}14a, WCKH10, WT10, wZfG15, ZMSM13]. Low-Area-Power-Delay [AK12]. Low-complexity [AUB11]. Low-Cost [IEBS19]. Low-Power [RDB^{+}14a, WT10, wZfG15]. Low-Storage [FEDHL16]. Low-Stretch [BF19]. Lower [LJ15, RMR^{+}15a]. LP [LS14]. LP-Based [LS14]. LTE [TL19]. LTE-A [TL19]. LTFs [ZYY19]. Lukasz [Gaz10]. Luo [RSD19]. LWE [XY18].

M [BV15, Ver17, YC19, BV15]. M/M/1 [BV15]. M2M [OKA17]. MAC [GH17, OB18, PR11, PA15, SM16, VN16, WCXZ17, YT11, ZTBW11]. MAC-REALM [PA15]. Machine [BY14, EFY16, For12, NL19, PXG^{+}17, PCC^{+}16, PP17, SL14, SK18a, TA16b, VMF^{+}14, WL18, WGL^{+}18, WXLL18, WLZ^{+}18, Zam19, ZHL^{+}17, LJWL19, PW19]. Machines [Do11, EFYS19, HT15, IJM14, JS15, LMR18, LZL^{+}15, TV12]. Macroscopically [HK15]. Mad [LCLL12]. Magic [KÖ14, KO15]. Mahalanobis [ZZ17]. Maimon [Lev10a]. Main [JYP^{+}15]. Maintaining [BCC^{+}19]. Maintenance [WL18, WXLL18]. Makespan [WLWL18]. Making [Ano10, DG15a, OS18, SS10b]. Malicious [BL15b, BL16, CL15, CL18, MS11, VGA15, XLXZ17]. Malicious-Resilient [CL15]. Malware [ATS15, CLC^{+}19, CLJ^{+}22, GAF^{+}15]. Management [BEG^{+}16, BKFP19, CLLH13, CP16, DP13, DFG10, Elg15, GST15, GK17, GLL^{+}13, HHV17, HuRH^{+}15, JYP^{+}15, JG15, LfHmXjL11, LZL^{+}17, LTL10, LTW10, Lop15b, Mit10, MHMSGH16, NK14, NM19, NB17, PANH10, RRDC^{+}18, Ros14, SBBB12, TMOO11, VKZ^{+}10, WYL^{+}13, ZNQR15]. Managing [Har10a, KBMA12]. Manets [FM11, AOS^{+}15, GVVL12, WS10]. Manhattan [ZX16]. Manipulation [Con12]. Many [ABH15, CCL^{+}19, CXF^{+}15, EDH^{+}18, FS18, GMS^{+}12, PHM^{+}12]. Many-Core [EDH^{+}18, GMS^{+}12, PHM^{+}12]. Many-Objective [FS18]. Manycore [LWDZ16]. Mapping [CFM17, GV16, KOA15, PCC^{+}16, VO16]. MapReduce [LLZY15, LWDZ16, LLpC16, MDY15]. Maps [ZWFW15]. Margin [Yil12]. Maritime [KS16]. Marked [CCY10]. Markers [BL15a]. Market [CZLY19]. Marking [FEDHL16]. Markov [AZHASD14, VM14, WL13, XHTH13, XXW11]. Markov-Based [AZHASD14]. Marquardt [BMRS11]. Mart [SMM^{+}19]. Massive [ABS12, GLK^{+}16, OR12, PWY^{+}13]. Matching [CFM17, CHL14, DE10, FP18, Jun12, Kha16, KS12, KVX12, LA12, Meg16, OR12, SY13, SLZ14, VWR11, WXZ^{+}12]. Matching-Based [DE10]. Materialized [ZZZ14]. Mathematical [AKA15, SZB19]. Mathematics [Ham12]. Matrix [LYY^{+}18b, OVGG14, VGF11, WWXH18, XZW^{+}17]. Maximal [HM17, Meg16, PGBFW14, Sin12, ZLL^{+}14]. Maximally [ZM19]. Maximization [CTD18, KAAE11, RASM17, Tam18, YHGL17]. Maximized [IEBS19]. Maximizing [Alm19, ABS13, HK15]. Maximum [DDG^{+}15, LPV10, Yil12]. May [Lar10]. MBST [ZYWW13]. MC [AAZ13, HIDFGPC15]. MC-2D [HIDFGPC15]. MC64 [EDH^{+}18]. MC64-Cluster [EDH^{+}18]. MD [NYT^{+}11].

MD-GRAPE-3 [NYT^{+}11]. MDC [LHYW12]. MDedup [VB16]. MDS [XHC^{+}15]. Mead [CGVP15]. Meadows [BBP13]. Mean [GYDX12, TZ11]. Mean-Based [GYDX12]. Meaningful [LTC^{+}15]. Means [BFCRH14, KRDH13]. Measure [BS16, ZZ17]. Measurement [MKK15]. Measures [GRK13]. Measuring [AHH13, BSK19, DSBB19, NC16]. Mechanical [Gra12]. Mechanism [BSK19, CLH^{+}14, CLLH13, CL18, KTTRJ10, LL15, LQZ^{+}10, LJYL13, LS17, YGLW15]. Mechanisms [KL14, LJA13, LZWY18, NSRP15, WBS15]. Media [HY15, KHYC15, SVG^{+}15, TY14, XLM^{+}14, YIUH14, YNP15]. Medical [AJ17]. Medium [STBB14, ZTBW11]. Medoids [EA17]. Meet [KH10, LJ18a, LYD^{+}18, LSG^{+}19]. Meet-in-the [LYD^{+}18]. Meet-in-the-Middle [KH10, LJ18a, LSG^{+}19]. MEI [DLL^{+}13]. Mel [CC11]. Mel- [CC11]. Mellin [CC11]. Mellin-cepstral [CC11]. Membership [GSAS12, mAYL10]. Memoization [MS14]. Memories [Whi12b]. Memory [BGM^{+}11, BMG12, CFJ^{+}10, DCLN11, Gra12, JYP^{+}15, KS19, KTA12, LWDZ16, LCMC11, LHCN11, MH11, PBH^{+}13, SSS16, VB16, ZZX16, HXQ^{+}19]. Merged [KM14]. Merging [MM17]. Mesh [BACD13, CLSV15, HH14, NSRP15, RJS^{+}17, SKK^{+}12, ZX16]. Message [FYF^{+}18, GTM15, HLLC11, Jia17, KTTRJ18, MPH14, VMF^{+}14, YGH^{+}14]. Message-Passing [VMF^{+}14, YGH^{+}14]. Messages [GST15, UKW^{+}18, YLL^{+}12]. Messaging [CQL10]. Meta [dLGCML14]. Meta-Modelling [dLGCML14]. Metadata [GLBS13, ISH13]. Metadata-Based [ISH13]. Metaheuristics [LZN^{+}16, Gon07, Mar10a]. Metamorphic [ATS15]. Metamorphosis [KV15b]. Metaphor [RCTK18]. Metaphors [Lev11b]. Meter [XLXZ17]. Method [CZ19, CZL^{+}18, DD19, FSMT19, FSGS15, GBBK14, HHS^{+}15, HC15, KTM19, LZL^{+}15, LYPL17, LSW10, LZWY18, MKK15, MDSF12, Ni16, QF19, SY15, SZW^{+}18, SP15, TSK17, WZXL12, WZCC18, WJ19, WCW^{+}14, YYO15, ZDM^{+}15, ZSX10, ZZM17b]. Methodological [CFJ^{+}10, PRG^{+}10]. Methodologies [AAA19, BBK11]. Methodology [GTL13, KOA15, NC16, PRJS11]. Methods [BGM^{+}13, BKBK14, FP19, GBBK11, PS15, TB10]. Metric [SM12, TRY16, YGFL15]. Metrics [MKK15, Meg18, TKM11, TA16b, Trc10, WXP^{+}10]. MFA [TL19]. MIBS [CWZ19]. Microblogging [GJQG14]. Microblogs [SIK14]. Microcosmic [WWC^{+}11]. Microgrid [LZZ^{+}17]. Micropayment [RM19]. Middle [KH10, LJ18a, LYD^{+}18, LSG^{+}19]. Middleware [BL11, LPL14, RTE^{+}13, dFHP^{+}11]. Migration [CK15, LZL^{+}15, NNN^{+}14, SL14, SYH11, YWR^{+}14]. Military [GTB10]. Miller [LL11a]. MILP [CWZ19]. MILP-based [CWZ19]. Min [NM19]. Min-uMax [NM19]. Mind [Lev11b]. Miniature [HWS^{+}19]. Minimal [BH10, Dun11, GAFP^{+}14]. Minimization [Chi14, GGZC11, WXLL18, WVGP11, WLWL18]. Minimize [PHB15]. Minimized [Ni16]. Minimizing [ABM12]. Minimum [BPBRT16, MRPR15, Xie11, Yan19, YWSH10]. Mining [CZ19, Gaz10, GBA18, GTL13, HY15, Lev10a, MBBA16, MDSF12, OKT^{+}16, PCLU12, PZL12, SCKH18, VvdAMG17, WZCC18, XLM^{+}14, YMS^{+}15, YNP15, ZW15, ZCX^{+}16, ZH14, ZWFW15, CPSK07, RM08]. Mins [APW11]. Misbehavior [KKPB14]. Missing [ACB17, ÖKA11]. Mission [Sta18, SNG^{+}10, dFHP^{+}11]. Mission-Critical [Sta18]. Mission-Driven [dFHP^{+}11]. Mitigating [AZHASD14, CWCS14]. Mitigation [DHT^{+}19]. Mixed [BD16, ST16, WLZ^{+}15]. MLC [JYP^{+}15]. MLH [GBBK14]. MLH-IDS [GBBK14]. Mobile [ABCG11, BCH^{+}15, CL13, CCC^{+}10, CL18, CL16, CMY17, DG15a, EOIH15, FZCL18, FT11, GVVL12, GdM16, GTM15, HB11, HK13, HLC10b, HC15, HLKL15, JAAA^{+}17, KAAE11, LWKB15, LCLL12, LS17, LZWY18, Meg18, MHW10, MK19, NK14, NRZQ15, OKT^{+}16, PL18, RHH12, RHG^{+}11, SCKH18, SZB19, SYH11, SJS12, TY14, TB11, WCKH10, WT10, ZTBW11, ZWC^{+}19]. Mobility [BDC11, GPK11, HK13, NK14, WB16]. Möbius [CFJ^{+}13]. Modal [CKP^{+}11, Möl13]. Model [Abd15, BK08, BFF^{+}15, BFMT15, BS10a, BP10, CBJX19, CCUA14, CK15, Das17, DLL^{+}13, GN19, GA18, GK17, GJJ15, HZX15, HMM11, HK13, HSZS17, HSZS18, HFP^{+}19, ISD15, IA15, JLDJ19, Kap11, KV15a, KHR^{+}19, KLA^{+}15, Lar10, LK18, LMA^{+}15, LDZ16, LZL^{+}19, LHM^{+}15, LCXZ16, LLS17, LKG10, MDS15, MK11, MPP15, MKW11, NM19, NB12, PYM^{+}15, PBH^{+}13, PTOM18, QLZ18, RSD19, RJV13, SZS14, Sin12, SLP11, SZL16, SK18a, Tra12, VBBR16, WWC^{+}11, WM19, WXP^{+}10, XHTH13, YWY10, YT16a, mAYL10, YT11, ZC10, ZX16, ZYY19, ZDCZ18, ZHL15, ZDZ^{+}15b, dLGCML14, TCL15]. Model-Based [CCUA14, RJV13, GN19, IA15]. Model-Development [ZC10]. Model-Driven [BFMT15, BS10a, GK17, LMA^{+}15, dLGCML14]. Modeled [ZM19]. Modeling [BFCRH14, BGM^{+}13, BL16, Cha11, CWRZ18, CCHL18, ISH13, IAG^{+}14, KS16, LZL^{+}19, MKN13, Mar10b, NHMI13, ZYY^{+}13]. Modelling [AAZ13, BL15b, DD10b, GB15, HJL16, Jar12, LDK11, LBZ19, PA15, RHF^{+}15, RMB15, RMR15b, SL10b, Vel10, WNNZ17, dLGCML14, DBC18, Kon10]. Models [BCK^{+}11, CLM16, DH12b, GÁVRRL16, HMS^{+}12, KMSM15, LR10, LH11, LNBFPA13, MBBA16, MEdJMGE^{+}19, OS16, Pek12, PGBFW14, SVP13, SRD^{+}12, TKB11, VN16, WDW12, XXW11, ZLCW14, TZ11]. Modern [NTSA16]. Modernizing [BFMT15]. Modification [LLSW16]. Modified [KV16, TPV18]. Modifying [WL18, ZHL^{+}17]. Modulated [MPP15]. Modulation [JYL18]. Module [OBA16]. Modules [PiLCH11]. Moduli [AJ15, BG15, HS19]. Modulo [HS19]. MoG [CLC^{+}19, CLJ^{+}22]. MOLAR [LGHD15]. Monitoring [BEG^{+}16, BDL^{+}13, CCC^{+}10, HM17, MGM12, NHMI13, SPRR^{+}17, TAC^{+}18, VKZ^{+}10, WCL15]. Monitors [Cha10b, IF16]. Monomial [Nil10]. Monotonic [ZZZ14]. Monte [WL13]. Moppet [BS10a]. Morphing [MBC15]. Morphology [IA15]. Morris [Mil10]. Most [CFS13]. Motif [FGR17, GÁVRRL16]. Motion [BY16, CDYC11, GIB12]. Motorcycle [SHR^{+}11]. MotorcycleSim [SHR^{+}11]. Movable [ACW13]. Movement [ZYR^{+}13]. Moves [MM17]. Movie [ÖKA11]. Moving [BDL^{+}13]. Mp2P [MK11]. MPEG [AAZ13, Ang13, GLBS13, HM13, YYO15]. MPEG-4 [YYO15]. MPEG-7 [AAZ13]. MPI [CRGM14, WT12]. MPI/OpenMP [WT12]. MPSoC [CK10]. MRC [BG15]. MS [VWR11]. MS-DFA [VWR11]. MSC [DH12a]. MSC-Based [DH12a]. MSCs [DH14]. Muller [WLZ^{+}15]. Multi [ABL^{+}18, ASS15, BHR10, BS16, BLRT10, CFMR14, Chi12, CLL10, DGFGHZ13, ELS11, GF17, GTS^{+}11, GA18, GBBK14, GB15, GGZC11, HM14, HNAS18, HLZ15, HFP^{+}19, ISST19, LH13, LBZ19, LR14, LV17, LZN^{+}16, LGC19, MZHY15, MZW^{+}18, MEH19, NGAuHQ16, OR12, OKG^{+}12, PT13, PB12, PW12, PXG^{+}17, RAKJ17, RLTZ17, RTE^{+}13, RG14, RA14, SL14, SU18, SCD15, TS19, Wan14, WOLP15, WHSW15, WLZ^{+}18, XZY^{+}10, YGH^{+}14, YYK^{+}17, YT11, YLLS16, jZ18, ZDL^{+}17, ZLG15, dFHP^{+}11]. Multi-Agent [CFMR14, LR14, PXG^{+}17, RA14, ZLG15, dFHP^{+}11, LBZ19]. Multi-Armed [LV17]. Multi-Authority [ZDL^{+}17]. Multi-Biometric [NGAuHQ16, YYK^{+}17]. Multi-Cloud [GB15]. Multi-Core [RTE^{+}13, XZY^{+}10, YGH^{+}14, CLL10]. Multi-criteria [BHR10]. Multi-Dimensional [LGC19]. Multi-Dividing [GF17]. Multi-floors [ISST19]. Multi-groups [BLRT10]. Multi-Hop [YT11]. Multi-Label [PT13]. Multi-layer [HNAS18]. Multi-Layered [HFP^{+}19]. Multi-Level [GBBK14, MZW^{+}18, TS19]. Multi-Machine [WLZ^{+}18]. Multi-Measure [BS16]. Multi-Objective [GA18, HLZ15, LZN^{+}16, RLTZ17, SU18, SL14]. Multi-Party [ABL^{+}18, SCD15]. Multi-Path [GTS^{+}11, GGZC11]. Multi-phase [ELS11]. Multi-pose [MZW^{+}18]. Multi-Proxy [ASS15]. Multi-Query [WHSW15]. Multi-Receiver [Wan14, Chi12]. Multi-Robot [MEH19]. Multi-Scale [LH13]. Multi-Segment [WOLP15]. Multi-Signature [ASS15]. Multi-source [PB12]. Multi-state [jZ18]. Multi-Stride [PW12]. Multi-task [RG14]. Multi-Threading [OR12, YLLS16]. Multi-Tier [RAKJ17]. Multi-Two-Party [DGFGHZ13]. Multi-User [MZHY15, OKG^{+}12]. Multi-Way [HM14]. Multi-Weight [BS16]. Multiagent [CCUA14, VRAC11]. Multicarrier [KV16]. Multicast [AUB11, LTW10, WF10]. Multicore [CWCS14, KLA^{+}15, OBA16, SPdGPM18, WT12]. Multicoupon [HIDFGPC15]. Multicriteria [DG15a]. Multidimension [AJA16]. Multihomed [HLC10b]. Multilevel [NSA15, YMS^{+}15]. Multilingual [Jun12]. Multimedia [AGP10, GIP^{+}12a, GIP^{+}12b, HLC10a, NSA15, OB18, PZPS15, PYM^{+}15, WLY^{+}15, ZW15, Zha15]. Multiobjective [ARVR15, GÁVRRL16, KNHK12, RLVRGÁ15]. Multiparty [CCL^{+}13, vdALM^{+}10]. Multipath [BAFF11, HLC10b, WS10, XS11, HCL15]. Multiple [ABCG11, DTFT11, DTFT12, DSB15, GSS19, HHL10, JDAZN16, LTC^{+}15, Ma17, OR12, PSS10, SVG^{+}15, TB11, VRD10, VWR11, WWB17]. Multiple-Collision [HHL10]. Multiple-Stride [VWR11]. Multiplication [ABS12, CW11]. Multiplicative [PB14]. Multiprocessor [NL19]. Multiprocessors [CCCS11, KW11]. Multiresolution [PABD10, SK18b]. Multiset [MSTA17]. Multisets [AC14]. Multisite [URHK19]. Multispectral [TS17]. Multithreaded [Lop13]. Multivalued [BD14]. Multivariate [CLND19, ST16, YT16b, YL17]. MultiView [URHK19]. Mum [SR10]. Mumford [Bra10]. Mutation [aSPW^{+}17]. Mutual [DLL^{+}13, WT10]. mvSERS [HLKL15].

N [Ver17, KV16]. Named [MDB^{+}18, PWY^{+}13]. Narrow [PRJS11]. Narrowing [HJP15]. NAS [WT12]. Nash [SPJA11]. Natural [Gel12, LHCN11, MMB13, PH15, Sab11, Whi12a]. Nature [Par15]. Navigational [ZV15]. NdRFT [BFCRH14]. Near [AFKT12, GTN10, TY14]. Near-Optimum [GTN10]. Near-Perfect [TY14]. Nearest [GYDX12, JD12, WCL15]. Necessary [LJC11]. Negotiation [DB15]. Neighbor [GYDX12, HLJ^{+}15, JD12, XZL17, ZLX^{+}19]. Neighborhood [GY13, KSA12, LCLL12, XLXZ17]. Neighbors [WCL15]. Neighbourhood [Dan11]. Nelder [CGVP15]. Nested [XHTH13, yZdZhZ18]. Nested-Stacking [yZdZhZ18]. Net [YHS^{+}17]. Nets [Abd15, BFF^{+}15, HJL16, LJC11, PTOM18, jZ18]. Network [AFG^{+}17, AS11, BHAC10, BSK19, BS16, BDL^{+}13, CFM17, CCF11, CGE^{+}14, CCUA14, CBA18, DJAJ15, DAOG14, FFH17, FLZC15, FYF^{+}18, GHMP18, GBBK11, GLL^{+}13, GV16, HMZ15, HNAS18, HM13, HLZ15, HXZ12, HLC10a, HMH18, JC10, JYL18, KSPR15, KZY16, LR12, LFLJ18, LQZ^{+}10, LTL10, MP18, MMH18, MDN^{+}11, MT11, Meg16, Meg18, MMPB10, MK19, NG17, NRZQ15, PBL14, RDB^{+}14a, RHH12, Sak10, SBV19, SSK19, Tam18, TST^{+}11, Tim10, VKZ^{+}10, VGA15, VKC15, WN11, WLI^{+}14, WF10, WCXZ17, WCW^{+}18, XHTH13, YZLC15, ZH14, ZL15, TYL^{+}18, SKK^{+}12]. Network-Based [RHH12]. Network-on-Chip [AS11, DAOG14]. Network-Scale [CCUA14]. Networking [CL13, CZL^{+}18, dRFMD^{+}17, GRVD^{+}15, HGRV15, MDB^{+}18, RJS^{+}17, ZHL15]. Networks [AG12, ABM12, ABG^{+}12, AKL^{+}19, AFGG11, ADML^{+}13, AHM15, Alm19, ABH15, AAH10, Ano17, BN14, BBM10, BL11, BEG^{+}16, BCC^{+}19, BMRS11, BAFF11, BPK10, BS10a, BK12a, BK14, CLSV15, CCF11, CLRJ14, CWS^{+}10, CL17, CL18, DE10, DA14, DA18, DSTC12, DLL^{+}13, EA17, ER14, ELS11, FWC13, FLWL19, GPK11, GN10, GM11, GH17, GTS^{+}11, GTB10, GTL13, GGZC11, HJS^{+}13, HLJ^{+}15, HB11, HWCZ16, HLC10b, HZHC11, Hua14, HC15, HLZ^{+}17, HH14, IAG^{+}14, Jay12, JYL18, KSA12, KL14, KNHK12, KAAE11, KXS^{+}10, KCC15, KMNA^{+}16, Koç10, KLT^{+}15, KLS18, Kon10, LH13, LH11, LHYW12, LL15, LZZZ13, Lev10b, LYPL17, LZL^{+}17, LBZ19, jLbLzH18, LWPZ13, LHM^{+}15, LZ19, LC14, LCLL12, LWW13, LLDL17, LSCG10, LPV10, MDN^{+}11, Mar10b, Meg18, Meg19, MS11, NSRP15, NSA15, NK14, NB12, OKG^{+}12, OKA17]. Networks [OB18, PB12, PR11, PYM^{+}15, PTP10, QLZ18, RDB^{+}14a, RMP^{+}16, RASM17, RG14, RL11, RKBY15, Rog11, RLVRGÁ15, SJ14, SJA17, SM12, SAKOK11, ŞLV^{+}11, SJ18a, ST17, SM16, SZL16, SLW^{+}17, SVG^{+}15, SJ18b, SYH11, SJS12, SGG^{+}13, TPG^{+}15, TL19, THP^{+}11, THP^{+}12, TB11, UKW^{+}18, VRAC11, WZ17, WOV^{+}10, WHYH12, WCL15, WEFJ15, WWW16, WB16, WCKH10, WSR11, WCW^{+}14, WLY^{+}15, XLXZ17, XS11, XZL17, YC14a, YWSH10, YDE11, YT11, YLX^{+}11, YZJH12, YNN11, ZTBW11, ZWJ^{+}14, ZYY^{+}13, ZYR^{+}13, ZLX^{+}15, ZW15, Zha15, ZJHJ17, ZZLL18, ZJHJ19, ZM19, ZLYX10, ZHY^{+}14, ZJH^{+}15, dFHP^{+}11, DTFT11, DTFT12]. Networks-On-Chip [ADML^{+}13]. Neumann [GDKP10]. Neural [BHAC10, BMRS11, BPK10, EA17, GV16, HNAS18, KNHK12, Koç10, jLbLzH18, NG17, NM19, RHH12, RG14, Tim10, Whi12b, WF10, TYL^{+}18]. Neuro [ALA19]. Neuro-Fuzzy [ALA19]. NIPSOM [VBMH10]. NMF [MP18]. NMR [AÇPD11]. No [TXJ^{+}19]. No-wait [TXJ^{+}19]. Node [AKL^{+}19, ABH15, BKP11, BK12a, CL17, DA18, LC14, OKG^{+}12, PK18, ZWJ^{+}14]. Node-Disjoint [ABH15]. Node-Link-Based [ZWJ^{+}14]. Node-Pancyclic [CL17]. Node-to-Set [BKP11, BK12a, LC14]. Nodes [ACG^{+}11, KTTRJ10, MS11, RHG^{+}11, VGA15, YT11]. Noise [FET17]. Noisy [Cao10, HZW^{+}14, YZLC15]. Non [AG12, AKA15, Ana10, BACD13, Cao14, EFYS19, ER14, GZXA19, HBDJ13, HWS^{+}19, LK18, LWYZ17, LZ19, PS15, RHH12, SKS19, SGG^{+}13, WCW10, WXP^{+}10, XHC^{+}15]. Non-abelian [HWS^{+}19]. Non-Archimedean [Ana10]. Non-Backtracking [LZ19]. Non-Cooperation [SGG^{+}13]. Non-Determinism [HBDJ13]. Non-Deterministic [PS15, EFYS19].

Non-Functional [AKA15, SKS19].

Non-homogeneous [AG12]. Non-IIDness [Cao14]. Non-interactive [LK18]. Non-MDS [XHC^{+}15]. Non-polynomial [LWYZ17]. Non-recursive [BACD13]. Non-regular [WCW10]. Non-Spatial [ER14]. Non-time [WXP^{+}10]. Non-transferable [GZXA19]. Non-uniform [RHH12]. Nonce [MPLDV13]. Nonlinear [ZL19]. Nonzero [KAAE11]. Normal [KMZ16, LJC11, PABD10, RiCH10, WVGP11]. Normalized [YGFL15]. Note [CGVP15, HWCZ16]. Noting [SDN15]. Novel [CLH^{+}14, DD19, DB13, HZJS17, KRDH13, LYY^{+}18a, LCMC11, MPP15, NM19, NC16, RR16, RATB^{+}13, RiCH10, VBVP14, VN16, WZXL12, ZZM17a]. Nuclear [AÇPD11]. null [BL15a]. Number [AJ15, Erg11, MMAY19, STW^{+}18, WJ19]. Numeral [Sta18]. Numerically [DH12b]. Nurmi [Jas10]. NVM [CP16]. NVR [AÇPD11]. NVR-BIP [AÇPD11].

O [AD11, DCLN11, GFPC16, LMR18, WHP^{+}13]. Obfuscated [ZM16]. Obfuscation [CLC^{+}19, CLJ^{+}22]. Obfuscators [PSD15]. Object [CLM16, KLL14, KS16, LTL10, PiLCH11, WSR11]. Object-Tracking [LTL10]. Objective [FS18, GA18, HLZ15, LZN^{+}16, RLTZ17, SU18, SL14]. Objects [AFKT12, BDL^{+}13, DGV17, GHXW16]. Oblivious [HSMY12]. Observations [ZC10]. Obtaining [PB14]. OCCI [YT16a]. Occurrence [GÁVRRL16, SK18b]. Occurring [LLZY15]. Oceanic [NHMI13]. OCLoptimizer [FAFD15]. Oded [Lev10a]. OFDM [LZZZ13, OJSO14]. Off [Tan11, YMWS11]. Off-Line [YMWS11, HHL10, Tan11]. Official [Küp15]. Offline [JMG^{+}16]. Offloading [ZWC^{+}19]. Offs [DDLM17, JLS11]. Offshore [SR10]. Offutt [Maj10]. Okey [EÇGK16]. OLAP [BFMT15, MH11]. On-Chip [ADML^{+}13, JC10]. On-Demand [CDYC11]. On-line [HHL10]. On-line/Off-line [HHL10]. One [ABH15, CBJX19, WCXZ17, XZLW15, YLL^{+}17]. One-Round [XZLW15, YLL^{+}17]. One-to-Many [ABH15]. One-Way [CBJX19, WCXZ17]. Online [FXV13, JMG^{+}16, LZZ^{+}17, YMS^{+}15, ZC10, ZHY^{+}14, ZHL15]. Online/Offline [JMG^{+}16]. OntCAAC [KHC15]. Onto [OJSO14]. Ontologies [AJ17, DBC18, TA16a]. Ontology [AAA19, AHH13, CvdT10, GF17, KHC15, MHW10, NNF19, SVS15, TK15, TSC^{+}17, mAYL10]. Ontology-Based [KHC15]. Ontology-Oriented [AAA19]. OP2 [GMS^{+}12]. Open [AFG^{+}17, BCH^{+}15, BY14, GTK^{+}19, KZY16, RJV13, WM19, AJBTT19]. Open-Edge [BCH^{+}15]. OpenCL [FAFD15]. Opening [Den12b, GDCC16, LLH18]. OpenMP [MKW11, WT12]. OpenStreetMap [AJBTT19]. Operating [HXZ^{+}16]. Operation [MT11, Whi12b, YLSL19]. Operational [TK15]. Operations [Ano10, GSS19, ZDZ^{+}15a]. Operators [JZ13]. Opponent [SL10b]. Opportunistic [BBM10, CCF11, MK13, WBS15]. Opportunities [WRSV12]. Opposing [SKS19]. Optical [NSRP15, PTWB14, RLVRGÁ15]. Optimal [BFCRH14, CLW11, FZCL18, FGS15, GA18, KSA12, LYC11, LLN^{+}15, PDNH15, RJ18, SHL^{+}15, TL19, WYL^{+}13, YC14b, ZZ17, yZdZhZ18, AUB11]. Optimistic [DGFGHZ13, WSA15]. Optimization [ADML^{+}13, ARVR15, BBM10, BBGM14, CFMR14, GMS^{+}12, GA18, GTM15, HGZ10, HXQ^{+}19, KNHK12, LWDZ16, LJYL13, LZN^{+}16, MK19, NM19, OJSO14, RA14, SU18, SC11, SBV19, SJ18b, WLZ^{+}15, WLZ^{+}18, XXW11, ZJHJ17, ZJHJ19, ZLYX10, ZZZ14, ZDZ^{+}15b]. Optimization-Based [SBV19, SJ18b]. Optimizations [NdMCdMM16]. Optimize [AGR15, Ort11, XYL^{+}11]. Optimized [FAFD15, HHHC16, PKM18, SJA17, dAEN^{+}18, MDN^{+}11]. Optimizing [HZQ^{+}19, Kha16, WWXH18]. Optimum [GTN10]. Options [PHB15]. OPV [HHHC16]. OPV-SLA [HHHC16]. Oracle [CBJX19, DH14]. Oracles [GSW^{+}16, GMS11, YLA^{+}13, ZYM18]. Orange4Ws [PZL12]. Orchestrations [GSS14]. Order [DCA18, KH18, WCCL13, WLZ^{+}18, ZZ17]. Order-Hiding [DCA18]. Orders [SV15]. Organization [HZJS17, LCMC11]. Organizational [PXG^{+}17]. Organized [GV16]. Organizing [DBHC15, FM11]. Orientated [TJZF12]. Oriented [AAA19, Bro10, Cal11a, HXLX18, HXLX22, MK13, MGM12, NdMCdMM16, OLF^{+}17, PZL12, MK11]. Origin [vdH15]. Orthogonal [BBM14, BZS^{+}16]. Oscillation [RDB14b]. Oscillation-Free [RDB14b]. Oscillator [Erg11]. Oscillatory [Whi12b]. OSNs [PZPS15]. Osteoporosis [SBV19]. Outerplanar [HM17]. Outlier [GBBK11, PCLU12]. Output [Kap11, SLP11]. Outsourceable [QZZ18]. Outsourced [RDZ^{+}16, ZDL^{+}17]. Outsourcing [MK11, MK13, SR10, WWJ18]. Overhead [BKFP19, WCKH10]. Overlapping [CMKJ10, JS15, WWB17]. Overtaking [EMB19]. Overview [YIUH14]. OWL [SVS15].

P2P [CLL14, EOIH15, ISH13, LY10, LHFF13, LFHF14, YLX^{+}11, ZWJ^{+}14]. PaaS [YT16a]. Package [LWKB15]. Packages [PiLCH11]. Packet [AG12, CC19, FEDHL16, GM11, HCL15, Sak10, TPG^{+}15, TY14, VWR11, ZYWW13]. Packet-Level [TY14, ZYWW13]. Packets [IS13]. Packing [PTWB14]. Page [LHCN11, VB16]. Page-Sharing [VB16]. Pages [WCCL17]. Pairing [HH14, LSQL18a, LGPRH14, YY17, ZM18]. Pairing-Based [LGPRH14]. Pairing-Free [LSQL18a, YY17, ZM18]. Pairings [ASS15, IL15]. Pairs [BBB^{+}15]. Pairwise [CFS13, CFS14, JS15]. Palsy [PKM18]. Pan [YLSL19]. Pancake [SZL16]. Pancycle [Fan10]. Pancyclic [CL17]. Pancyclicity [CKH18]. PAPR [OJSO14]. Paradigm [RATB^{+}13]. Paradox [Sta18]. Parallel [ABS12, CHL14, DN16, GLK^{+}16, GJ16, yHRT^{+}12, KK18, KTA12, OLL15, PB12, RLVRGÁ15, SHL^{+}15, VBMH10, VMF^{+}14, WHP^{+}13, WT12, ZHY^{+}14]. Parallelism [HK15]. Parallelization [NdMCdMM16, YLLS16]. Parallelized [PBL14]. Parameter [CGVP15, HBS^{+}19]. Parameter-free [HBS^{+}19]. Parameterized [SLL15]. Parameters [Gur15]. Paraphrase [TA16b]. Paravirtualization [AD11]. Parsing [SLZ14, ZTTM18]. Part [KKMG15]. Partial [CBJX19, KH18, Meg19, PS17, SHL^{+}15, YWY10]. Participatory [PZZ^{+}17]. Particle [NdMCdMM16, OJSO14, SJ18b, WLZ^{+}18]. Parties [YCR16]. Partition [GSRM17]. Partitioned [FVS17, KTA12, LAP11]. Partitioning [JC10, KAS13, MSH^{+}11, QZXR15, RR16]. Parts [DD19, FFH17]. Party [ABL^{+}18, BBKL19, DGFGHZ13, NSMS14, SCD15, ZM16]. Passage [BBDF11]. Passing [KTTRJ18, MPH14, VMF^{+}14, YGH^{+}14]. Password [FVS17, Lop15a, Lop15b]. Password-Based [FVS17]. Path [ADBPLV13, BKP11, BK12a, BK12b, DCLN11, FGS15, GTS^{+}11, KLS18, LR14, LKG10, LLF17, MBRM15, SPJA11, WVGP11, Xie11, ZW15, ZFZ12, GGZC11]. Path-Classification [SPJA11].

Path-Consistency [KLS18].

Path-Planning [MBRM15]. Paths [ABM12, ABH15, BK14, LC14, MMH18]. Patient [ZVG16]. Patient-Centric [ZVG16]. Pattern [BBM17, Cai12, CFM17, CAV17, DPZ11, DCA18, KVX12, LA12, LJA15, NPTZ16, NK14, OR12, OKT^{+}16, SK18b, VWR11, WCW^{+}14]. Pattern-Based [BBM17, WCW^{+}14]. Pattern-Matching [KVX12]. Patterns [FET17, GIP^{+}12a, GIP^{+}12b, HK13, Kha16, WOLP15]. Paul [Maj10]. Paxos [MPSP17]. Payment [DG15b]. PC [SPJA11]. PC-Nash [SPJA11]. PCM [RRCC^{+}15]. PCM-Based [RRCC^{+}15]. PeakGraph [LCXZ16]. Pedrycz [Gaz10]. Peer [GIP^{+}12a, GIP^{+}12b, LY10, YIUH14]. Peer-to-Peer [GIP^{+}12a, GIP^{+}12b, YIUH14]. Peering [KSPR15]. Peers [RMB11]. Penalties [WXLL18]. People [PRG^{+}10]. Pepa [TZ11]. Perceptions [SR10]. Perceptual [NS16, THY^{+}18]. Perfect [BKP11, BK12b, TY14]. Performability [EMTSM18]. Performability-Based [EMTSM18]. Performance [AKL^{+}19, ASCTFP16, AD11, Awa13, BBM10, BKFP19, BGM^{+}11, BS10a, BCK^{+}11, Bra11, CTIAP12, CXF^{+}15, DHT^{+}19, DN16, ECL15, EDH^{+}18, ETR^{+}16, GH17, GMS^{+}12, GFPC16, GB15, Har10b, HYZ17, yHRT^{+}12, HJL16, Jar11, Jar12, Jay12, KV16, KAS13, LZZZ13, LMR18, LJYL13, LGHD15, MK13, Mit10, MGM12, NZ14, Pek12, RMGT11, SPRR^{+}17, SWG13, WT12, WWZ^{+}17, YZLC15, ZYY^{+}13, EB12]. Performance-impacting [RMGT11]. Performance-Oriented [MGM12]. Perimeter [PL16]. Permission [VN16]. Permissions [CK15]. Permutation [LJ16]. Person [FFH17]. Personal [RSD19, Wet10]. Personalization [LNBFPA13]. Personalized [BDC11, GJQG14, HHHC16, VGA19].

Personnel [HSZS17, HSZS18]. persons [NK19]. Perspective [VRAC11, HHH^{+}18]. PerTiMo [CK15]. Pervasive [HHCL10, LSW10, MHW10, SVP13, ZSX10]. Peter [GG10]. Petri [Abd15, HJL16, PTOM18, YHS^{+}17, jZ18]. Petri-Net-Based [YHS^{+}17]. PGR [Cha10b]. Phase [DTFT11, DTFT12, ELS11]. Phenylene [YLW^{+}17]. Philosophical [JHHC15]. Phishing [EA17]. Phone [OKT^{+}16, SCKH18]. Photo [WHSW15]. Photographic [YSC^{+}15]. Phrase [KMZ16]. Phrase-Structure [KMZ16]. Physical [JYL18, NLDH11, SPS^{+}18, WYL^{+}13, YAQ12]. Physical-Layer [JYL18]. Physics [Bac12]. Pieter [Lar10]. Piggybacked [YC11]. Pixel [LTC^{+}15]. PKC [Ma17]. PKC-Based [Ma17]. PKE [GWW^{+}13, HTC^{+}15]. PKE-AET [HTC^{+}15]. PKI [YCR16]. Placement [RAKJ17, RR16, TPV18, WCW^{+}18, yZdZhZ18]. Places [BBDF11]. Plagiarism [– DG13, YWFQ18]. Plan [ISST19]. Planar [BD16, DDL^{+}15, DDLM17, ZQ13]. Planarity [AD16, RH17]. Plane [MMH18, YLL^{+}12]. Planning [LR14, MBRM15, SCT18a, SNG^{+}10]. Plant [KYU11]. Platform [AFGG11, CLL14, Mur10a]. Platforms [BHR10, EFV15, GTK^{+}19, PHB15]. Play [SL10b]. Playgrounds [URHK19]. PMC [SZL16]. PMCTrack [SPRR^{+}17]. PMIPv6 [CHDP17]. PMR [SC10]. Point [AAHTH10, ABS12, PB14, ZM16]. Point/Polynomial [ZM16]. Pointers [Lop12]. Points [PL16]. Poisoning [HLAZ15]. polar [YKK18]. Polarity [WLZ^{+}15]. Policies [BH10, JCSZ13, RRCC^{+}15, TAC^{+}18, TEP^{+}16]. Policy [CP16, CHH^{+}19, GSW^{+}16, HGRV15, KHR^{+}19, LLpC16, SL14, SL10a, SNG^{+}10, ZSX10, YC19]. Policy-Based [SL10a]. Policy-Driven [SNG^{+}10, ZSX10].

Pollutant [LWKB15]. Polygon [Xie11].

Polygons [BCH^{+}15]. Polynomial

[WT18, ZM16, LWYZ17]. Pools [WRSV12]. Population [MKL18]. Populations [HTG12]. Port [BBK11]. Portable [DLM^{+}14, DCLN11]. Portfolio [NM19]. Portrait [HCZ^{+}19]. Pose [KLL14, MZW^{+}18]. Position [PK18]. Positioning [Hua14]. positions [JJ18]. Positive [WVGP11]. Possible [Fra15, HJL10]. Potential [CMKJ10, ZW15]. Potentials [NYT^{+}11]. Power [AK12, BGM^{+}13, Cor11, Dim13, yHRT^{+}12, LBIC14, MSWI^{+}12, RDB^{+}14a, TMOO11, WLZ^{+}15, WCKH10, WT10, YAM^{+}15, ZH15, wZfG15, ZV15, ZJHJ19, ZJH^{+}15]. Power-Aware [Cor11, WCKH10]. Power-Proportional [LBIC14]. Powered [PL18]. Powerful [IF16]. pp [Gaz10, Jas10, Joh10, Lar10, Lev10a, Maj10]. Practical [FT11, HH17, JYL18, LW16, WHLH16, dLGCML14]. Practice [JRC^{+}10, Jun12, CMS10, Lev11a]. PRAM [JYP^{+}15]. Precedence [BHR10]. Precise [FEDHL16, Hie16]. Precondition [YWY10]. Predicate [LNWZ19, NMS14, ZYT13]. Predicting [LWKB15, OKT^{+}16, RHH12, SLL15, ZFZ12]. Prediction [ALH17, AKL^{+}19, CBA18, GOR^{+}10, IK17, KNHK12, ÖKA11, ZZM17b, ZHY^{+}14, ZH14]. Predictive [ALA19]. Predictors [JDAZN16]. Predistribution [ELS11]. Preface [Bra11]. Preference [LBZ19, WWJ18]. Preference-Based [LBZ19]. Preferences [CZLY19]. Prefetching [BBM17, Kha11, NTSA16]. Prefilter [RT12]. Prefix [LSTC11]. Prefix/Carry [LSTC11]. Prefix/Carry-Select [LSTC11]. Pregel [ZCX^{+}16]. PrescADE [NNF19]. Prescribed [WCW10]. Prescription [NNF19]. Presence [GOR^{+}10, Sin12, Tim11]. PRESENT [LJ16]. PRESENT-like [LJ16]. presented [GG10]. Preserved [TS17]. Preserving [EKOS19, HLLC11, IJY^{+}14, LMGC17, LSY^{+}16, LS17, LLLW17, LZWY18, NSMS14, RJ18, SCD15, SJ18a, WZCC18, ZM16, ZTL15]. Preset [HT17]. Press [Lar10, Maj10]. Prevent [HLAZ15, HHHC16]. Pricing [JBM^{+}19, LZZ^{+}17]. Principled [TB10]. Principles [BK08, Lar10]. Prior [YHGL17]. Prior-Free [YHGL17]. Priority [CCUA14, Dim13, LHFF13]. Privacy [BN14, EKOS19, HWY11, IJY^{+}14, JLS11, KKMG15, LMGC17, LSY^{+}16, LS17, LLLW17, LZWY18, NH19, NSMS14, PZPS15, PSD15, RJ18, RL11, SCD15, SJ18a, SWLZ12, Wak17, WZCC18, YYK^{+}17, ZM16, ZHL15, ZTL15]. Privacy-Aware [Wak17]. Privacy-Preserving [EKOS19, IJY^{+}14, LMGC17, LSY^{+}16, LS17, LLLW17, LZWY18, NSMS14, RJ18, SCD15, WZCC18, ZM16, ZTL15]. Private [BBKL19, Jia14, LSQX19, MV19, ST17, TSK17, ZMW16]. Private-Key [MV19]. Probabilistic [EA17, GRVD^{+}15, GTB10, PC12, RHF^{+}15, Rig14, WP17]. Probability [ACB17, Pek12, VM14]. Probably [DHW10, Mal10]. Probe [BS16]. Probing [WBS15]. Problem [AZHASD14, APW11, BK12b, Cal11a, Cal11b, CKH18, CL15, CK10, DH14, DDG^{+}15, HWS^{+}19, KS12, LWPZ13, NG17, NdMCdMM16, PTWB14, SU18, TMOO11, KT18]. Problems [DH12a, Fra15, KRDH13, Kot11, LZN^{+}16, NM19, OLL15, Tra12]. Procedure [Day11]. Procedures [RCS16, vdH15]. Process [DH12b, Fra12, HTG12, IAG^{+}14, LKG10, MZW^{+}18, NNN^{+}14, OLF^{+}17, PGBFW14, QS15, dMRGAS18, RMR15b, VvdAMG17, WDW12]. Process-Driven [QS15]. Processes [Cha10b, DBC18, KM14, KLL14, LTW10, vdALM^{+}10]. Processing [Ano10, BL11, FP19, HS19, LR12, NdMCdMM16, Pyl19, SCD15, WGL^{+}18, YC14a, ZHL^{+}17, JJ18, WLWL18]. Processor [Nur07, XZY^{+}10, YT16b, YS15, Jas10, VRD10]. Processors [HWXD14, OBA16, SPdGPM18]. Product [AK12, FSMT19, HM14, OVGG14, PS17, QLZ18, SAK16, WZF18, YC14b]. Product-Forms [HM14]. Products [LMG^{+}18]. Prof. [BTHS12, Kon10]. Professional [BT18]. Profile [AGF15, Hsu12]. Profiles [GSS14, YLSL19]. Profiling [HK13, PKM18]. Profit [Mit10]. Prognostics [BEG^{+}16]. Program [DHW10, Har11, HY11, Mal10, YLLS16]. Programmer [Bar11]. Programming [AAA19, AKA15, AFGG11, AÇPD11, CCY10, DP13, GA18, Gur15, HGZ10, KTTRJ10, LLDL17, PL18, RAJ15, SZB19, TLRE11, Wet10, ZH14]. Programming-Based [SZB19]. Programs [MKW11, Rig14, aSPW^{+}17, YWY10, ZLCW14]. Progressive [DG15a, ZZZ14]. Project [HSZS17, HSZS18, Zam19]. Projection [OJSO14, WHS^{+}16]. Proof [CZLC14, JLDJ19, LYY^{+}18a, LYY^{+}16, ZZM17a]. Proof-of-Knowledge [LYY^{+}16]. Proofs [NBN14, RB17]. propagating [GDKP10]. Propagation [AD11, Ni16, WWC^{+}11, ZFZ12]. Properties [ABG^{+}12, CL17, MMAY19, MGM12, OLL15, QYZ19, SAKOK11, WEFJ15]. Property [BPBRT16]. Proportional [LBIC14]. Prospects [PANH10]. Protected [NSRP15]. Protecting [LLPY19]. Protection [Lop12, Lop13, LÖ10, NGAuHQ16]. Protocol [AOS^{+}15, ABCG11, BBKL19, CL15, DSTC12, DGFGHZ13, GH17, HLC10a, HH14, KSA12, KSPR15, LLN^{+}15, LYY^{+}16, MBC15, PR11, SM16, SJS12, SWG13, TB10, WCKH10, WT10, YC11, YT11, YNN11, ZTBW11, ZXZ^{+}11, ZYR^{+}13, ZJHJ17, Mit19, OB18, HCL15]. Protocols [CZCD18, Fra15, HLLC11, LWL^{+}17, MV19, MPLDV13, NSMS14, RB17, WOV^{+}10]. Prototype [KV16]. Provably [BCKM17, IL15, WMS^{+}12]. Proven [Har10a]. Provider [HHHC16, HHH^{+}18]. Provider-Based [HHHC16]. Providers [LWS^{+}14]. Providing [KS18]. Proving [WSY19]. Provisioning [HZWT15, LCH16, NSRP15, SPJA11, YT16a, ÁHFE18]. Proxy [ASS15, DHT^{+}19, GSW^{+}16, GJJ15, GZXA19, HCZ^{+}19, HZX15, LK18, LSLW15, MBC15, Tan11, WYML16, YMWS11]. PRS [GLK^{+}16]. PRSDs [BMG12]. Pruning [LLTY13, STW^{+}18, YLC15]. PS [Hua14]. Pseudoknots [SLL15]. Pseudorandom [Ana10]. PSO [LH11, NdMCdMM16, TPV18]. Psychological [NC16]. Public [BVS^{+}13, Che15a, CLND19, EKOS19, ET19, ETR^{+}16, HWS^{+}19, HTC^{+}15, LLSW16, LSQL18b, LLH18, MZHY15, PDNH15, SLY^{+}16, TMC15, TT12, WP17, YL17, ZY17]. Public-Key [BVS^{+}13, LLH18, PDNH15, ZY17]. Publication [CMSML16, ZTL15]. Publishing [MP18]. Pull [YC11]. Pull-Based [YC11]. Pulse [Erg11]. Pulse-Excited [Erg11]. Pupil [IEBS19]. Pure [KKM19]. Purpose [HWXD14]. Push [GTM15, YC11]. Pushdown [KK18]. Pyramids [WW17].

Q [CCUA14]. Q-Learning [CCUA14]. QARMA [LJ18a]. QARMA-64 [LJ18a]. QARMA-64/128 [LJ18a]. QoS [BZS^{+}16, GN10, KSPR15, KCZJ14, LQZ^{+}10, MPH14, OB18, SPJA11]. QoS-Aware [BZS^{+}16, KSPR15, OB18]. QoS-Security [MPH14]. QPRR [KSPR15]. Quadratic [KRDH13]. Qualities [TU17]. Quality [AKL^{+}19, CXH14, GTB10, GLL^{+}13, HL15, KCZJ14, LCH16, LMA^{+}15, THY^{+}18, VKZ^{+}10]. Quantified [HSZS17, HSZS18]. Quantify [HS11].

Quantitative [BL15b, BL16, GRK13]. Quantum [CCL^{+}13, DP13, Mel13]. Quantum-Inspired [DP13]. Quasi [BD16, DDL^{+}15, PZ18, PWY^{+}13]. Quasi-Automatic [PWY^{+}13]. Quasi-Planar [DDL^{+}15]. Quasi-Upward [BD16]. Quaternion [HZAZ18]. Queries [Chi14, ER14, MH11, SC11, SCT18a, SCT18b, TST^{+}11, WVGP11, ZZQ^{+}19, ZV15, ZLL^{+}14, ZZZ14]. Query [CCC^{+}10, DCA18, HLC10a, LLZY15, MGZ18, SCD15, TSC^{+}17, WHSW15, WWJ18, XLX17, YC14a, YCL15]. Query-Based [CCC^{+}10]. Query-Driven [LLZY15]. Querying [AJBTT19, AAH10, BdBG^{+}17]. Question [CJYY17, Fre12]. Queue [ASCTFP16, BBDF11, BV15, Dim13]. Queueing [IAG^{+}14, YC19]. Queuing [HJM12]. Quicksort [GK16]. Quorums [Kuo10].

R [LJF19]. RAC [DDG^{+}15]. Race

[GAFP^{+}14, YB16]. Radar [Pyl19]. Radial [SAKOK11]. Radio [Alm19, DA18, KV16, KTC^{+}11, NB12, RASM17, SJA17, SJ18b, ZJHJ17, ZJHJ19, ZJH^{+}15]. Radiosity [PABD10]. Radius [YWSH10]. RAID [LDK11, WWZ^{+}17, XHC^{+}15, XHQX18]. RAID-6 [XHC^{+}15]. RAID-Coded [XHQX18]. Rainforest [KTC^{+}11]. Rallying [Day11]. RAM [RRDC^{+}18]. Ran [Har11]. Random [BHAC10, BMRS11, CBJX19, Erg11, GSW^{+}16, GN10, GHXW16, GMS11, Koç10, LPL15, LZ19, PZ18, PBL14, Tim10, YLA^{+}13, ZYM18, ZFZ12]. Randomized [KAS13]. Randomness [BWLA16]. Range [ABM12, BÜ11, DB13, DCA18, MH11]. Rank [JHBA17]. Ranking [BS16, WCCL13]. Raster [BdBG^{+}17]. Rate [FGS15, RASM17, WM19, ZHL^{+}17, ZLYX10]. Rate-Modifying [ZHL^{+}17]. Rating [ZZM17b]. Ratings [NB17]. Ratio [MS14]. Rational [KOTY17, LWYZ17]. Ray [SBV19]. RBAC [VN16]. RC [Mar10b]. Re [EFV15, FFH17, GSW^{+}16, GN10, GZXA19, LK18, LSLW15, LLpC16, ML13]. Re-Configurable [EFV15]. Re-Encryption [GSW^{+}16, LSLW15, GZXA19]. Re-Execution [LLpC16]. Re-identification [FFH17]. Re-Routing [GN10]. Re-Signature [LK18]. RE-UML [ML13]. Reachability [CCY10]. Read [BBM17, LLPY19]. Read/Write [LLPY19]. Readaheads [XXW11]. Readiness [HJL10]. Ready [WRSV12, JJ18, WLWL18]. Real [ALH17, ASCTFP16, AFKT12, ABL^{+}18, CDYC11, CAV17, FXV13, FGS15, GJQG14, GIB12, IMS10, JBM^{+}19, KW11, LZN^{+}16, MSH^{+}11, Meg19, NH19, NL19, YGH^{+}14, wZfG15]. Real-Time [ALH17, CDYC11, CAV17, FXV13, IMS10, JBM^{+}19, KW11, MSH^{+}11, NH19, NL19, YGH^{+}14, wZfG15, FGS15, GJQG14, GIB12]. Real-World [ABL^{+}18, LZN^{+}16, Meg19]. Realistic [CXF^{+}15, dRFMD^{+}17, GB14]. Reality [ZZLL18]. Realization [JHHC15]. Reallocation [LWZ^{+}18]. REALM [PA15]. Realtime [KXS^{+}10]. Reasoning [BPK10, JHHC15, SKS19, ZFZ12]. Receiver [Wan14, Chi12]. Recognition [BY16, Cai12, CC11, CLLH13, GB10, IMS10, IA15, JHBA17, jLbLzH18, MZW^{+}18, STBB14, WWHL12]. Recognize [Tah11]. Recognizing [BY14, TKM11]. Recollections [Wet10]. Recommendation [CLL14, MGBD15, ÖKA11, SMLM14, WLW^{+}18, YMS^{+}15, ZZM17b]. Recommender [CZ19, DJAJ15, TBBH18, YGLW15, TYL^{+}18]. Reconfigurable [ASG15, BHAC10, FWC13]. Reconstruction [ED09, ED10, KOTY17, KLT^{+}15, PG11, VGF11, XHC^{+}15]. Records [NHMI13, ZVG16]. Recovery [LYY^{+}18b, NRZQ15, XHQX18]. Rectangle [CWZ19]. Rectangular [BBM14]. Recurrent [KNHK12]. Recursion [MGZ18]. Recursive [CC19, Day11, SLZ14, vdH15, BACD13]. Redesign [HMZ15]. Reduce [Kuo10, MMH18]. Reduced [LJF16, LJ18a, LYD^{+}18, LSG^{+}19, THY^{+}18]. Reduced-Reference [THY^{+}18]. Reduced-Round [LJ18a, LYD^{+}18, LSG^{+}19]. Reducing [CSS16, RATB^{+}13, STW^{+}18, ZC10]. Reduction [GMS11, KH18, KMNA^{+}16, KMZ16, LZHS14, OJSO14, PT13, WSY19]. Reductions [Bla13]. Reed [WLZ^{+}15]. Reference [FS18, KL10, TK15, THY^{+}18]. Reference-Inspired [FS18]. References [PK18]. Refinement [BACD13, LYPL17, WJ16, ZWFW15]. Refinements [LL11a]. Reflected [SV15]. Reflections [Den12c]. Refutations [FSGS15]. Region [HZW^{+}14, Ros14]. Regions [SDW13]. Register [HYZ17, KL10]. Register-File [HYZ17]. Registers [ZH15, ZL19]. Regular [Cal11a, CYTP18, GJ16, KV15b, XLC19, WCW10]. Regularization [ED09, ED10]. Regulatory [LH11]. Rehabilitation [PKM18]. Rekeying [DT13, LTW10]. Related [CWZ19, CMA14, GV16, HLLG18, MEdJMGE^{+}19, NHMI13, ZH14]. Related-Key [CWZ19, CMA14, HLLG18]. Relating [CGVP15]. Relation [Hie16]. Relational [BP10, LLZY15, LZL^{+}19, WP17]. Relations [HLL11, Hie13, TA16a]. Relationship [CZ19, LCX14]. Relationships [GRK13, KCC15, YWDW12]. Relative [CXH14, FGN^{+}18]. Relaxation [GLK^{+}16]. Relay [JYL18]. Relays [Gra12]. Relevance [dMRGAS18]. Relevant [LH13]. Reliability [HXZ^{+}16, KSPR15, WM19, XZL17, jZ18]. Reliable [ABCG11, Elg15, KSA12, LS14, MK11, MS12]. Relocatable [RHG^{+}11].

Remembers [Bar11]. Remote [FYMY15, HM17]. Renewable [LZZ^{+}17]. Repair [BFCRH14, HC15]. Repairs [YC19]. Repeatable [RHG^{+}11]. Repetition [YLX^{+}11, ZSL19]. Replacement [AÇPD11, RRCC^{+}15]. Replay [BMG12]. Replica [HZQ^{+}19, TPG^{+}15, WCW^{+}18]. Replica-Based [HZQ^{+}19]. Replicas [CNV13]. Replication [LFHF14, WWB17]. Representation [Tam18, ZDM^{+}15]. Representing [DH12b, KBMA12]. Reputable [RMB11]. Reputation [BL15b, FM11, LHM^{+}15, MS11, SF17]. Request [LWS^{+}14]. Requirement [YWR^{+}14]. Requirements [AKA15, CvdT10, Cha10b, GK17, LMA^{+}15, ML13, MBBA16, OS16, SKK18, SKS19, SWLZ12]. Requirements-Based [GK17]. Resampling [FLCT10]. Rescue [RFMJ10]. Research [ZJLC16, ZDZ^{+}15a]. Reservation [CLLL17, LSY^{+}16]. Reshaping [Uli11, BS10b]. Residential [ZLG15]. Residue [AJ15]. Resilience [SM12]. Resiliency [BEG^{+}16, ELS11]. Resilient [CL15, EOIH15, HHS18, KSA12, LTZY16, PSD15, SJ18a, ZYT13, ZWTM15, ZZM17a, ZYY19, ZY17, ZYM18, ZYH^{+}19]. Resistance [CLC^{+}19, CLJ^{+}22]. Resistant [FET17, YLLS16, FM11]. Resistor [Mar10b]. Resolution [ABS13, BT18, ED09, ED10, HFP^{+}19, KOA15, ZJLC16]. Resource [AGP10, BKFP19, CLH^{+}14, Cha10b, CTD18, CK10, CMY17, CWCS14, Das17, EOIH15, JMG^{+}16, KV15a, KLT^{+}15, KCZJ14, LZL^{+}17, LS14, MK13, NNN^{+}14, RAJ15, SDN15, ŞLV^{+}11, TL19, TPV18, WXLL18, ZDCZ18, ZDZ^{+}15b, JJ18, LJWL19, TXJ^{+}19]. Resource-Aware [AGP10]. Resource-Dependent [WXLL18, JJ18]. Resource-Efficient [KLT^{+}15]. Resource-Sharing [MK13]. Resources [KHR^{+}19, LfHmXjL11, PCC^{+}16, YT16a, wZfG15]. Respect [ABS14]. Respective [VM14]. Response [BTHS12, HQL17, HLKL15, Kon10, Mil10, Pen10, Sin12]. Responses [Sin12]. Restricted [SV15, SJS12, WZF18]. Result [DLM^{+}14, ZWC^{+}19]. Results [BCH^{+}15, BLRT10, LJF16, RSW14, SLP11, Xie11]. Rethinking [MV16]. Retrial [Dim13]. Retrieval [ACB17, CJYY17, CMSML16, CW12b, IJM14, JMG^{+}16, KYU11, KAS13, Lev11a, LL11b, LGC19, Mel13, PB12, RKBY15, SA11, SK18b, TSK17, VBVP14, WHSW15, ZBY^{+}10, ZXZ^{+}11, CMS10]. Reuse [RRDC^{+}18]. Reusing [WLH15a, ZWC^{+}19]. Revenue [YHGL17]. Reverse [BG15, WCL15]. Reversible [HHS^{+}15, JDAZN16]. Review [Gaz10, Jas10, Joh10, Lar10, Lev10a, Lev11a, Lop15b, Maj10, Mar10a, Uli11, WHS^{+}16, ZJLC16]. Reviews [Kam10, Kam11a, Kam11b, Kam11c, Kam11d, Kam11e, Kam11f, Kam11g, Kam11h, Kam11i, Kam11j, Kam12a, Kam12b, Kam12c, Kam12d, Kam12e, Kam12f, Kam12g, Kam12h, Kam12i, Kam12j, Kam12k, Kam13]. Revision [MEdJMGE^{+}19]. Revisited [GWWC15, LL11a, Lop15a]. Revisiting [IAG^{+}14, RSD19, WSA15]. Revocability [WHLH16]. Revocable [AEHS15, CD16, IDVGMP^{+}13, LNWZ19, LLLW17, QZZ18, SZS14, TCL15, TT12]. Revocation [AEHS15, BP19, CGE^{+}14, LW16, Lop15b, RDZ^{+}16, ZDL^{+}17]. Reward [CLRJ14]. Reweighting [Kot11]. Rewriting [AC14, KS19, TSC^{+}17]. Rewritings [ZLL^{+}14]. RFID [BL11]. RGB [ST16]. Rich [MKW11]. Rider [SHR^{+}11]. Right [Tra12]. Rigorous [MBRM15]. Ring [CZCD18, KR14, LYY^{+}16, MPSP17, PTWB14, XY18, YLA^{+}13, ZJ14]. Ring-LWE [XY18]. Risk [Buz12, Cha11, HHH^{+}18, OS18, SR10, TKB18, Zha15]. Risk-based [HHH^{+}18]. RKA [SLY^{+}16]. RKA-Secure [SLY^{+}16]. RL [SVS15].

RNA [Mar10b, SLL15]. RNN [TST^{+}11]. RNS [ABS12, HS19, Par15]. RNS-Based [ABS12]. Road [ZWFW15]. Roadmap [PW19]. RoboCup [RFMJ10]. Robot [MEH19, SZB19]. Robotic [OLF^{+}17]. Robots [Ros14, WCL^{+}11]. Robust [ACW13, BCG12, Cai12, LYPL17, LY10, MKK15, WLHH18, YYO15]. Rokach [Lev10a]. Role [ZVH11, ZVG16]. Role-Based [ZVH11, ZVG16]. Roles [TKM11]. Roman [Gaz10]. Rooms [BBM14]. Rotation [CLW11, KJ11, LMMP16]. Rotations [LYC11]. Round [KOTY17, LJ18a, LYD^{+}18, LSG^{+}19, XZLW15, YLL^{+}17]. Rounding [KJ11]. Route [HC15, WLW^{+}18, WCKH10]. Router [AS11]. Routing [AOS^{+}15, ABH15, BBM10, BF19, BK12a, BK14, DSTC12, GN10, GM11, GTS^{+}11, GHMP18, GGZC11, HK15, HH14, KSA12, KSPR15, LLN^{+}15, MK19, SJS12, WCKH10, WS10, WF10, WLY^{+}15, YWSH10, ZYR^{+}13, ZX16, ZLYX10]. Row [Sta18]. RP [AOS^{+}15]. RSA [CCL^{+}19, MV19]. RSD [ZX16]. RTCC [WW17]. RTCC-Pyramids [WW17]. RTDB [EMB19]. Rule [CLH^{+}14, JDAS12, LSW10]. Rule-Based [JDAS12, LSW10]. Rules [GF13, GBA18, TS19]. Run [BJY11, IF16, LHCN11, LWC15]. Run-Length-Encoded [LWC15]. Run-Time [BJY11, IF16]. Running [NL19]. Runtime [SVP13, ZDCZ18].

S [ASG15, Cha10b, LH11, NHC13, RMP10, SC10, WJ19]. S-box [RMP10]. S-Boxes [WJ19]. S-DIRECT [ASG15]. S-system [LH11]. SaaS [LWS^{+}14]. Safe [CK15]. SAFER [YCL17]. Safety [BCK^{+}11, CLLL17, FYF^{+}18, GH17, GB14, KSH^{+}14, OS16, ZLCW14]. Safety-Critical [ZLCW14]. Said [Den12a]. Sale [CC14]. Salient [AFKT12, FFH17]. Samplable [Yas19]. Sampling [DB13, RDMRM12].

Sanitization [RJ18]. SAR [WLW^{+}18]. SAT [AGR15]. Satisfaction [BBGM14]. Satisfiability [Wak17]. Satisfiability-Based [Wak17]. Satisfying [SWLZ12]. Save [ZJHJ17]. Save-Then-Transmit [ZJHJ17]. Saving [ARVR15, Dim13, LSCG10, SDN15]. Scalability [ER14]. Scalable [ASG15, DT13, HIDFGPC15, yHRT^{+}12, RTE^{+}13, SBBB12, VWR11]. Scale [BPFK19, CCUA14, KTTRJ18, LH13, LLDL17, LPV10, MDY15, NS16, OS18, WSR11, WT12, WCW^{+}14, ZHY^{+}14, ZWFW15]. Scale-Invariant [NS16]. Scaling [SSK12, TS19]. Scanners [BÜ11]. Scans [BBK11, DB13]. Scenario [ADBPLV13]. Scenarios [DSB15, SS10b]. Scene [SA11]. Schedulability [CLSV15, NL19]. Scheduler [SDN15, SPRR^{+}17, XZY^{+}10]. Schedules [MK11]. Scheduling [BHR10, CL15, CK10, CMY17, CWCS14, EMTSM18, EV16, GA18, HCL15, HJM12, JJ18, KL10, KCZJ14, KV19, LHFF13, LWS^{+}14, PB12, RR16, SU18, SM16, SLW^{+}17, SHL^{+}15, SW14, TB11, VMF^{+}14, WS15, WL18, WGL^{+}18, WXLL18, WLZ^{+}18, ZLX^{+}15, wZfG15, LJWL19, TXJ^{+}19, WLWL18]. Schema [ABS14, KKM^{+}15, MDY15]. Schema-Free [MDY15]. Scheme [ASS15, BP19, CLLL17, CLG^{+}19, CHDP17, CL16, Dim13, HSMY14, HIDFGPC15, HWS^{+}19, HP17, IL15, LCH16, LTH^{+}15, LTZY16, LSQZ17, LSQL18a, LHFF13, LTC^{+}15, LYY^{+}18b, LGPRH14, MK13, NK14, NG17, RSD19, RR16, ST16, SPJA11, Tan11, TPG^{+}15, TL19, UKW^{+}18, WYL^{+}13, WLH15b, WWB17, WF10, XTH11, YWR^{+}14, YL17, YMWS11, YY17, ZZQ^{+}19, ZM18, ZY17, DT13]. Schemes [BVS^{+}13, BF19, CZCD18, CLND19, Do11, HLLG18, HHL10, HMS^{+}12, HCL15, LWL10, MS11, PDNH15, QS15, THP^{+}11, THP^{+}12]. Science [ET19, Ham12, MP17, Suz13]. Scientific [Lev10a, NP16, SMLM14, Tra12, WS15, YMS^{+}15]. Scores [WCCL17]. Scoring [CXH14]. Screening [LP14]. Scripting [DSB15]. Scrum [AAA19]. SDM [VL13]. SDM-Toolkit [VL13]. SDN [DHT^{+}19, URHK19]. SDN-Based [DHT^{+}19]. SDN-Enabled [URHK19]. SDWN [AFG^{+}17]. SE [Pop11]. SE-Compression [Pop11]. Sea [Cro10]. Search [Cha10a, Che15a, CMSML16, CMS10, DCA18, EB12, FP18, GJQG14, GN19, HQL17, IJY^{+}14, LSQL18b, TMC15, WDCL18, XLM^{+}14, KAZ18, Lev11a]. Searchable [ZZQ^{+}19]. Searches [EDH^{+}18]. Searching [LCXZ16, NZ14, PW12, YGLW15]. Secondary [SLL15]. Secret [CCL^{+}13, DD10a, KS18, KOTY17, LPL15, LTC^{+}15, LJ16]. Secure [ABL^{+}18, BVS^{+}13, BWLA16, BCG12, BFMT15, BP19, CC14, CZLC14, Che15a, CMA14, DM18, DG15b, GWW^{+}13, HJS^{+}13, HLLG18, HH17, HLKL15, IDVGMP^{+}13, IL15, KSA12, Küp15, LL15, LTH^{+}15, LTZY16, LYPL17, LSLW15, LSQL18b, LHL16, MK19, NMS14, NSMS14, QZZ18, RMP10, RLJ15, SZS14, SKK^{+}12, SGH15, SLY^{+}16, TCL15, WDCL18, WLH15b, WMS^{+}12, WWJ18, YLL^{+}17, YAM^{+}15, YY17, YNN11, ZXZ^{+}11, ZVH11, ZVG16, OKG^{+}12]. Secure-TWS [OKG^{+}12]. Securing [ZDL^{+}17]. Security [BKFP19, BJY11, Cha11, CBJX19, GMS11, HXZ12, HMS^{+}12, Jay12, JSP13, Jia14, KL14, KS18, LE13, LLPY19, LSQZ17, LYL^{+}18, LSQL18b, LWL^{+}17, LDB^{+}15, LLH18, MKN13, MKK15, MV16, Mit19, MPH14, MHMSGH16, MGM12, OS16, OS18, PZPS15, PDNH15, RB17, RMFM15, RMB15, RMR15b, TD12, TAC^{+}18, Trc10, TV12, TV15, TKB18, WSA15, YYK^{+}17, Zha15, ZM18, ZYY19, ZY17, ZYH^{+}19]. Seed [XZW^{+}17]. SEFE [AD16]. Segment [WOLP15, ZCL^{+}12]. Segment-Based [ZCL^{+}12]. Segmentation [CLM16, KS16, MPP15, RDMRM12, SCKH18, SLZ14].

Segmented [NPTZ16]. Select [LSTC11]. Selecting [MPLDV13]. Selection [AAH10, GRK13, HPG^{+}15, KJ11, KÖ14, KO15, LS17, SBBB12, SZW^{+}18, TNWT14]. Selective [CL18, DSZZ15, GDCC16, LLH18]. Selective-Opening [LLH18]. Selectivity [IS13]. Self [AZHASD14, BPBRT16, DM18, DLV10, FM11, FXV13, GDKP10, HJK13, HB11, LL14, LLPY19, MDN^{+}11, PRJS11, THP^{+}11, THP^{+}12, WCL^{+}11]. Self-Adaptive [FXV13]. Self-assembly [WCL^{+}11]. Self-Coexistence [AZHASD14]. Self-enriching [PRJS11]. Self-Healing [THP^{+}11, THP^{+}12]. Self-optimized [MDN^{+}11]. Self-organizing [FM11]. Self-propagating [GDKP10]. Self-Stabilizing [BPBRT16, DLV10, HJK13, LL14]. Self-Synchronized [DM18, HB11]. Self-Updatable [LLPY19]. Semantic [CW12b, DGV17, Hsu12, IJM14, JK12, VL13, ZDM^{+}15, vDBvEW10, FLZC15]. Semantically [MKW11]. Semantics [Chi14, GF13]. Semi [KV15b, SMM^{+}19, XYL^{+}11, XZLW15, XHTH13]. Semi-Extended [KV15b]. semi-Markov [XHTH13]. Semi-structured [SMM^{+}19]. Semi-Track [XYL^{+}11]. Semi-trusted [XZLW15]. Semirings [Möl13]. Sensational [YGFL15]. Sensing [Alm19, Ano10, DA18, FZCL18, LZWY18, PZZ^{+}17, SJA17, TMOO11, ZJHJ19, KT18]. Sensitive [KSPR15, KS19, LV17, SSS16, Tan15]. Sensor [ACG^{+}11, ABM12, AKL^{+}19, AFGG11, Alm19, AAH10, BN14, BL11, BEG^{+}16, BY14, BSK19, BS10a, CCF11, CLRJ14, CWS^{+}10, CDYC11, Cor11, DE10, DA14, DSTC12, DBHC15, ER14, ELS11, FLWL19, FT11, GPK11, GN10, GTS^{+}11, GTB10, GLL^{+}13, HJS^{+}13, HLJ^{+}15, HB11, HZHC11, KAAE11, KTTRJ10, KXS^{+}10, KLT^{+}15, LYPL17, LTL10, LWPZ13, LCLL12, LSCG10, MT11, Meg18, MMPB10, MGBD15, MS11, OKG^{+}12, OB18, PANH10, RL11, Rog11, SJ14, ŞLV^{+}11, SM16, SJ18b, SYH11, SJS12, TLRE11, TB11, VRAC11, WN11, WZ17, WOV^{+}10, WBS15, YC14a, YWSH10, YDE11, YNN11, ZTBW11, ZBY^{+}10, ZLX^{+}15, ZJHJ19, ZLYX10, dFHP^{+}11]. Sensor-Aided [Alm19]. Sensor-Based [MGBD15]. Sensor-Instrumented [FT11]. Sensor-Network [MMPB10]. Sensornet [TB10]. Sensors [BY16, yZdZhZ18]. Sentiment [FP19]. Separation [XZW^{+}17]. Sequence [CZL^{+}18, STW^{+}18, SC10, VRD10, WCW10]. Sequences [Bla13, HT15, HT16, HT17, SV15, SLL15, WWHL12, YWFQ18]. Sequential [Cha10b, GOR^{+}10, LR10, OKT^{+}16, Vel10]. Sequentially [CFJ^{+}10]. Series [EKOS19, KNHK12, NHMI13, SCKH18, SZL15, ZCL^{+}12]. Servants [ET19]. Server [Che15a, Do11, GMSV14, JWCZ13, LSQL18b, LNWZ19, MV19, Mit10, RDB14b]. Server-Aided [GMSV14, LNWZ19, MV19]. Server-Designation [Che15a, LSQL18b]. Service [AAZ13, BZS^{+}16, BKFP19, BDC11, BKBK14, BCKM17, CWS^{+}10, CCHL18, CL16, DTFT11, DTFT12, EV16, GVVL12, HMM11, HuRH^{+}15, KCZJ14, LCH16, LP14, LWS^{+}14, LDB^{+}15, LÖ10, MDS15, NRZQ15, OLF^{+}17, PZL12, PP17, WXP^{+}10, ZSX10, dAEN^{+}18, ÁHFE18, HHH^{+}18]. Service-Based [LP14, LDB^{+}15]. Service-Oriented [OLF^{+}17, PZL12]. Services [Ang13, BV15, DBC18, Elg15, ET19, FLZC15, GLBS13, HLC10a, HJM12, IDVGMP^{+}13, JSP13, KHC15, LPL14, NB17, SBBB12, SVP13, SSY15, WWJ18, ZHL15]. Session [HLC10a]. Set [AJ15, BG15, BKP11, BK12a, BK12b, BK14, CLW11, LC14, LHL16, MSH^{+}11, PH15, RCS16, YCL15]. Set-to-Set [BK12b, BK14]. Sets [HJK13, HS19, OJSO14]. Setting [MZHY15, Ma17, ZHL15]. Setup [HJM12]. Seven [CFS13]. SFP [HGRV15]. SGAC [HFP^{+}19]. Shadow [HZAZ18, KS16]. Shadows [YSC^{+}15]. Shape [CLM16, KYU11, NLDH11, SY13]. Share [LTC^{+}15]. Shared [CFJ^{+}10, NSRP15, NHC13, OKG^{+}12, OBA16, WWZ^{+}17, ZC10, wZfG15, PZPS15]. Sharing [CLG^{+}19, CK10, CCL^{+}13, DD10a, EOIH15, KOTY17, LPL15, LY10, LTC^{+}15, LZZ^{+}17, MK13, NH19, QZZ18, VB16, YC11, EFV15]. Shearlet [TS17]. Shell [WZCC18]. Shift [ZH15, ZL19]. Shih [Joh10]. Shilling [CZ19, TYL^{+}18]. Short [GMS11, LZL^{+}19, PRJS11, XGLM14]. Short-Text [LZL^{+}19]. shortening [WLWL18]. Shot [BPK10]. Shuffle [GÁVRRL16]. Side [KH10, RDB14b, YL17]. Side-Channel [KH10, YL17]. Sign [IMS10, LL15, jLbLzH18, ZHY^{+}14]. Sign-On [LL15]. Signal [CCUA14]. Signature [ASS15, AEHS15, CZCD18, CLND19, GJJ15, GMSV14, GHY18, HHL10, HZX15, HP17, LK18, LTH^{+}15, LDZ16, LYY^{+}16, LGPRH14, LLS17, OBA16, ST16, Tan11, TTH15, WZXL12, WLH15b, WYML16, WHLH16, XGLM14, YMWS11, YLA^{+}13, ZJ14]. Signatures [GdM16, GMS11, HMS^{+}12, HHS18, Ver17, WCD19, WLI^{+}14, YT16b]. Signcryption [CMA14, HWY11, IL15, LSQZ17, LSQL18a, RSD19, YY17, ZCL13, ZM18]. Significance [BPK10]. Significant [KTM19]. Significantly [YZLC15]. Signing [DGFGHZ13, YAM^{+}15]. Signposting [Thi11]. SIMD [HWXD14]. Similar [ZDCZ18]. Similarity [Cha10a, – DG13, HPG^{+}15, NZ14, ÖKA11, TA16b, ZZ17]. Similarity-Based [HPG^{+}15]. Simple [Cha10b, EKOS19, LYY^{+}18b, Xie11, ZH15]. SimpleLock [YB16]. Simpler [YLL^{+}17]. Simplicity [Yas19]. Simplifications [ZTTM18]. Simplified [RHF^{+}15]. Simulated [HGZ10]. Simulation [GLK^{+}16, GB15, yHRT^{+}12, Jar12, KOA15, LDK11, LLH18, TKB11, WXP^{+}10]. Simulation-Based [LLH18]. Simulator [DFG10, GFPC16, SHR^{+}11]. Simultaneous [DDL^{+}15, LJWL19, VGF11]. Since [Har11]. Single [ED09, ED10, HZAZ18, LL15, LJWL19, RH17, WGL^{+}18, WXLL18, XHQX18, ZHL^{+}17]. Single-Image [HZAZ18]. Single-Machine [ZHL^{+}17, LJWL19]. Singular [NS16]. Sink [KAAE11]. Sinkhole [HLJ^{+}15]. Sinks [ABCG11, TB11]. SIP [PP17]. SIP-Based [PP17]. Sirt [VGF11]. Site [DSB15]. Situation [KBMA12, ZFZ12]. Situations [KHYC15, STBB14]. Six [GTK^{+}19]. Size [AEHS15, LSQX19, WCXZ17, YWSH10, ZMW16, ZSL19]. Sized [GHXW16]. Sizes [ZL15]. Skeleton [YGLW15]. Skip [FP18]. Skip-Search [FP18]. Skyline [SCT18b]. SLA [HHHC16, DB15, HHHC16, HHH^{+}18, NSRP15]. SLA-Based [NSRP15]. Slacking [SHL^{+}15]. SLC [JYP^{+}15]. SLC/MLC [JYP^{+}15]. Slices [VBVP14]. Slicing [LCX14, YWY10]. Sliding [MDY15]. SLK [WGL^{+}18]. SLLCs [RRDC^{+}18]. Sloman [Mil10]. Small [ABG^{+}12, ARVR15, HXQ^{+}19, OS18, QYZ19, WCXZ17, YZLC15, YTV16, YT16b]. Small-Scale [OS18]. Small-World [ABG^{+}12, ARVR15, QYZ19, YZLC15]. Smart [DSTC12, DFG10, GTK^{+}19, MHW10, SSK19, ZNQR15, JBM^{+}19, SKK^{+}12, XLXZ17]. Smart-Context [MHW10]. Smartphone [BDC11, LWKB15]. Smartphones [PZZ^{+}17]. SmartRec [XHQX18]. SmartX [URHK19]. SmiDCA [SK18a]. Smishing [SK18a]. SMML [Dow15]. Smooth [LYY^{+}18a]. SMS4 [LYL^{+}18]. SMT [AGR15]. SOAs [QS15]. Social [Cao14, DLL^{+}13, EÇGK16, FMRS17, HY15, Hsu12, HLZ^{+}17, HZJS17, KHYC15, KCC15, KKBF12, Lev10b, LBZ19, MP18, NSA15, NRZQ15, PYM^{+}15, RKBY15, ST17, SCT18b, SVG^{+}15, SMLM14, STBB14, UKW^{+}18, WLW^{+}18, WCW^{+}14, WLY^{+}15, XLM^{+}14, XZA14, YNP15, ZW15, Zha15, ZHY^{+}14, ZL15, ZHL15, TYL^{+}18]. Social-Aware [SCT18b, WLW^{+}18]. Socially [CZL^{+}18, MK13]. Societies [LLV10, SS10a]. Society [GG10]. Socio [MK11]. Socio-oriented [MK11]. Soclake [KKBF12]. SoD [VN16]. Soft [DN16, MSH^{+}11, WLI^{+}14]. Soft-Failures [WLI^{+}14]. Software [AFG^{+}17, AAA19, AO08, Ano17, Bro10, dRFMD^{+}17, FSMT19, GHMP18, Ham12, Har10a, JAAA^{+}17, KHC15, LWKB15, LMA^{+}15, LZL^{+}17, LLDL17, Llo13, LCX14, Maj10, MV16, OLF^{+}17, PS17, Pyl19, RJS^{+}17, RASM17, SKKM15, TKM11, TLRE11, WB16, WM19, YHS^{+}17, YWFQ18]. Software-Defined [AFG^{+}17, Ano17, dRFMD^{+}17, GHMP18, JAAA^{+}17, LZL^{+}17, LLDL17, RJS^{+}17, RASM17, WB16]. Solution [DHT^{+}19, Fra15, HLKL15, KT18]. Solutions [KKMG15, PANH10]. Solve [OLL15]. Solvers [AGR15]. Solving [KLS18, NG17, NM19, WJ16]. SOM [VBMH10]. Some [BCH^{+}15, BBP13, LWL10, LCMC11, SV15, Xie11]. Sort [WWW16, ZHW19, WW19]. Sorting [LHCN11, TKM11, Tah11]. Source [BY14, – DG13, GRK13, JLS11, MBC15, RH17, RL11, Tah11, WM19, PB12]. Source-Based [MBC15]. Source-Location [RL11]. Sources [JLS11, SMM^{+}19]. SP [WT12, PW19]. Space [Abd15, BBM10, BUB13, BGM^{+}11, BWR12, PT13, SZB15, ZLG15]. Space-Filling [BWR12]. Spaces [BP10, NZ14, SH15]. Spammer [WHS^{+}16]. Spanning [BCC^{+}19, BPBRT16, CFJ^{+}13, WW17, YC14b, YLC15]. SPARQL [ZV15]. Sparse [DB13, OVGG14, ZDM^{+}15]. Spatial [ACW13, CK15, ER14, FGG13, JYL18, LPD13, MCT19, RDMRM12, SDW13, SCT18b, TST^{+}11, VB16, WCL15]. Spatiotemporal [NHC13]. Special [Ano17, CQL10, Jay12, Llo13, OLL15, Pek12, RA14, RLJ15, SS10a, Suz13, XZA14, Ano10]. Specific [DAOG14, TEP^{+}16]. Specification [JCSZ13, KW11, RJV13]. Specifications [SLP11, SZL15]. Spectral [QYZ19]. Spectrum [Alm19, DA18, SJA17, YHGL17, ZJHJ19]. Speed [ASG15, PW12, GIB12, MDSF12]. Speeding [KTTRJ18, Rig14]. SPEKS [Che15a]. Spheres [SAKOK11]. Spiking [WF10]. Splicing [YSC^{+}15]. spline [RT12]. Splitting [MV19]. SPMD [BMG12]. Spoof [SP15]. Sporadic [wZfG15]. Sports [BY14]. Springer [Gaz10, Jas10]. Squares [KÖ14, KO15]. SRN [YT11]. SSA [QO17]. SSD [LJYL13, LGHD15]. SSD-Based [LGHD15]. SST [yHRT^{+}12]. Stability [ZYWW13]. Stabilizing [BPBRT16, DLV10, HJK13, KSA12, LL14, SJ14]. Stable [Meg18]. Stacking [yZdZhZ18]. stage [KT18]. Standalone [CAV17]. Standard [GJJ15, HZX15, LK18, LDZ16, LLS17, RSD19, SZS14, TCL15, XLC19]. Standards [GLBS13]. Star [WWW16, WW19]. Stars [CYTP18, YLC15]. Start [SWG13]. Start-up [SWG13]. starting [JJ18]. State [Abd15, BVS^{+}13, CCL^{+}13, EFY16, EFYS19, HT15, KH18, KLA^{+}15, MMH18, RSW14, Ros14, TV12, VM14, jZ18]. State-Based [RSW14, TV12]. State/Event [KH18]. Stateful [BVS^{+}13]. Stateless [KLA^{+}15]. Statement [Den12a, Den12b, HXZ12]. Static [IF16, PiLCH11]. Statically [Ort11]. Station [LSY^{+}16]. Stationary [CTIAP12]. Stations [ISST19]. Statistical [CMSML16, FNP12, HGRV15, Hey17, WLI^{+}14]. Statistically [MBRM15]. Status [RJS^{+}17]. Steady [VM14]. Steady-State [VM14]. Steganalysis [YLL^{+}12]. Steganographic [HHS^{+}15]. Steganography [BCG12, HZW^{+}14, Joh10, KTM19, LLY^{+}12, Shi08, TJZF12]. Stego [YLL^{+}12].

Stego-Image [YLL^{+}12]. Stemming [SVG^{+}15]. Step [HJS^{+}13]. Stochastic [ASCTFP16, BBM10, DH12b, HTG12, Kon10, NB12, PL16]. Storage [BBM17, DCLN11, FEDHL16, HZQ^{+}19, HMH18, Küp15, LPL14, LDLJ15, LGHD15, LBIC14, WS15, XHQX18, ZVH11, ZVG16]. Stores [HXQ^{+}19]. Storing [Mer13]. Straight [DDLM17]. Straight-Line [DDLM17]. Strand [SH15]. Strategies [BFCRH14, Har10a, NdMCdMM16, SM12, SZL15, TJZF12]. Strategy [BACD13, DB15, FM11, FYF^{+}18, GTM15, HL15, HFP^{+}19, LFHF14, NRZQ15, YGLW15]. Stream [Abd15, DM18, DG12, DJG^{+}15, Hey17, JZ13, LR12, MK11, ZH15]. Streaming [AGF15, AAH10, DSBB19, HZWT15, HXLX18, HXLX22, ISH13, LHYW12, LHFF13, TY14, YWDW12]. Streams [ALH17, MDSF12, YIUH14, ZWFW15]. Stress [GSS14]. Stretch [BF19]. Stride [PW12, VWR11]. String [CHL14, Kha16, KS12, LK14, PW12]. String-Matching [Kha16]. String-to-Dictionary [KS12]. Strings [LWC15, Mar10b]. Stripe [BÜ11]. Stripe-Based [BÜ11]. Strong [GSAS12, PYM^{+}15, WW19]. Stronger [ZYY19]. Strongly [LK14, YLL^{+}17]. Structural [DD19, KAZ18, LYY^{+}18b, LZN^{+}16, MGZ18, Yas19]. Structure [AÇPD11, ECL15, GRK13, JYP^{+}15, KMZ16, LJ15, LLF17, MPP15, Mur10a, SLL15, WJ19, WLC^{+}19, Yan19]. Structure- [Yan19]. Structure-Activity [GRK13]. Structure-Based [AÇPD11]. Structured [YLX^{+}11, SMM^{+}19]. Structures [CLG^{+}19, FMRS17, LPL15, Lop13, TD12]. Structuring [PGBFW14]. STT [RRDC^{+}18]. STT-RAM [RRDC^{+}18]. Study [Bla13, DCLN11, FLZC15, GÁVRRL16, HK15, LMR18, MK15, MCT19, NZ14, OS18, PXG^{+}17, PRG^{+}10, RDB^{+}14a, RCTK18, RMGT11, SR10, SY15, aSPW^{+}17, TKB18, WCCL17, WVGP11]. Subgraph [DDG^{+}15, LCXZ16, ZCX^{+}16]. Subgraphs [DP16]. Subgroup [CCL^{+}19, LPP^{+}13, VL13]. Subject [BY16]. Subnet [ZFZ12]. Subnetworks [CTIAP12]. Subsequences [LWC15]. Substitution [KTM19]. Substructure [Yan19]. Substructure-Cuts [Yan19]. Subsystem [HLC10a]. Subtraction [HL15]. Subtrees [YLW^{+}17]. Subtyping [DGV17]. Sufficient [LJC11]. Suffix [FGN^{+}18, OR12]. Suites [CSS16]. Sum [APW11, RASM17]. Sum-Rate [RASM17]. Summaries [HM13]. Summarization [KCC10, SIK14]. Super [ED09, ED10, LY10, WZF18]. Super-Peer-Based [LY10]. Super-Resolution [ED09, ED10]. Support [CLL10, FGG13, GAFP^{+}14, IJM14, JS15, JAAA^{+}17, KCZJ14, LQZ^{+}10, MMPB10, PP17, SPRR^{+}17, SNG^{+}10, VBBR16, WB16, XHC^{+}15, Zam19, ZZQ^{+}19, dFHP^{+}11]. Supporting [ET19, LPL14, PSS10, SAPS19, VvdAMG17, Ver17, WXP^{+}10]. Suppression [KS16]. Surface [CQS13, SAKOK11]. Surfaces [JHBA17]. Survey [Cal11b, GTK^{+}19, GBBK11, JHHC15, KL14, LBIC14, LÖ10, Sak10, THP^{+}11, THP^{+}12, Tim10, VRAC11, WGS17]. Surveying [BBK11]. Surveys [NPTZ16]. Survivability [CCHL18, RMB15]. Surviving [YAQ12]. Suspicious [FXV13]. Sustainability [JG15]. Sustainable [LZZ^{+}17]. SVC [LHYW12, CLL14]. SVC-Based [CLL14]. Swap [FP18]. Swarm [OJSO14, SJ18b, WCL^{+}11, WLZ^{+}18]. Swiniarski [Gaz10]. Switch [LGC19]. Switch-Centric [LGC19]. Sybil [FM11]. Sybil-resistant [FM11]. Symbol [Con12]. Symbolic [LZHS14]. Symmetric [BFF^{+}15, DTFT11, DTFT12]. Symmetries [BFF^{+}15]. Symmetry [LZHS14, Win11]. Symposium [Den12c]. SYN [DHT^{+}19].

Synchronization [HBS^{+}19, WHYH12]. Synchronizations [HM14]. Synchronized [DM18, HB11]. Synchronous [YEFVJ15]. Synopsis [Lav12]. Syntax [XLC19]. Synthesis [WLZ^{+}15]. Synthesized [AAHTH10]. System [AAA19, AJ15, ALA19, ALZ^{+}17, CGE^{+}14, CLL14, CC14, CWRZ18, CZLC14, CLLH13, CMSML16, DJAJ15, DP13, DG15b, – DG13, EMB19, GJQG14, GHXW16, HXZ^{+}16, HYZ17, HHS^{+}15, HXZ12, Hua14, HJM12, IK17, ISH13, IDVGMP^{+}13, JDAS12, JAAA^{+}17, Jas10, JG15, Kap11, KTC^{+}11, LP14, LDLJ15, LBZ19, LL11b, LSY^{+}16, Lop12, ML13, MGM12, MS14, NNF19, NM19, NSMS14, Nur07, PWY^{+}13, Pyl19, RLTZ17, dMRGAS18, SPS^{+}18, Sta18, SSS^{+}12b, SSK19, TAC^{+}18, WLW^{+}18, WHP^{+}13, YKK18, ZZM17a, jZ18, ZMSM13, ZVG16, LH11, TYL^{+}18]. System-on-Chip [Jas10, Nur07]. Systematical [OLL15]. Systems [AC14, Awa13, BL11, BJY11, BL15b, BL16, Bro10, BMG12, CZ19, CFMR14, Cha10b, Che15b, CLL10, Cro10, CWCS14, Dim13, DCLN11, DN16, FM11, Fra11, GB14, HGZ10, HS11, Hsu12, HHCL10, Jar12, JRC^{+}10, JK12, JMB12, KAS13, KSH^{+}14, KV19, LYY^{+}18a, LDK11, LE13, LWDZ16, LSW10, LY10, LWYZ17, LSTC11, LFHF14, LBIC14, Llo13, Lop13, MSH^{+}11, MK11, MSWI^{+}12, Nil10, NL19, OLF^{+}17, PABD10, RTE^{+}13, RSW14, RA14, RRCC^{+}15, RJV13, RLJ15, RMR15b, SL14, SU18, SF17, SL10a, Sta18, TBBH18, TD12, TKB18, VL13, WYL^{+}13, WNNZ17, XHQX18, YGH^{+}14, YC19, YIUH14, YDHW18, YHS^{+}17, YGLW15, wZfG15].

TABEMS [JG15]. Table

[CCL^{+}13, HLC10a]. Table-Based [HLC10a]. Tables [RMB11]. Tactical [SS10b]. Tactile [YLSL19]. Tag [Hsu12, LBD^{+}19, XLM^{+}14]. Tag-Based [XLM^{+}14, Hsu12]. Tagged [BBDF11].

Tagging [Hsu12]. Tags [Jun12]. Taiwan [WCCL17]. Target [CZ19, TNWT14]. Targeting [AK12, LCMC11]. Targets [BEG^{+}16]. Tariff [JG15]. Tariff-Aware [JG15]. Task [BHR10, CMKJ10, CK10, GA18, HGZ10, KZY16, LLpC16, MSH^{+}11, MEH19, PSS10, PZZ^{+}17, PCC^{+}16, SU18, SSK12, SZB15, Tim11, wZfG15, ZWC^{+}19, RG14]. Tasks [GTN10, SAPS19, SHL^{+}15, ZJLC16, ZDCZ18]. Taxonomy [AJA16, GAF^{+}15, KMSM15]. Taylor [Joh10]. TCAM [ASG15]. TCP [DHT^{+}19]. TDMA [CLSV15, SM16]. Team [HSZS17, HSZS18, SS10b]. Teams [HLZ^{+}17]. Technical [CGVP15]. Technique [BBM17, FEDHL16, KH10, SK18b, SJ18b, VGF11]. Techniques [ABG^{+}12, AK12, CWRZ18, DN16, ISST19, OKT^{+}16, PKM18, RCS16, SU18, Shi08, SKKM15, VO16, WGS17, Joh10]. Technologies [BT18, CvdT10, ET19, JSP13, KJR15, KHC14, Rog11]. Technology [Ano10, Uli11, BS10b, WLWL18]. Telecom [IK17]. Telemetry [KTC^{+}11]. Template [NGAuHQ16, YYK^{+}17]. Templates [SLZ14]. Temporal [CWWK14, KH18, VB16]. Tenant [TV15]. Tensor [MPP15]. Term [KTA12, LLZY15]. Term-Partitioned [KTA12]. Terminal [HM13]. Terminals [GTM15]. Terminating [BCKM17]. Ternary [VBVP14]. Tessellations [DKB^{+}14]. Test [AGR15, CHH^{+}19, CSS16, EFY16, EFYS19, FSMT19, HTC^{+}15, ISD15, LLSW16, MZHY15, NL19, SP10, Sin12]. Testable [RMP10]. Testbed [RHG^{+}11]. Testing [AO08, DH12a, DH14, Hie13, HT16, Hie16, HT17, PS15, RH17, SZL15, aSPW^{+}17, TAC^{+}18, YHS^{+}17, Maj10]. Tests [PABD10]. Text [DLL^{+}13, FNP12, GdM16, KVX12, KCC10, LZL^{+}19, SPS^{+}18]. Text-Augmented [DLL^{+}13].

Text-dependent [GdM16]. Texts [PRJS11]. Textual [NC16, SHH^{+}15]. Texture [FET17, KYU11, LL11b, MPP15]. TFRC [SWG13]. Their [BBK11, CL17, VM14, ZYY19, CGVP15, JHHC15, JJ18]. Theorem [BBP13, MEdJMGE^{+}19]. Theorems [KMNA^{+}16]. Theoretic [Meg18, TNWT14]. Theories [Roc12]. Theory [Ana10, HY11, HHCL10, JRC^{+}10, Lev10a, SH10, SKK18, SKS19, TKB11, RM08]. Theory-Based [SKK18, SKS19]. Thermal [ADML^{+}13, yHRT^{+}12, XZY^{+}10]. Thermal-Aware [XZY^{+}10]. Thickness [DDLM17]. Thin [Chi16]. Things [MDB^{+}18, NNF19, PZ18, SSK19]. Think [NH19]. Thinking [Aho12]. Thinning [Cai12]. Threading [OR12, YLLS16]. Threats [AJA16]. Three [ADML^{+}13, GB15, ZH19]. Three-Connectivity [ZH19]. Three-Dimensional [ADML^{+}13]. Three-Tier [GB15]. Threshold [DD10a, GWW^{+}13, LWL10, LYY^{+}16, WLH15b, YLA^{+}13]. Throughput [BBM17, MPSP17, SPdGPM18, ZJHJ17]. Ticket [CC14]. Ticket-Sale [CC14]. Tide [NHMI13]. Tie [CLS15]. Tie-Breaking [CLS15]. Tier [GB15, RAKJ17]. Ties [PYM^{+}15]. Tight [GDCC16, LLH18, ZYH^{+}19]. Tighter [GMS11]. Tightly [HLLG18]. Time [Abd15, ALH17, AFKT12, Alm19, ARR^{+}16, BBM10, BBDF11, BJY11, BUB13, CDYC11, CAV17, DJAJ15, DB15, DLV10, EKOS19, FXV13, GF13, Har10b, HJK13, IMS10, IF16, IDVGMP^{+}13, JBM^{+}19, JMB12, KNHK12, KAAE11, KW11, MSH^{+}11, NSRP15, NHMI13, NH19, NL19, OKT^{+}16, SCKH18, SZB15, WGL^{+}18, Whi12a, WXP^{+}10, YGH^{+}14, ZC10, ZCL^{+}12, wZfG15, ZHL^{+}17, ZH14, FGS15, GJQG14, GIB12]. Time-Aware [DJAJ15, NSRP15]. Time-Based [IDVGMP^{+}13].

Time-Branching [GF13]. Time-Dependent [DB15, ZHL^{+}17]. Time-Related [ZH14]. Time-Series [EKOS19, KNHK12, NHMI13, SCKH18]. Timed [Tan15]. Timed-Ephemerizer [Tan15]. Times [DTFT11, DTFT12, ZHL^{+}17, JJ18, WLWL18]. Timing [CK10, GB14]. tisements [NK19]. Token [ZM16]. Token-Leakage [ZM16]. Tolerance [CNV13, HZHC11, WZF18, WLC^{+}19, ZMSM13, ZM19]. Tolerant [Fan11, LTL10, WCD19, YWR^{+}14, ZM16, ZX16]. Tomographic [PG11]. Tool [RJV13, VvdAMG17]. Toolkit [VL13]. Tools [BKBK14, HM13]. Top [ALH17, BGM^{+}13, GN19, SCT18b, WWJ18]. Top- [ALH17, GN19, WWJ18]. Top-Down [BGM^{+}13]. Top-k [SCT18b]. Topic [DLL^{+}13, LZL^{+}19]. Topological [BP10, HPG^{+}15, QYZ19, SDW13, SAKOK11]. Topology [BHAC10, JC10, KNHK12, LWW13, YZLC15, YDE11]. Total [ABS13, MMAY19, WXLL18]. Touch [YLSL19]. Touchscreen [YLSL19]. TPR [NHC13]. Trace [BMG12, PiLCH11, WJ16]. Traceability [BJY11, WYML16, WHLH16, WSR11]. Traceback [FEDHL16]. Traces [ASCTFP16]. Tracing [LW16, PBH^{+}13, WHP^{+}13]. Track [XYL^{+}11]. Tracking [KTC^{+}11, LTL10, WL13, WSR11]. Tractable [QLZ18]. Trade [DDLM17, JLS11]. Trade-Offs [DDLM17, JLS11]. Tradeoffs [MPH14]. Trading [CZLY19]. Traffic [ASCTFP16, CLSV15, FGS15, GIP^{+}12a, GIP^{+}12b, HM16, KKPB14, OB18, RLVRGÁ15, SAPS19, XHTH13, ZYWW13, ZH14]. Training [BMRS11, KNHK12]. Traitor [LW16]. Trajectory [LZHS14]. Transactional [LM17]. Transactions [DG15b, TV15]. Transducer [KK18]. Transductive [KLL14]. Transfer [GRK13, HSMY14, HLC10b]. transferable [GZXA19]. Transform [BCPV11, KTM19, NS16, TS17]. Transformation [Kha16, VM14]. Transformations [BPFK19, QO17, RCS16]. Transformed [MZW^{+}18]. Transient [AKL^{+}19, CTIAP12, LJA13]. Transit [CCUA14]. Translation [EFY16, TA16b]. Transmission [Alm19, CCUA14, GIB12, RMP^{+}16, SGG^{+}13, TB11, YWSH10]. Transmit [ZJHJ17, ZJHJ19]. Transparency [TJZF12]. Transparency-Orientated [TJZF12]. Transport [CLLL17]. Transportation [Alh19, BDL^{+}13, LE13, WLW^{+}18]. Trapdoor [CCL^{+}19, CBJX19, HHL10]. Traveling [KAAE11]. Tree [BPBRT16, EDH^{+}18, Kuo10, LV17, MCT19, Tah11, WCW^{+}18, Yil12, NHC13]. Trees [CFJ^{+}13, FGN^{+}18, Kor11, Lev10a, LYC11, LHL16, MMH18, Meg18, SSS16, TRY16, WW17, WCW10, WCCL13, YC14b, YLC15, YTV16, RM08, XZLL18]. Trench [NHMI13]. Trend [WM19]. Trends [ZYR^{+}13]. Tri [WWHL12]. Tri-view [WWHL12]. Triangles [ST17]. Triangulations [BCH^{+}15]. Trie [EB12, ECL15]. Trigger [KR14]. Triggers [ZYF17]. Trip [SCT18a]. Trojan [Sta18]. Tropical [KTC^{+}11]. True [YDHW18]. True-Concurrency [YDHW18]. Truly [Erg11]. Trust [BL16, FMRS17, KMSM15, MDS15, MK19, NB17, PYM^{+}15, RHH12, TV15, WLY^{+}15, YGLW15, Zha15]. Trust-Aware [YGLW15]. Trust-Based [MK19]. Trusted [FPY15, YCR16, XZLW15]. TS [TL19]. Tuning [TLRE11, TB10]. Turing [Lav12, For12]. Turkish [KCC10]. Turning [GF13, Har10b]. Turtle [WZCC18]. Tutorial [CFMR14]. TWINE [LYD^{+}18]. TWINE-128 [LYD^{+}18]. Twitter [Alh19, GST15, SIK14, WGS17]. Two [BY14, DP16, DGFGHZ13, HJS^{+}13, JYL18, Koç10, NSMS14, SV15, SAK16, SZB19, ZM16, ZH19, ZL19, KT18]. Two-Party [NSMS14, ZM16]. two-stage [KT18]. Two-Step [HJS^{+}13]. Two-Way [JYL18]. TWS [OKG^{+}12]. Type [DTFT11, DTFT12, HWS^{+}19, LM17, Ort11, QO17, SH15]. Type-Flaw [SH15]. Typed [Ort11, QO17]. Types [ALH17]. Typicality [mAYL10].

U [VBBR16]. U-Library [VBBR16]. Ubiquitous [JAAA^{+}17, OS16, PSP14, RiCH10, SL10a, VBBR16]. UK [Ano10, GTK^{+}19, Jar11, OS18]. Ultra [LZL^{+}17]. Ultrametric [Mur10b]. uMax [NM19]. UML [ML13]. Unattended [BN14]. Uncertainty [Buz12, GSS14, Tim11]. Unconditional [Jia14]. Uncovering [Mur10a]. Understand [HY15]. Understanding [Sab11, VKZ^{+}10]. Underwater [LCLL12, SM16]. Undisturbed [YCL17]. Unicyclic [TRY16]. Unified [HY11]. Uniform [GF17, PZ19, RHH12]. Unit [LR12, PXG^{+}17]. Unitable [CCCS11]. Units [BY14, HS19, NdMCdMM16]. Univariate [Yil12]. Universal [LK14, WF10]. Universe [LW16]. University [Maj10]. Unknown [CLRJ14, GSAS12, Sin12]. Unranking [WCCL13]. Unreliable [dMRGAS18]. Unsupervised [CZ19, JBM^{+}19]. Untangling [LYC11]. Updatable [LLPY19]. Update [GdJ13, WCW^{+}18]. Updated [Cal11b]. Updates [HXQ^{+}19, LM17, MH11]. Upward [BD16, RH17]. Urgent [CCF11]. urgMAC [OB18]. US/UK [Ano10]. Usage [AHH13, HYZ17]. Use [AGR15, HY15, JG15, JLS11, PS17, SHH^{+}15, VO16]. Use-Cases [SHH^{+}15]. Usefulness [KBN10]. User [CMY17, GdM16, HCZ^{+}19, HK13, HXLX18, HZJS17, HuRH^{+}15, KHYC15, MZHY15, MBC15, PRG^{+}10, SCKH18, Swa11, WLI^{+}14, XTH11, XLM^{+}14, OKG^{+}12, HXLX22].

User-Based [HuRH^{+}15]. User-Experience-Oriented [HXLX18, HXLX22]. Users [IDVGMP^{+}13, NK14, OKT^{+}16]. Using [AJ17, ATA19, BT18, BY14, BY16, BCPV11, BP10, BMG12, CW11, Cha11, CMKJ10, CHDP17, CCL^{+}13, CAV17, CRGM14, Das17, DBHC15, DBC18, DD19, FAFD15, FET17, FLCT10, FMRS17, GSS14, GTS^{+}11, GK16, HM16, HNAS18, HHS^{+}15, HJL16, ISST19, IS13, ISD15, IJM14, IL15, JDAZN16, JBM^{+}19, JHBA17, JC10, KTM19, KYU11, KAAE11, KÖ14, KO15, KTTRJ10, KH10, KCZJ14, KS16, LR10, LLZY15, LWYZ17, MDY15, MBC15, Mar10b, MDS15, MK11, MK13, MGBD15, MK19, MKL18, NGAuHQ16, NSMS14, NLDH11, OKT^{+}16, PT13, PZ19, PRG^{+}10, PP17, QZXR15, RDB14b, RJ18, SBV19, SZW^{+}18, SHL^{+}15, TKM11, Tah11, TB10, TS17, TA16b, TEP^{+}16, VBVP14, WSY19, WLZ^{+}18, YCL15, YTV16, YCL17, ZH15, ZC10, ZYR^{+}13, ZFZ12, ABG^{+}12, AAHTH10, AÇPD11, BBM17, CC19, ET19, GHMP18, GA18, HZAZ18]. using [IK17, KOA15, LZL^{+}19, PHM^{+}12, PG11, SRD^{+}12, SK18b, SKS19, TPV18, Yil12, ZH14]. Utility [Tam18, TL19, UKW^{+}18]. Utility-Based [TL19, UKW^{+}18]. Utilization [HWXD14, NL19, SA11, TPV18]. Utilization-Based [NL19].

V2V [EMB19]. v3 [ZFL18]. Vacation [Dim13]. Validation [FPY15]. Value [ECL15, NS16, TZ11]. Value-Coded [ECL15]. Values [CGVP15, RHH12]. VANET [BP19]. VANETs [ALZ^{+}17]. Variable [GHXW16, JJO^{+}17, LR10]. Variable-to-Fixed [JJO^{+}17]. Variables [TKM11]. Variants [ASS15]. Variation [CK10]. Variation-Aware [CK10]. Variations [BY16]. Vector [AÇPD11, BdBG^{+}17, IJM14, JS15, LYPL17, PG11, PP17, TEP^{+}16, ZYT13, ZM16]. Vectors [OLL15, JJO^{+}17]. Vehicle [ALZ^{+}17, WYL^{+}13, XYL^{+}11]. Vehicles [CLLL17, LSY^{+}16]. Vehicular [FYF^{+}18, GH17, HLKL15, LHYW12, TPG^{+}15, UKW^{+}18, ZYR^{+}13]. Venues [SCT18a]. Verifiability [EKOS19]. Verifiable [RDZ^{+}16, SJ18a, YCR16]. Verification [BL15b, BL16, GdM16, GMSV14, JCSZ13, KSH^{+}14, LWYZ17, MV16, SHH^{+}15, XTH11]. Verifier [BDT10]. Verifying [AAHTH10, WNNZ17]. Version [KK18, KÖ14, KO15]. Vertex [DA14, KMNA^{+}16]. Vertex-Weighted [KMNA^{+}16]. Vertices [CFS13]. Very [YT16b]. Vese [MPP15]. Via [LHCN11, NHC13, PBH^{+}13, CCY10, CWCS14, KMNA^{+}16, KVX12, LYC11, LMMP16, OJSO14, PC12, WL13, WHS^{+}16]. Viable [HHHC16]. Video [BPK10, BWR12, CDYC11, DSBB19, GIB12, HZWT15, HM13, ISH13, LHYW12, LNBFPA13, PA15, RDB14b, WWHL12, ZDM^{+}15]. Video-Streaming [ISH13]. Videos [AFKT12, MMB13]. View [ZGC16, WWHL12]. Views [VGA15, ZLL^{+}14, ZZZ14]. Violation [HHHC16, HHH^{+}18]. Virtual [AD11, CFM17, Do11, FMRS17, HM16, HHV17, HLZ15, LP14, LMR18, LZL^{+}15, LWZ^{+}18, NL19, PXG^{+}17, PCC^{+}16, QZXR15, SL14, TV12, WRSV12, YAQ12, YHS^{+}17, ZZLL18]. Virtualization [AFG^{+}17, JWCZ13]. Virtualized [CWRZ18, LfHmXjL11, LJYL13, RAKJ17]. Virus [WOLP15]. Visibility [URHK19]. Visible [FFH17]. Vision [NLDH11, WRSV12]. Visual [ALA19, DD10a, HHS^{+}15, KTTRJ10, LPL15, LWL10, LTC^{+}15, WL13, ZXZ^{+}11, ZZLL18, NK19]. Visualization [URHK19]. Visually [MMB13]. VLIW [HYZ17]. VM [HXZ^{+}16, TPV18, YWR^{+}14]. VoD [LHFF13, LFHF14]. Voice [TJZF12]. Voice-over-IP [TJZF12]. Voltage [SSK12]. Volume [SAKOK11]. Volumetric [SBV19]. Voting [IEBS19, LGPRH14, TAC^{+}18, ZSJ10]. Voting-Based [ZSJ10]. vs [BGM^{+}13]. Vulnerabilities [CZC10]. Vulnerability [HS11, MKN13, VKC15]. VxWorks [YGH^{+}14].

W [Gaz10]. WaaS [EV16]. wait [TXJ^{+}19]. Walk [LZ19]. WalkSAT [CLS15]. Wan [RSD19]. Warning [HuRH^{+}15, UKW^{+}18]. Watermarking [BCPV11, Fra15, Joh10, NGAuHQ16, Shi08, YKK18, YYO15, ZXZ^{+}11]. Watson [ZTTM18]. Waveforms [YLSL19]. Wavefront [PHM^{+}12]. Wavelet [IS13, KTM19]. Way [CBJX19, HM14, JYL18, WCXZ17]. WBAN [KS18]. Weak [HJL10, PYM^{+}15]. Weaker [YLL^{+}17]. Weakest [YWY10]. Weakness [LLS17]. Wearable [BY16, CAV17]. web [KAZ18, AJ17, AHH13, BDC11, CDYC11, DG15a, Elg15, FLZC15, GSS14, QF19, SBBB12, SSY15, SP15, TU17, WXP^{+}10, WHSW15, YZJH12, ZGC16]. Web-Based [BDC11]. Web-Orchestrations [GSS14]. Websites [EA17]. Weierstrass [LL11a]. Weight [BS16, YWFQ18]. Weighted [KMNA^{+}16, QLZ18, THY^{+}18, WHYH12, YLL^{+}12, ZDM^{+}15, ZSJ10]. Weil [HH14]. WG [DJG^{+}15]. Where [Sab11]. White [BW16, LYL^{+}18]. White-Box [BW16, LYL^{+}18]. Whitewashing [SF17]. Wicked [SGH15]. Wide [CYTP18]. Width [DP16]. Wikipedia [SSS12a, WXZ^{+}12]. Window [MDY15, ZHL^{+}17]. Wired [BS10b, Uli11]. Wireless [AFG^{+}17, ACG^{+}11, ABG^{+}12, AKL^{+}19, AFGG11, AUB11, Ano17, BN14, BBM10, BSK19, BS10a, CLSV15, CCF11, CCC^{+}10, CL18, Cor11, DA14, DSTC12, Dim13, ER14, ELS11, FLWL19, dRFMD^{+}17, GPK11, HJS^{+}13, HLJ^{+}15, HLC10b, HZHC11, Hua14, HC15, HH14, IAG^{+}14, KAAE11, KTTRJ10, Kon10, LYPL17, LZL^{+}17, LTL10, LHM^{+}15, LWW13, LLDL17, LSCG10, MK13, OB18, RJS^{+}17, RDB^{+}14a, RSD19, RL11, SJ14, SM16, SZB19, SYH11, THP^{+}11, THP^{+}12, WZ17, WOV^{+}10, YWSH10, YDE11, YNN11, ZBY^{+}10, ZLYX10]. Within [GK17, Ano10]. Without [ASS15, CCL^{+}13, GSW^{+}16, GMS11, LTC^{+}15, YLA^{+}13, DCA18, LTW10, ZYM18]. Witold [Gaz10]. WIVET [TU17]. WLANs [IAG^{+}14, KKPB14]. Wolf [GA18]. Word [FNP12, LZL^{+}19]. Word-Based [FNP12]. Words [GdM16]. Work [HSZS17, HSZS18, NTSA16]. Workflow [EMTSM18, EV16, PB12, WLH15a]. Workflow-as-a-Service [EV16]. Workflows [EV16, VL13, WS15]. Working [YC19]. Workload [HSZS17, HSZS18, KV19, RAKJ17]. Workload-Aware [RAKJ17]. Workloads [NTSA16]. Workshop [Jar11]. World [ABG^{+}12, ABL^{+}18, ARVR15, Lav12, Lev10a, LZN^{+}16, Meg19, QYZ19, YZLC15]. Worm [WWC^{+}11, ZFZ12]. Worms [GDKP10]. Worn [BY14]. Wrangling [BPFK19, SAPS19]. WRANs [AZHASD14]. Wrapper [IK17]. Write [LLPY19, RRCC^{+}15, ZDZ^{+}15a]. Write-Aware [RRCC^{+}15]. Writes [GHXW16]. Writing [CXH14]. WS [MK15, aSPW^{+}17]. WS-BPEL [MK15, aSPW^{+}17]. WSN [AV16, RHG^{+}11]. WSNS [ABCG11, HL15, HK15]. WWW [JHHC15].

X [SBV19]. X-Ray [SBV19]. XML [ABS14, KKM^{+}15]. XPath [ZLL^{+}14].

Years [EEK17, Har11].

Zero [ST17, SW14, YCL17]. Zero-Clairvoyant [SW14]. Zero-correlation [YCL17]. Zero-Knowledge-Private [ST17]. ZIDS [NSMS14]. ZigBee [YNN11]. Zoned [LDK11]. Zooming [HNAS18].

References

Adnan:2019:OOS

[AAA19] Muhammad Adnan, Muhammad Afzal, and Khadim Hussain Asif. Ontology-oriented software effort estimation system for e-commerce applications based on extreme programming and scrum methodologies. The Computer Journal, 62(11):1605– 1624, November 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                                       academic.oup.com/comjnl/
                                                       article/62/11/1605/5319151.

Anagnostopoulos:2010:SCS

[AAH10] Christoforos Anagnostopoulos, Niall M. Adams, and David J. Hand. Streaming covariance selection with applications to adaptive querying in sensor networks. The Computer Journal, 53(9):1401–1414, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                                       comjnl.oxfordjournals.
                                                       org/content/53/9/1401.
                                                       full.pdf+html.

Akbarpour:2010:VSI

[AAHTH10] Behzad Akbarpour, Amr T.

Abdel-Hamid, Sofiène Tahar, and John Harrison. Verifying a synthesized implementation of IEEE-754 floatingpoint exponential function using HOL. The Computer Journal, 53(4):465– 488, May 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/4/465;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/4/465.

Agius:2013:MFS

[AAZ13] Harry Agius, Marios C. Angelides, and Damon Daylamani Zad. MC 2: a framework and service for MPEG-7 content-modelling communities. The Computer Journal, 56(5):593– 616, May 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/5/593.full.
pdf+html.

Anastasi:2011:HAP

[ABCG11] Giuseppe Anastasi, Eleonora Borgia, Marco Conti, and Enrico Gregori. A hybrid adaptive protocol for reliable data delivery in WSNS with multiple mobile sinks. The Computer Journal, 54(2):213–229, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/2/213.full.
                                        pdf+html.

Abdelli:2015:ISS

[Abd15] Abdelkrim Abdelli. Improving the state space computation of the time stream Petri nets model. The Computer Journal, 58(7):1607– 1627, July 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/7/1607.

Agarwal:2012:ASW

[ABG^{+}12] Rachit Agarwal, Abhik Banerjee, Vincent Gauthier, Monique Becker, Chai Kiat Yeo, and Bu Sung Lee. Achieving small-world properties using bio-inspired techniques in wireless networks. The Computer Journal, 55(8):909–931, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/909.full.
                                        pdf+html.

Alsaleh:2015:OMN

[ABH15] Omar Alsaleh, Bella Bose, and Bechir Hamdaoui. Oneto-many node-disjoint paths routing in dense Gaussian networks. The Computer Journal, 58(2):173– 187, February 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/2/173.

Archer:2018:KDR

[ABL^{+}18] David W. Archer, Dan Bogdanov, Yehuda Lindell, Liina Kamm, Kurt Nielsen, Jakob Illeborg Pagter, Nigel P. Smart, and Rebecca N. Wright. From keys to databases — realworld applications of secure multi-party computation. The Computer Journal, 61(12):1749–1771, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1749/5095655.

Abellanas:2012:MRC

[ABM12] Manuel Abellanas, Antonio Leslie Bajuelos, and Inês Matos. Minimizing the range for k-covered paths on sensor networks. The Computer Journal, 55(1):69–81, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

                                        org/content/55/1/69.full.
                                        pdf+html.

Antao:2012:RBE

[ABS12] Samuel Antão, Jean-Claude Bajard, and Leonel Sousa. RNS-based elliptic curve point multiplication for massive parallel architectures. The Computer Journal, 55(5):629–647, May 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/5/629.full.
                                        pdf+html.

Argyriou:2013:MTR

[ABS13] E. N. Argyriou, M. A. Bekos, and A. Symvonis. Maximizing the total resolution of graphs. The Computer Journal, 56(7):887– 900, July 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/7/887.full.
                                        pdf+html.

Amavi:2014:CXD

[ABS14] Joshua Amavi, Béatrice Bouchou, and Agata Savary. On correcting XML documents with respect to a schema. The Computer Journal, 57(5):639– 674, May 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/5/639.full.
pdf+html.

Agrigoroaiei:2014:RSI

[AC14] Oana Agrigoroaiei and Gabriel Ciobanu. Rewriting systems over indexed multisets. The Computer Journal, 57(1):165–179, January 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/165.full.
pdf+html.

Abdulahhad:2017:LLP

[ACB17] Karam Abdulahhad, Jean-Pierre Chevallet, and Catherine Berrut. Logics, lattices and probability: The missing links to information retrieval. The Computer Journal, 60(7):995– 1018, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/995/2608043.

Abdelhak:2011:EAD

[ACG^{+}11] Sherine Abdelhak, Rabi S.

Chaudhuri, Chandra S. Gurram, Soumik Ghosh, and Magdy Bayoumi. Energyaware distributed QR decomposition on wireless sensor nodes. The Computer Journal, 54(3):373– 391, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/3/373.full.
                                        pdf+html.

Apaydin:2011:NBN

[AÇPD11] Mehmet Serkan Apaydin, Bülent Çatay, Nicholas Patrick, and Bruce R. Donald. NVR-BIP: Nuclear vector replacement using binary integer programming for NMR structure-based assignments. The Computer Journal, 54(5):708– 716, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/5/708.full.
                                        pdf+html.

Ackley:2013:MAR

[ACW13] David H. Ackley, Daniel C. Cannon, and Lance R. Williams. A movable architecture for robust spatial computing. The Computer Journal, 56(12):1450–1468, December 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/12/1450.
                                        full.pdf+html.

Armstrong:2011:PIC

[AD11] Django Armstrong and Karim Djemame. Performance issues in clouds: an evaluation of virtual image propagation and I/O paravirtualization. The Computer Journal, 54(6):836– 849, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/836.full.
pdf+html.

Angelini:2016:SCP

[AD16] P. Angelini and G. Da Lozzo. SEFE = Cplanarity? The Computer Journal, 59(12):1831–1838, December 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/12/1831.

Abellanas:2013:LCP

[ADBPLV13] Manuel Abellanas, Jose M.

Díaz-Báñez, Pablo Pérez-Lantero, and Inmaculada Ventura. Locating a communication path in a competitive scenario. The Computer Journal, 56(7):819– 826, July 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/7/819.full.
pdf+html.

Al-Dujaily:2013:DCT

[ADML^{+}13] Ra’ed Al-Dujaily, Terrence Mak, Kai-Pui Lam, Fei Xia, Alex Yakovlev, and Chi-Sang Poon. Dynamic on-chip thermal optimization for three-dimensional networks-on-chip. The Computer Journal, 56(6): 756–770, June 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/6/756.full.
                                        pdf+html.

Attrapadung:2015:RGS

[AEHS15] Nuttapong Attrapadung, Keita Emura, Goichiro Hanaoka, and Yusuke Sakai. Revocable group signature with constant-size revocation list. The Computer Journal, 58(10):2698–2715, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2698.

Abdelaziz:2017:SDW

[AFG^{+}17] Ahmed Abdelaziz, Ang Tan Fong, Abdullah Gani, Suleman Khan, Faiz Alotaibi, and Muhammad Khurram Khan. On softwaredefined wireless network (SDWN) network virtualization: Challenges and open issues. The Computer Journal, 60(10):1510–1519, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1510/4321712.

Aiello:2011:JBA

[AFGG11] Francesco Aiello, Giancarlo Fortino, Raffaele Gravina, and Antonio Guerrieri. A Java-based agent platform for programming wireless sensor networks. The Computer Journal, 54(3):439– 454, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/439.full.
pdf+html.

Akamine:2012:FAE

[AFKT12] Kazuma Akamine, Ken Fukuchi, Akisato Kimura, and Shigeru Takagi. Fully automatic extraction of salient objects from videos in near real time. The Computer Journal, 55(1):3– 14, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/1/3.full.
pdf+html.

Abdelrahman:2012:PDE

[AG12] Omer H. Abdelrahman and Erol Gelenbe. Packet delay and energy consumption in non-homogeneous networks. The Computer Journal, 55(8):950–964, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/950.full.
                                        pdf+html.

Abbasoglu:2015:APC

[AGF15] Mehmet Alİ Abbasoğlu, Buğra Gedİk, and Hakan Ferhatosmanoğlu. Aggregate profile clustering for streaming analytics. The Computer Journal, 58(9):2092–2108, September 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/2092.

Araujo:2016:EEC

[AGM^{+}16] J. Araujo, F. Giroire, J. Moulierac, Y. Liu, and R. Modrzejewski. Energy efficient content distribution. The Computer Journal, 59(2):192–207, February 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/2/192.

Allani:2010:RAM

[AGP10] Mouna Allani, Benoît Garbinato, and Fernando Pedone.

Resource-aware multimedia content delivery: a gambling approach. The Computer Journal, 53(2):234– 248, February 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/2/234;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/2/234.

Arcaini:2015:HOU

[AGR15] Paolo Arcaini, Angelo Gargantini, and Elvinia Riccobene. How to optimize the use of SAT and SMT solvers for test generation of Boolean expressions. The Computer Journal, 58(11): 2900–2920, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Alvarez:2018:CDP

[ÁHFE18] Pedro Álvarez, Sergio Hernández, Javier Fabra, and Joaquín Ezpeleta. Cost-driven provisioning and execution of a computing-intensive service on the Amazon EC2. The Computer Journal, 61 (9):1407–1421, September 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1407/4835632.

Ashraf:2013:FMO

[AHH13] Jamshaid Ashraf, Omar Khadeer Hussain, and Farookh Khadeer Hussain. A framework for measuring ontology usage on the Web. The Computer Journal, 56(9):1083–1101, September 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/9/1083.
                                        full.pdf+html.

Aliakbary:2015:FED

[AHM15] Sadegh Aliakbary, Jafar Habibi, and Ali Movaghar. Feature extraction from degree distribution for comparison and analysis of complex networks. The Computer Journal, 58(9):2079– 2091, September 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/9/2079.

Aho:2012:CCT

[Aho12]  Alfred V. Aho.  Com-

putation and computational thinking. The Computer Journal, 55(7):832– 835, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/7/832.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Ahmadifar:2015:NRN

[AJ15] H. Ahmadifar and G. Jaberipur.

A new residue number system with 5-moduli set: 2^{2q}, 2^{q} ± 3, 2^{q} ± 1. The Computer Journal, 58(7):1548– 1565, July 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/7/1548.

AlFayez:2017:ULD

[AJ17] Reem Q. Al Fayez and Mike Joy. Using linked data for integrating educational medical Web databases based on BioMedical ontologies. The Computer Journal, 60(3):60– ??, March 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/3/369/2632623/Using-
Linked-Data-for-Integrating-
Educational.

Alhanahnah:2016:MTI

[AJA16] Mohannad J. Alhanahnah, Arshad Jhumka, and Sahel Alouneh. A multidimension taxonomy of insider threats in cloud computing. The Computer Journal, 59(11):1612–1622, November 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/11/1612.

Almendros-Jimenez:2019:IQO

[AJBTT19] Jesús M. Almendros-Jiménez, Antonio Becerra-Terón, and Manuel Torres. Integrating and querying Open-StreetMap and Linked Geo Open data. The Computer Journal, 62(3):321– 345, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/321/4210213.

Aggarwal:2012:DTT

[AK12] Supriya Aggarwal and Kavita Khare. Design techniques targeting low-area-power-delay product in hyperbolic CORDIC algorithm. The Computer Journal, 55(5): 616–628, May 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/5/616.full.
                                        pdf+html.

Affleck:2015:NFR

[AKA15] Amy Affleck, Aneesh Krishna, and Narasimaha R. Achuthan. Non-functional requirements framework: a mathematical programming approach. The Computer Journal, 58(5):1122–1139, May 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1122.

Ahmed:2019:AWS

[AKL^{+}19] Faisal Ahmed, Corentin Kervadec, Yannick Le Moullec, Gert Tamberg, and Paul Annus. Autonomous wireless sensor networks: Implementation of transient computing and energy prediction for improved node performance and link quality. The Computer Journal, 62(6):820– 837, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/820/5123536.

Amri:2019:PVA

[ALA19] Saber Amri, Hela Ltifi, and Mounir Ben Ayed. A predictive visual analytics evaluation approach based on adaptive neuro-fuzzy inference system. The Computer Journal, 62(7):977– 1000, July 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/7/977/5110549.

Acharya:2017:CPT

[ALH17] Saurav Acharya, Byung Suk Lee, and Paul Hines. Causal prediction of top-k event types over real-time event streams. The Computer Journal, 60(11):1561–1581, November 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/11/1561/3051819.

Alhumoud:2019:TAI

[Alh19] Sarah Alhumoud. Twitter analysis for intelligent transportation. The Computer Journal, 62(11):1547–1556, November 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/11/1547/5253756.

Almasaeid:2019:MAT

[Alm19] Hisham M Almasaeid. Maximizing achievable transmission time in cognitive radio networks under sensoraided crowdsourced spectrum sensing. The Computer Journal, 62(10):1477– 1489, October 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/10/1477/5487027.

Au:2017:AAS

[ALZ^{+}17] Man Ho Au, Joseph K. Liu, Zhenfei Zhang, Willy Susilo, Jin Li, and Jianying Zhou. Anonymous Announcement System (AAS) for electric vehicle in VANETs. The Computer Journal, 60(4):588–599, March 23, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/588/2433260.

Anashin:2010:NAE

[Ana10] Vladimir Anashin. Non-Archimedean ergodic theory and pseudorandom generators. The Computer Journal, 53(4):370– 392, May 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/4/370;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/4/370.

Angelides:2013:EMA

[Ang13] Marios C. Angelides. Editorial: MPEG applications and services. The Computer Journal, 56(5):527– 528, May 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/5/527.full.
pdf+html.

Anonymous:2010:ISI

[Ano10] Anonymous. Introduction to the Special Issue on Advances in Sensing, Information Processing and Decision Making for Coalition Operations within the US/ UK International Technology Alliance. The Computer Journal, 53(5):491– 492, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/5/491.

Anonymous:2017:SIS

[Ano17] Anonymous. Special issue on software-defined wireless networks. The Computer Journal, 60(10):1415–1416, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1415/4191408.

Ammann:2008:IST

[AO08] Paul Ammann and Jeff Offutt. Introduction to Software Testing. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88038-6 (hardback). xxii + 322 pp. LCCN QA76.76.T48 A56

2008. URL http://www.
loc.gov/catdir/enhancements/
fy0803/2007035077-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0803/2007035077-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0803/2007035077-t.html.

Abid:2015:RDB

[AOS^{+}15] S. A. Abid, Mazliza Othman, Nadir Shah, Mazhar Ali, and A. R. Khan. 3D-RP: a DHT-based routing protocol for MANETs. The Computer Journal, 58(2):258–279, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/258.

Asghar:2011:HSM

[APW11] Hassan Jameel Asghar, Josef Pieprzyk, and Huaxiong Wang. On the hardness of the sum of k mins problem. The Computer Journal, 54(10):1652–1660, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1652.
full.pdf+html.

Andrade:2016:AIF

[ARR^{+}16] Rodrigo Andrade, Márcio Ribeiro, Henrique Rebêlo, Paulo Borba, Vaidas Gasiunas, and Lucas Satabin. Assessing idioms for a flexible feature binding time. The Computer Journal, 59(1):1– 32, January 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/1/1.

Arsuaga-Rios:2015:MSW

[ARVR15] María Arsuaga-Ríos and Miguel A. Vega-Rodríguez. Multiobjective small-world optimization for energy saving in Grid environments. The Computer Journal, 58(3):432–447, March 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/3/432.

Ahmadinia:2011:HAE

[AS11] Ali Ahmadinia and Alireza Shahrabi. A highly adaptive and efficient router architecture for network-on-chip. The Computer Journal, 54(8):1295–1307, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/8/1295.
                                        full.pdf+html.

Ait-Salaht:2016:PAQ

[ASCTFP16] Farah Aït-Salaht, Hind Castel-Taleb, Jean-Michel Fourneau, and Nihal Pekergin. Performance analysis of a queue by combining stochastic bounds, real traffic traces and histograms. The Computer Journal, 59(12):1817–1830, December 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/12/1817.

Ayyildiz:2015:DSD

[ASG15] Nizam Ayyıldız, Ece Güran Schmidt, and Hasan Cengiz Güran. S-DIRECT: Scalable and dynamically reconfigurable TCAM architecture for high-speed IP lookup. The Computer Journal, 58(6):1443–1455, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1443.

Sun:2017:ESM

[aSPW^{+}17] Chang ai Sun, Lin Pan, Qiaoling Wang, Huai Liu, and Xiangyu Zhang. An empirical study on mutation testing of WS-BPEL programs. The Computer Journal, 60(1):143–158, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Asaar:2015:IBM

[ASS15] Maryam Rajabzadeh Asaar, Mahmoud Salmasizadeh, and Willy Susilo. An identity-based multi-proxy multi-signature scheme without bilinear pairings and its variants. The Computer Journal, 58(4):1021– 1039, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/1021.

Angulu:2019:AGE

[ATA19] Raphael Angulu, Jules R. Tapamo, and Aderemi O. Adewumi. Age-group estimation using feature and decision level fusion. The Computer Journal, 62(3):346–358, March 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/346/4995614.

Alam:2015:ACF

[ATS15] Shahid Alam, Issa Traore, and Ibrahim Sogukpinar. Annotated control flow graph for metamorphic malware detection. The Computer Journal, 58(10):2608– 2621, October 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/10/2608.

Akyurek:2011:DOL

[AUB11] A. Sinan Akyurek and Elif Uysal-Biyikoglu. A depthoptimal low-complexity distributed wireless multicast algorithm. The Computer Journal, 54(6):988–1003, June 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/988.full.
pdf+html.

Akila:2016:FBE

[AV16] I. S. Akila and R. Venkatesan. A fuzzy based energyaware clustering architecture for cooperative communication in WSN. The Computer Journal, 59(10):1551– 1562, October 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/10/1551.

Awan:2013:EPE

[Awa13]  Irfan Awan.  Editorial:

Performance engineering of communication systems and applications. The Computer Journal, 56(2):139– 140, February 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/2/139.full.
pdf+html.

Al-Zubi:2014:MBD

[AZHASD14] Raed Al-Zubi, Mohammed Hawa, Ghazi Al-Sukkar, and Khalid A. Darabkh. Markov-based distributed approach for mitigating self-coexistence problem in IEEE 802.22 WRANs. The Computer Journal, 57(12):1765–1775, December 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/12/1765.

Bacon:2012:CFP

[Bac12] Dave Bacon. Computation and fundamental physics. The Computer Journal, 55 (7):826–829, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/7/826.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Boo:2013:EAD

[BACD13] M. Bóo, M. Amor, R. Concheiro, and M. Doggett. Efficient adaptive and dynamic mesh refinement based on a nonrecursive strategy. The Computer Journal, 56(7): 843–851, July 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/7/843.full.
                                        pdf+html.

Ben-Asher:2011:DMA

[BAFF11] Yosi Ben-Asher, Sharoni Feldman, and Moran Feldman. Dynamic multipath allocation in ad hoc networks. The Computer Journal, 54(2):197–212, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/2/197.full.
pdf+html.

Bajcsy:2012:CI

[Baj12] Ruzena Bajcsy. Computation and information. The Computer Journal, 55(7): 825, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/825.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Barron:2011:EPR

[Bar11] David Barron. EDSAC:

a programmer remembers. The Computer Journal, 54(1):139–142, January 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/139.full.
pdf+html.

Beerenwinkel:2015:CPD

[BBB^{+}15] Niko Beerenwinkel, Stefano Beretta, Paola Bonizzoni, Riccardo Dondi, and Yuri Pirola. Covering pairs in directed acyclic graphs. The Computer Journal, 58(7): 1673–1686, July 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/7/1673.

Balbo:2011:FPT

[BBDF11] Gianfranco Balbo, Marco Beccuti, Massimiliano De Pierro, and Giuliana Franceschinis. First passage time computation in tagged GSPNs with queue places. The Computer Journal, 54 (5):653–673, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/5/653.full.
pdf+html.

Bessiere:2014:GCD

[BBGM14] Christian Bessiere, Ismel Brito, Patricia Gutierrez, and Pedro Meseguer. Global constraints in distributed constraint satisfaction and optimization. The Computer Journal, 57(6):906– 923, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/6/906.full.
pdf+html.

Bhuyan:2011:SPS

[BBK11] Monowar H. Bhuyan, D. K.

Bhattacharyya, and J. K. Kalita. Surveying port scans and their detection methodologies. The Computer Journal, 54(10):1565–1581, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1565.
full.pdf+html.

Bingol:2019:EPP

[BBKL19] Muhammed Ali Bingöl, Osman Biçer, Mehmet Sabir Kiraz, and Albert Levi. An efficient 2-party private function evaluation protocol based on half gates. The Computer Journal, 62(4): 598–613, April 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/598/5259181.

Baccelli:2010:TSO

[BBM10] François Baccelli, Bartlomiej Blaszczyszyn, and Paul Mühlethaler. Time– space opportunistic routing in wireless ad hoc networks: Algorithms and performance optimization by stochastic geometry. The Computer Journal, 53(5): 592–609, June 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/5/592;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/5/592.

Bajuelos:2014:GOG

[BBM14] Antonio L. Bajuelos, Sergey Bereg, and Mafalda Martins. Guarding orthogonal galleries with rectangular rooms. The Computer Journal, 57(11):1668–1673, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/11/1668.

Balasundaram:2017:IRT

[BBM17] Prabavathy Balasundaram, Chitra Babu, and Subha Devi M. Improving read throughput of deduplicated cloud storage using frequent patternbased prefetching technique. The Computer Journal, 60 (3):60–??, March 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/3/444/2609375/
                                        Improving-Read-Throughput-
                                        of-Deduplicated-Cloud.

Bergstra:2013:GE

[BBMW13] Jan Bergstra, Jens Blanck, Faron Moller, and Stan Wainer. Guest editorial. The Computer Journal, 56 (1):2, January 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/1/2.full.
pdf+html.

Bergstra:2013:CMG

[BBP13] Jan A. Bergstra, Inge Bethke, and Alban Ponse. Cancellation meadows: a generic basis theorem and some applications. The Computer Journal, 56(1):3– 14, January 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/1/3.full.
pdf+html.

Barjon:2019:MDS

[BCC^{+}19] Matthieu Barjon, Arnaud Casteigts, Serge Chaumette, Colette Johnen, and Yessin M. Neggaz. Maintaining a distributed spanning forest in highly dynamic networks. The Computer Journal, 62(2):231–246, February 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/231/5051848.

Bahi:2012:SCS

[BCG12] Jacques M. Bahi, Jean-François Couchot, and Christophe Guyeux. Steganography: a class of secure and robust algorithms. The Computer Journal, 55(6): 653–666, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/6/653.full.
pdf+html.

Bajuelos:2015:SRO

[BCH^{+}15] Antonio Leslie Bajuelos, Santiago Canales, Gregorio Hernández, Mafalda Martins, and Inês Matos. Some results on open-edge and open mobile guarding of polygons and triangulations. The Computer Journal, 58(1):160– 171, January 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/160.

Bozzano:2011:SDP

[BCK^{+}11] Marco Bozzano, Alessandro Cimatti, Joost-Pieter Katoen, Viet Yen Nguyen, Thomas Noll, and Marco Roveri. Safety, dependability and performance analysis of extended AADL models. The Computer Journal, 54(5):754–775, May 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

org/content/54/5/754.full.
pdf+html.

Bonnaire:2017:APT

[BCKM17] Xavier Bonnaire, Rudyar Cortés, Fabrice Kordon, and Olivier Marin. ASCENT: a provably terminating decentralized logging service. The Computer Journal, 60 (12):1889–1911, December 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1889/4107197.

Basso:2011:BWC

[BCPV11] Alessandro Basso, Davide Cavagnino, Victor Pomponiu, and Annamaria Vernone. Blind watermarking of color images using Karhunen–Loève transform keying. The Computer Journal, 54(7):1076– 1090, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1076.
full.pdf+html.

Babaee:2014:DMC

[BD14] Arta Babaee and Moez Draief. Distributed multivalued consensus. The Computer Journal, 57(8):1132– 1140, August 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/8/1132.

Binucci:2016:CQU

[BD16] Carla Binucci and Walter Didimo. Computing quasiupward planar drawings of mixed graphs. The Computer Journal, 59(1):133– 150, January 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/1/133.

Brisaboa:2017:EQV

[BdBG^{+}17] Nieves R. Brisaboa, Guillermo de Bernardo, Gilberto Gutiérrez, Miguel R. Luaces, and José R. Paramá. Efficiently querying vector and raster data. The Computer Journal, 60(9): 1395–1413, September 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/9/1395/2996415.

Bayir:2011:WBP

[BDC11] Murat Ali Bayir, Murat Demirbas, and Ahmet Cosar. A Web-based personalized mobility service for Smartphone applications. The Computer Journal, 54 (5):800–814, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/5/800.full.
pdf+html.

Both:2013:DMM

[BDL^{+}13] Alan Both, Matt Duckham, Patrick Laube, Tim Wark, and Jeremy Yeoman. Decentralized monitoring of moving objects in a transportation network augmented with checkpoints. The Computer Journal, 56(12):1432– 1449, December 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/12/1432.
full.pdf+html.

Beal:2013:E

[BDMS13] Jacob Beal, Stefan Dulman, Olivier Michel, and Antoine Spicher. Editorial. The Computer Journal, 56(12):1397–1398, December 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/12/1397.
full.pdf+html.

Barbuti:2010:AIA

[BDT10] Roberto Barbuti, Nicoletta De Francesco, and Luca Tesei. An abstract interpretation approach for enhancing the Java Bytecode Verifier. The Computer Journal, 53 (6):679–700, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/679;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/679.

Burgin:2012:EAE

[BE12] Mark Burgin and Eugene Eberbach. Evolutionary automata: Expressiveness and convergence of evolutionary computation. The Computer Journal, 55(9):1023– 1029, September 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/9/1023.
full.pdf+html.

Bahi:2016:RDS

[BEG^{+}16] J. Bahi, W. Elghazel, C. Guyeux, M. Haddad, M. Hakem, K. Medjaher, and N. Zerhouni. Resiliency in distributed sensor networks for prognostics and health management of the monitoring targets. The Computer Journal, 59(2):275–284, February 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/2/275.

Balliu:2019:CCL

[BF19] Alkida Balliu and Pierre Fraigniaud. Certification of compact low-stretch routing schemes. The Computer Journal, 62(5):730– 746, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/730/5085053.

Beccuti:2014:COR

[BFCRH14] Marco Beccuti, Giuliana Franceschinis, Daniele Codetta-Raiteri, and Serge Haddad. Computing optimal repair strategies by means of NdRFT modeling and analysis. The Computer Journal, 57(12):1870–1892, December 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/12/1870.

Beccuti:2015:SND

[BFF^{+}15] Marco Beccuti, Chiara Fornari, Giuliana Franceschinis, Sami M. Halawani, Omar Ba-Rukab, Ab Rahman Ahmad, and Gianfranco Balbo. From symmetric nets to differential equations exploiting model symmetries. The Computer Journal, 58(1):23–39, January 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/1/23.

Blanco:2015:MSO

[BFMT15] Carlos Blanco, Eduardo Fernández-Medina, and Juan Trujillo. Modernizing secure OLAP applications with a model-driven approach. The Computer Journal, 58(10):2351–2367, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2351.

Bankas:2015:NMA

[BG15] Edem Kwedzo Bankas and Kazeem Alagbe Gbolagade. New MRC adder-based reverse converter for the moduli set 2^{n}, 2^{2n+1} − 1, 2^{2n+2} − 1. The Computer Journal, 58(7):1566–1572, July 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/7/1566.

Berl:2010:EEC

[BGD^{+}10] Andreas Berl, Erol Gelenbe, Marco Di Girolamo, Giovanni Giuliani, Hermann De Meer, Minh Quan Dang, and Kostas Pentikousis. Energy-efficient cloud computing. The Computer Journal, 53(7):1045–1051, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/1045;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1045.

Bertran:2011:LMD

[BGM^{+}11] Ramon Bertran, Marc Gonzàlez, Xavier Martorell, Nacho Navarro, and Eduard Ayguadé. Local memory design space exploration for high-performance computing. The Computer Journal, 54(5):786– 799, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/5/786.full.
pdf+html.

Bertran:2013:CBP

[BGM^{+}13] Ramon Bertran, Marc Gonzàlez, Xavier Martorell, Nacho Navarro, and Eduard Ayguadé. Counterbased power modeling methods: Top-down vs. bottomup. The Computer Journal, 56(2):198–213, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/2/198.full.
pdf+html.

Broda:2010:DEP

[BH10] Krysia Broda and Christopher J. Hogger. Designing effective policies for minimal agents. The Computer Journal, 53(8):1184–1209, October 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/8/1184.
                                        full.pdf+html.

Badaroglu:2010:CRN

[BHAC10] Mustafa Badaroğlu, Ugur Halici, Isik Aybay, and Cuneyt Cerkez. A cascadable random neural network chip with reconfigurable topology. The Computer Journal, 53(3):289– 303, March 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/3/289;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/3/289.

Benoit:2010:MCS

[BHR10] Anne Benoit, Mourad Hakem, and Yves Robert. Multicriteria scheduling of precedence task graphs on heterogeneous platforms. The Computer Journal, 53(6): 772–785, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/772;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/772.

Bauer:2011:RTS

[BJY11] Andreas Bauer, Jan Jürjens, and Yijun Yu. Run-time security traceability for evolving systems. The Computer Journal, 54(1):58–87, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/58.full.
pdf+html.

Baier:2008:PMC

[BK08] Christel Baier and Joost-Pieter Katoen. Principles of Model Checking. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-02649-X (hardcover). xvii + 975 pp. LCCN QA76.76.V47 B35 2008. URL http://

www.loc.gov/catdir/toc/
ecip0727/2007037603.html.

Bossard:2012:NSD

[BK12a] Antoine Bossard and Keiichi Kaneko. Node-to-set disjoint-path routing in hierarchical cubic networks. The Computer Journal, 55(12):1440–1446, December 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/12/1440.
                                        full.pdf+html.

Bossard:2012:SSD

[BK12b] Antoine Bossard and Keiichi Kaneko. The setto-set disjoint-path problem in perfect hierarchical hypercubes. The Computer Journal, 55(6):769– 775, June 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/6/769.full.
                                        pdf+html.

Bossard:2014:SSD

[BK14] Antoine Bossard and Keiichi Kaneko. Set-to-set disjoint paths routing in hierarchical cubic networks. The Computer Journal, 57(2):332– 337, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/2/332.full.
                                        pdf+html.

Bhuyan:2014:DDD

[BKBK14] Monowar H. Bhuyan, H. J. Kashyap, D. K. Bhattacharyya, and J. K. Kalita. Detecting distributed denial of service attacks: Methods, tools and future directions. The Computer Journal, 57(4):537– 556, April 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/4/537.full.
pdf+html.

Batista:2019:SOS

[BKFP19] Bruno Guazzelli Batista, Bruno Tardiole Kuehne, Dionisio Machado Leite Filho, and Maycon Leone Maciel Peixoto. Security overhead on a service with automatic resource management: a performance analysis. The Computer Journal, 62(2):161–173, February 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/161/4964852.

Bossard:2011:NNS

[BKP11] Antoine Bossard, Keiichi Kaneko, and Shietung Peng. A new node-to-set disjointpath algorithm in perfect hierarchical hypercubes. The Computer Journal, 54(8):1372–1381, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/8/1372.
full.pdf+html.

Bekos:2010:AFB

[BKPS10] Michael A. Bekos, Michael Kaufmann, Katerina Potika, and Antonios Symvonis. Area-feature boundary labeling. The Computer Journal, 53(6):827–841, July 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/6/827;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/6/827.

Bade:2011:ABE

[BL11] Dirk Bade and Winfried Lamersdorf. An agent-based event processing middleware for sensor networks and RFID systems. The Computer Journal, 54(3):321– 331, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/3/321.full.
                                        pdf+html.

Badia:2015:FDN

[BL15a] Antonio Badia and Daniel Lemire. Functional dependencies with null markers. The Computer Journal, 58(5):1160–1168, May 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1160.

Bidgoly:2015:MQV

[BL15b] Amir Jalaly Bidgoly and Behrouz Tork Ladani. Modelling and quantitative verification of reputation systems against malicious attackers. The Computer Journal, 58(10):2567–2582, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2567.

Bidgoly:2016:MQV

[BL16] Amir Jalaly Bidgoly and Behrouz Tork Ladani. Modeling and quantitative verification of trust systems against malicious attackers. The Computer Journal, 59(7):1005–1027, July 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/7/1005.

Blanck:2013:IDC

[Bla13] Jens Blanck. Interval domains and computable sequences: a case study of domain reductions. The Computer Journal, 56(1):45–52, January 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/1/45.full.
                                        pdf+html.

Blin:2010:HRH

[BLRT10] Lélia Blin, Christian Laforest, Stephane Rovedakis, and Nicolas Thibault. Hardness results and heuristic for multi-groups interconnection. The Computer Journal, 53(9):1497–1507, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/9/1497.
                                        full.pdf+html.

Brognoli:2016:IDE

[BLS16] Simone Brognoli, Gianfranco Lamperti, and Michele Scandale. Incremental determinization of expanding automata. The Computer Journal, 59(12):1872–1899, December 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/12/1872.

Budanur:2012:MTC

[BMG12] Sandeep Budanur, Frank Mueller, and Todd Gamblin. Memory trace compression and replay for SPMD systems using extended PRSDs. The Computer Journal, 55(2):206– 217, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/2/206.full.
pdf+html.

Basterrech:2011:LMT

[BMRS11] Sebastián Basterrech, Samir Mohammed, Gerardo Rubino, and Mostafa Soliman. Levenberg–Marquardt training algorithms for random neural networks. The Computer Journal, 54(1): 125–135, January 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/1/125.full.
pdf+html.

Babamir:2014:AKP

[BN14] Faezeh Sadat Babamir and Ali Norouzi. Achieving key privacy and invisibility for unattended wireless sensor networks in healthcare. The Computer Journal, 57(4): 624–635, April 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/4/624.full.
pdf+html.

Bradley:2010:URM

[BP10] Patrick Erik Bradley and Norbert Paul. Using the relational model to capture topological information of spaces. The Computer Journal, 53(1):69–89, January 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/69;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/69.

BrijilalRuban:2019:CBS

[BP19] C. BrijilalRuban and B. Paramasivan. Cluster-based secure communication and certificate revocation scheme for VANET. The Computer Journal, 62(2):263– 275, February 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/263/5068285.

Blin:2016:NSS

[BPBRT16] Lélia Blin, Maria Potop-Butucaru, Stéphane Rovedakis, and Sébastien Tixeuil. A new self-stabilizing minimum spanning tree construction with loop-free property. The Computer Journal, 59(2):225– 243, February 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/2/225.

Bogatu:2019:TAD

[BPFK19] Alex Bogatu, Norman W.

Paton, Alvaro A. A. Fernandes, and Martin Koehler. Towards automatic data format transformations: Data wrangling at scale. The Computer Journal, 62(7): 1044–1060, July 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/7/1044/5224764.

Benois-Pineau:2010:SDR

[BPK10] Jenny Benois-Pineau and Andrei Khrennikov. Significance delta reasoning with p-adic neural networks: Application to shot change detection in video. The Computer Journal, 53(4):417– 431, May 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/4/417;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/4/417.

Bradley:2010:MD

[Bra10] Patrick Erik Bradley. Mumford dendrograms. The Computer Journal, 53(4): 393–404, May 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/

                                        53/4/393;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/4/393.

Bradley:2011:CPE

[Bra11] Jeremy T. Bradley. Computer performance evaluation: Preface. The Computer Journal, 54(5):641– 642, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/5/641.full.
                                        pdf+html.

Broy:2010:LBC

[Bro10] Manfred Broy. A logical basis for component-oriented software and systems engineering. The Computer Journal, 53(10):1758–1782, December 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/10/1758.
                                        full.pdf+html.

Boonma:2010:MMD

[BS10a] Pruet Boonma and Junichi Suzuki. Moppet: a model-driven performance engineering framework for wireless sensor networks. The Computer Journal, 53(10):1674–1690, December 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/10/1674.
full.pdf+html.

Brynjolfsson:2010:WIH

[BS10b] Erik Brynjolfsson and Adam Saunders. Wired for innovation: how information technology is reshaping the economy. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01366-5 (hardcover). xvii + 154 pp. LCCN HC79.T4 B79 2010.

Bhattacharya:2016:MMM

[BS16] Sangeeta Bhattacharya and S. Selvakumar. Multimeasure multi-weight ranking approach for the identification of the network features for the detection of DoS and probe attacks. The Computer Journal, 59(6): 923–943, June 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/6/923.

Bayoumi:2019:IMD

[BSK19] Mohamed A Bayoumi, Tarek M Salem, and Samir M Koriem. Improving the mechanism of detecting and measuring holes in ad hoc wireless sensor network. The Computer Journal, 62(10):1505–1514, October 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/10/1505/5540369.

Barnett:2018:ADR

[BT18] Jeremy Barnett and Philip Treleaven. Algorithmic dispute resolution — the automation of professional dispute resolution using AI and blockchain technologies. The Computer Journal, 61(3):399–408, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/3/399/4608879.

Bradley:2012:IRC

[BTHS12] Jeremy T. Bradley, Nigel Thomas, Richard A. Hayden, and Anton Stefanek. Invited response to Computer Journal Lecture by Prof. Jane Hillston. The Computer Journal, 55(7): 882–886, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/7/882.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Benveniste:2011:CIL

[BÜ11] Rıfat Benveniste and Cem Ünsalan. A color invariant for line stripe-based range scanners. The Computer Journal, 54(5):738– 753, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/5/738.full.
pdf+html.

Beal:2013:EST

[BUB13] Jacob Beal, Kyle Usbeck, and Brett Benyo. On the evaluation of space–time functions. The Computer Journal, 56(12):1500–1517, December 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/12/1500.
full.pdf+html.

Buzen:2012:CUR

[Buz12] Jeffrey P. Buzen. Computation, uncertainty and risk. The Computer Journal, 55 (7):838–847, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/838.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Balbo:2015:AMM

[BV15] Gianfranco Balbo and Maria Grazia Vigliotti. On the analysis of a M/M/1 queue with bulk services. The Computer Journal, 58(1):57–74, January 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/1/57.

Baek:2013:SPK

[BVS^{+}13] Joonsang Baek, Quang Hieu Vu, Abdulhadi Shoufan, Andrew Jones, and Duncan S. Wong. Stateful public-key encryption schemes forward-secure against state exposure. The Computer Journal, 56(4):497– 507, April 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/4/497.full.
                                        pdf+html.

Bai:2016:ALC

[BW16] Kunpeng Bai and Chuankun Wu. An AES-like cipher and its white-box implementation. The Computer Journal, 59(7):1054–1065, July 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/7/1054.

Baek:2016:EGC

[BWLA16] Joonsang Baek, Duncan S. Wong, Jin Li, and Man Ho Au. Efficient generic construction of CCA-secure identity-based encryption from randomness extraction. The Computer Journal, 59(4):508–521, April 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/508.

Bhatnagar:2012:IVE

[BWR12] Gaurav Bhatnagar, Q. M. Jonathan Wu, and Balasubramanian Raman. Image and video encryption based on dual space-filling curves. The Computer Journal, 55(6): 667–685, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/6/667.full.
pdf+html.

Barshan:2014:RDS

[BY14] Billur Barshan and Murat Cihan Yüksek. Recognizing daily and sports activities in two open source machine learning environments using body-worn sensor units. The Computer Journal, 57(11):1649–1667, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1649.

Barshan:2016:IIS

[BY16] Billur Barshan and Aras Yurtman. Investigating inter-subject and interactivity variations in activity recognition using wearable motion sensors. The Computer Journal, 59(9):1345–1362, September 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/9/1345.

Bao:2016:OGA

[BZS^{+}16] Liang Bao, Fen Zhao, Mengqing Shen, Yutao Qi, and Ping Chen. An orthogonal genetic algorithm for QoS-aware service composition. The Computer Journal, 59(12):1857–1871, December 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/12/1857.

Cai:2012:RFB

[Cai12] Jinhai Cai. Robust filteringbased thinning algorithm for pattern recognition. The Computer Journal, 55(7): 887–896, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/887.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Calamoneri:2011:LPO

[Cal11a] Tiziana Calamoneri. The L(2, 1)-labeling problem on oriented regular grids. The Computer Journal, 54(11): 1869–1875, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Calamoneri:2011:LPU

[Cal11b] Tiziana Calamoneri. The L(h, k)-labelling problem: an updated survey and annotated bibliography. The Computer Journal, 54(8): 1344–1371, August 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/54/8/1344.
full.pdf+html.

Cao:2010:HBE

[Cao10] Yongzhi Cao. A hierarchy of behavioral equivalences in the π-calculus with noisy channels. The Computer Journal, 53(1):3–20, January 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/3;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/3.

Cao:2014:NIL

[Cao14] Longbing Cao. Non-IIDness learning in behavioral and social data. The Computer Journal, 57(9):1358–1370, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1358.

Cola:2017:RTI

[CAV17] Guglielmo Cola, Marco Avvenuti, and Alessio Vecchio. Real-time identification using gait pattern analysis on a standalone wearable accelerometer. The Computer Journal, 60(8):1173–1186, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/8/1173/2861346.

Chehreghani:2018:DDB

[CBA18] Mostafa Haghir Chehreghani, Albert Bifet, and Talel Abdessalem. Discriminative distance-based network indices with application to link prediction. The Computer Journal, 61(7):998– 1014, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/998/4985553.

Chang:2019:KCS

[CBJX19] Jinyong Chang, Genqing Bian, Yanyan Ji, and Maozhi Xu. On the KDM– CCA security from partial trapdoor one-way family in the random oracle model. The Computer Journal, 62(8):1232–1245, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1232/5492772.

Cakir:2011:MMC

[CC11] Serdar Cakir and A. Enis Cetin. Mel- and Mellincepstral feature extraction algorithms for face recognition. The Computer Journal, 54(9):1526–1534, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1526.
full.pdf+html.

Chang:2014:SDT

[CC14] Chin-Chen Chang and Ting-Fang Cheng. A secure diverse ticket-sale system in a distributed cloud environment. The Computer Journal, 57(10):1441–1459, October 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/10/1441.

Chang:2019:FPC

[CC19] Yeim-Kuan Chang and Han-Chen Chen. Fast packet classification using recursive endpoint-cutting and bucket compression on FPGA. The Computer Journal, 62(2):198– 214, February 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/2/198/5026282.

Cheng:2010:EQB

[CCC^{+}10] Long Cheng, Yimin Chen, Canfeng Chen, Jian Ma, Lei Shu, Athanasios V. Vasilakos, and Naixue Xiong. Efficient query-based data collection for mobile wireless monitoring applications. The Computer Journal, 53(10):1643–1657, December 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/10/1643.
                                        full.pdf+html.

Chiu:2011:UCA

[CCCS11] Jih-Ching Chiu, Yu-Liang Chou, Po-Kai Chen, and Ding-Siang Su. A unitable computing architecture for chip multiprocessors. The Computer Journal, 54(12): 2033–2052, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/54/12/2033.
full.pdf+html.

Cardone:2011:CNO

[CCF11] Giuseppe Cardone, Antonio Corradi, and Luca Foschini. Cross-network opportunistic collection of urgent data in wireless sensor networks. The Computer Journal, 54(12):1949–1962, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/12/1949.
full.pdf+html.

Chen:2018:SMA

[CCHL18] Zhi Chen, Xiaolin Chang, Zhen Han, and Lin Li. Survivability modeling and analysis of cloud service in distributed data centers. The Computer Journal, 61 (9):1296–1305, September 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1296/4683984.

Chou:2013:UGS

[CCL^{+}13] Yao-Hsin Chou, Shuo-Mao Chen, Yu-Ting Lin, Chi-Yuan Chen, and Han-Chieh Chao. Using GHZ-state for multiparty quantum secret sharing without code table. The Computer Journal, 56(10):1167–1175, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/10/1167.
full.pdf+html.

Cao:2019:AML

[CCL^{+}19] Nanyuan Cao, Zhenfu Cao, Zhen Liu, Xiaolei Dong, and Xiaopeng Zhao. All-but-many lossy trapdoor functions under decisional RSA subgroup assumption and application. The Computer Journal, 62(8):1148– 1157, August 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1148/5369686.

Chanloha:2014:CTM

[CCUA14] Pitipong Chanloha, Jatuporn Chinrungrueng, Wipawee Usaha, and Chaodit Aswakul. Cell transmission modelbased multiagent Q-learning for network-scale signal control with transit priority. The Computer Journal, 57(3):451–468, March 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/3/451.full.
pdf+html.

Chen:2010:RAA

[CCY10] Chien-Liang Chen, Shao-Chi Chin, and Hsu-Chun Yen. Reachability analysis of augmented marked graphs via integer linear programming. The Computer Journal, 53(6):623– 633, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/623;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/623.

Cui:2016:RDA

[CD16] Hui Cui and Robert H.

Deng. Revocable and decentralized attribute-based encryption. The Computer Journal, 59(8):1220–1235, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1220.

Chen:2011:RTD

[CDYC11] Zeqiang Chen, Liping Di, Genong Yu, and Nengcheng Chen. Real-time on-demand motion video change detection in the sensor Web environment. The Computer Journal, 54(12):2000–2016, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/12/2000.
                                        full.pdf+html.

Cholvi:2010:MCE

[CFJ^{+}10] Vicent Cholvi, Antonio Fernández, Ernesto Jiménez, Pilar Manzano, and Michel Raynal. A methodological construction of an efficient sequentially consistent distributed shared memory. The Computer Journal, 53(9):1523–1534, November 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/9/1523.
                                        full.pdf+html.

Cheng:2013:CAI

[CFJ^{+}13] Baolei Cheng, Jianxi Fan, Xiaohua Jia, Shukui Zhang, and Bangrui Chen. Constructive algorithm of independent spanning trees on Möbius cubes. The Computer Journal, 56(11):1347– 1362, November 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/11/1347.
                                        full.pdf+html.

Cao:2017:VNM

[CFM17] Yang Cao, Wenfei Fan, and Shuai Ma. Virtual network mapping in cloud computing: a graph pattern matching approach. The Computer Journal, 60(3): 60–??, March 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/3/287/2608063/Virtual-
Network-Mapping-in-Cloud-
Computing-A-Graph.

Cerquides:2014:TOM

[CFMR14] Jesus Cerquides, Alessandro Farinelli, Pedro Meseguer, and Sarvapali D. Ramchurn. A tutorial on optimization for multiagent systems. The Computer Journal, 57(6):799– 824, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/6/799.full.
pdf+html.

Calamoneri:2013:AGM

[CFS13] Tiziana Calamoneri, Dario Frascaria, and Blerina Sinaimeri. All graphs with at most seven vertices are pairwise compatibility graphs. The Computer Journal, 56(7):882– 886, July 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/7/882.full.
pdf+html.

Calamoneri:2014:PCG

[CFS14] Tiziana Calamoneri, Antonio Frangioni, and Blerina Sinaimeri. Pairwise compatibility graphs of caterpillars. The Computer Journal, 57(11):1616–1623, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/11/1616.

Caubet:2014:CRL

[CGE^{+}14] Juan Caubet, Carlos Gañán, Oscar Esparza, Jose L. Muñoz, Jorge Mata-Díaz, and Juanjo Alins. Certificate revocation list distribution system for the KAD network. The Computer Journal, 57(2):273– 280, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/2/273.full.
                                        pdf+html.

Corominas:2015:TNR

[CGVP15] Albert Corominas, Alberto García-Villoria, and Rafael Pastor. Technical note: Relating to the parameter values given by Nelder and Mead in their algorithm. The Computer Journal, 58(1):157–159, January 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/157.

Cha:2010:CAS

[Cha10a] Guang-Ho Cha. A contextaware similarity search for a handwritten digit image database. The Computer Journal, 53(8):1291–1301, October 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1291.
full.pdf+html.

Chao:2010:FMM

[Cha10b] Daniel Y. Chao. Fewer monitors and more efficient controllability for deadlock control in S^{3}PGR^{2} (systems of simple sequential processes with general resource requirements). The Computer Journal, 53(10):1783–1798, December 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/10/1783.
full.pdf+html.

Chan:2011:ISR

[Cha11] Chien-Lung Chan. Information security risk modeling using Bayesian index. The Computer Journal, 54(4): 628–638, April 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/4/628.full.
                                        pdf+html.

Chiang:2017:GPG

[CHDP17] Meng-Shu Chiang, Chung-Ming Huang, Duy-Tuan Dao, and Binh-Chau Pham. GB-PMIPv6: a group-based handover control scheme for PMIPv6 using the ‘hitch on’ concept. The Computer Journal, 60(6):822– 834, June 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/6/822/3044151.

Chehreghani:2014:EAA

[Che14] Mostafa Haghir Chehreghani. An efficient algorithm for approximate betweenness centrality computation. The Computer Journal, 57(9): 1371–1382, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/57/9/1371.

Chen:2015:SSS

[Che15a] Yu-Chi Chen. SPEKS: Secure server-designation public key encryption with keyword search against keyword guessing attacks. The Computer Journal, 58(4): 922–933, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/4/922.

Chen:2015:CSA

[Che15b] Zhe Chen. Control systems on automata and grammars. The Computer Journal, 58(1):75–94, January 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/75.

Cui:2019:CPA

[CHH^{+}19] Yuzhao Cui, Qiong Huang, Jianye Huang, Hongbo Li, and Guomin Yang. Ciphertext-policy attributebased encrypted data equality test and classification. The Computer Journal, 62(8):1166–1177, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1166/5480373.

Chien:2012:IAM

[Chi12]  Hung-Yu Chien.  Im-

proved anonymous multireceiver identity-based encryption. The Computer Journal, 55(4):439– 446, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/4/439.full.

pdf+html. See comment on insecurity [Wan14].

Chirkova:2014:CSE

[Chi14] Rada Chirkova. Combinedsemantics equivalence and minimization of conjunctive queries. The Computer Journal, 57(5):775– 795, May 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/5/775.full.
                                        pdf+html.

Chien:2016:GAI

[Chi16] Hung-Yu Chien. A generic approach to improving Diffie–Hellman key agreement efficiency for thin clients. The Computer Journal, 59(4):592–601, April 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/4/592.

Chen:2014:BPA

[CHL14] Kuei-Hao Chen, Guan-Shieng Huang, and Richard Chia-Tung Lee. Bit-parallel algorithms for exact circular string matching. The Computer Journal, 57(5):731– 743, May 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/5/731.full.
pdf+html.

Chen:2017:HAQ

[CJYY17] Long Chen, Joemon M.

Jose, Haitao Yu, and Fajie Yuan. A hybrid approach for question retrieval in community question answerin. The Computer Journal, 60(7):1019–1031, July 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/1019/2608045.

Chon:2010:RSP

[CK10] Haneul Chon and Taewhan Kim. Resource sharing problem of timing variation-aware task scheduling and binding in MPSoC. The Computer Journal, 53(7):883– 894, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/883;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/883.

Ciobanu:2015:PMS

[CK15] Gabriel Ciobanu and Maciej Koutny. PerTiMo: a model of spatial migration with safe access permissions. The Computer Journal, 58(5): 1041–1060, May 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/5/1041.

Chen:2018:APP

[CKH18] Hon-Chan Chen, Tzu-Liang Kung, and Lih-Hsing Hsu. An augmented pancyclicity problem of crossed cubes. The Computer Journal, 61(1):54–62, January 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/1/54/3111549.

Cirstea:2011:MLC

[CKP^{+}11] Corina Cîrstea, Alexander Kurz, Dirk Pattinson, Lutz Schröder, and Yde Venema. Modal logics are coalgebraic. The Computer Journal, 54(1):31–41, January 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/1/31.full.
                                        pdf+html.

Chen:2013:GMN

[CL13] Min Chen and Victor C. M. Leung. Green mobile networking and communications. The Computer Journal, 56(8):923–925, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/8/923.full.
pdf+html.

Cheng:2015:MRP

[CL15] Chien-Fu Cheng and Hsien-Chun Liao. A maliciousresilient protocol for consistent scheduling problem in the cloud computing environment. The Computer Journal, 58(2):315– 330, February 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/2/315.

Chunlin:2016:CAI

[CL16] Li Chunlin and Li LaYuan.

Context-aware integrated scheme for mobile cloud service allocation. The Computer Journal, 59(1):47–63, January 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/1/47.

Chen:2017:NPP

[CL17] Weidong Chen and Shan Ling. Node-pancyclic properties of biswapped networks based on cycles in their factor networks. The Computer Journal, 60(1):1– 12, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Chuang:2018:DMM

[CL18] Yung-Ting Chuang and Yi-Fan Lee. Defense mechanism for malicious and selective forwarding attacks in large and mobile wireless networks. The Computer Journal, 61(12):1862–1879, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/12/1862/5105707.

Cheng:2019:MBO

[CLC^{+}19] Binlin Cheng, Jinjun Liu, Jiejie Chen, Shudong Shi, Xufu Peng, Xingwen Zhang, and Haiqing Hai. MoG: Behavior-obfuscation resistance malware detection. The Computer Journal, 62(12):1734–1747, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/12/1734/5510727.

See erratum [CLJ^{+}22].

Chen:2019:EAB

[CLG^{+}19] Yang Chen, Wenmin Li, Fei Gao, Wei Yin, Kaitai Liang, Hua Zhang, and Qiaoyan Wen. Efficient attributebased data sharing scheme with hidden access structures. The Computer Journal, 62(12):1748–1760, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/12/1748/5518308.

Chang:2014:NAM

[CLH^{+}14] Hong-Yi Chang, Hsin-Che Lu, Yu-Huei Huang, Yuan-Wei Lin, and Yih-Jou Tzang. Novel auction mechanism with factor distribution rule for cloud resource allocation. The Computer Journal, 57(2):255– 262, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/255.full.
pdf+html.

Cheng:2022:EMB

[CLJ^{+}22] Binlin Cheng, Jinjun Liu, Chen Jiejie, Shi Shudong, Peng Xufu, Zhang Xingwen, and Haiqing Hai. Erratum to: MoG: Behaviorobfuscation resistance malware detection. The Computer Journal, 65(10):2846, October 2022. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/

                                        article/65/10/2846/6154265.

See [CLC^{+}19].

Chung:2010:ASE

[CLL10] Sung Woo Chung, Hsien-Hsin S. Lee, and Woo Hyong Lee. Architecture/OS support for embedded multicore systems. The Computer Journal, 53(8):1134– 1135, October 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/53/8/1134.
                                        full.pdf+html.

Chang:2014:FSB

[CLL14] Hong-Yi Chang, Chih-Chun Lai, and Yuan-Wei Lin. A fast SVC-based channelrecommendation system for an IPTV on a cloud and P2P hybrid platform. The Computer Journal, 57(12):1776–1789, December 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/12/1776.

Cho:2013:AAA

[CLLH13] Wei-Ting Cho, Ying-Xun Lai, Chin-Feng Lai, and Yueh-Min Huang. Applianceaware activity recognition mechanism for IoT energy management system. The Computer Journal, 56(8):1020–1033, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/8/1020.
full.pdf+html.

Chen:2017:GBR

[CLLL17] Yen-Hung Chen, Yuan-Cheng Lai, Ching-Neng Lai, and Yang-Chi Li. A group bandwidth reservation scheme to enhance the Driver’s safety in vehicles and transport environment. The Computer Journal, 60(2):60– ??, February 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/2/253/2754559/A-Group-
Bandwidth-Reservation-
Scheme-to-Enhance.

Chopin:2016:IOS

[CLM16] Josh Chopin, Hamid Laga, and Stanley J. Miklavcic. The influence of object shape on the convergence of active contour models for image segmentation. The Computer Journal, 59(5): 603–615, May 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/5/603.

Chen:2019:IBS

[CLND19] Jiahui Chen, Jie Ling, Jianting Ning, and Jintai Ding. Identity-based signature schemes for multivariate public key cryptosystems. The Computer Journal, 62(8):1132–1147, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/8/1132/5369678.

Chapman:2014:LUR

[CLRJ14] Archie C. Chapman, David S. Leslie, Alex Rogers, and Nicholas R. Jennings. Learning in unknown reward games: Application to sensor networks. The Computer Journal, 57(6):875– 892, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/6/875.full.
                                        pdf+html.

Cai:2015:IWE

[CLS15] Shaowei Cai, Chuan Luo, and Kaile Su. Improving WalkSAT by effective tiebreaking and efficient implementation. The Computer Journal, 58(11):2864–2875, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Cappanera:2015:SDC

[CLSV15] P. Cappanera, A. Lori, G. Stea, and G. Vaglini.

On the schedulability of deadline-constrained traffic in TDMA wireless mesh networks. The Computer Journal, 58(2):215–233, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/215.

Chen:2011:ORD

[CLW11] Yen-Ju Chen, Jia-Jie Liu, and Yue-Li Wang. An optimal rotation distance set. The Computer Journal, 54 (5):824–830, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/5/824.full.
pdf+html.

Cui:2014:SSA

[CMA14] Hui Cui, Yi Mu, and Man Ho Au. Signcryption secure against linear related-key attacks. The Computer Journal, 57(10):1472–1483, October 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/10/1472.

Chapman:2010:DDT

[CMKJ10] Archie C. Chapman, Rosa Anna Micillo, Ramachandra Kota, and Nicholas R. Jennings.

Decentralized dynamic task allocation using overlapping potential games. The Computer Journal, 53(9):1462– 1477, November 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/9/1462.
full.pdf+html.

Croft:2010:SEI

[CMS10] W. Bruce Croft, Donald Metzler, and Trevor Strohman. Search engines: information retrieval in practice. Pearson Education, Boston, MA, USA, 2010. ISBN 0-13-136489-8 (paperback). xxv + 524 pp. LCCN TK5105.884 CRO 2010.

Chuang:2016:SED

[CMSML16] Yung-Ting Chuang, P. M. Melliar-Smith, Louise E. Moser, and Isaií Michel Lombera. Statistical estimation and dynamic adaptation algorithms for the iTrust publication, search and retrieval system. The Computer Journal, 59(10): 1492–1510, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/59/10/1492.

Chunlin:2017:ELB

[CMY17] Li Chunlin, Zhou Min, and Luo Youlong. Efficient loadbalancing aware cloud resource scheduling for mobile user. The Computer Journal, 60(6):925–939, June 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/925/3746509.

Correia:2013:BIT

[CNV13] Miguel Correia, Nuno Ferreira Neves, and Paulo Verissimo. BFT-TO: Intrusion tolerance with less replicas. The Computer Journal, 56(6):693–715, June 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/6/693.full.
pdf+html.

Conery:2012:CSM

[Con12] John S. Conery. Computation is symbol manipulation. The Computer Journal, 55 (7):814–816, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/814.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Corkill:2011:DPA

[Cor11] Daniel D. Corkill. Deploying power-aware, wireless sensor agents. The Computer Journal, 54(3):392– 405, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/392.full.
pdf+html.

Choi:2016:DAN

[CP16] Ju-Hee Choi and Gi-Ho Park. Demand-aware NVM capacity management policy for hybrid cache architecture. The Computer Journal, 59(5):685– 700, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/5/685.

Cios:2007:DMK

[CPSK07] Krzysztof J. Cios, Witold Pedrycz, Roman W. Swiniarski, and Lukasz A. Kurgan. Data mining: a knowledge discovery approach. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-33333-9 (hardcover), 0-387-36795-0 (ebook). xv + 606 pp. LCCN QA76.9.D343.D365

2007. URL http://www.
loc.gov/catdir/enhancements/
fy0824/2007921581-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0824/2007921581-
t.html.

Cutting:2010:SIM

[CQL10] Daniel Cutting, Aaron Quigley, and Björn Landfeldt. Special interest messaging: a comparison of IGM approaches. The Computer Journal, 53(1):50–68, January 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/50;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/50.

Cheng:2013:GFA

[CQS13] Eddie Cheng, Ke Qiu, and Zhizhang Shen. A generating function approach to the edge surface area of the arrangement graphs. The Computer Journal, 56(7): 871–881, July 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/7/871.full.
pdf+html.

Cores:2014:FAM

[CRGM14] Iván Cores, Gabriel Rodríguez, Patricia González, and María J. Martín. Failure avoidance in MPI applications using an applicationlevel approach. The Computer Journal, 57(1):100– 114, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/1/100.full.
                                        pdf+html.

Crowcroft:2010:IFE

[Cro10] Jon Crowcroft. Internet failures: an emergent sea of complex systems and critical design errors? The Computer Journal, 53(10):1752– 1757, December 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/53/10/1752.
                                        full.pdf+html.

Cutigi:2016:RFB

[CSS16] Jorge Francisco Cutigi, Adenilso Simao, and Simone R. S. Souza. Reducing FSM-based test suites with guaranteed fault coverage. The Computer Journal, 59(8):1129–1143, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/8/1129.

Chaurasia:2018:REE

[CTD18] Nisha Chaurasia, Shashikala Tapaswi, and Joydip Dhar. A resource efficient expectation maximization clustering approach for cloud. The Computer Journal, 61(1):95–104, January 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/95/3861965.

Castel-Taleb:2012:BAT

[CTIAP12] H. Castel-Taleb, I. Ismael-Aouled, and N. Pekergin. Bounding aggregations for transient and stationary performance analysis of subnetworks. The Computer Journal, 55(5):564– 576, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/5/564.full.
pdf+html.

Caire:2010:CAT

[CvdT10] Patrice Caire and Leendert van der Torre. Convivial ambient technologies: Requirements, ontology and design. The Computer Journal, 53(8):1229–1256, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1229.
full.pdf+html.

Cavagnino:2011:AAD

[CW11] D. Cavagnino and A. E.

Werbrouck. An analysis of associated dividends in the DBM algorithm for division by constants using multiplication. The Computer Journal, 54(1):148– 156, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/1/148.full.
                                        pdf+html.

Chung:2012:CBI

[CW12a] Yu-Fang Chung and Zhen-Yu Wu. Casting ballots over Internet connection against bribery and coercion. The Computer Journal, 55(10):1169–1179, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/10/1169.
                                        full.pdf+html.

Conilione:2012:FAS

[CW12b] Paul Conilione and Dianhui Wang. Fuzzy approach for semantic face image retrieval. The Computer Journal, 55(9):1130–1145, September 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/9/1130.
                                        full.pdf+html.

Cui:2014:MRC

[CWCS14] Yan Cui, Yingxin Wang, Yu Chen, and Yuanchun Shi. Mitigating resource contention on multicore systems via scheduling. The Computer Journal, 57(8):1178–1194, August 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/8/1178.

Chang:2018:MAH

[CWRZ18] Xiaolin Chang, Tianju Wang, Ricardo J. Rodríguez, and Zhenjiang Zhang. Modeling and analysis of high availability techniques in a virtualized system. The Computer Journal, 61(2): 180–198, February 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/61/2/180/
3863077.

Chen:2010:DSE

[CWS^{+}10] Lei Chen, Zijian Wang, Boleslaw Szymanski, Joel W. Branch, Dinesh Verma, Raju Damarla, and John Ibbotson. Dynamic service execution in sensor networks. The Computer Journal, 53(5):513– 527, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/513;   http://comjnl.

                                        oxfordjournals.org/cgi/
                                        reprint/53/5/513.

Chasin:2014:EDT

[CWWK14] Rachel Chasin, Daryl Woodward, Jeremy Witmer, and Jugal Kalita. Extracting and displaying temporal and geospatial entities from articles on historical events. The Computer Journal, 57(3): 403–426, March 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/3/403.full.
                                        pdf+html.

Chen:2019:MBR

[CWZ19] Lele Chen, Gaoli Wang, and GuoYan Zhang. MILPbased related-key rectangle attack and its application to GIFT, Khudra, MIBS. The Computer Journal, 62(12):1805–1821, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/12/1805/5587703.

Che:2015:RPC

[CXF^{+}15] Yonggang Che, Chuanfu Xu, Jianbin Fang, Yongxian Wang, and Zhenghua Wang. Realistic performance characterization of CFD applications on Intel many integrated core architecture. The Computer Journal, 58(12):3279–3294, December 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/12/3279.

Chen:2014:AES

[CXH14] Hongbo Chen, Jungang Xu, and Ben He. Automated essay scoring by capturing relative writing quality. The Computer Journal, 57(9):1318–1330, September 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1318.

Chang:2018:WDR

[CYTP18] Jou-Ming Chang, Jinn-Shyong Yang, Shyue-Ming Tang, and Kung-Jui Pai. The wide diameters of regular hyper-stars and folded hyper-stars. The Computer Journal, 61(1):121– 128, January 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/121/3861969.

Cai:2019:UMD

[CZ19] Hongyun Cai and Fuzhi Zhang. An unsupervised method for detecting shilling attacks in recommender systems by mining item relationship and identifying target items. The Computer Journal, 62(4): 579–597, April 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/579/5255729.

Chen:2010:CFC

[CZC10] Zhongqiang Chen, Yuan Zhang, and Zhongrong Chen. A categorization framework for common computer vulnerabilities and exposures. The Computer Journal, 53(5):551– 580, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/551;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/5/551.

Chen:2018:ERS

[CZCD18] Siyuan Chen, Peng Zeng, Kim-Kwang Raymond Choo, and Xiaolei Dong. Efficient ring signature and group signature schemes based on q-ary identification protocols. The Computer Journal, 61(4):545–560, April 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/545/4656252.

Cheng:2018:DDM

[CZL^{+}18] Jieren Cheng, Jinghe Zhou, Qiang Liu, Xiangyan Tang, and Yanxiang Guo. A DDoS detection method for socially aware networking based on forecasting fusion feature sequence. The Computer Journal, 61(7):959– 970, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/959/4953373.

Chen:2014:CSI

[CZLC14] Yu Chen, Zongyang Zhang, Dongdai Lin, and Zhenfu Cao. CCA-secure IB– KEM from identity-based extractable hash proof system. The Computer Journal, 57(10):1537–1556, October 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/10/1537.

Cai:2019:DAD

[CZLY19] Hui Cai, Yanmin Zhu, Jie Li, and Jiadi Yu. Double auction for a data trading market with preferences and conflicts of interest. The Computer Journal, 62(10):1490–1504, October 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/10/1490/5516469.

Dagdeviren:2014:EED

[DA14] Orhan Dagdeviren and Vahid Khalilpour Akram. An energy-efficient distributed cut vertex detection algorithm for wireless sensor networks. The Computer Journal, 57(12):1852– 1869, December 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/12/1852.

Das:2018:FND

[DA18] Soumya Das and Tamaghna Acharya. Faulty node detection in HMM-based cooperative spectrum sensing for cognitive radio networks. The Computer Journal, 61(10):1468–1478, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/10/1468/4791880.

deAzevedo:2018:OSL

[dAEN^{+}18] Leonildo J. M. de Azevedo, Júlio C. Estrella, Luis H. V. Nakamura, Marcos J. Santana, Regina H. C. Santana, Cláudio F. Motta Toledo, Bruno G. Batista, and Stephan Reiff-Marganiec. Optimized service level agreement establishment in cloud computing. The Computer Journal, 61(10):1429–1442, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1429/4210212.

Dantchev:2011:DNC

[Dan11] Stefan Dantchev. Dynamic neighbourhood cellular automata. The Computer Journal, 54(1):26–30, January 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/26.full.
pdf+html.

Demirbas:2014:ASH

[DAOG14] Dilek Demirbas, Ismail Akturk, Ozcan Ozturk, and Uğur Güdükbay. Applicationspecific heterogeneous networkon-chip design. The Computer Journal, 57(8):1117– 1131, August 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/8/1117.

Daszczuk:2017:CRD

[Das17] Wiktor B. Daszczuk. Communication and resource deadlock analysis using IMDS formalism and model checking. The Computer Journal, 60(5):729– 750, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/5/729/2725498.

Daylight:2011:DRC

[Day11] Edgar G. Daylight. Dijkstra’s rallying cry for generalization: the advent of the recursive procedure, late 1950s–early 1960s. The Computer Journal, 54(11): 1756–1772, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Dobrucali:2013:NCA

[DB13] Oğuzcan Dobrucali and Billur Barshan. Novel compression algorithm based on sparse sampling of 3-D laser range scans. The Computer Journal, 56(7):852– 870, July 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/7/852.full.
                                        pdf+html.

Dastjerdi:2015:ATD

[DB15] Amir Vahid Dastjerdi and Rajkumar Buyya. An autonomous time-dependent SLA negotiation strategy for cloud computing. The Computer Journal, 58(11): 3202–3216, November 2015.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Derguech:2018:UOB

[DBC18] Wassim Derguech, Sami Bhiri, and Edward Curry. Using ontologies for business capability modelling: Describing what services and processes achieve. The Computer Journal, 61(7): 1075–1098, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://academic.oup.com/
comjnl/article/61/7/1075/
4924485.

Derguech:2015:UFC

[DBHC15] Wassim Derguech, Sami Bhiri, Souleiman Hasan, and Edward Curry. Using formal concept analysis for organizing and discovering sensor capabilities. The Computer Journal, 58(3): 356–367, March 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/3/356.

Dou:2018:OHR

[DCA18] Yi Dou, Henry C. B.

Chan, and Man Ho Au. Order-hiding range query over encrypted data without search pattern leakage. The Computer Journal, 61(12):1806–1824, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/12/1806/5065094.

Doh:2011:ESD

[DCLN11] In Hwan Doh, Jongmoo Choi, Donghee Lee, and Sam H. Noh. An empirical study of deploying storage class memory into the I/O path of portable systems. The Computer Journal, 54(8):1267–1281, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/8/1267.
                                        full.pdf+html.

DePrisco:2010:CIT

[DD10a] Roberto De Prisco and Alfredo De Santis. Cheating immune threshold visual secret sharing. The Computer Journal, 53(9):1485–1496, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/9/1485.
                                        full.pdf+html.

Dragovich:2010:AMG

[DD10b] Branko Dragovich and Alexandra Dragovich. padic modelling of the genome and the genetic code. The Computer Journal, 53(4):432–442, May 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/4/432;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/4/432.

Dhivya:2019:DSD

[DD19] M. Dhivya and M. Renuka Devi. Detection of structural defects in fabric parts using a novel edge detection method. The Computer Journal, 62(7):1036– 1043, July 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/7/1036/5224763.

DiGiacomo:2015:HML

[DDG^{+}15] Emilio Di Giacomo, Walter Didimo, Luca Grilli, Giuseppe Liotta, and Salvatore Agostino Romeo. Heuristics for the maximum 2-layer RAC subgraph problem. The Computer Journal, 58(5):1085–1098, May 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1085.

DiGiacomo:2015:PQP

[DDL^{+}15] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, Henk Meijer, and Stephen K. Wismath. Planar and quasi-planar simultaneous geometric embedding. The Computer Journal, 58(11):3126–3140, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Giacomo:2017:ATT

[DDLM17] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Fabrizio Montecchiani. Area-thickness trade-offs for straight-line drawings of planar graphs. The Computer Journal, 60 (1):135–142, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Dagdeviren:2010:GMB

[DE10] Orhan Dagdeviren and Kayhan Erciyes. Graph matching-based distributed clustering and backbone formation algorithms for sensor networks. The Computer Journal, 53(10):1553–1575, December 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/10/1553.
                                        full.pdf+html.

Denning:2012:CSW

[Den12a] Peter J. Denning. Closing statement: What have we said about computation? The Computer Journal, 55 (7):863–865, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/863.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Denning:2012:OSW

[Den12b] Peter J. Denning. Opening statement: What is computation? The Computer Journal, 55(7):805– 810, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/805.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Denning:2012:RSC

[Den12c] Peter J. Denning. Reflections on a symposium on computation. The Computer Journal, 55(7):799– 802, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/799.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Dimakis:2010:DBE

[DFG10] Nikolaos Dimakis, Avgoustinos Filippoupolitis, and Erol Gelenbe. Distributed building evacuation simulator for smart emergency management. The Computer Journal, 53(9):1384–1400, November 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/9/1384.
                                        full.pdf+html.

deFreitas:2011:MAS

[dFHP^{+}11] Edison Pignaton de Freitas, Tales Heimfarth, Carlos Eduardo Pereira, Armando Morado Ferreira, Flávio Rech Wagner, and Tony Larsson. Multi-agent support in a middleware for mission-driven heterogeneous sensor networks. The Computer Journal, 54(3): 406–420, March 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/3/406.full.
                                        pdf+html.

Ding:2012:CLS

[DG12] Lin Ding and Jie Guan. Cryptanalysis of Loiss stream cipher. The Computer Journal, 55(10):1192–1201, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1192.
full.pdf+html.

Desruelle:2015:CDP

[DG15a] Heiko Desruelle and Frank Gielen. Context-driven progressive enhancement of mobile Web applications: a multicriteria decision-making approach. The Computer Journal, 58(8):1732–1746, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/8/1732.

Djuric:2015:FSF

[DG15b] Zoran Djuric and Dragan Gasevic. FEIPS: a secure fair-exchange payment system for Internet transactions. The Computer Journal, 58(10):2537–2556, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2537.

Draper-Gil:2013:AOP

[DGFGHZ13] Gerard Draper-Gil, Josep-Lluís Ferrer-Gomila, M. Francisca Hinarejos, and Jianying Zhou. An asynchronous optimistic protocol for atomic multitwo-party contract signing. The Computer Journal, 56(10):1258–1267, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/10/1258.
                                        full.pdf+html.

Dardha:2017:SSO

[DGV17] Ornela Dardha, Daniele Gorla, and Daniele Varacca. Semantic subtyping for objects and classes. The Computer Journal, 60(5):636– 656, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/5/636/2632620.

Dan:2012:CPM

[DH12a] Haitao Dan and Robert M. Hierons. Controllability problems in MSC-based testing. The Computer Journal, 55(11):1270–1287, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/11/1270.
                                        full.pdf+html.

Ding:2012:NRS

[DH12b] Jie Ding and Jane Hillston. Numerically representing stochastic process algebra models. The Computer Journal, 55(11):1383–1397, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1383.
full.pdf+html.

Dan:2014:OPW

[DH14] Haitao Dan and Robert M.

Hierons. The oracle problem when testing from MSCs. The Computer Journal, 57(7):987–1001, July 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/7/987.full.
pdf+html.

Dang:2019:SBS

[DHT^{+}19] Van Tuyen Dang, Truong Thu Huong, Nguyen Huu Thanh, Pham Ngoc Nam, Nguyen Ngoc Thanh, and Alan Marshall. SDN-based SYN proxy — a solution to enhance performance of attack mitigation under TCP SYN flood. The Computer Journal, 62(4): 518–534, April 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/518/5183521.

DiPierro:2010:PAP

[DHW10] Alessandra Di Pierro, Chris Hankin, and Herbert Wiklicky. Program analysis probably counts. The Computer Journal, 53(6):871– 880, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/871;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/871.

Dimitriou:2013:APR

[Dim13] Ioannis Dimitriou. Analysis of a priority retrial queue with dependent vacation scheme and application to power saving in wireless communication systems. The Computer Journal, 56(11):1363–1380, November 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/11/1363.
full.pdf+html.

Daneshmand:2015:TAR

[DJAJ15] Seyed Mohammadhadi Daneshmand, Amin Javari, Seyed Ebrahim Abtahi, and Mahdi Jalili. A time-aware recommender system based on dependency network of items. The Computer Journal, 58(9):1955–1966, September 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/1955.

Ding:2015:CWF

[DJG^{+}15] Lin Ding, Chenhui Jin, Jie Guan, Shaowu Zhang, Ting Cui, Dong Han, and Wei Zhao. Cryptanalysis of WG family of stream ciphers. The Computer Journal, 58(10):2677–2685, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2677.

Duan:2014:ILT

[DKB^{+}14] Qibin Duan, Dirk P. Kroese, Tim Brereton, Aaron Spettl, and Volker Schmidt. Inverting Laguerre tessellations. The Computer Journal, 57(9):1431–1440, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1431.

deLara:2014:EDM

[dLGCML14] Juan de Lara, Esther Guerra, Ruth Cobos, and Jaime Moreno-Llorena. Extending deep meta-modelling for practical model-driven engineering. The Computer Journal, 57(1):36–58, January 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/1/36.full.
                                        pdf+html.

Duan:2013:MME

[DLL^{+}13] Dongsheng Duan, Yuhua Li, Ruixuan Li, Zhengding Lu, and Aiming Wen. MEI: Mutual enhanced infinite community-topic model for analyzing text-augmented social networks. The Computer Journal, 56(3):336– 354, March 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/3/336.full.
                                        pdf+html.

Danalis:2014:BPH

[DLM^{+}14] Anthony Danalis, Piotr Luszczek, Gabriel Marin, Jeffrey S. Vetter, and Jack Dongarra. BlackjackBench: Portable hardware characterization with automated results’ analysis. The Computer Journal, 57(7):1002– 1016, July 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/7/1002.
                                        full.pdf+html.

Datta:2010:SST

[DLV10] Ajoy K. Datta, Lawrence L. Larmore, and Priyanka Vemula. A self-stabilizing O(k)-time k-clustering algorithm. The Computer Journal, 53(3):342–350, March 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/3/342;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/3/342.

Daneshgar:2018:SSS

[DM18] Amir Daneshgar and Fahimeh Mohebbipoor. A secure self-synchronized stream cipher. The Computer Journal, 61(8):1180–1201, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/8/1180/5005423.

Rossetto:2018:IFU

[dMRGAS18] Anubis Graciela de Moraes Rossetto, Cláudio F. R. Geyer, Luciana Arantes, and Pierre Sens. Impact FD: an unreliable failure detector based on process relevance and confidence in the system. The Computer Journal, 61(10):1557–1576, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1557/4990390.

Dorronsoro:2016:PSC

[DN16] Bernabe Dorronsoro and Sergio Nesmachnow. Parallel soft computing techniques in high-performance computing systems. The Computer Journal, 59(6): 775–776, June 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/6/775.

Do:2011:CAS

[Do11] Tien Van Do. Comparison of allocation schemes for virtual machines in energyaware server farms. The Computer Journal, 54(11): 1790–1797, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Dowty:2015:SED

[Dow15] James G. Dowty. SMML estimators for 1-dimensional continuous data. The Computer Journal, 58(1):126– 133, January 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/1/126.

Dias:2013:QIL

[DP13] Douglas Mota Dias and Marco Aurélio C. Pacheco. Quantum-inspired linear genetic programming as a knowledge management system. The Computer Journal, 56(9):1043–1062, September 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/9/1043.
full.pdf+html.

Dabrowski:2016:CWG

[DP16] Konrad K. Dabrowski and Daniël Paulusma. Cliquewidth of graph classes defined by two forbidden induced subgraphs. The Computer Journal, 59(5):650– 666, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/5/650.

Dong:2011:IIC

[DPZ11] Jing Dong, Tu Peng, and Yajing Zhao. On instantiation and integration commutability of design pattern. The Computer Journal, 54(1):164–184, January 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/164.full.
pdf+html.

Fontes:2017:HFC

[dRFMD^{+}17] Ramon dos Reis Fontes, Mohamed Mahfoudi, Walid Dabbous, Thierry Turletti, and Christian Rothenberg. How far can we go? Towards realistic softwaredefined wireless networking experiments. The Computer Journal, 60(10):1458–1471, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1458/3093135.

Das:2015:DCS

[DSB15] Debasish Das, Utpal Sharma, and D. K. Bhattacharyya. Detection of cross-site scripting attack under multiple scenarios. The Computer Journal, 58(4):808– 822, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/808.

Darwich:2019:CEC

[DSBB19] Mahmoud Darwich, Mohsen Amini Salehi, Ege Beyazit, and Magdy Bayoumi. Costefficient cloud-based video streaming through measuring hotness. The Computer Journal, 62(5):641– 656, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/641/5043541.

De:2012:EEA

[DSTC12] Debraj De, Wen-Zhan Song, Shaojie Tang, and Diane Cook. EAR: an energy and activity-aware routing protocol for wireless sensor networks in smart environments. The Computer Journal, 55(12):1492–1506, December 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/12/1492.
full.pdf+html.

Du:2015:SDE

[DSZZ15] Lei Du, Qinbao Song, Lei Zhu, and Xiaoyan Zhu. A selective detector ensemble for concept drift detection. The Computer Journal, 58(3): 457–471, March 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/3/457.

Dini:2013:HHS

[DT13] Gianluca Dini and Marco Tiloca. HISS: a HIghly Scalable Scheme for group rekeying. The Computer Journal, 56(4):508– 525, April 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/4/508.full.
pdf+html.

Dao-Thi:2011:MCS

[DTFT11] Thu-Ha Dao-Thi, Jean-Michel Fourneau, and Minh-Anh Tran. Multiple class symmetric G-networks with phase type service times. The Computer Journal, 54(2):274–284, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/2/274.full.

pdf+html. See erratum [DTFT12].

Dao-Thi:2012:EMC

[DTFT12] Thu-Ha Dao-Thi, Jean-Michel Fourneau, and Minh-Anh Tran. Erratum: Multiple class symmetric Gnetworks with phase type service times. The Computer Journal, 55(5):577, May 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/5/577.full.

pdf+html. See [DTFT11].

Dunne:2011:CMF

[Dun11]  Paul E. Dunne.  On

constructing minimal formulae. The Computer Journal, 54(7):1067–1075, July 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.

org/content/54/7/1067.
full.pdf+html.

Denning:2012:IWC

[DW12] Peter J. Denning and Peter Wegner. Introduction to what is computation. The Computer Journal, 55(7): 803–804, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/803.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

El-Alfy:2017:DPW

[EA17] El-Sayed M. El-Alfy. Detection of phishing Websites based on probabilistic neural networks and K-medoids clustering. The Computer Journal, 60(12):1745–1759, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1745/3738789.

Erdem:2012:HPI

[EB12] Oguzhan Erdem and Cüneyt F.

Bazlamaçci. High-performance IP lookup engine with compact clustered trie search. The Computer Journal, 55(12):1447–1466, December 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/12/1447.
                                        full.pdf+html.

Ekmekci:2016:EOS

[EÇGK16] Ümit Ekmekçi, Zehra Çataltepe, Güvenç Güngör, and Demir Can Kahya. EigenBots for the okey social board game. The Computer Journal, 59(9):1426–1432, September 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/9/1426.

Erdem:2015:VCT

[ECL15] Oğuzhan Erdem, Aydin Carus, and Hoang Le. Value-coded trie structure for high-performance IPv6 lookup. The Computer Journal, 58(2):204– 214, February 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/2/204.

Elad:2009:EBR

[ED09] Michael Elad and Dmitry Datsenko. Example-based regularization deployed to super-resolution reconstruction of a single image. The Computer Journal, 52(1): 15–30, January 2009. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
52/1/15;   http://comjnl.
oxfordjournals.org/cgi/
content/full/52/1/15;
http://comjnl.oxfordjournals.
org/cgi/reprint/52/1/15.

See corrigendum [ED10].

Elad:2010:CEB

[ED10] Michael Elad and Dmitry Datsenko. Corrigendum: “Example-Based Regularization Deployed to Super-Resolution Reconstruction of a Single Image”. The Computer Journal, 53(7): 1131, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/full/53/
7/1131;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1131. See

[ED09].

Esteban:2018:MCM

[EDH^{+}18] Francisco José Esteban, David Díaz, Pilar Hernández, Juan Antonio Caballero, Gabriel Dorado, and Sergio Gálvez. MC64-Cluster: Many-core CPU cluster architecture and performance analysis in B-tree searches. The Computer Journal, 61(6):912– 925, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/6/912/4769286.

Edelkamp:2017:HCY

[EEK17] Stefan Edelkamp, Amr Elmasry, and Jyrki Katajainen. Heap construction — 50 years later. The Computer Journal, 60(5):657– 674, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/5/657/2632621.

Esteves:2015:CPR

[EFV15] Sérgio Esteves, Paulo Ferreira, and Luís Veiga. C^{3}P: a re-configurable framework to design cycle-sharing computing cloud platforms. The Computer Journal, 58(12):3217–3241, December 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/12/3217.

El-Fakih:2016:TTE

[EFY16] Khaled El-Fakih and Nina Yevtushenko. Test translation for embedded finite state machine components. The Computer Journal, 59(12):1805–1816, December 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/12/1805.

El-Fakih:2019:IHA

[EFYS19] Khaled El-Fakih, Nina Yevtushenko, and Ayat Saleh. Incremental and heuristic approaches for deriving adaptive distinguishing test cases for non-deterministic finite-state machines. The Computer Journal, 62(5): 757–768, May 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/757/5102233.

Emura:2019:PPA

[EKOS19] Keita Emura, Hayato Kimura, Toshihiro Ohigashi, and Tatsuya Suzuki. Privacypreserving aggregation of time-series data with public verifiability from simple assumptions and its implementations. The Computer Journal, 62(4):614– 630, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/614/5263983.

Elgedawy:2015:CRF

[Elg15] Islam Elgedawy. CRES-CENT: a reliable framework for durable composite Web services management. The Computer Journal, 58(2):280–299, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/2/280.

Ergun:2011:IRM

[ELS11] Murat Ergun, Albert Levi, and Erkay Savas. Increasing resiliency in multi-phase wireless sensor networks: Generationwise key predistribution approach. The Computer Journal, 54(4): 602–616, April 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/4/602.full.
                                        pdf+html.

Elleuch:2019:COA

[EMB19] Islam Elleuch, Achraf Makni, and Rafik Bouaziz. Cooperative overtaking assistance system based on V2V communications and RTDB. The Computer Journal, 62(10):1426–1449, October 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/10/1426/5436924.

Entezari-Maleki:2018:PBW

[EMTSM18] Reza Entezari-Maleki, Kishor S. Trivedi, Leonel Sousa, and Ali Movaghar. Performabilitybased workflow scheduling in Grids. The Computer Journal, 61(10):1479–1495, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1479/4796923.

Elgazzar:2015:RPA

[EOIH15] Khalid Elgazzar, Sharief M. A. Oteafy, Walid M. Ibrahim, and Hossam S. Hassanein. A resilient P2P architecture for mobile resource sharing. The Computer Journal, 58(8):1689– 1700, August 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/8/1689.

Enigo:2014:ESN

[ER14] V. S. Felix Enigo and V. Ramachandran. Enhancing the scalability of nonspatial concurrent queries in wireless sensor networks. The Computer Journal, 57(12):1912–1924, December 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/12/1912.

Ergun:2011:TRN

[Erg11]  Salih Ergün.  A truly

random number generator based on a pulse-excited cross-coupled chaotic oscillator. The Computer Journal, 54(10):1592–1602, October 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/10/1592.
                                        full.pdf+html.

Engin:2019:AGA

[ET19] Zeynep Engin and Philip Treleaven. Algorithmic government: Automating public services and supporting civil servants in using data science technologies. The Computer Journal, 62(3):448–460, March 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/448/5070384.

Exposito:2016:PED

[ETR^{+}16] Roberto R. Expósito, Guillermo L. Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Performance evaluation of data-intensive computing applications on a public IaaS cloud. The Computer Journal, 59(3): 287–307, March 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/3/287.

Esteves:2016:WWS

[EV16] Sérgio Esteves and Luís Veiga. WaaS: Workflowas-a-service for the cloud with scheduling of continuous and data-intensive workflows. The Computer Journal, 59(3):371– 383, March 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/3/371.

Fabeiro:2015:AGO

[FAFD15] Jorge F. Fabeiro, Diego Andrade, Basilio B. Fraguela, and Ramón Doallo. Automatic generation of optimized OpenCL codes using OCLoptimizer. The Computer Journal, 58(11): 3057–3073, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Fang:2010:BCP

[Fan10]  Jywe-Fei Fang.  The

bipancycle-connectivity and the m-pancycle-connectivity of the k-ary n-cube. The Computer Journal, 53(6): 667–678, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/667;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/667.

Fang:2011:EFT

[Fan11] Jywe-Fei Fang. The edgefault-tolerant bipancyclicity of the even k-ary ncube. The Computer Journal, 54(2):255–262, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/2/255.full.
pdf+html.

Fadel:2016:LSP

[FEDHL16] Magdy M. Fadel, Ali I. El-Desoky, Amira Y. Haikel, and Labib M. Labib. A low-storage precise IP traceback technique based on packet marking and logging. The Computer Journal, 59(11):1581–1592, November 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1581.

Fekri-Ershad:2017:INR

[FET17] Shervan Fekri-Ershad and Farshad Tajeripour. Impulsenoise resistant color-texture classification approach using hybrid color local binary patterns and Kullback– Leibler divergence. The Computer Journal, 60(11): 1633–1648, November 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/11/1633/3738788.

Fendri:2017:APR

[FFH17] Emna Fendri, Mayssa Frikha, and Mohamed Hammami. Adaptive person re-identification based on visible salient body parts in large camera network. The Computer Journal, 60(11):1590–1608, November 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/11/1590/2970361.

Filippoupolitis:2013:SCE

[FGG13] Avgoustinos Filippoupolitis, Gokce Gorbil, and Erol Gelenbe. Spatial computers for emergency support. The Computer Journal, 56(12):1399–1416, December 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/12/1399.
full.pdf+html.

Farruggia:2018:RST

[FGN^{+}18] Andrea Farruggia, Travis Gagie, Gonzalo Navarro, Simon J. Puglisi, and Jouni Sirén. Relative suffix trees. The Computer Journal, 61(5):773–788, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/5/773/4643569.

Furfaro:2017:MBA

[FGR17] Angelo Furfaro, Maria Carmela Groccia, and Simona E. Rombo. 2D motif basis applied to the classification of digital images. The Computer Journal, 60(7):1096– 1109, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/7/1096/2608076.

Frangioni:2015:OJP

[FGS15] Antonio Frangioni, Laura Galli, and Giovanni Stea. Optimal joint path computation and rate allocation for real-time traffic. The Computer Journal, 58(6): 1416–1430, June 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/6/1416.

Feng:2010:RHI

[FLCT10] Jen-Bang Feng, Iuon-Chang Lin, Yen-Ping Chu, and Shyh-Chang Tsaur. Resampling halftone images using interpolation and error-diffusion. The Computer Journal, 53(6):802– 813, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/802;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/802.

Fan:2019:ILJ

[FLWL19] Jianhua Fan, Tao Liang, Tongxiang Wang, and Jianwei Liu. Identification and localization of the jammer in wireless sensor networks. The Computer Journal, 62(10):1515–1527, October 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/10/1515/5525445.

Feng:2015:SSW

[FLZC15] Zhiyong Feng, Bo Lan, Zhen Zhang, and Shizhan Chen. A study of Semantic Web services network. The Computer Journal, 58(6):1293– 1305, June 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1293.

Fallah:2011:GBS

[FM11] Mehran S. Fallah and Maryam Mouzarani. A game-based Sybil-resistant strategy for reputation systems in self-organizing manets. The Computer Journal, 54(4):537–548, April 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/4/537.full.
                                        pdf+html.

Fotia:2017:ULT

[FMRS17] Lidia Fotia, Fabrizio Messina, Domenico Rosaci, and Giuseppe M. L. Sarné. Using local trust for forming cohesive social structures in virtual communities. The Computer Journal, 60(11):1717–1727, November 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/11/1717/4091311.

Farina:2012:BTC

[FNP12] Antonio Fariña, Gonzalo Navarro, and José R. Paramá. Boosting text compression with word-based statistical encoding. The Computer Journal, 55(1): 111–131, January 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/1/111.full.
                                        pdf+html.

Fortnow:2012:ELT

[For12] Lance Fortnow. The enduring legacy of the Turing machine. The Computer Journal, 55(7):830– 831, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/830.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Faro:2018:ESS

[FP18] Simone Faro and Arianna Pavone. An efficient skipsearch approach to swap matching. The Computer Journal, 61(9):1351–1360, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1351/4779883.

Ferreira:2019:BDP

[FP19] Roger Santos Ferreira and Denilson Alves Pereira. BigFeel — a distributed processing environment for the integration of sentiment analysis methods. The Computer Journal, 62(11):1671–1683, November 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/11/1671/5389528.

Fu:2015:TVG

[FPY15] Dong Lai Fu, Xin Guang Peng, and Yu Li Yang.

Trusted validation for geolocation of cloud data. The Computer Journal, 58(10):2595–2607, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2595.

Francois:2011:AFD

[Fra11] Alexandre R. J. François. An architectural framework for the design, analysis and implementation of interactive systems. The Computer Journal, 54(7):1188– 1204, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/7/1188.
                                        full.pdf+html.

Frailey:2012:CP

[Fra12] Dennis J. Frailey. Computation is process. The Computer Journal, 55(7):817– 819, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/7/817.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Frattolillo:2015:WPP

[Fra15] Franco Frattolillo. Watermarking protocols: Problems, challenges and a possible solution. The Computer Journal, 58(4):944– 960, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/944.

Freeman:2012:CQW

[Fre12] Peter A. Freeman. Consideration of the question “What Is Computation” considered harmful. The Computer Journal, 55(7): 861–862, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/861.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Fu:2018:RIM

[FS18] Xiaogang Fu and Jianyong Sun. Reference-inspired many-objective evolutionary algorithm based on decomposition. The Computer Journal, 61(7):1015– 1037, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/1015/4160532.

Fuentes:2015:MFF

[FSGS15] Joel Fuentes, Pablo Sáez, Gilberto Gutiérrez, and Isaac D. Scherson. A method to find functional dependencies through refutations and duality of hypergraphs. The Computer Journal, 58(5):1186–1198, May 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1186.

Fragal:2019:EHT

[FSMT19] Vanderson Hafemann Fragal, Adenilso Simao, Mohammad Reza Mousavi, and Uraz Cengiz Turker. Extending HSI test generation method for software product lines. The Computer Journal, 62(1):109–129, January 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/109/4993054.

Ferranti:2011:PID

[FT11] Ettore Ferranti and Niki Trigoni. Practical issues in deploying mobile agents to explore a sensor-instrumented environment. The Computer Journal, 54(3):309– 320, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/309.full.
pdf+html.

Fiore:2017:PGP

[FVS17] Dario Fiore, María Isabel González Vasco, and Claudio Soriente. Partitioned group passwordbased authenticated key exchange. The Computer Journal, 60(12):1912–1922, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1912/4210211.

Fan:2013:DAF

[FWC13] Hongbing Fan, Yu-Liang Wu, and Ray C. C. Cheung. Design automation framework for reconfigurable interconnection networks. The Computer Journal, 56(2):258–269, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/2/258.full.
pdf+html.

Ford:2013:RTS

[FXV13] Benjamin J. Ford, Haiping Xu, and Iren Valova. A realtime self-adaptive classifier for identifying suspicious bidders in online auctions. The Computer Journal, 56 (5):646–663, May 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.

                                        org/content/56/5/646.full.
                                        pdf+html.

Feng:2018:SMB

[FYF^{+}18] Dan Feng, Ma Yajie, Zhou Fengxing, Wang Xiaomao, and He Kai. A safety message broadcast strategy in hybrid vehicular network environment. The Computer Journal, 61(6):789– 797, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/6/789/4080206.

Fan:2015:IRD

[FYMY15] Xinyu Fan, Guomin Yang, Yi Mu, and Yong Yu. On indistinguishability in remote data integrity checking. The Computer Journal, 58(4): 823–830, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/4/823.

Feng:2018:ODA

[FZCL18] Zhenni Feng, Yanmin Zhu, Hui Cai, and Pingyi Luo. Optimal distributed auction for mobile crowd sensing. The Computer Journal, 61(10):1443–1459, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1443/4725105.

Duric:2013:SCS

[– DG13] Zoran – Durić and Dragan Gašević. A source code similarity system for plagiarism detection. The Computer Journal, 56(1):70–86, January 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/1/70.full.
pdf+html.

Gobalakrishnan:2018:NMO

[GA18] N. Gobalakrishnan and C. Arun. A new multiobjective optimal programming model for task scheduling using genetic gray wolf optimization in cloud computing. The Computer Journal, 61(10):1523–1536, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1523/4953372.

Gregio:2015:TTM

[GAF^{+}15] André Ricardo Abed Grégio, Vitor Monte Afonso, Dario Simões Fernandes Filho, Paulo Lício de Geus, and Mario Jino. Toward a taxonomy of malware behaviors. The Computer Journal, 58(10):2758–2777, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2758.

Gonzalez-Alberquilla:2014:DRD

[GAFP^{+}14] Rekai González-Alberquilla, Fernando Emmanuel Frati, Luis Piñuel, Karin Strauss, and Luis Ceze. Data race detection with minimal hardware support. The Computer Journal, 57(5):675– 692, May 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/5/675.full.
                                        pdf+html.

Gonzalez-Alvarez:2016:CSD

[GÁVRRL16] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Álvaro Rubio-Largo. A comparative study of different motif occurrence models applied to a hybrid multiobjective shuffle frog leaping algorithm. The Computer Journal, 59(3):384– 402, March 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/3/384.

Gazi:2010:BRK

[Gaz10]  Boran Gazi.  Book re-

view: Krzysztof J. Cios, Witold Pedrycz, Roman W. Swiniarski and Lukasz A.

Kurgan, Data Mining: A Knowledge Discovery Approach. Springer (2007). ISBN-13 978-0-387-33333-5. £55.99. 606 pp. Hardcover. The Computer Journal, 53 (4):489–490, May 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/reprint/53/4/489-
a.

Gunlu:2010:FRD

[GB10] Göksel Günlü and Hasan S.

Bilge. Face recognition with discriminating 3D DCT coefficients. The Computer Journal, 53(8):1324–1337, October 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1324.
full.pdf+html.

Graydon:2014:RSC

[GB14] Patrick Graydon and Iain Bate. Realistic safety cases for the timing of systems. The Computer Journal, 57 (5):759–774, May 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/5/759.full.
pdf+html.

Grozev:2015:PMS

[GB15] Nikolay Grozev and Rajkumar Buyya. Performance modelling and simulation of three-tier applications in cloud and multi-cloud environments. The Computer Journal, 58(1):1–22, January 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/1.

Goyal:2018:EFM

[GBA18] Lalit Mohan Goyal, M. M. Sufyan Beg, and Tanvir Ahmad. An efficient framework for mining association rules in the distributed databases. The Computer Journal, 61(5):645– 657, May 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/5/645/3978075.

Gogoi:2011:SOD

[GBBK11] Prasanta Gogoi, D. K. Bhattacharyya, B. Borah, and Jugal K. Kalita. A survey of outlier detection methods in network anomaly identification. The Computer Journal, 54(4):570– 588, April 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/4/570.full.
pdf+html.

Gogoi:2014:MIM

[GBBK14] Prasanta Gogoi, D. K. Bhattacharyya, B. Borah, and Jugal K. Kalita. MLH-IDS: a multi-level hybrid intrusion detection method. The Computer Journal, 57(4): 602–623, April 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/4/602.full.
pdf+html.

Gong:2016:ATI

[GDCC16] Junqing Gong, Xiaolei Dong, Zhenfu Cao, and Jie Chen. Almost-tight identity based encryption against selective opening attack. The Computer Journal, 59(11):1669–1688, November 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1669.

Gierasimczuk:2013:CCU

[GdJ13] Nina Gierasimczuk and Dick de Jongh. On the complexity of conclusive update. The Computer Journal, 56(3):365–377, March 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/3/365.full.
pdf+html.

Giannetsos:2010:ACI

[GDKP10] Thanassis Giannetsos, Tassos Dimitriou, Ioannis Krontiris, and Neeli R. Prasad. Arbitrary code injection through self-propagating worms in von Neumann architecture devices. The Computer Journal, 53(10): 1576–1593, December 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/53/10/1576.
                                        full.pdf+html.

Guerin:2016:TDU

[GdM16] Nilson Donizete Guerin, Jr., Flavio de Barros Vidal, and Bruno Macchiavello. Text-dependent user verification of handwritten words and signatures on mobile devices. The Computer Journal, 59(9):1415–1425, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/9/1415.

Gelenbe:2010:E

[Gel10] Erol Gelenbe. Editorial. The Computer Journal, 53(1):1– 2, January 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/reprint/53/1/1.

Gelenbe:2012:NC

[Gel12]  Erol Gelenbe.  Natural

computation. The Computer Journal, 55(7):848– 851, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/848.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Gazda:2013:TGR

[GF13] Maciej Gazda and Wan Fokkink. Turning GSOS rules into equations for linear time-branching time semantics. The Computer Journal, 56(1):34–44, January 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/1/34.full.
pdf+html.

Gao:2017:GBU

[GF17] Wei Gao and Mohammad Reza Farahani. Generalization bounds and uniform bounds for multidividing ontology algorithms with convex ontology loss function. The Computer Journal, 60(9):1289–1299, September 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/9/1289/2882686.

Gonzalez-Ferez:2016:IPT

[GFPC16] Pilar González-Férez, Juan Piernas, and Toni Cortes. Improving I/O performance through an in-kernel disk simulator. The Computer Journal, 59(10):1433–1452, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/10/1433.

Gelenbe:2010:DCB

[GG10] Erol Gelenbe and Stephen Gilmore. Discussants’ comments on the Computer Journal Lecture by Peter Harrison presented at the British Computer Society on 24th February 2009. The Computer Journal, 53(6): 869–870, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/reprint/53/6/869.

Gurcan:2011:BEC

[GGZC11] Mustafa K. Gurcan, Hadhrami Ab Ghani, Jihai Zhou, and Anusorn Chungtragarn. Bit energy consumption minimization for multi-path routing in ad hoc networks. The Computer Journal, 54(6):944–959, June 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/944.full.
pdf+html.

Ghahramani:2017:GBP

[GH17] Seyyed Amir Ali Ghafourian Ghahramani and Ali Mohammad Afshin Hemmatyar. A graph-based performance analysis of the 802.11p MAC protocol for safety communications in highway vehicular ad hoc networks. The Computer Journal, 60(2):60– ??, February 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/2/185/2754557/A-Graph-
based-Performance-Analysis-
of-the-802-11p.

Gu:2018:ECF

[GHFY18] Mei-Mei Gu, Rong-Xia Hao, Yan-Quan Feng, and Ai-Mei Yu. The 3-extra connectivity and faulty diagnosability. The Computer Journal, 61(5):672–686, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/5/672/4371609.

Giroire:2018:EAR

[GHMP18] Frédéric Giroire, Nicolas Huin, Joanna Moulierac, and Truong Khoa Phan. Energy-aware routing in software-defined network using compression. The Computer Journal, 61(10): 1537–1556, October 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1537/4953376.

Gong:2016:DFS

[GHXW16] Yili Gong, Chuang Hu, Yanyan Xu, and Wenjie Wang. A distributed file system with variable sized objects for enhanced random writes. The Computer Journal, 59(10):1536–1550, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/10/1536.

Guo:2018:AFH

[GHY18] Qingwen Guo, Qiong Huang, and Guomin Yang. Authorized function homomorphic signature. The Computer Journal, 61(12):1897–1908, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1897/5158246.

Goel:2012:HSM

[GIB12] Sumeer Goel, Yasser Ismail, and Magdy Bayoumi. Highspeed motion estimation architecture for real-time video transmission. The Computer Journal, 55(1): 35–46, January 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/1/35.full.
pdf+html.

Gomes:2012:CEB

[GIP^{+}12a] João V. Gomes, Pedro R. M. Inácio, Manuela Pereira, Mário M. Freire, and Paulo P. Monteiro. Corrigendum: Exploring behavioral patterns through entropy in multimedia peerto-peer traffic. The Computer Journal, 55(10):1265, October 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1265.

full.pdf+html. See [GIP^{+}12b].

Gomes:2012:EBP

[GIP^{+}12b] João V. Gomes, Pedro R. M. Inácio, Manuela Pereira, Mário M. Freire, and Paulo P. Monteiro. Exploring behavioral patterns through entropy in multimedia peer-to-peer traffic. The Computer Journal, 55(6): 740–755, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/6/740.full.

pdf+html. See corrigendum [GIP^{+}12a].

Garhwal:2016:PFR

[GJ16] Sunita Garhwal and Ram Jiwari. Parallel fuzzy regular expression and its conversion to epsilon-free fuzzy automaton. The Computer Journal, 59(9):1383–1391, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/9/1383.

Gu:2015:EIB

[GJJ15] Ke Gu, Weijia Jia, and Chunlin Jiang. Efficient identity-based proxy signature in the standard model. The Computer Journal, 58(4):792–807, April 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/792.

Gao:2014:RTP

[GJQG14] Ming Gao, Cheqing Jin, Weining Qian, and Xueqing Gong. Real-time and personalized search over a microblogging system. The Computer Journal, 57(9):1281–1295, September 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1281.

Gueron:2016:FQI

[GK16] Shay Gueron and Vlad Krasnov. Fast Quicksort implementation using AVX instructions. The Computer Journal, 59(1):83–90, January 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/1/83.

Goncalves:2017:ICM

[GK17] Joshua Goncalves and Aneesh Krishna. Incorporating change management within dynamic requirements-based model-driven agent development. The Computer Journal, 60(7):1044–1077, July 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/1044/2608053.

Gibbon:2013:ACM

[GLBS13] D. C. Gibbon, Z. Liu, A. Basso, and B. Shahraray. Automated content metadata extraction services based on MPEG standards. The Computer Journal, 56 (5):628–645, May 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/5/628.full.
                                        pdf+html.

Gao:2016:PPR

[GLK^{+}16] Jianliang Gao, Ping Liu, Xuedan Kang, Lixia Zhang, and Jianxin Wang. PRS: Parallel relaxation simulation for massive graphs. The Computer Journal, 59(6): 848–860, June 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/6/848.

Gong:2013:QIS

[GLL^{+}13] Wei Gong, Kebin Liu, Xiaoxu Li, Xin Miao, and Tong Zhu. Quality of interaction for sensor network energy-efficient management. The Computer Journal, 56(8):926–937, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/8/926.full.
                                        pdf+html.

Gelenbe:2011:FEA

[GM11] Erol Gelenbe and Christina Morfopoulou. A framework for energy-aware routing in packet networks. The Computer Journal, 54(6):850– 859, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/850.full.
pdf+html.

Guo:2011:SST

[GMS11] Fuchun Guo, Yi Mu, and Willy Susilo. Short signatures with a tighter security reduction without random oracles. The Computer Journal, 54(4):513– 524, April 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/4/513.full.
pdf+html.

Giles:2012:PAO

[GMS^{+}12] M. B. Giles, G. R. Mudalige, Z. Sharif, G. Markall, and P. H. J. Kelly. Performance analysis and optimization of the OP2 framework on many-core architectures. The Computer Journal, 55(2):168–180, February 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/2/168.full.
pdf+html.

Guo:2014:SAS

[GMSV14] Fuchun Guo, Yi Mu, Willy Susilo, and Vijay Varadharajan. Server-aided signature verification for lightweight devices. The Computer Journal, 57(4): 481–493, April 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/4/481.full.
                                        pdf+html.

Gelenbe:2010:ARR

[GN10] Erol Gelenbe and Edith Ngai. Adaptive random re-routing for differentiated QoS in sensor networks. The Computer Journal, 53(7):1052–1061, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/7/1052;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/7/1052.

Ghanbarpour:2019:MBK

[GN19] Asieh Ghanbarpour and Hassan Naderi. A modelbased keyword search approach for detecting top-k effective answers. The Computer Journal, 62(3):377– 393, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/377/5033366.

Gonzalez:2007:HAA

[Gon07] Teofilo F. Gonzalez, editor. Handbook of approximation algorithms and metaheuristics. Chapman and Hall/CRC computer and information science series. Chapman and Hall, Ltd., London, UK, 2007. ISBN 1-58488-550-5. xxi + 1427 pp. LCCN QA76.9.A43 H36

2007. URL http://www.
loc.gov/catdir/enhancements/
fy0707/2007002478-d.html;
 http://www.loc.gov/catdir/
toc/ecip079/2007002478.
html.

Garnett:2010:SBP

[GOR^{+}10] Roman Garnett, Michael A.

Osborne, Steven Reece, Alex Rogers, and Stephen J. Roberts. Sequential Bayesian prediction in the presence of changepoints and faults. The Computer Journal, 53(9):1430–1446, November 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1430.
full.pdf+html.

Gavalas:2011:MWS

[GPK11] Damianos Gavalas, Grammati Pantziou, and Charalampos Konstantopoulos. Mobility in wireless sensor networks. The Computer Journal, 54(12):1928–1930, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/12/1928.
                                        full.pdf+html.

Grassmann:2012:CBR

[Gra12] Winfried K. Grassmann. A computer built with relays and a mechanical memory, and Algol. The Computer Journal, 55(11):1309–1316, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/11/1309.
                                        full.pdf+html.

Girschick:2013:ATD

[GRK13] Tobias Girschick, Ulrich Rückert, and Stefan Kramer. Adapted transfer of distance measures for quantitative structure-activity relationships and data-driven selection of source datasets. The Computer Journal, 56(3): 274–288, March 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/3/274.full.
                                        pdf+html.

Garcia-Reinoso:2015:AEP

[GRVD^{+}15] Jaime Garcia-Reinoso, Ivan Vidal, David Diez, Daniel Corujo, and Rui L. Aguiar. Analysis and enhancements to probabilistic caching in content-centric networking. The Computer Journal, 58(8):1842–1856, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/8/1842.

Greve:2012:ESF

[GSAS12] Fabíola Greve, Pierre Sens, Luciana Arantes, and Véronique Simon. Eventually strong failure detector with unknown membership. The Computer Journal, 55(12): 1507–1524, December 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/55/12/1507.
full.pdf+html.

Grigorious:2017:PDC

[GSRM17] Cyriac Grigorious, Sudeep Stephen, Bharati Rajan, and Mirka Miller. On the partition dimension of circulant graphs. The Computer Journal, 60(2):60– ??, February 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/2/180/2354605/On-the-
Partition-Dimension-of-
Circulant-Graphs.

Gabarro:2014:AWO

[GSS14] Joaquim Gabarro, Maria Serna, and Alan Stewart. Analysing Weborchestrations under stress using uncertainty profiles. The Computer Journal, 57(11):1591–1615, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/11/1591.

Gonen:2019:EDM

[GSS19] Mira Gonen, Dana Shapira, and James A. Storer. Edit distance with multiple block operations. The Computer Journal, 62(5):657– 669, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/5/657/5048917.

Ghahremanlou:2015:GTM

[GST15] Lida Ghahremanlou, Wanita Sherchan, and James A. Thom. Geotagging Twitter messages in crisis management. The Computer Journal, 58(9):1937–1954, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/1937.

Ge:2016:KPA

[GSW^{+}16] Chunpeng Ge, Willy Susilo, Jiandong Wang, Zhiqiu Huang, Liming Fang, and Yongjun Ren. A keypolicy attribute-based proxy re-encryption without random oracles. The Computer Journal, 59(7):970– 982, July 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/7/970.

Gillies:2010:PAE

[GTB10] Duncan Gillies, David Thornley, and Chatschik Bisdikian. Probabilistic approaches to estimating the quality of information in military sensor networks. The Computer Journal, 53(5):493– 502, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/493;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/5/493.

Ghahremanlou:2019:SOD

[GTK^{+}19] Lida Ghahremanlou, Abdel-Rahman H. Tawil, Paul Kearney, Hossein Nevisi, Xia Zhao, and Ali Abdallah. A survey of open data platforms in six UK smart city initiatives. The Computer Journal, 62(7):961– 976, July 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/7/961/5075438.

Grcar:2013:MMD

[GTL13] Miha Grčar, Nejc Trdin, and Nada Lavrač. A methodology for mining documentenriched heterogeneous information networks. The Computer Journal, 56(3): 321–335, March 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/3/321.full.
                                        pdf+html.

Guo:2015:AAM

[GTM15] Kehua Guo, Biao Tian, and Jianhua Ma. AMPS: an adaptive message push strategy for the energy efficiency optimization in mobile terminals. The Computer Journal, 58(6):1243– 1253, June 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1243.

Gelenbe:2010:FDN

[GTN10] Erol Gelenbe, Stelios Timotheou, and David Nicholson. Fast distributed nearoptimum assignment of assets to tasks. The Computer Journal, 53(9):1360–1369, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1360.
full.pdf+html.

Ghica:2011:CMP

[GTS^{+}11] Oliviu Ghica, Goce Trajcevski, Peter Scheuermann, Nikolay Valtchanov, and Zachary Bischof. Controlled multi-path routing in sensor networks using Bézier curves. The Computer Journal, 54(2):230–254, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/2/230.full.
pdf+html.

Gurski:2015:LPF

[Gur15] Frank Gurski. Linear programming formulations for computing graph layout parameters. The Computer Journal, 58(11):2921–2927, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gorman:2016:GGO

[GV16] Chris Gorman and Iren Valova. GORMANN: Gravitationally organized related mapping artificial neural network. The Computer Journal, 59(6):875– 888, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/6/875.

Gonzalez-Valenzuela:2012:LSD

[GVVL12] Sergio González-Valenzuela, Son T. Vuong, and Victor C. M. Leung. Leveraging service discovery in MANETs with mobile directories. The Computer Journal, 55(2):218–231, February 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/2/218.full.
                                        pdf+html.

Gan:2013:ECC

[GWW^{+}13] Yuanju Gan, Lihua Wang, Licheng Wang, Ping Pan, and Yixian Yang. Efficient construction of CCA-secure threshold PKE based on hashed Diffie–Hellman assumption. The Computer Journal, 56(10):1249–1257, October 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/10/1249.
                                        full.pdf+html.

Gao:2015:GCC

[GWWC15] Wei Gao, Guilin Wang, Xueli Wang, and Kefei Chen. Generic construction of certificate-based encryption from certificateless encryption revisited. The Computer Journal, 58(10):2747–2757, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2747.

Gou:2013:LBD

[GY13] Jianping Gou and Zhang Yi. Locality-based discriminant neighborhood embedding. The Computer Journal, 56(9):1063–1082, September 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/9/1063.
full.pdf+html.

Gou:2012:LMB

[GYDX12] Jianping Gou, Zhang Yi, Lan Du, and Taisong Xiong. A local mean-based knearest centroid neighbor classifier. The Computer Journal, 55(9):1058–1071, September 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/9/1058.
full.pdf+html.

Guo:2019:NTP

[GZXA19] Hui Guo, Zhenfeng Zhang, Jing Xu, and Ningyu An. Non-transferable proxy reencryption. The Computer Journal, 62(4):490– 506, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/490/5146175.

Hamlet:2012:SMC

[Ham12]  Dick Hamlet.  Science,

mathematics, computer science, software engineering. The Computer Journal, 55(1):99–110, January 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/1/99.full.
pdf+html.

Harrin:2010:GSP

[Har10a] Elizabeth Harrin. Growing software: Proven strategies for managing software engineers. The Computer Journal, 53(7):1129–1130, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/7/1129-
a.

Harrison:2010:TBT

[Har10b] Peter G. Harrison. Turning back time — what impact on performance? The Computer Journal, 53(6):860– 868, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/860;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/860.

Hartley:2011:ECY

[Har11] David Hartley. EDSAC at 60—a celebration of 60 years since the first program ran on the EDSAC at Cambridge. The Computer Journal, 54(1):136– 138, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/136.full.
pdf+html.

Hernandez:2011:FAS

[HB11] Hugo Hernández and Christian Blum. Foundations of Antcycle: Selfsynchronized duty-cycling in mobile sensor networks. The Computer Journal, 54(9):1427–1448, September 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1427.
full.pdf+html.

Han:2019:ABI

[HBC^{+}19] Jinguang Han, Maoxuan Bei, Liqun Chen, Yang Xiang, Jie Cao, Fuchun Guo, and Weizhi Meng. Attribute-based information flow control. The Computer Journal, 62(8):1214– 1231, August 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1214/5488733.

Hayes:2013:CDN

[HBDJ13] Ian J. Hayes, Alan Burns, Brijesh Dongol, and Cliff B. Jones. Comparing degrees of non-determinism in expression evaluation. The Computer Journal, 56(6):741– 755, June 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/6/741.full.
pdf+html.

Huang:2019:PFC

[HBS^{+}19] Jianbin Huang, Qingquan Bian, Heli Sun, Yaming Yang, and Yu Zhou. Parameter-free community detection through distance dynamic synchronization. The Computer Journal, 62(11):1625–1638, November 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/11/1625/5364033.

Huang:2015:BRR

[HC15] Shih-Chang Huang and Hong-Yi Chang. A bidirectional route repair method for wireless mobile ad hoc networks. The Computer Journal, 58(2):338– 353, February 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/2/338.

Huang:2015:PSC

[HCL15] Chung-Ming Huang, Yih-Chung Chen, and Shih-Yang Lin. Packet scheduling and congestion control schemes for Multipath Datagram Congestion Control Protocol. The Computer Journal, 58(2):188–203, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/188.

Han:2019:DPU

[HCZ^{+}19] Zhen-Hui Han, Xing-Shu Chen, Xue-Mei Zeng, Yi Zhu, and Ming-Yong Yin. Detecting proxy user based on communication behavior portrait. The Computer Journal, 62(12):1777–1792, December 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/12/1777/5543066.

Heys:2017:SCF

[Hey17] Howard M. Heys. Statistical cipher feedback of stream ciphers. The Computer Journal, 60(12):1839–1851, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/12/1839/3959607.

Huynh:2019:SML

[HFP^{+}19] Nghi Huynh, Marc Frappier, Herman Pooda, Amel Mammar, and Régine Laleau. SGAC: a multi-layered access control model with conflict resolution strategy. The Computer Journal, 62(12):1707–1733, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/12/1707/5488734.

Hernandez:2015:SSF

[HGRV15] David Diez Hernandez, Jaime Garcia-Reinoso, and Ivan Vidal. SFP: Statistical filtering policy for caching in content-centric networking. The Computer Journal, 58(8):1763–1775, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1763.

He:2010:TAO

[HGZ10] Xiuqiang He, Zonghua Gu, and Yongxin Zhu. Task allocation and optimization of distributed embedded systems with simulated annealing and geometric programming. The Computer Journal, 53(7):1071–1091, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/1071;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1071.

Hwang:2014:ARP

[HH14] Ren-Junn Hwang and Yu-Kai Hsiao. Anonymous routing protocol based on Weil pairing for wireless mesh networks. The Computer Journal, 57(10):1557– 1569, October 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/10/1557.

Harn:2017:PHG

[HH17] Lein Harn and Ching-Fang Hsu. A practical hybrid group key establishment for secure group communications. The Computer Journal, 60(11):1582–1589, November 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/11/1582/2962044.

Hu:2010:CTS

[HHCL10] Bo Hu, Bin Hu, Victor Callaghan, and Zongkai Lin. Combining theory and systems building in pervasive computing. The Computer Journal, 53(2):129– 130, February 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/reprint/53/2/129.

Hussain:2018:RBF

[HHH^{+}18] Walayat Hussain, Farookh Khadeer Hussain, Omar Hussain, Ravindra Bagia, and Elizabeth Chang. Risk-based framework for SLA violation abatement from the cloud service provider’s perspective. The Computer Journal, 61(9):1306–1322, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/9/1306/4791156.

Hussain:2016:PBO

[HHHC16] Walayat Hussain, Farookh Khadeer Hussain, Omar Khadeer Hussain, and Elizabeth Chang. Provider-based optimized personalized viable SLA (OPV-SLA) framework to prevent SLA violation. The Computer Journal, 59(12):1760–1783, December 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/12/1760.

Harn:2010:ELL

[HHL10] Lein Harn, Wen-Jung Hsin, and Changlu Lin. Efficient on-line/off-line signature schemes based on multiple-collision trapdoor hash families. The Computer Journal, 53(9):1478– 1484, November 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/9/1478.
full.pdf+html.

Hong:2015:RSM

[HHS^{+}15] Wien Hong, Gwoboa Horng, Chih-Wei Shiu, Tung-Shou Chen, and Yu-Chi Chen. Reversible steganographic method using complexity control and human visual system. The Computer Journal, 58(10):2583–2594, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2583.

Huang:2018:LRD

[HHS18] Jianye Huang, Qiong Huang, and Willy Susilo. Leakageresilient dual-form signatures. The Computer Journal, 61(8):1216–1227, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1216/5035762.

Higgins:2017:ADM

[HHV17] Joshua Higgins, Violeta Holmes, and Colin Venters. Autonomous discovery and management in virtual container clusters. The Computer Journal, 60(2): 60–??, February 2017. CO-DEN CMPJA6. ISSN ????

                                        URL https://academic.
                                        oup.com/comjnl/article/
                                        60/2/240/2715225/Autonomous-
                                        Discovery-and-Management-
                                        in-Virtual.

Hinarejos:2015:MES

[HIDFGPC15] M. Francisca Hinarejos, Andreu Pere Isern-Deyà, Josep-Lluís Ferrer-Gomila, and Magdalena Payeras-Capellà. MC-2D: an efficient and scalable multicoupon scheme. The Computer Journal, 58(4):758– 778, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/758.

Hierons:2013:IRT

[Hie13] Robert M. Hierons. Implementation relations for testing through asynchronous channels. The Computer Journal, 56(11):1305–1319, November 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/11/1305.
full.pdf+html.

Hierons:2016:MPI

[Hie16] Robert M. Hierons. A more precise implementation relation for distributed testing. The Computer Journal, 59(1):33–46, January 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/1/33.

Hedetniemi:2013:LTS

[HJK13] Stephen T. Hedetniemi, David P. Jacobs, and K. E. Kennedy. Linear-time selfstabilizing algorithms for disjoint independent sets. The Computer Journal, 56(11):1381–1387, November 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/11/1381.
full.pdf+html.

Huang:2010:CFW

[HJL10] Xiaowei Huang, Li Jiao, and Weiming Lu. Congruence formats for weak readiness equivalence and weak possible future equivalence. The Computer Journal, 53(1): 21–36, January 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/21;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/21.

Hu:2016:MPA

[HJL16] Xiang Hu, Li Jiao, and Zhijia Li. Modelling and performance analysis of IEEE 802.11 DCF using coloured Petri nets. The Computer Journal, 59(10):1563–1580, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/10/1563.

Hyon:2012:SSQ

[HJM12] Emmanuel Hyon and Alain Jean-Marie. Scheduling services in a queuing system with impatience and setup costs. The Computer Journal, 55(5):553– 563, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

org/content/55/5/553.full.
pdf+html.

Huang:2015:NCG

[HJP15] Shenwei Huang, Matthew Johnson, and Daniël Paulusma. Narrowing the complexity gap for colouring (C_{s}, P_{t})- free graphs. The Computer Journal, 58(11):3074–3088, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Han:2013:TSS

[HJS^{+}13] Guangjie Han, Jinfang Jiang, Lei Shu, Mohsen Guizani, and Shojiro Nishio. A two-step secure localization for wireless sensor networks. The Computer Journal, 56(10):1154–1166, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/10/1154.
full.pdf+html.

Hong:2013:DMM

[HK13] Jinpyo Hong and Hwangnam Kim. A dual mobility model with user profiling: Decoupling user mobile patterns from association patterns. The Computer Journal, 56(6):771– 784, June 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/6/771.full.
                                        pdf+html.

Huang:2015:SMP

[HK15] Chun-Chieh Huang and Ren-Song Ko. A study on maximizing the parallelism of macroscopically derived routing algorithms for WSNs. The Computer Journal, 58(12):3306–3324, December 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/12/3306.

Hu:2015:EHS

[HL15] Yanling Hu and Anfeng Liu. An efficient heuristic subtraction deployment strategy to guarantee quality of event detection for WSNs. The Computer Journal, 58(8):1747–1762, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1747.

Hmood:2015:ACA

[HLAZ15] Haider Salim Hmood, Zhitang Li, Hasan Khalaf Abdulwahid, and Yang Zhang. Adaptive caching approach to prevent DNS cache poisoning attack. The Computer Journal, 58(4):973– 985, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/973.

Huang:2010:DHT

[HLC10a] Chung-Ming Huang, Jian-Wei Li, and Chun-Ta Chen. Distributed hash tablebased interrogating-call session control function network in the Internet protocol multimedia subsystem for efficient query services. The Computer Journal, 53(7):918–933, September 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/918;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/918.

Huang:2010:DMC

[HLC10b] Chung-Ming Huang, Ming-Sian Lin, and Lik-Hou Chang. The design of mobile concurrent multipath transfer in multihomed wireless mobile networks. The Computer Journal, 53(10):1704– 1718, December 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/10/1704.
full.pdf+html.

Han:2015:IDA

[HLJ^{+}15] Guangjie Han, Xun Li, Jinfang Jiang, Lei Shu, and Jaime Lloret. Intrusion detection algorithm based on neighbor information against sinkhole attack in wireless sensor networks. The Computer Journal, 58(6):1280–1292, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1280.

Huang:2015:MSE

[HLKL15] Chanying Huang, Hwaseong Lee, Hyoseung Kim, and Dong Hoon Lee. mvSERS: a secure emergency response solution for mobile healthcare in vehicular environments. The Computer Journal, 58(10):2461–2475, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2461.

Hartmann:2011:CFK

[HLL11] Sven Hartmann, Uwe Leck, and Sebastian Link. On Codd families of keys over incomplete relations. The Computer Journal, 54(7): 1166–1180, July 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/7/1166.
full.pdf+html.

Harn:2011:FDM

[HLLC11] Lein Harn, Chia-Yin Lee, Changlu Lin, and Chin-Chen Chang. Fully deniable message authentication protocols preserving confidentiality. The Computer Journal, 54(10):1688–1699, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1688.
full.pdf+html.

Han:2018:TSE

[HLLG18] Shuai Han, Shengli Liu, Lin Lyu, and Dawu Gu. Tightly secure encryption schemes against related-key attacks. The Computer Journal, 61 (12):1825–1844, December 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1825/5067538.

Houidi:2015:EMO

[HLZ15] Ines Houidi, Wajdi Louati, and Djamal Zeghlache. Exact multi-objective virtual network embedding in cloud environments. The Computer Journal, 58(3):403– 415, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/3/403.

Huang:2017:FGT

[HLZ^{+}17] Jianbin Huang, Ze Lv, Yu Zhou, He Li, Heli Sun, et al. Forming grouped teams with efficient collaboration in social networks. The Computer Journal, 60 (11):1545–1560, November 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/11/1545/2527543.

Herranz:2013:CMT

[HM13] Luis Herranz and José María Martínez. Combining MPEG tools to generate video summaries adapted to the terminal and network. The Computer Journal, 56 (5):529–553, May 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/5/529.full.
                                        pdf+html.

Harrison:2014:PFM

[HM14] Peter G. Harrison and Andrea Marin. Productforms in multi-way synchronizations. The Computer Journal, 57(11):1693–1710, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1693.

Hameed:2016:TBT

[HM16] Abdul Hameed and Adnan Noor Mian. Towards better traffic localization of virtual LANs using genetic algorithm. The Computer Journal, 59(2):178– 191, February 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/2/178.

Hernandez:2017:RME

[HM17] Gregorio Hernández and Mafalda Martins. Remote monitoring by edges and faces of maximal outerplanar graphs. The Computer Journal, 60(9):1279–1288, September 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/9/1279/2770525.

Hussain:2018:ESA

[HMH18] Tahani H. Hussain, Paulvanna N. Marimuthu, and Sami J. Habib. Exploration of storage architectures for enterprise network. The Computer Journal, 61(2):233–247, February 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/2/233/3861970.

Hnetynka:2011:CSC

[HMM11] Petr Hnetynka, Liam Murphy, and John Murphy. Comparing the service component architecture and fractal component model. The Computer Journal, 54 (7):1026–1037, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/54/7/1026.
                                        full.pdf+html.

Huang:2012:CSN

[HMS^{+}12] Xinyi Huang, Yi Mu, Willy Susilo, Duncan S. Wong, and Wei Wu. Certificateless signatures: New schemes and security models. The Computer Journal, 55(4): 457–474, April 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/4/457.full.
                                        pdf+html.

Habib:2015:CAE

[HMZ15] Sami Habib, Paulvanna N. Marimuthu, and Naser Zaeri. Carbon-aware enterprise network through redesign. The Computer Journal, 58(2):234–245, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/234.

Hassanpour:2018:IZU

[HNAS18] H. Hassanpour, N. Nowrozian, M. M. AlyanNezhadi, and N. Samadiani. Image zooming using a multi-layer neural network. The Computer Journal, 61(11):1737–1748, November 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/11/1737/5106602.

Hyla:2017:HLS

[HP17] Tomasz Hyla and Jerzy Pejaś. A Hess-like signature scheme based on implicit and explicit certificates. The Computer Journal, 60(4):457–475, March 23, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/457/2608054.

Han:2015:TSB

[HPG^{+}15] Yongkoo Han, Kisung Park, Donghai Guan, Sajal Halder, and Young-Koo Lee. Topological similarity-based feature selection for graph classification. The Computer Journal, 58(9):1884–1893, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/1884.

Han:2017:BFG

[HQL17] XiaoHong Han, Yan Qiang, and Yuan Lan. A bird flock gravitational search algorithm based on the collective response of birds. The Computer Journal, 60 (11):1687–1716, November 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/11/1687/4110360.

Homaei:2011:CAQ

[HS11] Hossein Homaei and Hamid Reza Shahriari. Compositional approach to quantify the vulnerability of computer systems. The Computer Journal, 54(10):1616–1631, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/10/1616.
                                        full.pdf+html.

Hiasat:2019:DRI

[HS19] Ahmad Hiasat and Leonel Sousa. On the design of RNS inter-modulo processing units for the arithmeticfriendly moduli sets 2^{n+k}, 2^{n} − 1, 2^{n+1} − 1. The Computer Journal, 62(2):292– 300, February 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/292/5224762.

Han:2012:ABO

[HSMY12] Jinguang Han, Willy Susilo, Yi Mu, and Jun Yan. Attribute-based oblivious access control. The Computer Journal, 55(10):1202– 1215, October 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/10/1202.
full.pdf+html.

Han:2014:ABD

[HSMY14] Jinguang Han, Willy Susilo, Yi Mu, and Jun Yan. Attribute-based data transfer with filtering scheme in cloud computing. The Computer Journal, 57(4):579– 591, April 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/4/579.full.
pdf+html.

Hsu:2012:STB

[Hsu12]  I-Ching Hsu.  Semantic

tag-based profile framework for social tagging systems. The Computer Journal, 55(9):1118–1129, September 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/9/1118.
                                        full.pdf+html.

Huang:2017:TFM

[HSZS17] Jianbin Huang, Xiaojing Sun, Yu Zhou, and Heli Sun. A team formation model with personnel work hours and project workload quantified. The Computer Journal, 60(9):1382–1394, September 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/9/1382/3051820.

See corrigendum [HSZS18].

Huang:2018:CTF

[HSZS18] Jianbin Huang, Xiaojing Sun, Yu Zhou, and Heli Sun. Corrigendum: A team formation model with personnel work hours and project workload quantified. The Computer Journal, 61(1): 158, January 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/1/158/3852072.

See [HSZS17].

Hierons:2015:IDS

[HT15] Robert M. Hierons and Uraz Cengiz Türker. Incomplete distinguishing sequences for finite state machines. The Computer Journal, 58(11):3089–3113, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Hierons:2016:DSD

[HT16] Robert M. Hierons and Uraz Cengiz Türker. Distinguishing sequences for distributed testing: Adaptive distinguishing sequences. The Computer Journal, 59(8):1186–1206, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1186.

Hierons:2017:DSD

[HT17] Robert M. Hierons and Uraz Cengiz Türker. Distinguishing sequences for distributed testing: Preset distinguishing sequences. The Computer Journal, 60 (1):110–125, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Huang:2015:PAP

[HTC^{+}15] Kaibin Huang, Raylin Tso, Yu-Chi Chen, Sk Md Mizanur Rahman, Ahmad Almogren, and Atif Alamri. PKE-AET: Public key encryption with authorized equality test. The Computer Journal, 58(10):2686–2697, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2686.

Hillston:2012:SPA

[HTG12] Jane Hillston, Mirco Tribastone, and Stephen Gilmore. Stochastic process algebras: from individuals to populations. The Computer Journal, 55(7):866– 881, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/7/866.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Huang:2014:DPD

[Hua14] Shih-Chang Huang. D^{2}PS: Direction and distance positioning system in wireless networks. The Computer Journal, 57(6):939– 951, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/6/939.full.
                                        pdf+html.

Hussain:2015:UBE

[HuRH^{+}15] Omar Khadeer Hussain, Zia ur Rahman, Farookh Khadeer Hussain, Jaipal Singh, Naeem Khalid Janjua, and Elizabeth Chang. A userbased early warning service management framework in cloud computing. The Computer Journal, 58(3):472– 496, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/3/472.

Hou:2016:NEC

[HWCZ16] Rui Hou, Jigang Wu, Yawen Chen, and Haibo Zhang. Note on edge-colored graphs for networks with homogeneous faults. The Computer Journal, 59(10):1470–1478, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/10/1470.

Hong:2019:MCP

[HWS^{+}19] Haibo Hong, Licheng Wang, Jun Shao, Jianhua Yan, Haseeb Ahmad, Guiyi Wei, Mande Xie, and Yixian Yang. A miniature CCA public key encryption scheme based on nonabelian factorization problem in finite groups of Lie type. The Computer Journal, 62(12):1840–1848, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/12/1840/5627776.

Huang:2014:EUS

[HWXD14] Libo Huang, Zhiying Wang, Nong Xiao, and Qiang Dou. Efficient utilization of SIMD engines for general-purpose processors. The Computer Journal, 57(8):1141–1154, August 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/8/1141.

Huang:2011:HSK

[HWY11] Qiong Huang, Duncan S. Wong, and Guomin Yang. Heterogeneous signcryption with key privacy. The Computer Journal, 54(4):525– 536, April 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/4/525.full.
                                        pdf+html.

Hou:2018:UEO

[HXLX18] Yonghong Hou, Lin Xue, Shuo Li, and Jiaming Xing. User-experience-oriented fuzzy logic controller for adaptive streaming. The Computer Journal, 61(7):1064–1074, July 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/1064/4843991.

See corrigendum [HXLX22].

Hou:2022:CUE

[HXLX22] Yonghong Hou, Lin Xue, Shuo Li, and Jiaming Xing. Corrigendum: Userexperience-oriented fuzzy logic controller for adaptive streaming. The Computer Journal, 65(7):1937, July 2022. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/65/7/1937/5498465.

See [HXLX18].

Huang:2019:OSU

[HXQ^{+}19] Jianzhong Huang, Jie Xia, Xiao Qin, Qiang Cao, and Changsheng Xie. Optimization of small updates for erasure-coded inmemory stores. The Computer Journal, 62(6):869– 883, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/869/5310128.

Hu:2012:ANS

[HXZ12] Jiankun Hu, Yang Xiang, and Wanlei Zhou. Advances in network and system security: Editorial statement. The Computer Journal, 55(4):382– 383, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/4/382.full.
                                        pdf+html.

Hao:2016:IRO

[HXZ^{+}16] Zheng Hao, Dong Xiaoshe, Zhu Zhengdong, Chen Baoke, Bai Xiuxiu, Zhang Xingjun, and Wang Endong. Improving the reliability of the operating system inside a VM. The Computer Journal, 59(5):715– 740, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/5/715.

Honda:2011:UTP

[HY11] Kohei Honda and Nobuko Yoshida. A unified theory of program logics: an approach based on the π-calculus. The Computer Journal, 54(1): 88–107, January 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/1/88.full.
                                        pdf+html.

He:2015:MBF

[HY15] Wu He and Gongjun Yan. Mining blogs and forums to understand the use of social media in customer cocreation. The Computer Journal, 58(9):1909–1920, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/1909.

He:2017:IPE

[HYZ17] Hu He, Xu Yang, and Yanjun Zhang. On improving performance and energy efficiency for registerfile connected clustered VLIW architectures for embedded system usage. The Computer Journal, 60 (9):1338–1352, September 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/9/1338/2937754.

Hiary:2018:SIS

[HZAZ18] Hazem Hiary, Rawan Zaghloul, and Moh’d Belal Al-Zoubi. Single-image shadow detection using quaternion cues. The Computer Journal, 61(3):459–468, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/3/459/4827073.

Huang:2011:FTD

[HZHC11] Guangyan Huang, Yanchun Zhang, Jing He, and Jinli Cao. Fault tolerance in data gathering wireless sensor networks. The Computer Journal, 54(6):976– 987, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/6/976.full.
                                        pdf+html.

Huang:2017:NSE

[HZJS17] Jianbin Huang, Yu Zhou, Xiaolin Jia, and Heli Sun. A novel social event organization approach for diverse user choices. The Computer Journal, 60(7):1078– 1095, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/7/1078/2608060.

Huang:2019:OEC

[HZQ^{+}19] Jianzhong Huang, Panping Zhou, Xiao Qin, Yanqun Wang, and Changsheng Xie. Optimizing erasurecoded data archival for replica-based storage clusters. The Computer Journal, 62(2):247–262, February 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/2/247/5065104.

Han:2014:ATS

[HZW^{+}14] Tao Han, Weiming Zhang, Chao Wang, Nenghai Yu, and Yuefei Zhu. Adaptive ±1 steganography in extended noisy region. The Computer Journal, 57(4): 557–566, April 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/4/557.full.
pdf+html.

He:2015:CAC

[HZWT15] Huajun He, Yang Zhao, Jinfu Wu, and Ye Tian. Cost-aware capacity provisioning for Internet video streaming CDNs. The Computer Journal, 58(12):3255– 3270, December 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/12/3255.

He:2015:IEI

[HZX15] Debiao He, Mingwu Zhang, and Baowen Xu. Insecurity of an efficient identity-based proxy signature in the standard model. The Computer Journal, 58(10):2507–2508, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2507.

Islam:2015:MBA

[IA15] Md Saiful Islam and Naif Alajlan. Model-based alignment of heartbeat morphology for enhancing human recognition capability. The Computer Journal, 58(10):2622–2635, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2622.

Iradat:2014:RAB

[IAG^{+}14] Faisal Iradat, Sergey Andreev, Sayeed Ghani, Syed Irfan Nabi, and Waseem Arain. Revisiting assumptions in backoff process modeling and queueing analysis of wireless local area networks (WLANs). The Computer Journal, 57 (6):924–938, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/57/6/924.full.
pdf+html.

Isern-Deya:2013:SAF

[IDVGMP^{+}13] Andreu Pere Isern-Deyà, Arnau Vives-Guasch, Macià Mut-Puigserver, Magdalena Payeras-Capellà, and Jordi Castellà-Roca. A secure automatic fare collection system for time-based or distance-based services with revocable anonymity for users. The Computer Journal, 56(10):1198–1215, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/10/1198.
full.pdf+html.

Ince:2019:LCP

[IEBS19] Ibrahim Furkan Ince, Yusuf Sait Erdem, Faruk Bulut, and Md Haidar Sharif. A lowcost pupil center localization algorithm based on maximized integral voting of circular hollow kernels. The Computer Journal, 62(7): 1001–1015, July 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/7/1001/5116110.

Imanimehr:2016:HPR

[IF16] Fatemeh Imanimehr and Mehran S. Fallah. How powerful are run-time monitors with static information? The Computer Journal, 59(11):1623–1636, November 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1623.

Irtaza:2014:SIR

[IJM14] Aun Irtaza, M. Arfan Jaffar, and Muhammad Tariq Mahmood. Semantic image retrieval in a Grid computing environment using support vector machines. The Computer Journal, 57(2):205– 216, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/205.full.
pdf+html.

Ibrahim:2014:TEY

[IJY^{+}14] Ayad Ibrahim, Hai Jin, Ali A. Yassin, Deqing Zou, and Peng Xu. Towards efficient yet privacy-preserving approximate search in cloud computing. The Computer Journal, 57(2):241– 254, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/241.full.
pdf+html.

Idris:2017:CPS

[IK17] Adnan Idris and Asifullah Khan. Churn prediction system for telecom using filter–wrapper and ensemble classification. The Computer Journal, 60(3): 60–??, March 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/3/410/3063777/Churn-
Prediction-System-for-
Telecom-using-Filter.

Islam:2015:LFP

[IL15] Sk Hafizul Islam and Fagen Li. Leakage-free and provably secure certificateless signcryption scheme using bilinear pairings. The Computer Journal, 58(10):2636– 2648, October 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/10/2636.

Ibarguren:2010:LAR

[IMS10] Aitor Ibarguren, Iñaki Maurtua, and Basilio Sierra. Layered architecture for real-time sign recognition. The Computer Journal, 53(8):1169–1183, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1169.
full.pdf+html.

Ilic:2013:UWP

[IS13] S. S. Ilić and P. Spalević. Using wavelet packets for selectivity estimation. The Computer Journal, 56(7):827– 842, July 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/7/827.full.
pdf+html.

Ipate:2015:MLT

[ISD15] Florentin Ipate, Alin Stefanescu, and Ionut Dinca. Model learning and test generation using cover automata. The Computer Journal, 58(5):1140–1159, May 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1140.

Iqbal:2013:MEM

[ISH13] R. Iqbal, S. Shirmohammadi, and B. Hariri. Modeling and evaluation of a metadata-based adaptive P2P video-streaming system. The Computer Journal, 56(5):554–572, May 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/5/554.full.
pdf+html.

Ibrahim:2019:UCT

[ISST19] Lamiaa F. Ibrahim, Hesham A. Salman, Sara Y. Sery, and Zaki Taha. Using clustering techniques to plan indoor femtocell base stations layout in multi-floors. The Computer Journal, 62(6):919– 930, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/919/5370971.

Jararweh:2017:SDS

[JAAA^{+}17] Yaser Jararweh, Mohammad Alsmirat, Mahmoud Al-Ayyoub, Elhadj Benkhelifa, Ala’ Darabseh, Brij Gupta, and Ahmad Doulat. Software-defined system support for enabling ubiquitous mobile edge computing. The Computer Journal, 60(10):1443–1457, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1443/3051824.

Jarvis:2011:UPE

[Jar11] S. A. Jarvis. UK Performance Engineering Workshop 2010. The Computer Journal, 54(6):960– 961, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/960.full.
pdf+html.

Jarvis:2012:EPM

[Jar12] S. A. Jarvis. Editorial performance modelling, benchmarking and simulation of high-performance computing systems. The Computer Journal, 55(2):136– 137, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/2/136.full.
                                        pdf+html.

Jasinski:2010:BRJ

[Jas10] Ricardo Jasinski. Book review: Jari Nurmi, Processor Design — System-on-Chip Computing for ASICs and FPGAs. Springer (2007). ISBN-13 978-1-4020-5529-4. 525 pp. Hardcover. The Computer Journal, 53(1): 127, January 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/reprint/53/1/127.

Jayaraman:2012:SIS

[Jay12] Bharat Jayaraman. Special issue on security and performance of networks and clouds: Guest Editor’s introduction. The Computer Journal, 55(8):907–908, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/907.full.
                                        pdf+html.

Javed:2019:FRT

[JBM^{+}19] Hafiz Tayyeb Javed, Mirza Omer Beg, Hasan Mujtaba, Hammad Majeed, and Muhammad Asim. Fairness in real-time energy pricing for Smart Grid using unsupervised learning. The Computer Journal, 62(3):414– 429, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/3/414/5053260.

Jun:2010:DCC

[JC10] Minje Jun and Eui-Young Chung. Design of onchip crossbar network topology using chained edge partitioning. The Computer Journal, 53(7):904– 917, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/904;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/904.

Janicke:2013:DAC

[JCSZ13] H. Janicke, A. Cau, F. Siewe, and H. Zedan. Dynamic access control policies: Specification and verification. The Computer Journal, 56(4): 440–463, April 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/4/440.full.
pdf+html.

James:2012:NNC

[JD12] Alex Pappachen James and Sima Dimitrijev. Nearest neighbor classifier based on nearest feature decisions. The Computer Journal, 55(9):1072–1087, September 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/9/1072.
full.pdf+html.

Jafar:2012:RBF

[JDAS12] Iyad F. Jafar, Khalid A. Darabkh, and Ghazi M. Al-Sukkar. A rule-based fuzzy inference system for adaptive image contrast enhancement. The Computer Journal, 55(9):1041–1057, September 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/9/1041.
full.pdf+html.

Jafar:2016:ERD

[JDAZN16] Iyad F. Jafar, Khalid A. Darabkh, Raed T. Al-Zubi, and Rami A. Al Na’mneh. Efficient reversible data hiding using multiple predictors. The Computer Journal, 59(3): 423–438, March 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/3/423.

Javidi:2015:TTA

[JG15] Hamed Javidi and Maziar Goudarzi. TABEMS: Tariffaware building energy management system for sustainability through better use of electricity. The Computer Journal, 58(6):1384– 1398, June 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1384.

Jemaa:2017:FSR

[JHBA17] Salma Ben Jemaa, Mohamed Hammami, and Hanene Ben-Abdallah. Finger surfaces recognition using rank level fusion. The Computer Journal, 60(7): 969–985, July 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/969/2608037.

Janjua:2015:PLB

[JHHC15] Naeem Khalid Janjua, Omar Khadeer Hussain, Farookh Khadeer Hussain, and Elizabeth Chang. Philosophical and logic-based argumentation-driven reasoning approaches and their realization on the WWW: a survey. The Computer Journal, 58(9):1967–1999, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/1967.

Jiang:2014:UIS

[Jia14] Shaoquan Jiang. On unconditional μ-security of private key encryption. The Computer Journal, 57(10):1570– 1579, October 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/10/1570.

Jiang:2017:BMA

[Jia17] Shaoquan Jiang. Bounds for message authentication with distortion. The Computer Journal, 60(4):497– 506, March 23, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/4/497/2608061.

Jin:2018:SJR

[JJ18] Jian Jin and Ping Ji. Scheduling jobs with resourcedependent ready times and processing times depending on their starting times and positions. The Computer Journal, 61(9):1323–1328, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/9/1323/4779881.

Jo:2017:CBV

[JJO^{+}17] Seungbum Jo, Stelios Joannou, Daisuke Okanohara, Rajeev Raman, and Srinivasa Rao Satti. Compressed bit vectors based on variable-to-fixed encodings. The Computer Journal, 60(5):761–775, April 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/5/761/2754558.

Jung:2012:EEK

[JK12] Jason J. Jung and Dariusz Król. Editorial: Engineering knowledge and semantic systems. The Computer Journal, 55(3):256– 257, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/3/256.full.
pdf+html.

Liang:2018:CNN

[jLbLzH18] Zhi jie Liang, Sheng bin Liao, and Bing zhang Hu. 3D convolutional neural networks for dynamic sign language recognition. The Computer Journal, 61 (11):1724–1736, November 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/11/1724/4995616.

Jiang:2019:TCM

[JLDJ19] Ying Jiang, Jian Liu, Gilles Dowek, and Kailiang Ji. Towards combining model checking and proof checking. The Computer Journal, 62(9):1365–1402, September 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/9/1365/5210380.

Jhumka:2011:UFS

[JLS11] Arshad Jhumka, Matthew Leeke, and Sambid Shrestha. On the use of fake sources for source location privacy: Trade-offs between energy and privacy. The Computer Journal, 54(6):860– 874, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/6/860.full.
                                        pdf+html.

Jung:2012:EJA

[JMB12] Dong-Heon Jung, Soo-Mook Moon, and Sung-Hwan Bae. Evaluation of a Java aheadof-time compiler for embedded systems. The Computer Journal, 55(2):232– 252, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/2/232.full.
pdf+html.

Jiang:2016:OOC

[JMG^{+}16] Peng Jiang, Yi Mu, Fuchun Guo, Xiaofen Wang, and Qiaoyan Wen. Online/ offline ciphertext retrieval on resource constrained devices. The Computer Journal, 59(7):955–969, July 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/7/955.

Johnson:2010:BRF

[Joh10] Neil F. Johnson. Book review: Frank Y. Shih, Digital Watermarking and Steganography: Fundamentals and Techniques. CRC/ Taylor & Francis (2008). ISBN-13 978-1-4200-4757-8. £46.99. 180 pp. Hardcover. The Computer Journal, 53(5):616–617, June 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/5/616.

Jennings:2010:DDI

[JRC^{+}10] N. R. Jennings, A. Rogers, S. Case, R. Johnston, and D. Philpot. Decentralized data and information systems: Theory and practice. The Computer Journal, 53(9):1341–1343, November 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1341.
full.pdf+html.

Janicki:2015:CPC

[JS15] Ryszard Janicki and Mohammad Hadi Soudkhah. On classification with pairwise comparisons, support vector machines and feature domain overlapping. The Computer Journal, 58(3): 416–431, March 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/3/416.

Jeong:2013:AST

[JSP13] Young-Sik Jeong, Damien Sauveron, and Jong Hyuk Park. Advanced security technologies and services for future computing environments. The Computer Journal, 56(10):1151–1153, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/10/1151.
full.pdf+html.

Jung:2012:DCL

[Jun12] Jason J. Jung. Discovering community of lingual practice for matching multilingual tags from folksonomies. The Computer Journal, 55(3):337– 346, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/3/337.full.
pdf+html.

Jin:2013:EII

[JWCZ13] Yichao Jin, Yonggang Wen, Qinghua Chen, and Zuqing Zhu. An empirical investigation of the impact of server virtualization on energy efficiency for green data center. The Computer Journal, 56(8):977– 990, August 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/8/977.full.
pdf+html.

Jung:2018:PPL

[JYL18] Bang Chul Jung, Jae Sook Yoo, and Woongsup Lee. A practical physical-layer network coding with spatial modulation in two-way relay networks. The Computer Journal, 61(2):264– 272, February 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/2/264/3978074.

Jang:2015:DCM

[JYP^{+}15] Sung-In Jang, Su-Kyung Yoon, Kihyun Park, Gi-Ho Park, and Shin-Dug Kim. Data classification management with its interfacing structure for hybrid SLC/ MLC PRAM main memory. The Computer Journal, 58(11):2852–2863, November 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

James:2013:CCS

[JZ13] Nick D. James and Jeffery Zucker. A class of contracting stream operators. The Computer Journal, 56(1): 15–33, January 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/1/15.full.
                                        pdf+html.

Zhang:2018:MSS

[jZ18] Xin ju Zhang. Multi-state system reliability analysis based on fuzzy colored Petri nets. The Computer Journal, 61(1):1–13, January 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/1/2981969.

Keskin:2011:LMW

[KAAE11] M. Emre Keskin, I. Kuban Altinel, Necati Aras, and Cem Ersoy. Lifetime maximization in wireless sensor networks using a mobile sink with nonzero traveling time. The Computer Journal, 54(12):1987–1999, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/12/1987.
full.pdf+html.

Kamareddine:2010:CR

[Kam10] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 53(6):619– 622, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/619;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/619.

Kamareddine:2011:CRa

[Kam11a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(2):185– 186, February 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.

                                        org/content/54/2/185.full.
                                        pdf+html.

Kamareddine:2011:CRb

[Kam11b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(4):503– 504, April 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/4/503.full.
                                        pdf+html.

Kamareddine:2011:CRc

[Kam11c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(5):639– 640, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/5/639.full.
                                        pdf+html.

Kamareddine:2011:CRd

[Kam11d] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(6):833– 835, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/6/833.full.
                                        pdf+html.

Kamareddine:2011:CRe

[Kam11e] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(7):1005– 1009, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1005.
full.pdf+html.

Kamareddine:2011:CRf

[Kam11f] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(8):1233– 1234, August 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/8/1233.
full.pdf+html.

Kamareddine:2011:CRg

[Kam11g] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(9):1413–1415, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1413.
full.pdf+html.

Kamareddine:2011:CRh

[Kam11h] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(10):1561–1564, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1561.
full.pdf+html.

Kamareddine:2011:CRi

[Kam11i] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(11):1737–1740, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Kamareddine:2011:CRj

[Kam11j] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 54(12):1925–1927, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/12/1925.
                                        full.pdf+html.

Kamareddine:2012:CRa

[Kam12a] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(1):1–2, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/1/1.full.
                                        pdf+html.

Kamareddine:2012:CRb

[Kam12b] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(2):133– 135, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.

org/content/55/2/133.full.
pdf+html.

Kamareddine:2012:CRc

[Kam12c] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(3):253– 255, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/3/253.full.
pdf+html.

Kamareddine:2012:CRd

[Kam12d] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(4):379– 381, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/4/379.full.
pdf+html.

Kamareddine:2012:CRe

[Kam12e] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(5):509– 511, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/5/509.full.
pdf+html.

Kamareddine:2012:CRf

[Kam12f] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(6):649– 650, June 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/6/649.full.
pdf+html.

Kamareddine:2012:CRg

[Kam12g] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(7):777– 778, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/777.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Kamareddine:2012:CRh

[Kam12h] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(9):1021–1022, September 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/9/1021.
full.pdf+html.

Kamareddine:2012:CRi

[Kam12i] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(10):1147, October 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

org/content/55/10/1147.
full.pdf+html.

Kamareddine:2012:CRj

[Kam12j] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(11):1267–1269, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1267.
full.pdf+html.

Kamareddine:2012:CRk

[Kam12k] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 55(12):1419, December 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/12/1419.
full.pdf+html.

Kamareddine:2013:CRa

[Kam13] Fairouz Kamareddine. Capsule reviews. The Computer Journal, 56(1):1, January 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/1/1.full.
pdf+html.

Kapus:2011:CSD

[Kap11]  Tatjana Kapus.  Clos-

ing a system in the dynamic input/output automata model. The Computer Journal, 54(7):1038– 1048, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1038.
full.pdf+html.

Khan:2013:RPA

[KAS13] Malik Jahan Khan, Mian Muhammad Awais, and Shafay Shamail. A randomized partitioning approach for CBRbased autonomic systems to improve retrieval performance. The Computer Journal, 56(2):175–183, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/2/175.full.
pdf+html.

Kolias:2018:SAC

[KAZ18] Vasilis Kolias, Ioannis Anagnostopoulos, and Sherali Zeadally. Structural analysis and classification of search interfaces for the deep web. The Computer Journal, 61(3):386– 398, March 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/3/386/4539942.

Krishnamoorthy:2012:RMC

[KBMA12] Shivsubramani Krishnamoorthy, Preeti Bhargava, Matthew Mah, and Ashok Agrawala. Representing and managing the context of a situation. The Computer Journal, 55(8):1005–1019, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/8/1005.
full.pdf+html.

Klein:2010:UFC

[KBN10] Shmuel T. Klein and Miri Kopel Ben-Nissan. On the usefulness of Fibonacci compression codes. The Computer Journal, 53(6): 701–716, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/701;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/701.

Kutlu:2010:GTS

[KCC10] Mücahid Kutlu, Celal Cığır, and Ilyas Cicekli. Generic text summarization for Turkish. The Computer Journal, 53(8):1315–1323, October 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

                                        org/content/53/8/1315.
                                        full.pdf+html.

Kim:2015:FFC

[KCC15] Sang-Yeon Kim, Dong-Wan Choi, and Chin-Wan Chung. Finding a friendly community in social networks considering bad relationships. The Computer Journal, 58(6):1469–1481, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1469.

Kumar:2014:AQS

[KCZJ14] Neeraj Kumar, Naveen Chilamkurti, Sherali Zeadally, and Young-Sik Jeong. Achieving quality of service (QoS) using resource allocation and adaptive scheduling in cloud computing with Grid support. The Computer Journal, 57(2):281– 290, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/2/281.full.
                                        pdf+html.

Kim:2010:SCA

[KH10] Jongsung Kim and Seokhie Hong. Side-channel attack using meet-in-the-middle technique. The Computer Journal, 53(7):934– 938, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/934;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/934.

Kan:2018:POR

[KH18] Shuanglong Kan and Zhiqiu Huang. Partial order reduction for the full class of state/event linear temporal logic. The Computer Journal, 61(5):629– 644, May 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/5/629/3920737.

Khan:2011:DCP

[Kha11] Minhaj Ahmad Khan. Data cache prefetching with dynamic adaptation. The Computer Journal, 54(5): 815–823, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/5/815.full.
pdf+html.

Khan:2016:TOS

[Kha16] Minhaj Ahmad Khan. A transformation for optimizing string-matching algorithms for long patterns. The Computer Journal, 59(12):1749–1759, December 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/12/1749.

Kim:2014:ACT

[KHC14] Jongsung Kim, Bo Hong, and Naveen Chilamkurti. Advanced computer technologies and applications in Grid and cloud computing. The Computer Journal, 57(2):181–182, February 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/2/181.full.
                                        pdf+html.

Kayes:2015:OOB

[KHC15] A. S. M. Kayes, Jun Han, and Alan Colman. Ont-CAAC: an ontology-based approach to context-aware access control for software services. The Computer Journal, 58(11):3000–3034, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Kayes:2019:PMF

[KHR^{+}19] A. S. M. Kayes, Jun Han, Wenny Rahayu, Tharam Dillon, Md Saiful Islam, and Alan Colman. A policy model and framework for context-aware access control to information resources. The Computer Journal, 62 (5):670–705, May 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/670/5055357.

Kabir:2015:IUS

[KHYC15] Muhammad Ashad Kabir, Jun Han, Jian Yu, and Alan Colman. Inferring user situations from interaction events in social media. The Computer Journal, 58(9):2026–2043, September 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/2026.

Kaivani:2011:DCR

[KJ11] Amir Kaivani and Ghassem Jaberipur. Decimal CORDIC rotation based on selection by rounding: Algorithm and architecture. The Computer Journal, 54(11): 1798–1809, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Kallel:2015:ETI

[KJR15] Slim Kallel, Mohamed Jmaiel, and Sumitra Reddy. Enabling technologies: Infrastructure for collaborative enterprises. The Computer Journal, 58(3):355, March 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/3/355.

Kalra:2018:DDP

[KK18] Nidhi Kalra and Ajay Kumar. Deterministic deep pushdown transducer and its parallel version. The Computer Journal, 61(1): 63–73, January 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/1/63/3746508.

Kukla:2012:SSL

[KKBF12] Grzegorz Kukla, Przemyslaw Kazienko, Piotr Bródka, and Tomasz Filipowski. Soclake: Social latent knowledge explorator. The Computer Journal, 55(3):258– 276, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/3/258.full.
                                        pdf+html.

Klempa:2015:JFX

[KKM^{+}15] Michal Klempa, Michal Kozak, Mário Mikula, Robert Smetana, Jakub Starka, Michal Švirec, Matej Vitásek, Martin Nečaský, and Irena Holubova (Mlýnková). jInfer: a framework for XML schema inference. The Computer Journal, 58(1):134– 156, January 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/134.

Krivka:2019:JPG

[KKM19] Zbyněk Křivka, Jiří Kučera, and Alexander Meduna. Jumping pure grammars. The Computer Journal, 62(1):30–41, January 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/30/4942085.

Kavakli:2015:PIP

[KKMG15] Evangelia Kavakli, Christos Kalloniatis, Haralambos Mouratidis, and Stefanos Gritzalis. Privacy as an integral part of the implementation of cloud solutions. The Computer Journal, 58(10):2213–2224, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2213.

Kim:2014:DMT

[KKPB14] Hwangnam Kim, Hwantae Kim, Wonkyun Park, and Mungyu Bae. Disabling misbehavior with traffic constraints in WLANs. The Computer Journal, 57(12):1817–1833, December 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/12/1817.

Kim:2010:FGR

[KL10] Dae-Hwan Kim and Hyuk-Jae Lee. Fine-grain register allocation and instruction scheduling in a reference flow. The Computer Journal, 53(6):717– 740, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/717;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/717.

Karaoglan:2014:SDS

[KL14] Duygu Karaoğlan and Albert Levi. A survey on the development of security mechanisms for body area networks. The Computer Journal, 57(10):1484–1512, October 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/10/1484.

Kong:2015:FMB

[KLA^{+}15] Weiqiang Kong, Leyuan Liu, Takahiro Ando, Hirokazu Yatsu, Kenji Hisazumi, and Akira Fukuda. Facilitating multicore bounded model checking with stateless explicit-state exploration. The Computer Journal, 58(11):2824–2840, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Kim:2014:TGP

[KLL14] Hyun-Chul Kim, Jaewook Lee, and Daewon Lee. Transductive Gaussian processes with applications to object pose estimation. The Computer Journal, 57(3): 339–346, March 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/3/339.full.
pdf+html.

Kong:2018:EDP

[KLS18] Shufeng Kong, Sanjiang Li, and Michael Sioutis. Exploring directional pathconsistency for solving constraint networks. The Computer Journal, 61(9): 1338–1350, September 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/9/1338/4779882.

Kong:2015:RED

[KLT^{+}15] Linghe Kong, Xiao-Yang Liu, Meixia Tao, Min-You Wu, Yu Gu, Long Cheng, and Jianwei Niu. Resourceefficient data gathering in sensor networks for environment reconstruction. The Computer Journal, 58(6): 1330–1343, June 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/6/1330.

Khomenko:2014:DCC

[KM14] Victor Khomenko and Andrey Mokhov. Direct construction of complete merged processes. The Computer Journal, 57(5): 693–707, May 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/5/693.full.
                                        pdf+html.

Klavzar:2016:ADI

[KMNA^{+}16] Sandi Klavžar, Paul Manuel, M. J. Nadjafi-Arani, R. Sundara Rajan, Cyriac Grigorious, and Sudeep Stephen. Average distance in interconnection networks via reduction theorems for vertexweighted graphs. The Computer Journal, 59(12):1900– 1910, December 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/12/1900.

Kanwal:2015:TTM

[KMSM15] Ayesha Kanwal, Rahat Masood, Muhammad Awais Shibli, and Rafia Mumtaz. Taxonomy for trust models in cloud computing. The Computer Journal, 58(4): 601–626, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/4/601.

Krivka:2016:PSG

[KMZ16] Zbyněk Křivka, Alexander Meduna, and Petr Zemek. Phrase-structure grammars: Normal forms and reduction. The Computer Journal, 59(8):1180–1185, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1180.

Katagiri:2012:MEO

[KNHK12] Hideki Katagiri, Ichiro Nishizaki, Tomohiro Hayashida, and Takanori Kadoma. Multiobjective evolutionary optimization of training and topology of recurrent neural networks for time-series prediction. The Computer Journal, 55(3):325– 336, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/3/325.full.
                                        pdf+html.

Kheiri:2014:CCV

[KÖ14] Ahmed Kheiri and Ender Özcan. Constructing constrained-version of magic squares using selection hyper-heuristics. The Computer Journal, 57(3): 469–479, March 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/3/469.full.

pdf+html. See correction [KO15].

Kheiri:2015:CCC

[KO15] Ahmed Kheiri and Ender Ozcan. Corrigendum: Constructing constrainedversion of magic squares using selection hyper-heuristics. The Computer Journal, 58 (8):1857, August 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/8/1857.

See [KÖ14].

Kara:2015:MRM

[KOA15] Ahmet Kara, Halit Oğuztüzün, and M. Nedim Alpdemir. A methodology for resolution mapping for crossresolution simulation using Event-B. The Computer Journal, 58(11):2804–2823, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Kocak:2010:TDR

[Koç10]  Taskin Koçak.  Two

decades of random neural networks. The Computer Journal, 53(3):249– 250, March 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/3/249.

Konstantopoulos:2010:RPB

[Kon10] Takis Konstantopoulos. Response to Prof. Baccelli’s lecture on Modelling of Wireless Communication Networks by Stochastic Geometry. The Computer Journal, 53(5):612– 614, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/5/612.

Korsh:2011:FGA

[Kor11] James F. Korsh. Fast generation of t-ary trees. The Computer Journal, 54(5): 776–785, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/5/776.full.
                                        pdf+html.

Kotsiantis:2011:CGR

[Kot11]  S. B. Kotsiantis.  Cas-

cade generalization with reweighting data for handling imbalanced problems. The Computer Journal, 54(9):1547–1559, September 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/9/1547.
                                        full.pdf+html.

Kawachi:2017:GCR

[KOTY17] Akinori Kawachi, Yoshio Okamoto, Keisuke Tanaka, and Kenji Yasunaga. General constructions of rational secret sharing with expected constant-round reconstruction. The Computer Journal, 60(5):711– 728, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/5/711/2715224.

Karmakar:2014:IAD

[KR14] Sushanta Karmakar and A. Chandrakanth Reddy. An improved algorithm for distributed trigger counting in ring. The Computer Journal, 57(7):980– 986, July 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/7/980.full.
pdf+html.

Kannan:2013:NQF

[KRDH13] S. R. Kannan, S. Ramthilagam, R. Devi, and Yueh-Min Huang. Novel quadratic fuzzy c-means algorithms for effective data clustering problems. The Computer Journal, 56(3):393– 406, March 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/3/393.full.
pdf+html.

Klein:2012:SDM

[KS12] Shmuel T. Klein and Dana Shapira. The string-to-dictionary matching problem. The Computer Journal, 55(11):1347–1356, November 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1347.
full.pdf+html.

Kushwaha:2016:MOS

[KS16] Alok Kumar Singh Kushwaha and Rajeev Srivastava. Maritime object segmentation using dynamic background modeling and shadow suppression. The Computer Journal, 59(9):1303–1329, September 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/9/1303.

Karthiga:2018:PSA

[KS18] I Karthiga and Sharmila Sankar. Providing secret authentication in clustered security architecture for cloud-based WBAN. The Computer Journal, 61(2):223–232, February 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/2/223/3861967.

Klein:2019:CSR

[KS19] Shmuel T. Klein and Dana Shapira. Context sensitive rewriting codes for flash memory. The Computer Journal, 62(1):20–29, January 1, 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/20/4922802.

Karaata:2012:OIS

[KSA12] Mehmet Hakan Karaata, Ozgur Sinanoglu, and Bader AlBdaiwi. An optimal inherently stabilizing 2neighborhood crash resilient protocol for secure and reliable routing in hypercube networks. The Computer Journal, 55(5):578– 589, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/5/578.full.
pdf+html.

Kong:2014:NBC

[KSH^{+}14] Hui Kong, Xiaoyu Song, Dong Han, Ming Gu, and Jiaguang Sun. A new barrier certificate for safety verification of hybrid systems. The Computer Journal, 57(7): 1033–1045, July 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/7/1033.
full.pdf+html.

Khan:2015:QQA

[KSPR15] Zahoor A. Khan, Shyamala C. Sivakumar, William J. Phillips, and Bill Robertson. QPRR: QoS-aware peering routing protocol for reliability sensitive data in body area network communication. The Computer Journal, 58(8):1701–1716, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1701.

Kanti:2018:SSF

[KT18] Jyotshana Kanti and Geetam Singh Tomar. Solution of sensing failure problem: an improved twostage detector. The Computer Journal, 61(6):847– 855, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/6/847/4602864.

Kucukyilmaz:2012:PFM

[KTA12] Tayfun Kucukyilmaz, Ata Turk, and Cevdet Aykanat. A parallel framework for in-memory construction of term-partitioned inverted indexes. The Computer Journal, 55(11):1317–1330, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/11/1317.
                                        full.pdf+html.

Kays:2011:TAL

[KTC^{+}11] Roland Kays, Sameer Tilak, Margaret Crofoot, Tony Fountain, Daniel Obando, Alejandro Ortega, Franz Kuemmeth, Jamie Mandel, George Swenson, Thomas Lambert, Ben Hirsch, and Martin Wikelski. Tracking animal location and activity with an automated radio telemetry system in a tropical rainforest. The Computer Journal, 54(12):1931– 1948, November 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/12/1931.
full.pdf+html.

Kalita:2019:NSM

[KTM19] Manashee Kalita, Themrichon Tuithung, and Swanirbhar Majumder. A new steganography method using integer wavelet transform and least significant bit substitution. The Computer Journal, 62(11):1639–1655, November 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/11/1639/5369945.

Kho:2010:ABD

[KTTRJ10] Johnsen Kho, Long Tran-Thanh, Alex Rogers, and Nicholas R. Jennings. An agent-based distributed coordination mechanism for wireless visual sensor nodes using dynamic programming. The Computer Journal, 53(8):1277–1290, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

                                        org/content/53/8/1277.
                                        full.pdf+html.

Khan:2018:SGB

[KTTRJ18] Md Mosaddek Khan, Long Tran-Thanh, Sarvapali D. Ramchurn, and Nicholas R. Jennings. Speeding up GDL-based message passing algorithms for large-scale DCOPs. The Computer Journal, 61(11):1639–1666, November 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/11/1639/4942084.

Kuo:2010:GTC

[Kuo10]  Yu-Chen Kuo.  General

tree k-coteries to reduce the degradation of quorums. The Computer Journal, 53 (6):634–643, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/6/634;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/6/634.

Kupcu:2015:OAS

[Küp15] Alptekin Küpçü. Official arbitration with secure cloud storage application. The Computer Journal, 58(4): 831–852, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/4/831.

Kaushik:2015:GEM

[KV15a] Achal Kaushik and Deo Prakash Vidyarthi. A green energy model for resource allocation in computational grid. The Computer Journal, 58(7):1530–1547, July 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/7/1530.

Kumar:2015:IAM

[KV15b] Ajay Kumar and Anil Kumar Verma. An improved algorithm for the metamorphosis of semiextended regular expressions to deterministic finite automata. The Computer Journal, 58(3):448– 456, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/3/448.

Kang:2016:CPE

[KV16] A. S. Kang and Renu Vig.

Comparative performance evaluation of modified prototype filter bank multicarrier cognitive radio under constraints of L_{p}, K, N and D. The Computer Journal, 59(10):1479–1491, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/10/1479.

Kumar:2019:HHL

[KV19] Neetesh Kumar and Deo Prakash Vidyarthi. A hybrid heuristic for load-balanced scheduling of heterogeneous workload on heterogeneous systems. The Computer Journal, 62(2):276– 291, February 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/2/276/5074549.

Kulekci:2012:FPM

[KVX12] M. Oğuzhan Külekci, Jeffrey Scott Vitter, and Bojian Xu. Fast patternmatching via k-bit filtering based text decomposition. The Computer Journal, 55(1):62–68, January 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/1/62.full.
                                        pdf+html.

Kim:2011:MAE

[KW11] Minseong Kim and Andy Wellings. Multiprocessors and asynchronous event handling in the real-time specification for Java. The Computer Journal, 54(8):1308–1324, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/8/1308.
full.pdf+html.

Kiely:2010:ALF

[KXS^{+}10] Aaron Kiely, Mingsen Xu, Wen-Zhan Song, Renjie Huang, and Behrooz Shirazi. Adaptive linear filtering compression on realtime sensor networks. The Computer Journal, 53(10):1606– 1620, December 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/10/1606.
full.pdf+html.

Kebapci:2011:PIR

[KYU11] Hanife Kebapci, Berrin Yanikoglu, and Gozde Unal. Plant image retrieval using color, shape and texture features. The Computer Journal, 54(9):1475–1490, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1475.
full.pdf+html.

Kong:2016:ABA

[KZY16] Yan Kong, Minjie Zhang, and Dayong Ye. An auctionbased approach for group task allocation in an open network environment. The Computer Journal, 59(3): 403–422, March 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/3/403.

Lin:2012:AAA

[LA12] Jie Lin and Don Adjeroh. All-against-all circular pattern matching. The Computer Journal, 55(7):897– 906, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/897.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Lee:2011:PCA

[LAP11] Jia Lee, Susumu Adachi, and Ferdinand Peper. A partitioned cellular automaton approach for efficient implementation of asynchronous circuits. The Computer Journal, 54(7):1211– 1220, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1211.
full.pdf+html.

Laroussinie:2010:BRC

[Lar10] François Laroussinie. Book review: Christel Baier and Joost-Pieter Katoen, Principles of Model Checking. MIT Press (May 2008). ISBN-13 978-0-262-02649-9. £44.95. 975 pp. Hardcover. The Computer Journal, 53(5):615–616, June 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/5/615-
a.

Lavington:2012:SBB

[Lav12] Simon Lavington. A synopsis of the book Alan Turing and his Contemporaries: Building the World’s First Computers. The Computer Journal, 55(7):779– 787, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/779.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Li:2019:CTC

[LBD^{+}19] Chunlin Li, Jingpan Bai, Shaofeng Du, Chunguang Yang, and Youlong Luo. Combining tag correlation and interactive behaviors for community discovery. The Computer Journal, 62(5):

785–800, May 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/5/785/5110554.

Llopis:2014:SEE

[LBIC14] Pablo Llopis, Javier Garcia Blas, Florin Isaila, and Jesus Carretero. Survey of energy-efficient and power-proportional storage systems. The Computer Journal, 57(7):1017–1032, July 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/7/1017.
                                        full.pdf+html.

Li:2019:MAS

[LBZ19] Weihua Li, Quan Bai, and Minjie Zhang. A multiagent system for modelling preference-based complex influence diffusion in social networks. The Computer Journal, 62(3):430– 447, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/430/5067534.

Ling:2014:NSD

[LC14] Shan Ling and Weidong Chen. Node-to-set disjoint paths in biswapped networks. The Computer Journal, 57(7):953– 967, July 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/7/953.full.
pdf+html.

Layek:2016:ADD

[LCH16] Md. Abu Layek, TaeChoong Chung, and Eui-Nam Huh. Adaptive desktop delivery scheme for provisioning quality of experience in cloud desktop as a service. The Computer Journal, 59(2):260–274, February 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/2/260.

Liu:2012:IKD

[LCLL12] Chuan-Gang Liu, Chih-Hung Chao, Chih-Wen Leou, and Jung-Shian Li. Iterative key distribution based on mad neighborhood in underwater mobile sensor networks. The Computer Journal, 55(12):1467–1485, December 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/12/1467.
full.pdf+html.

Liu:2011:CCL

[LCMC11] Qiang Liu, George A. Constantinides, Konstantinos Masselos, and Peter Y. K. Cheung. Compiling Clike languages to FPGA hardware: Some novel approaches targeting data memory organization. The Computer Journal, 54(1):1– 10, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/1/1.full.
                                        pdf+html.

Lun:2014:RBF

[LCX14] Lijun Lun, Xin Chi, and Hui Xu. The relationship between forward slicing and backward slicing for software architecture. The Computer Journal, 57(5): 744–758, May 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/5/744.full.
                                        pdf+html.

Liu:2016:SIS

[LCXZ16] Huilin Liu, Chen Chen, Jun-Chang Xin, and Liyuan Zhang. Searching the informative subgraph based on the PeakGraph model. The Computer Journal, 59(8):1207–1219, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1207.

Lotz:2015:SCS

[LDB^{+}15] Volkmar Lotz, Francesco Di Cerbo, Michele Bezzi, Samuel Paul Kaluvuri, Antonino Sabetta, and Slim Trabelsi. Security certification for service-based business ecosystems. The Computer Journal, 58(4):709– 723, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/709.

Lebrecht:2011:ASM

[LDK11] Abigail S. Lebrecht, Nicholas J.

Dingle, and William J. Knottenbelt. Analytical and simulation modelling of zoned RAID systems. The Computer Journal, 54(5): 691–707, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/5/691.full.
pdf+html.

Li:2015:DBE

[LDLJ15] He Li, Mianxiong Dong, Xiaofei Liao, and Hai Jin. Deduplication-based energy efficient storage system in cloud environment. The Computer Journal, 58(6):

1373–1383, June 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/6/1373.

Li:2016:CBK

[LDZ16] Jiguo Li, Haiting Du, and Yichen Zhang. Certificatebased key-insulated signature in the standard model. The Computer Journal, 59(7):1028–1039, July 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/7/1028.

Lee:2013:ISI

[LE13] Jong-Hyouk Lee and Thierry Ernst. IPv6 security issues in cooperative intelligent transportation systems. The Computer Journal, 56(10):1189–1197, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/10/1189.
                                        full.pdf+html.

Levene:2010:BRR

[Lev10a] Mark Levene. Book review: Rokach Lior and Maimon Oded, Data Mining with Decision Trees: Theory and Applications. World Scientific (2008). ISBN-13 978-981-277-171-1. 244 pp. Hardcover. The Computer Journal, 53(4):489, May 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/4/489.

Levene:2010:SNI

[Lev10b]  Mark Levene.  Social

networks: An introduction. The Computer Journal, 53(7):1129, September 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/7/1129.

Levene:2011:BRS

[Lev11a] Mark Levene. Book review:

Search Engines: Information Retrieval in Practice. The Computer Journal, 54 (5):831–832, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/5/831.full.

pdf+html. See [CMS10].

Levene:2011:CMA

[Lev11b]  Mark Levene.  Chess

metaphors, artificial intelligence and the human mind. The Computer Journal, 54(9):1560, September 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/9/1560.
                                        full.pdf+html.

Liu:2014:BAB

[LFHF14] Pingshan Liu, Shengzhong Feng, Guimin Huang, and Jianping Fan. Bandwidthavailability-based replication strategy for P2P VoD systems. The Computer Journal, 57(8):1211–1229, August 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/8/1211.

Li:2011:IAA

[LfHmXjL11] Qiang Li, Qin fen Hao, Li min Xiao, and Zhou jun Li. An integrated approach to automatic management of virtualized resources in cloud environments. The Computer Journal, 54(6): 905–919, June 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/6/905.full.
                                        pdf+html.

Li:2018:DED

[LFLJ18] Xiaoyan Li, Jianxi Fan, Cheng-Kuan Lin, and Xiaohua Jia. Diagnosability evaluation of the data center network DCell. The Computer Journal, 61(1):129– 143, January 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/129/3867580.

Luo:2019:ADR

[LGC19] Xinjian Luo, Xiaofeng Gao, and Guihai Chen. Accelerate data retrieval by multi-dimensional indexing in switch-centric data centers. The Computer Journal, 62(2):301–320, February 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/301/5253758.

Liu:2015:MCE

[LGHD15] Yi Liu, Xiongzi Ge, Xiaoxia Huang, and David H. C. Du. MOLAR: a costefficient, high-performance SSD-based hybrid storage cache. The Computer Journal, 58(9):2061–2078, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/2061.

Lopez-Garcia:2014:PBB

[LGPRH14] Lourdes López-García, Luis J. Dominguez Perez, and Francisco Rodríguez-Henríquez. A pairing-based blind signature e-voting scheme.

The Computer Journal, 57(10):1460–1471, October 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/10/1460.

Lee:2011:AGP

[LH11] Wei-Po Lee and Yu-Ting Hsiao. An adaptive GA– PSO approach with gene clustering to infer S-system models of gene regulatory networks. The Computer Journal, 54(9):1449–1464, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/9/1449.
                                        full.pdf+html.

LeMartelot:2013:FMS

[LH13] Erwan Le Martelot and Chris Hankin. Fast multiscale detection of relevant communities in large-scale networks. The Computer Journal, 56(9):1136–1150, September 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/9/1136.
                                        full.pdf+html.

Liu:2011:ESF

[LHCN11] Yang Liu, Zhen He, Yi-Ping Phoebe Chen, and Thi Nguyen. External sorting on flash memory via natural page run generation. The Computer Journal, 54(11): 1882–1990, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Liu:2013:EDH

[LHFF13] Pingshan Liu, Guimin Huang, Shengzhong Feng, and Jianping Fan. Eventdriven high-priority first data scheduling scheme for P2P VoD streaming. The Computer Journal, 56(2):239–257, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/2/239.full.
pdf+html.

Liu:2016:CGS

[LHL16] Jia-Jie Liu, Cheng-Ju Hsu, and Chien-Hung Lin. Computing global secure set on trees. The Computer Journal, 59(5):616– 629, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/5/616.

Lin:2015:CND

[LHM^{+}15] Hui Lin, Jia Hu, Jianfeng Ma, Li Xu, and Li Yang. CRM: a new dynamic cross-layer reputation computation model in wireless networks. The Computer Journal, 58(4):656– 667, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/656.

Lee:2012:CSM

[LHYW12] Chao-Hsien Lee, Chung-Ming Huang, Chia-Ching Yang, and Tai-Hsiang Wang. CO-SVC-MDC-based cooperative video streaming over vehicular networks. The Computer Journal, 55(6): 756–768, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/6/756.full.
                                        pdf+html.

Liu:2015:LBD

[LJ15] Huacui Liu and Chenhui Jin. Lower bounds of differential and linear active S-boxes for 3D-like structure. The Computer Journal, 58(4):904– 921, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/904.

Liu:2016:LCP

[LJ16] Guo-Qiang Liu and Chen-Hui Jin. Linear cryptanalysis of PRESENT-like ciphers with secret permutation. The Computer Journal, 59(4):549– 558, April 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/549.

Li:2018:MMA

[LJ18a] Rongjia Li and Chenhui Jin.

Meet-in-the-middle attacks on reduced-round QARMA-64/128. The Computer Journal, 61(8):1158–1165, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/8/1158/4993053.

Linda:2018:HAD

[LJ18b] C. Harriet Linda and G. Wiselin Jiji. Hierarchical approach to detect fractures in CT DICOM images. The Computer Journal, 61(7):1099–1108, July 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/1099/4925399.

Liu:2019:ICA

[LJ19] Hanqiu Liu and Chenhui Jin. An improvement of the CS attack to DSC cipher. The Computer Journal, 62(8):1158–1165, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/8/1158/5476715.

Leeke:2013:TDE

[LJA13] Matthew Leeke, Arshad Jhumka, and Sarabjot Singh Anand. Towards the design of efficient error detection mechanisms for transient data errors. The Computer Journal, 56(6):674– 692, June 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/6/674.full.
                                        pdf+html.

Lin:2015:CPD

[LJA15] Jie Lin, Yue Jiang, and Don Adjeroh. Circular pattern discovery. The Computer Journal, 58(5):1061– 1073, May 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/5/1061.

Liu:2011:NSC

[LJC11] GuanJun Liu, ChangJun Jiang, and Daniel Y. Chao. A necessary and sufficient condition for the liveness of normal nets. The Computer Journal, 54(1):157– 163, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/157.full.
pdf+html.

Li:2016:IRI

[LJF16] Xinran Li, Chen-Hui Jin, and Fang-Wei Fu. Improved results of impossible differential cryptanalysis on reduced FOX. The Computer Journal, 59(4):541– 548, April 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/541.

Li:2019:IID

[LJF19] Rongjia Li, Chenhui Jin, and Ruya Fan. Improved integral distinguishers on compression function of GOST R hash function. The Computer Journal, 62(4): 535–544, April 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/535/5224765.

Liu:2019:SMS

[LJWL19] Weiwei Liu, Chong Jiang, Ji-Bo Wang, and Yuan-Yuan Lu. Single-machine scheduling with simultaneous considerations of resource allocation and deteriorating jobs. The Computer Journal, 62(1):81–89, January 1, 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/1/81/4953377.

Liao:2013:POM

[LJYL13] Xiaofei Liao, Hai Jin, Jia Yu, and Dingding Li. A performance optimization mechanism for SSD in virtualized environment. The Computer Journal, 56(8):991–1000, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/8/991.full.
                                        pdf+html.

Lemire:2014:SUS

[LK14] Daniel Lemire and Owen Kaser. Strongly universal string hashing is fast. The Computer Journal, 57(11):1624–1638, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/11/1624.

Lee:2018:NIC

[LK18] Eunsung Lee and Sang Woo Kim. Non-interactive conditional proxy re-signature in the standard model. The Computer Journal, 61 (12):1772–1782, December 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1772/4965847.

Luo:2010:ILP

[LKG10] Junzhou Luo, Weining Kong, and Liang Ge. Implementation of learning path in process control model. The Computer Journal, 53(2):131–141, February 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/2/131;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/2/131.

Le:2011:RMA

[LL11a] Duc-Phong Le and Chao-Liang Liu. Refinements of Miller’s algorithm over Weierstrass curves revisited. The Computer Journal, 54(10):1582–1591, October 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1582.
full.pdf+html.

Lin:2011:IRS

[LL11b] Chuen-Horng Lin and Wei-Chih Lin. Image retrieval system based on adaptive color histogram and texture features. The Computer Journal, 54(7):1136– 1147, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/7/1136.
                                        full.pdf+html.

Lee:2014:SSD

[LL14] Chia-Lin Lee and Tzong-Jye Liu. A self-stabilizing distance-2 edge coloring algorithm. The Computer Journal, 57(11):1639–1648, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/11/1639.

Lee:2015:TSS

[LL15] Cheng-Chi Lee and Yan-Ming Lai. Toward a secure single sign-on mechanism for distributed computer networks. The Computer Journal, 58(4):934– 943, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/934.

Liu:2017:LSP

[LLDL17] Xiao Liu, Anfeng Liu, Qingyong Deng, and Haolin Liu. Large-scale programming code dissemination for software-defined wireless networks. The Computer Journal, 60(10):1417–1442, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1417/3061103.

Lv:2017:HCP

[LLF17] Yali Lv, Cheng-Kuan Lin, and Jianxi Fan. Hamiltonian cycle and path embeddings in k-ary n-cubes based on structure faults. The Computer Journal, 60(2): 60–??, February 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/2/159/2608070/Hamiltonian-
Cycle-and-Path-Embeddings-
in-k-Ary-n.

Lyu:2018:PKE

[LLH18] Lin Lyu, Shengli Liu, and Shuai Han. Publickey encryption with tight simulation-based selectiveopening security. The Computer Journal, 61(2):288– 318, February 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/2/288/4259796.

Liu:2017:ARP

[LLLW17] Weiran Liu, Xiao Liu, Jianwei Liu, and Qianhong Wu. Auditing revocable privacy-preserving access control for EHRs in clouds. The Computer Journal, 60(12):1871–1888, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/12/1871/4080205.

Liu:2015:GEE

[LLN^{+}15] Jin Liu, Juan Li, Xiaoguang Niu, Xiaohui Cui, and Yunchuan Sun. GreenOCR: an energy-efficient optimal clustering routing protocol. The Computer Journal, 58(6):1344–1359, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1344.

Lloyd:2013:SID

[Llo13] John Lloyd. Special issue on dependable software systems: Associate Editor’s introduction. The Computer Journal, 56(6):671– 673, June 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/6/671.full.
                                        pdf+html.

Lin:2016:ITR

[LLpC16] Jia-Chun Lin, Fang-Yie Leu, and Ying ping Chen. Impacts of task re-execution policy on MapReduce jobs. The Computer Journal, 59 (5):701–714, May 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/5/701.

Lee:2019:CSS

[LLPY19] Kwangsu Lee, Dong Hoon Lee, Jong Hwan Park, and Moti Yung. CCA security for self-updatable encryption: Protecting cloud data when clients read/ write ciphertexts. The Computer Journal, 62(4):545– 562, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/545/5239643.

Lu:2017:WIC

[LLS17] Yang Lu, Jiguo Li, and Jian Shen. Weakness and improvement of a certificatebased key-insulated signature in the standard model. The Computer Journal, 60 (12):1729–1744, December 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1729/3091766.

Lee:2016:CAM

[LLSW16] Hyung Tae Lee, San Ling, Jae Hong Seo, and Huaxiong Wang. CCA2 attack and modification of Huang et al.’s public key encryption with authorized equality test. The Computer Journal, 59(11):1689–1694, November 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/11/1689.

Li:2013:EPD

[LLTY13] Jiuyong Li, Jixue Liu, Hannu Toivonen, and Jianming Yong. Effective pruning for the discovery of conditional functional dependencies. The Computer Journal, 56(3):378– 392, March 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/3/378.full.
                                        pdf+html.

Locatelli:2010:ASC

[LLV10] Marco P. Locatelli, Marco Loregian, and Giuseppe Vizzari. Artificial societies in a community-based approach to ambient intelligence. The Computer Journal, 53(8):1152–1168, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1152.
full.pdf+html.

Luo:2012:FSI

[LLY^{+}12] Xiangyang Luo, Fenlin Liu, Chunfang Yang, Shiguo Lian, and Daoshun Wang. On F5 steganography in images. The Computer Journal, 55(4):447–456, April 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/4/447.full.
pdf+html.

Li:2015:QDF

[LLZY15] Jianxin Li, Chengfei Liu, Rui Zhou, and Jeffrey Xu Yu. Query-driven frequent co-occurring term computation over relational data using MapReduce. The Computer Journal, 58(3):497– 513, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/3/497.

Liu:2017:TIT

[LM17] Yu Liu and Peter McBrien.

Transactional and incremental type inference from data updates. The Computer Journal, 60(3):60– ??, March 2017. CO-DEN CMPJA6. ISSN ????

                                        URL https://academic.
                                        oup.com/comjnl/article/
                                        60/3/347/2608051/Transactional-
                                        and-Incremental-Type-Inference-
                                        from.

Lengyel:2015:QAM

[LMA^{+}15] László Lengyel, Tamás Mészáros, Márk Asztalos, Péter Boros, Attila Máté, Gábor Madács, Péter Hudák, Kristóf Kovács, András Tresch, and Hassan Charaf. Quality assured model-driven requirements engineering and software development. The Computer Journal, 58(11):3171–3186, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Lai:2018:IBB

[LMG^{+}18] Jianchang Lai, Yi Mu, Fuchun Guo, Peng Jiang, and Sha Ma. Identity-based broadcast encryption for inner products. The Computer Journal, 61(8):1240–1251, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1240/5035766.

Lai:2017:FPP

[LMGC17] Jianchang Lai, Yi Mu, Fuchun Guo, and Rongmao Chen. Fully privacypreserving ID-based broadcast encryption with authorization. The Computer Journal, 60(12):1809–1821, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1809/3861972.

Luccio:2016:CBR

[LMMP16] Fabrizio Luccio, Bernard Mans, Luke Mathieson, and Linda Pagli. Complete balancing via rotation. The Computer Journal, 59(8):1252–1263, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1252.

Lettieri:2018:SPV

[LMR18] Giuseppe Lettieri, Vincenzo Maffione, and Luigi Rizzo. A study of I/O performance of virtual machines. The Computer Journal, 61(6): 808–831, June 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/808/4259797.

Lopez-Nores:2013:CBP

[LNBFPA13] Martín López-Nores, Yolanda Blanco-Fernández, and José Juan Pazos-Arias. Cloud-based personalization of new advertising and e-commerce models for video consumption. The Computer Journal, 56(5):573–592, May 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/5/573.full.
pdf+html.

Ling:2019:SAR

[LNWZ19] San Ling, Khoa Nguyen, Huaxiong Wang, and Juanyang Zhang. Server-aided revocable predicate encryption: Formalization and lattice-based instantiation. The Computer Journal, 62(12):1849–1862, December 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/12/1849/5628022.

Loukas:2010:PAD

[LÖ10] Georgios Loukas and Gülay Öke. Protection against denial of service attacks: a survey. The Computer Journal, 53(7):1020–1037, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/1020;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1020.

Lopriore:2012:EPP

[Lop12] Lanfranco Lopriore. Encrypted pointers in protection system design. The Computer Journal, 55(4): 497–507, April 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/4/497.full.
pdf+html.

Lopriore:2013:PSM

[Lop13] Lanfranco Lopriore. Protection structures in multithreaded systems. The Computer Journal, 56(4): 478–496, April 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/4/478.full.
pdf+html.

Lopriore:2015:PCR

[Lop15a] Lanfranco Lopriore. Password capabilities revisited. The Computer Journal, 58(4):782–791, April 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/782.

Lopriore:2015:PMD

[Lop15b] Lanfranco Lopriore. Password management: Distribution, review and revocation. The Computer Journal, 58(10):2557–2566, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2557.

Lee:2014:GSB

[LP14] Hwamin Lee and Doosoon Park. A Grid servicebased virtual screening system. The Computer Journal, 57(2):302–307, February 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/2/302.full.
pdf+html.

Liu:2013:GBD

[LPD13] Qingzhi Liu, Andrei Pruteanu, and Stefan Dulman. Gradientbased distance estimation for spatial computers. The Computer Journal, 56(12):1469–1499, December 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/12/1469.
full.pdf+html.

Lee:2014:CMS

[LPL14] Hong-Chang Lee, Jong-Eun Park, and Myung-Joon Lee. C3ware: a middleware supporting collaborative services over cloud storage. The Computer Journal, 57(2):217–224, February 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/2/217.full.
pdf+html.

Lian:2015:GRG

[LPL15] Chunfeng Lian, Liaojun Pang, and Jimin Liang. Generalized random gridbased visual secret sharing for general access structures. The Computer Journal, 58(10):2426–2442, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2426.

Langohr:2013:CSD

[LPP^{+}13] Laura Langohr, Vid Podpečan, Marko Petek, Igor Mozetič, Kristina Gruden, Nada Lavrač, and Hannu Toivonen. Contrasting subgroup discovery. The Computer Journal, 56(3):289– 303, March 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/3/289.full.
pdf+html.

Loz:2010:NBL

[LPV10] Eyal Loz and Guillermo Pineda-Villavicencio. New benchmarks for large-scale networks with given maximum degree and diameter. The Computer Journal, 53(7):1092–1105, September 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/7/1092;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/7/1092.

Liao:2010:EIH

[LQZ^{+}10] Jianxin Liao, Qi Qi, Xiaomin Zhu, Yufei Cao, and Tonghong Li. Enhanced IMS handoff mechanism for QoS support over heterogeneous network. The Computer Journal, 53(10):1719– 1737, December 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/53/10/1719.
                                        full.pdf+html.

Lee:2010:SDC

[LR10] Seung Min Lee and Stephen J. Roberts. Sequential dynamic classification using latent variable models. The Computer Journal, 53(9):1415–1429, November 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1415.
full.pdf+html.

Lee:2012:ANC

[LR12] Sangpil Lee and Won W.

Ro. Accelerated network coding with dynamic stream decomposition on graphics processing unit. The Computer Journal, 55(1):21–34, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/1/21.full.
pdf+html.

Lim:2014:CAM

[LR14] Sejoon Lim and Daniela Rus. Congestion-aware multi-agent path planning: Distributed algorithm and applications. The Computer Journal, 57(6):825– 839, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/6/825.full.
pdf+html.

Liao:2014:LBA

[LS14] Kewen Liao and Hong Shen.

LP-based approximation algorithms for reliable resource allocation. The Computer Journal, 57(1):154– 164, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/1/154.full.
                                        pdf+html.

Liu:2017:APP

[LS17] Yang Liu and Andrew Simpson. AdSelector: a privacypreserving advertisement selection mechanism for mobile devices. The Computer Journal, 60(8):1251–1270, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/8/1251/3100392.

Lloret:2010:SEW

[LSCG10] Jaime Lloret, Sandra Sendra, Hugo Coll, and Miguel Garcia. Saving energy in wireless local area sensor networks. The Computer Journal, 53(10):1658–1673, December 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/10/1658.
                                        full.pdf+html.

Liu:2019:IMM

[LSG^{+}19] Ya Liu, Yifan Shi, Dawu Gu, Zhiqiang Zeng, Fengyu Zhao, Wei Li, Zhiqiang Liu, and Yang Bao. Improved meet-in-the-middle attacks on reduced-round Kiasu-BC and Joltik-BC. The Computer Journal, 62(12):1761– 1776, December 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/12/1761/5525447.

Liang:2015:EFC

[LSLW15] Kaitai Liang, Willy Susilo, Joseph K. Liu, and Duncan S. Wong. Efficient and fully CCA secure conditional proxy re-encryption from hierarchical identitybased encryption. The Computer Journal, 58(10):2778– 2792, October 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/10/2778.

Lin:2018:CPF

[LSQL18a] Xi-Jun Lin, Lin Sun, Haipeng Qu, and Dongxiao Liu. Cryptanalysis of a pairing-free certificateless signcryption scheme. The Computer Journal, 61(4): 539–544, April 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/539/4608880.

Lin:2018:SSS

[LSQL18b] Xi-Jun Lin, Lin Sun, Haipeng Qu, and Dongxiao Liu. On the security of secure server-designation public key encryption with keyword search. The Computer Journal, 61(12):1791–1793, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1791/5055854.

Lin:2019:CCA

[LSQX19] Xi-Jun Lin, Lin Sun, Haipeng Qu, and He-Qun Xian. Cryptanalysis of a compact anonymous HIBE with constant size private keys. The Computer Journal, 62(8):1087–1091, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1087/5253748.

Lin:2017:ESF

[LSQZ17] Xi-Jun Lin, Lin Sun, Haipeng Qu, and Xiaoshuai Zhang. Editorial: On the security of the first leakagefree certificateless signcryption scheme. The Computer Journal, 60(4):491– 496, March 23, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/491/2608059.

Liu:2011:FAH

[LSTC11] Feng Liu, Xiaoyu Song, Qingping Tan, and Gang Chen. Formal analysis of hybrid prefix/carry-select arithmetic systems. The Computer Journal, 54(6): 894–904, June 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/6/894.full.
pdf+html.

Lin:2010:RBM

[LSW10] Jinjiao Lin, Chengxiang Song, and Haiyang Wang. A rule-based method for improving adaptability in pervasive systems. The Computer Journal, 53(2):177– 190, February 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/2/177;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/2/177.

Liu:2016:EPP

[LSY^{+}16] Joseph K. Liu, Willy Susilo, Tsz Hon Yuen, Man Ho Au, Junbin Fang, Zoe L. Jiang, and Jianying Zhou. Efficient privacy-preserving charging station reservation system for electric vehicles. The Computer Journal, 59(7): 1040–1053, July 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/7/1040.

Liu:2015:MSG

[LTC^{+}15] Chao-Liang Liu, Wang-Jui Tsai, Ting-Yi Chang, Chun-Cheng Peng, and Peng-Shiang Wong. Meaningful share generation for (2, 2)- multiple visual secret sharing scheme without pixel expansion. The Computer Journal, 58(7):1598–1606, July 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/7/1598.

Li:2015:FSC

[LTH^{+}15] Jiguo Li, Huiyun Teng, Xinyi Huang, Yichen Zhang, and Jianying Zhou. A forward-secure certificatebased signature scheme. The Computer Journal, 58(4): 853–866, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/4/853.

Lin:2010:ITL

[LTL10] Chih-Yu Lin, Yu-Chee Tseng, and Yung-Chih Liu. Imprecision-tolerant location management for object-tracking wireless sensor network. The Computer Journal, 53(3):351– 364, March 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/3/351;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/3/351.

Lin:2010:MKM

[LTW10] Iuon-Chang Lin, Shih-Shan Tang, and Chung-Ming Wang. Multicast key management without rekeying processes. The Computer Journal, 53(7):939– 950, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/939;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/939.

Li:2016:LRC

[LTZY16] Jiguo Li, Meilin Teng, Yichen Zhang, and Qihong Yu. A leakage-resilient CCA-secure identity-based encryption scheme. The Computer Journal, 59(7): 1066–1075, July 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/7/1066.

Lomax:2017:CSD

[LV17] Susan Lomax and Sunil Vadera. A cost-sensitive decision tree learning algorithm based on a multiarmed bandit framework. The Computer Journal, 60(7):941–956, July 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/7/941/2609376.

Liu:2016:PAB

[LW16] Zhen Liu and Duncan S. Wong. Practical attributebased encryption: Traitor tracing, revocation and large universe. The Computer Journal, 59(7):983– 1004, July 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/7/983.

Liu:2015:CLC

[LWC15] Jia-Jie Liu, Yue-Li Wang, and Yu-Shan Chiu. Constrained longest common subsequences with runlength-encoded strings. The Computer Journal, 58(5): 1074–1084, May 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/5/1074.

Li:2016:OMA

[LWDZ16] Liang Li, Endong Wang, Xiaoshe Dong, and Zhengdong Zhu. The optimization of memory access congestion for MapReduce applications on manycore systems. The Computer Journal, 59(3): 325–337, March 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/3/325.

Larkin:2015:DSS

[LWKB15] Andrew Larkin, David E.

Williams, Molly L. Kile, and William M. Baird. Developing a Smartphone software package for predicting atmospheric pollutant concentrations at mobile locations. The Computer Journal, 58(6):1431–1442, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1431.

Liu:2010:SET

[LWL10] Feng Liu, ChuanKun Wu, and XiJun Lin. Some extensions on threshold visual cryptography schemes. The Computer Journal, 53(1): 107–119, January 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/107;   http://comjnl.

                                        oxfordjournals.org/cgi/
                                        reprint/53/1/107.

Liu:2017:SRG

[LWL^{+}17] Jing Liu, Yunyun Wu, Xuezheng Liu, Yunchun Zhang, Gang Xue, Wei Zhou, and Shaowen Yao. On the (in)security of recent group key distribution protocols. The Computer Journal, 60(4):507–526, March 23, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/4/507/2608062.

Lin:2013:EEC

[LWPZ13] Kai Lin, Xiaofei Wang, Limei Peng, and Xuan Zhu. Energy-efficient Kcover problem in hybrid sensor networks. The Computer Journal, 56(8):957– 967, August 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/8/957.full.
                                        pdf+html.

Liu:2014:CAC

[LWS^{+}14] Zhipiao Liu, Shangguang Wang, Qibo Sun, Hua Zou, and Fangchun Yang. Costaware cloud service request scheduling for SaaS providers. The Computer Journal, 57(2):291– 301, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/291.full.
pdf+html.

Liu:2013:DTC

[LWW13] Jiaojiao Liu, Yige Wang, and Gang Wei. Distributed topology control based on coalition formation game in wireless networks. The Computer Journal, 56(8): 968–976, August 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/8/968.full.
pdf+html.

Lin:2017:VNP

[LWYZ17] Wang Lin, Min Wu, Zhengfeng Yang, and Zhenbing Zeng. Verification for non-polynomial hybrid systems using rational invariants. The Computer Journal, 60(5):675– 689, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/5/675/2632263.

Liu:2018:TBG

[LWZ^{+}18] Jialei Liu, Shangguang Wang, Ao Zhou, Xiang Xu, Sathish A. P. Kumar, and Fangchun Yang. Towards bandwidth guaranteed virtual cluster reallocation in the cloud. The Computer Journal, 61(9):1284–1295, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/9/1284/4779880.

Lin:2010:RSP

[LY10] Jenn-Wei Lin and Ming-Feng Yang. Robust superpeer-based P2P file-sharing systems. The Computer Journal, 53(7):951–968, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/7/951;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/7/951.

Liu:2011:OAU

[LYC11] Jia-Jie Liu, William Chung-Kung Yen, and Yen-Ju Chen. An optimal algorithm for untangling binary trees via rotations. The Computer Journal, 54(11):1838–1844, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Liu:2018:IMM

[LYD^{+}18] Ya Liu, Anren Yang, Bo Dai, Wei Li, Zhiqiang Liu, Dawu Gu, and Zhiqiang Zeng. Improved meetin-the middle attacks on reduced-round TWINE-128. The Computer Journal, 61(8):1252–1258, August 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/8/1252/5040794.

Lin:2018:SEI

[LYL^{+}18] Tingting Lin, Hailun Yan, Xuejia Lai, Yixin Zhong, and Yin Jia. Security evaluation and improvement of a white-box SMS4 implementation based on affine equivalence algorithm. The Computer Journal, 61 (12):1783–1790, December 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1783/5055352.

Li:2017:SRD

[LYPL17] Xiaoyin Li, Lianshan Yan, Wei Pan, and Bin Luo. Secure and robust DV-hop localization based on the vector refinement feedback method for wireless sensor networks. The Computer Journal, 60(6):810– 821, June 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/810/2962043.

Liu:2016:FIT

[LYY^{+}16] Joseph K. Liu, Sze Ling Yeo, Wun-She Yap, Sherman S. M. Chow, Duncan S. Wong, and Willy Susilo. Faulty instantiations of threshold ring signature from threshold proofof-knowledge protocol. The Computer Journal, 59(7): 945–954, July 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/7/945.

Lai:2018:NSH

[LYY^{+}18a] Qiqi Lai, Bo Yang, Yong Yu, Yuan Chen, and Jian Bai. Novel smooth hash proof systems based on lattices. The Computer Journal, 61(4):561–574, April 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/4/561/4725104.

Liu:2018:SKR

[LYY^{+}18b] Jinhui Liu, Yong Yu, Bo Yang, Jianwei Jia, Shijia Wang, and Houzhen Wang. Structural key recovery of simple matrix encryption scheme family. The Computer Journal, 61 (12):1880–1896, December 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1880/5110544.

Lin:2019:NBC

[LZ19] Yuan Lin and Zhongzhi Zhang. Non-backtracking centrality based random walk on networks. The Computer Journal, 62(1): 63–80, January 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/63/4953375.

Li:2014:CSR

[LZHS14] Yongjian Li, Naiju Zeng, William N. N. Hung, and Xiaoyu Song. Combining symmetry reduction with generalized symbolic trajectory evaluation. The Computer Journal, 57(1):115– 128, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/115.full.
pdf+html.

Li:2015:ITA

[LZL^{+}15] Jianxin Li, Jieyu Zhao, Yi Li, Lei Cui, Bo Li, Lu Liu, and John Panneerselvam. iMIG: Toward an adaptive live migration method for KVM virtual machines. The Computer Journal, 58(6):1227– 1242, June 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1227.

Li:2017:UDC

[LZL^{+}17] Shichao Li, Gang Zhu, Siyu Lin, Chao Shen, Qian Gao, Weiliang Xie, and Xiaoyu Qiao. Ultra dense cells management and resource allocation in green softwaredefined wireless networks. The Computer Journal, 60 (10):1472–1481, October 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/10/1472/3073685.

Li:2019:RBT

[LZL^{+}19] Ximing Li, Ang Zhang, Changchun Li, Lantian Guo, Wenting Wang, and Jihong Ouyang. Relational biterm topic model: Short-text topic modeling using word embeddings. The Computer Journal, 62(3):359–372, March 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/359/5005494.

Luna:2016:DMO

[LZN^{+}16] Francisco Luna, Gustavo R. Zavala, Antonio J. Nebro, Juan J. Durillo, and Carlos A. Coello Coello. Distributed multi-objective metaheuristics for realworld structural optimization problems. The Computer Journal, 59(6):777– 792, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/6/777.

Liu:2018:LPP

[LZWY18] Tong Liu, Yanmin Zhu, Ting Wen, and Jiadi Yu. Location privacy-preserving method for auction-based incentive mechanisms in mobile crowd sensing. The Computer Journal, 61(6): 937–948, June 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/937/4781603.

Liu:2017:OPE

[LZZ^{+}17] Tong Liu, Yanmin Zhu, Hongzi Zhu, Jiadi Yu, Yuanyuan Yang, and Fan Ye. Online pricing for efficient renewable energy sharing in a sustainable microgrid. The Computer Journal, 60(2):60– ??, February 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/2/268/2882689/Online-

                                        Pricing-for-Efficient-
                                        Renewable-Energy.

Lei:2013:FLA

[LZZZ13] Lei Lei, Jing Zhao, Zhangdui Zhong, and Kan Zheng. Flow-level analysis of energy efficiency performance for device-to-device communications in OFDM cellular networks. The Computer Journal, 56(8):1001– 1009, August 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/8/1001.
                                        full.pdf+html.

Ma:2017:AEJ

[Ma17] Sha Ma. Authorized equijoin for multiple data contributors in the PKC-based setting. The Computer Journal, 60(12):1822–1838, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/12/1822/3861973.

Majumdar:2010:BRP

[Maj10] Rupak Majumdar. Book review: Paul Ammann and Jeff Offutt, Introduction to Software Testing. Cambridge University Press (2008). ISBN-13 978-0-521-88038-1. £32.99. 322 pp. Hardcover. The Computer Journal, 53(5):615, June 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/5/615.

Malacaria:2010:PAP

[Mal10] Pasquale Malacaria. Program analysis probably counts: Discussant contribution for the Computer Journal Lecture by Chris Hankin. The Computer Journal, 53(6):881, July 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/6/881.

Marecek:2010:BRB

[Mar10a] Jakub Mareček. Book review: Handbook of Approximation Algorithms and Metaheuristics. The Computer Journal, 53(8):1338– 1339, October 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/8/1338.
full.pdf+html.

Marshall:2010:MDR

[Mar10b]  R. Marshall.  Modeling

DNA/RNA strings using resistor–capacitor (RC) ladder networks. The Computer Journal, 53(6):644– 660, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/6/644;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/6/644.

Yeung:2010:FMO

[mAYL10] Ching man Au Yeung and Ho-Fung Leung. A formal model of ontology for handling fuzzy membership and typicality of instances. The Computer Journal, 53(3): 316–341, March 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/3/316;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/3/316.

Mefteh:2016:MFM

[MBBA16] Mariem Mefteh, Nadia Bouassida, and Hanêne Ben-Abdallah. Mining feature models from functional requirements. The Computer Journal, 59(12):1784– 1804, December 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/12/1784.

Mao:2015:PUA

[MBC15] Qian Mao, K. Bharanitharan, and Chin-Chen Chang.

A proxy user authentication protocol using source-based image morphing. The Computer Journal, 58(7):1573– 1584, July 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/7/1573.

Munoz:2015:SRA

[MBRM15] Pablo Muñoz, David F.

Barrero, and María D. R-Moreno. A statistically rigorous analysis of 2D pathplanning algorithms. The Computer Journal, 58(11): 2876–2891, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Mou:2019:CSD

[MCT19] Chengcheng Mou, Shaoping Chen, and Yi-Cheng Tu. A comparative study of dual-tree algorithms for computing spatial distance histograms. The Computer Journal, 62(1):42–62, January 1, 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/42/4942083.

Meddeb:2018:CFN

[MDB^{+}18] Maroua Meddeb, Amine Dhraief, Abdelfettah Belghith, Thierry Monteil, Khalil Drira, and Saad AlAhmadi. Cache freshness in named data networking for the Internet of things. The Computer Journal, 61(10):1496–1511, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1496/4827074.

Manzalini:2011:SOC

[MDN^{+}11] A. Manzalini, P. H. Deussen, S. Nechifor, M. Mamei, R. Minerva, C. Moiso, A. Salden, T. Wauters, and F. Zambonelli. Selfoptimized cognitive network of networks. The Computer Journal, 54(2):189– 196, February 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/2/189.full.
pdf+html.

Marudhadevi:2015:TEM

[MDS15] D. Marudhadevi, V. Neelaya Dhatchayani, and V. S. Shankar Sriram. A trust evaluation model for cloud computing using service level agreement. The Computer Journal, 58(10):2225–2232, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2225.

Memar:2012:EFI

[MDSF12] Mina Memar, Mahmood Deypir, Mohammad Hadi Sadreddini, and Seyyed Mostafa Fakhrahmad. An efficient frequent itemset mining method over high-speed data streams. The Computer Journal, 55(11):1357– 1366, November 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/11/1357.
full.pdf+html.

Ma:2015:LSS

[MDY15] Kun Ma, Fusen Dong, and Bo Yang. Large-scale schema-free data deduplication approach with adaptive sliding window using MapReduce. The Computer Journal, 58(11):3187–3201, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Munoz-Escoi:2019:CTR

[MEdJMGE^{+}19] Francesc D. Muñoz-Escoí, Rubén de Juan-Marín, José-Ramón García-Escrivá, J. R. González de Mendívil, and José M. Bernabéu-Aubán. CAP theorem: Revision of its related consistency models. The Computer Journal, 62(6): 943–960, June 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/6/943/5381952.

Meghanathan:2016:MAM

[Meg16] Natarajan Meghanathan. Maximal assortative matching and maximal dissortative matching for complex network graphs. The Computer Journal, 59(5):667– 684, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/5/667.

Meghanathan:2018:CNA

[Meg18] Natarajan Meghanathan. Complex network analysisbased graph theoretic metrics to determine stable data gathering trees for mobile sensor networks. The Computer Journal, 61(2):199– 222, February 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/2/199/3861966.

Meghanathan:2019:CPC

[Meg19] Natarajan Meghanathan. Centrality and partial correlation coefficient-based assortativity analysis of realworld networks. The Computer Journal, 62(9):1247– 1264, September 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/9/1247/5112948.

Mohamed:2019:HDC

[MEH19] Khalil Mohamed, Ayman El Shenawy, and Hany Harb. A hybrid decentralized coordinated approach for multi-robot exploration task. The Computer Journal, 62(9):1284– 1300, September 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/9/1284/5146171.

Melucci:2013:DQI

[Mel13] Massimo Melucci. Deriving a quantum information retrieval basis. The Computer Journal, 56(11):1279–1291, November 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/11/1279.
full.pdf+html.

Merca:2013:BDS

[Mer13] Mircea Merca. Binary diagrams for storing ascending compositions. The Computer Journal, 56(11):1320– 1327, November 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/56/11/1320.
                                        full.pdf+html.

Movahedi:2015:ASB

[MGBD15] Zahra Movahedi, Walid Gaaloul, Sami Bhiri, and Bruno Defude. Assisting sensor-based application design and instantiation using activity recommendation. The Computer Journal, 58(3):368– 384, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/3/368.

Munoz:2012:POM

[MGM12] Antonio Muñoz, Javier Gonzalez, and Antonio Maña. A performanceoriented monitoring system for security properties in cloud computing applications. The Computer Journal, 55(8):979–994, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/979.full.
                                        pdf+html.

Meng:2018:EGQ

[MGZ18] Xiaodong Meng, Minyi Guo, and Jingyu Zhang. An efficient graph query framework with structural recursion. The Computer Journal, 61(1):144–157, January 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/144/3920736.

McCarthy:2011:EUO

[MH11] Mitzi McCarthy and Zhen He. Efficient updates for OLAP range queries on flash memory. The Computer Journal, 54(11):1773–1789, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Munoz-Hernandez:2016:EES

[MHMSGH16] Mario Diego Munoz-Hernandez, Miguel Morales-Sandoval, and Jose Juan Garcia-Hernandez. An end-to-end security approach for digital document management. The Computer Journal, 59(7):1076–1090, July 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/7/1076.

Moore:2010:SCC

[MHW10] Philip Moore, Bin Hu, and Jizheng Wan. Smartcontext: a context ontology for pervasive mobile computing. The Computer Journal, 53(2):191–207, February 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/2/191;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/2/191.

Milner:2010:DRB

[Mil10] Robin Milner. Discussant of response to the Computer Journal Lecture by Morris Sloman. The Computer Journal, 53(7):1128, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/reprint/53/7/1128.

Misra:2014:CIF

[Mis14] Janardan Misra. A calculus of incomplete fusions. The Computer Journal, 57(7): 1105–1116, July 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/7/1105.
                                        full.pdf+html.

Mitrani:2010:MSF

[Mit10] Isi Mitrani. Management of server farms for performance and profit. The Computer Journal, 53(7):1038– 1044, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/1038;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1038.

Mitchell:2012:BC

[Mit12] Melanie Mitchell. Biological computation. The Computer Journal, 55(7):852– 855, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/852.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Mitchell:2019:SIG

[Mit19]  Chris J. Mitchell.  Se-

curity issues in a group key establishment protocol. The Computer Journal, 62(3):373–376, March 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/3/373/5033364.

Mavromoustakis:2011:ESO

[MK11] Constandinos X. Mavromoustakis and Helen D. Karatza. Embedded sociooriented model for end-to-end reliable stream schedules by using collaborative outsourcing in Mp2P systems. The Computer Journal, 54(8):1235–1247, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/8/1235.
                                        full.pdf+html.

Mavromoustakis:2013:PEO

[MK13] Constandinos X. Mavromoustakis and Helen D. Karatza. Performance evaluation of opportunistic resource-sharing scheme using socially oriented outsourcing in wireless devices. The Computer Journal, 56(2):184–197, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/2/184.full.
                                        pdf+html.

Maalej:2015:SLW

[MK15] Afef Jmal Maâlej and Moez Krichen. Study on the limitations of WS-BPEL compositions under load conditions. The Computer Journal, 58(3):385– 402, March 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/3/385.

Mukhedkar:2019:TBS

[MK19] Moresh Madhukar Mukhedkar and Uttam Kolekar.

Trust-based secure routing in mobile ad hoc network using hybrid optimization algorithm. The Computer Journal, 62(10):1528–1545, October 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/10/1528/5544160.

Mazur:2015:RMM

[MKK15] Katarzyna Mazur, Bogdan Ksiezopolski, and Zbigniew Kotulski. The robust measurement method for security metrics generation. The Computer Journal, 58(10):2280–2296, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2280.

Mula:2018:FPC

[MKL18] Wojciech Muła, Nathan Kurz, and Daniel Lemire. Faster population counts using AVX2 instructions. The Computer Journal, 61 (1):111–120, January 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/111/3852071.

Marconato:2013:VLC

[MKN13] G. Vache Marconato, M. Kaâniche, and V. Nicomette. A vulnerability life cycle-based security modeling and evaluation approach. The Computer Journal, 56(4):422– 439, April 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/4/422.full.
pdf+html.

Mokbel:2011:ASR

[MKW11] Mohammed F. Mokbel, Robert D. Kent, and Michael Wong. An abstract semantically rich compiler collocative and interpretative model for OpenMP programs. The Computer Journal, 54(8):1325–1343, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/8/1325.
full.pdf+html.

Mahmood:2013:RUC

[ML13] Sajjad Mahmood and Richard Lai. RE-UML: a componentbased system requirements analysis language. The Computer Journal, 56(7): 901–922, July 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/7/901.full.
pdf+html.

Mergen:2017:DMF

[MM17] Sergio L. S. Mergen and Viviane P. Moreira. DuelMerge: Merging with fewer moves. The Computer Journal, 60(9):1271–1278, September 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/9/1271/3062794.

Martinez:2019:CCP

[MMAY19] Abel Cabrera Martínez, Frank A. Hernández Mira, José M. Sigarreta Almira, and Ismael G. Yero. On computational and combinatorial properties of the total co-independent domination number of graphs. The Computer Journal, 62(1):97–108, January 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/97/4969887.

Mittal:2013:VLH

[MMB13] Anish Mittal, Anush K.

Moorthy, and Alan C. Bovik. Visually lossless H.264 compression of natural videos. The Computer Journal, 56(5):617– 627, May 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

                                        org/content/56/5/617.full.
                                        pdf+html.

Mamede:2018:BAN

[MMH18] Margarida Mamede, José Legatheaux Martins, and João Horta. BOUQUET — aggregating network paths in trees to reduce data-plane forwarding state. The Computer Journal, 61(10):1512– 1522, October 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/10/1512/4883869.

Mileo:2010:LAH

[MMPB10] Alessandra Mileo, Davide Merico, Stefano Pinardi, and Roberto Bisiani. A logical approach to home healthcare with intelligent sensor-network support. The Computer Journal, 53(8):1257–1276, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/8/1257.
                                        full.pdf+html.

Moller:2013:MKG

[Möl13] Bernhard Möller. Modal knowledge and game semirings. The Computer Journal, 56(1):53–69, January 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/1/53.full.
pdf+html.

Maneth:2017:DS

[MP17] Sebastian Maneth and Alexandra Poulovassilis. Data science. The Computer Journal, 60(3):60– ??, March 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/3/285/2608072/Data-
Science.

Macwan:2018:NAS

[MP18] Kamalkumar R. Macwan and Sankita J. Patel. k-NMF anonymization in social network data publishing. The Computer Journal, 61(4):601–613, April 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/601/4843992.

Mostafa:2014:MPD

[MPH14] Hala Mostafa, Partha Pal, and Patrick Hurley. Message passing for distributed QoS-security tradeoffs. The Computer Journal, 57(6): 840–855, June 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/6/840.full.
pdf+html.

Mitrokotsa:2013:SNL

[MPLDV13] Aikaterini Mitrokotsa, Pedro Peris-Lopez, Christos Dimitrakakis, and Serge Vaudenay. On selecting the nonce length in distance-bounding protocols. The Computer Journal, 56(10):1216–1227, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/10/1216.
                                        full.pdf+html.

Mewada:2015:NST

[MPP15] Hiren Mewada, Rahul Patel, and Suprava Patnaik. A novel structure tensor modulated Chan–Vese model for texture image segmentation. The Computer Journal, 58(9):2044–2060, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/2044.

Marandi:2017:RPH

[MPSP17] Parisa Jalili Marandi, Marco Primi, Nicolas Schiper, and Fernando Pedone. Ring Paxos: High-throughput atomic broadcast. The Computer Journal, 60(6): 866–882, June 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/866/3058780.

Miller:2015:MLA

[MRPR15] Mirka Miller, R. Sundara Rajan, N. Parthiban, and Indra Rajasingh. Minimum linear arrangement of incomplete hypercubes. The Computer Journal, 58(2):331–337, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/331.

Mukherjee:2011:CRS

[MS11] Partha Mukherjee and Sandip Sen. Comparing reputation schemes for detecting malicious nodes in sensor networks. The Computer Journal, 54(3):482– 489, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/482.full.
pdf+html.

Mazurczyk:2012:TER

[MS12] Wojciech Mazurczyk and Krzysztof Szczypiorski. Toward effective and reliable digital forensics. The Computer Journal, 55(6):651– 652, June 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/6/651.full.
                                        pdf+html.

Mutlu:2014:IHR

[MS14] Alev Mutlu and Pinar Senkul. Improving hit ratio of ILP-based concept discovery system with memoization. The Computer Journal, 57(1):138– 153, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/1/138.full.
                                        pdf+html.

March:2011:NEA

[MSH^{+}11] José Luis March, Julio Sahuquillo, Houcine Hassan, Salvador Petit, and José Duato. A new energyaware dynamic task set partitioning algorithm for soft and hard embedded realtime systems. The Computer Journal, 54(8):1282– 1294, August 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/8/1282.
                                        full.pdf+html.

Maitin-Shepard:2017:ECM

[MSTA17]  Jeremy  Maitin-Shepard,

Mehdi Tibouchi, and Diego F. Aranha. Elliptic curve multiset hash. The Computer Journal, 60(4):476– 490, March 23, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/476/2608055.

McIntosh-Smith:2012:BEE

[MSWI^{+}12] Simon McIntosh-Smith, Terry Wilson, Amaurys Ávila Ibarra, Jonathan Crisp, and Richard B. Sessions. Benchmarking energy efficiency, power costs and carbon emissions on heterogeneous systems. The Computer Journal, 55(2):192– 205, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/2/192.full.
pdf+html.

Markham:2011:AED

[MT11] Andrew Markham and Niki Trigoni. The automatic evolution of distributed controllers to configure sensor network operation. The Computer Journal, 54(3): 421–438, March 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/3/421.full.
pdf+html.

Murtagh:2010:CAP

[Mur10a] Fionn Murtagh. The correspondence analysis platform for uncovering deep structure in data and information. The Computer Journal, 53(3):304– 315, March 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/3/304;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/3/304.

Murtagh:2010:UAI

[Mur10b] Fionn Murtagh. On ultrametric algorithmic information. The Computer Journal, 53(4):405– 416, May 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/4/405;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/4/405.

Min:2016:RSC

[MV16] Byungho Min and Vijay Varadharajan. Rethinking software component security: Software component level integrity and cross verification. The Computer Journal, 59(11):1735–1748, November 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1735.

Mefenza:2019:CSA

[MV19] Thierry Mefenza and Damien Vergnaud. Cryptanalysis of server-aided RSA protocols with private-key splitting. The Computer Journal, 62(8):1194–1213, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1194/5488732.

Ma:2015:PKE

[MZHY15] Sha Ma, Mingwu Zhang, Qiong Huang, and Bo Yang. Public key encryption with delegated equality test in a multi-user setting. The Computer Journal, 58(4): 986–1002, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/4/986.

Mao:2018:CML

[MZW^{+}18] Qirong Mao, Feifei Zhang, Liangjun Wang, Sidian Luo, and Ming Dong. Cascaded multi-level transformed Dirichlet process for multi-pose facial expression recognition. The Computer Journal, 61(11):1605–1619, November 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/11/1605/4930692.

Nguyen:2012:SGM

[NB12] Tien Viet Nguyen and François Baccelli. A stochastic geometry model for cognitive radio networks. The Computer Journal, 55 (5):534–552, May 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/5/534.full.
                                        pdf+html.

Nguyen:2017:ETM

[NB17] Tung Doan Nguyen and Quan Bai. Enhance trust management in composite services with indirect ratings. The Computer Journal, 60(11):1619–1632, November 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/11/1619/3073683.

Ngo:2014:APC

[NBN14] Long Ngo, Colin Boyd, and Juan González Nieto. Automated proofs for computational indistinguishability. The Computer Journal, 57(10):1513–1536, October 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/10/1513.

Neuman:2016:NMA

[NC16] Yair Neuman and Yochai Cohen. A novel methodology for automatically measuring psychological dimensions in textual data. The Computer Journal, 59(9):1408–1414, September 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/9/1408.

Nedjah:2016:PDC

[NdMCdMM16] Nadia Nedjah, Rogério de Moraes Calazan, and Luiza de Macedo Mourelle. Particle, dimension and cooperation-oriented PSO parallelization strategies for efficient high-dimension problem optimizations on graphics processing units. The Computer Journal, 59 (6):810–835, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/59/6/810.

Nazemi:2017:DNN

[NG17] Alireza Nazemi and Ozra Ghezelsofla. A dual neural network scheme for solving the assignment problem. The Computer Journal, 60 (3):60–??, March 2017. CO-DEN CMPJA6. ISSN ????

                                        URL https://academic.
                                        oup.com/comjnl/article/
                                        60/3/431/2609374/A-Dual-
                                        Neural-Network-Scheme-
                                        for-Solving-the.

Nafea:2016:HMB

[NGAuHQ16] Ohoud Nafea, Sanaa Ghouzali, Wadood Abdul, and Emad ul Haq Qazi. Hybrid multibiometric template protection using watermarking. The Computer Journal, 59(9):1392–1407, September 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/9/1392.

Naher:2019:TAE

[NH19] Nazmun Naher and Tanzima Hashem. Think ahead: Enabling continuous sharing of location data in real-time with privacy guarantee. The Computer Journal, 62(1):1– 19, January 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/1/1/4796925.

Nguyen:2013:TEB

[NHC13] Thi Nguyen, Zhen He, and Yi-Ping Phoebe Chen. S^{e}TPR^{∗}-tree: Efficient buffering for spatiotemporal indexes via shared execution. The Computer Journal, 56(1):115–137, January 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/1/115.full.
pdf+html.

Nagao:2013:TSM

[NHMI13] Hiromichi Nagao, Tomoyuki Higuchi, Satoshi Miura, and Daisuke Inazu. Timeseries modeling of tide gauge records for monitoring of the crustal activities related to oceanic trench earthquakes around Japan. The Computer Journal, 56(3):355– 364, March 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/3/355.full.
pdf+html.

Ni:2016:MEP

[Ni16] Wei Ni. Minimized error propagation location method based on error estimation. The Computer Journal, 59(9):1282–1288, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/9/1282.

Nicholson:2011:DAA

[Nic11]  David Nicholson.  De-

fence applications of agentbased information fusion. The Computer Journal, 54(2):263–273, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/2/263.full.
                                        pdf+html.

Nilsson:2010:CAM

[Nil10] Marcus Nilsson. Computational aspects of monomial dynamical systems. The Computer Journal, 53(4): 365–369, May 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/4/365;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/4/365.

Nath:2014:MMS

[NK14] Prem Nath and Chiranjeev Kumar. Mobility management scheme for fixed mobility pattern mobile users in IPv4 networks. The Computer Journal, 57(12):1893– 1911, December 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/12/1893.

Nengroo:2019:AAT

[NK19] Ab Shaqoor Nengroo and K. S. Kuppusamy. ‘Advertisements or adversetisements?’ — an accessibility barrier for persons with visual impairments. The Computer Journal, 62(6): 855–868, June 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/855/5126881.

Niyizamwiyitira:2019:UBS

[NL19] Christine Niyizamwiyitira and Lars Lundberg. A utilization-based schedulability test of real-time systems running on a multiprocessor virtual machine. The Computer Journal, 62(6): 884–904, June 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/884/5319152.

Nixon:2011:UPA

[NLDH11] Mark S. Nixon, Xin U.

Liu, Cem Direkoḡlu, and David J. Hurley. On using physical analogies for feature and shape extraction in computer vision. The Computer Journal, 54(1):11–25, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/11.full.
pdf+html.

Nazemi:2019:NCN

[NM19] Alireza Nazemi and Marziyeh Mortezaee. A novel collaborate neural dynamic system model for solving a class of min-umax optimization problems with an application in portfolio management. The Computer Journal, 62(7):1061–1085, July 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/7/1061/5239663.

Nieto:2014:FSH

[NMS14] Juan Manuel González Nieto, Mark Manulis, and Dongdong Sun. Forwardsecure hierarchical predicate encryption. The Computer Journal, 57(4):510– 536, April 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/4/510.full.
                                        pdf+html.

Nakhla:2019:PAD

[NNF19] Zina Nakhla, Kaouther Nouira, and Ahmed Ferchichi. Prescription Adverse Drug Events System (PrescADE) based on ontology and Internet of Things. The Computer Journal, 62(6):801– 805, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/801/5114357.

Nguyen:2014:IRA

[NNN^{+}14] Tien-Dung Nguyen, An Thuy Nguyen, Man Doan Nguyen, Mui Van Nguyen, and Eui-Nam Huh. An improvement of resource allocation for migration process in cloud environment. The Computer Journal, 57(2):308– 318, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/308.full.
pdf+html.

Nepal:2016:GEC

[NP16] Surya Nepal and Suraj Pandey. Guest editorial: Cloud Computing and Scientific Applications (CCSA)

— big data analysis in the cloud. The Computer Journal, 59(3):285– 286, March 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/3/285.

Napoli:2016:CDG

[NPTZ16] Christian Napoli, Giuseppe Pappalardo, Emiliano Tramontana, and Gaetano Zappalà. A cloud-distributed GPU architecture for pattern identification in segmented detectors big-data surveys. The Computer Journal, 59(3):338–352, March 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/3/338.

Niu:2015:CRS

[NRZQ15] Danmei Niu, Lanlan Rui, Cheng Zhong, and Xuesong Qiu. A composition and recovery strategy for mobile social network service in disaster. The Computer Journal, 58(4):700– 708, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/700.

Neelima:2016:PHF

[NS16] Arambam Neelima and Kh Manglem Singh. Perceptual hash function based on scale-invariant feature transform and singular value decomposition. The Computer Journal, 59(9): 1275–1281, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/59/9/1275.

Natarajan:2015:MAD

[NSA15] V. Natarajan, Shina Sheen, and R. Anitha. Multilevel analysis to detect covert social botnet in multimedia social networks. The Computer Journal, 58(4):679– 687, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/679.

Niksefat:2014:ZPP

[NSMS14] Salman Niksefat, Babak Sadeghiyan, Payman Mohassel, and Saeed Sadeghian. ZIDS: a privacy-preserving intrusion detection system using secure two-party computation protocols. The Computer Journal, 57(4): 494–509, April 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/4/494.full.
pdf+html.

Nafarieh:2015:SBT

[NSRP15] Alireza Nafarieh, Shyamala Sivakumar, William Robertson, and William Phillips. SLA-based time-aware provisioning mechanisms in shared mesh protected optical networks. The Computer Journal, 58(8):1717– 1731, August 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1717.

Naderan-Tahan:2016:WDD

[NTSA16] Mahmood Naderan-Tahan and Hamid Sarbazi-Azad. Why does data prefetching not work for modern workloads? The Computer Journal, 59(2):244–259, February 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/2/244.

Nurmi:2007:PDS

[Nur07] Jari Nurmi, editor. Processor Design: System-on-Chip Computing for ASICs and FPGAs. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5529-3 (hardcover), 1-4020-5530-7 (ebook). xix + 525 pp. LCCN TK7895.E42.P763 2007.

Narumi:2011:FCE

[NYT^{+}11] Tetsu Narumi, Kenji Yasuoka, Makoto Taiji, Francesco Zerbetto, and Siegfried Höfinger. Fast calculation of electrostatic potentials on the GPU or the ASIC MD-GRAPE-3. The Computer Journal, 54(7):1181– 1187, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1181.
full.pdf+html.

Novak:2014:PSI

[NZ14] David Novak and Pavel Zezula. Performance study of independent anchor spaces for similarity searching. The Computer Journal, 57(11):1741–1755, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1741.

Ozen:2018:UNT

[OB18] Yunus Ozen and Cuneyt Bayilmis. urgMAC: a new traffic and QoS-aware crosslayer MAC protocol for wireless multimedia sensor networks. The Computer Journal, 61(10):1460–1467, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/10/1460/4781604.

Orosa:2016:AAS

[OBA16] Lois Orosa, Javier D.

Bruguera, and Elisardo Antelo. Asymmetric allocation in a shared flexible signature module for multicore processors. The Computer Journal, 59(10):1453–1469, October 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/10/1453.

Ouqour:2014:PRO

[OJSO14] A. Ouqour, Y. Jabrane, B. Ait Es Said, and A. Ait Ouahman. PAPR reduction in OFDM via active constellation extension–projection onto convex sets combined with particle swarm optimization. The Computer Journal, 57(8):1230–1237, August 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/8/1230.

Ozbal:2011:CBC

[ÖKA11] Gözde Özbal, Hılal Karaman, and Ferda N. Alpaslan. A content-boosted collaborative filtering approach for movie recommendation based on local and global similarity and missing data prediction. The Computer Journal, 54(9):1535–1546, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/9/1535.
                                        full.pdf+html.

Ozcelik:2017:EEI

[OKA17] Ihsan Mert Ozcelik, Ibrahim Korpeoglu, and Ashok Agrawala. Energy efficient IP-connectivity with IEEE 802.11 for home M2M networks. The Computer Journal, 60(6):883–897, June 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/883/3058781.

Oliveira:2012:STA

[OKG^{+}12] Leonardo B. Oliveira, Aman Kansal, Conrado P. L. Gouvêa, Diego F. Aranha, Julio López, Bodhi Priyantha, Michel Goraczko, and Feng Zhao. Secure-TWS: Authenticating node to multi-user communication in shared sensor networks. The Computer Journal, 55(4):384–396, April 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/4/384.full.
pdf+html.

Ozer:2016:PLT

[OKT^{+}16] Mert Ozer, Ilkcan Keles, Hakki Toroslu, Pinar Karagoz, and Hasan Davulcu. Predicting the location and time of mobile phone users by using sequential pattern mining techniques. The Computer Journal, 59(6): 908–922, June 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/6/908.

Oliveira:2017:ASP

[OLF^{+}17] Lucas Bueno Ruas Oliveira, Elena Leroux, Katia Romero Felizardo, Flavio Oquendo, and Elisa Yumi Nakagawa. ArchSORS: a software process for designing software architectures of serviceoriented robotic systems. The Computer Journal, 60 (9):1363–1381, September 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/9/1363/2996411.

Ou:2015:SPA

[OLL15] Chia Shin Ou, Chin Lung Lu, and R. C. T. Lee. A systematical and parallel approach to solve problems involving special properties of bit-vectors. The Computer Journal, 58(5):1112– 1121, May 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/5/1112.

Oh:2012:MTS

[OR12] Doohwan Oh and Won W. Ro. Multi-threading and suffix grouping on massive multiple pattern matching algorithm. The Computer Journal, 55(11):1331–1346, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1331.
full.pdf+html.

Ortin:2011:TIO

[Ort11] Francisco Ortin. Type inference to optimize a hybrid statically and dynamically typed language. The Computer Journal, 54(11): 1901–1924, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Osborn:2016:SSR

[OS16] Emma Osborn and Andrew Simpson. On safety and security requirements in emerging ubiquitous computing models. The Computer Journal, 59(4):570– 591, April 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/570.

Osborn:2018:RSS

[OS18] Emma Osborn and Andrew Simpson. Risk and the small-scale cyber security decision making dialogue — a UK case study. The Computer Journal, 61(4):472– 495, April 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/4/472/4259798.

Ortega:2014:FEL

[OVGG14] Gloria Ortega, Francisco Vázquez, Inmaculada García, and Ester M. Garzón. Fast-SpMM: an efficient library for sparse matrix matrix product on GPUs. The Computer Journal, 57(7): 968–979, July 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/7/968.full.
                                        pdf+html.

Parmar:2015:MRV

[PA15] Minaz Parmar and Marios C. Angelides. MAC-REALM: a video content feature extraction and modelling framework. The Computer Journal, 58(9):2135– 2171, September 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/9/2135.

Padron:2010:HRM

[PABD10] Emilio J. Padrón, Margarita Amor, Montserrat Bóo, and Ramón Doallo. Hierarchical radiosity for multiresolution systems based on normal tests. The Computer Journal, 53(6):741– 752, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/741;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/741.

Pavlidis:2010:PBS

[PANH10] Nicos G. Pavlidis, Niall M.

Adams, David Nicholson, and David J. Hand. Prospects for bandit solutions in sensor management. The Computer Journal, 53(9):1370–1383, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1370.
full.pdf+html.

Parhami:2015:DAN

[Par15] Behrooz Parhami. Digital arithmetic in nature: Continuous-digit RNS. The Computer Journal, 58(5): 1214–1223, May 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/5/1214.

Pandey:2012:SWA

[PB12] Suraj Pandey and Rajkumar Buyya. Scheduling workflow applications based on multi-source parallel data retrieval in distributed computing networks. The Computer Journal, 55(11):1288–1308, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1288.
full.pdf+html.

Piso:2014:OAE

[PB14] D. Piso and J. D. Bruguera. Obtaining accurate error expressions and bounds for floating-point multiplicative algorithms. The Computer Journal, 57(2):319– 331, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/319.full.
pdf+html.

Perks:2013:TAM

[PBH^{+}13] O. F. J. Perks, D. A. Beckingsale, S. D. Hammond, I. Miller, J. A. Herdman, A. Vadgama, A. H. Bhalerao, L. He, and S. A. Jarvis. Towards automated memory model generation via event tracing. The Computer Journal, 56(2):156– 174, February 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/2/156.full.
pdf+html.

Park:2014:HPD

[PBL14] Joon-Sang Park, Seung Jun Baek, and Kyogu Lee. A highly parallelized decoder for random network coding leveraging GPGPU. The Computer Journal, 57(2):233–240, February 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/2/233.full.
pdf+html.

Paz:2012:CEG

[PC12] Azaria Paz and Jack W.

Carlyle. Chaotic evolution via generalized probabilistic automata (probabilistic arrays). The Computer Journal, 55(5):522– 533, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/5/522.full.
pdf+html.

Piraghaj:2016:VMC

[PCC^{+}16] Sareh Fotuhi Piraghaj, Rodrigo N. Calheiros, Jeffrey Chan, Amir Vahid Dastjerdi, and Rajkumar Buyya. Virtual machine customization and task mapping architecture for efficient allocation of cloud data center resources. The Computer Journal, 59(2):208– 224, February 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/2/208.

Petric:2012:ODC

[PCLU12] Ingrid Petric, Bojan Cestnik, Nada Lavrac, and Tanja Urbancic. Outlier detection in cross-context link discovery for creative literature mining. The Computer Journal, 55(1):47–61, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/1/47.full.
pdf+html.

Pereira:2015:PKE

[PDNH15] Mayana Pereira, Rafael Dowsley, Anderson C. A. Nascimento, and Goichiro Hanaoka. Public-key encryption schemes with bounded CCA security and optimal ciphertext length based on the CDH and HDH assumptions. The Computer Journal, 58(10):2738–2746, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2738.

Pekergin:2012:ISI

[Pek12] Nihal Pekergin. Introduction to the special issue on probability models in performance analysis. The Computer Journal, 55(5): 512–514, May 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/5/512.full.
pdf+html.

Penrose:2010:DRB

[Pen10] Mathew D. Penrose. Discussant of response to the Computer Journal Lecture by François Baccelli. The Computer Journal, 53(5):610– 611, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/reprint/53/5/610.

Petrou:2011:FTR

[PG11] Maria Petrou and Archontis Giannakidis. Full tomographic reconstruction of 2D vector fields using discrete integral data. The Computer Journal, 54(9):1491– 1504, September 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/9/1491.
full.pdf+html.

Polyvyanyy:2014:MSA

[PGBFW14] Artem Polyvyanyy, Luciano García-Bañuelos, Dirk Fahland, and Mathias Weske. Maximal structuring of acyclic process models. The Computer Journal, 57(1):12–35, January 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/12.full.
pdf+html.

Prochazka:2015:CSF

[PH15] Petr Procházka and Jan Holub. Compression of a set of files with natural language content. The Computer Journal, 58(5):1169– 1185, May 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1169.

Pallister:2015:ICO

[PHB15] James Pallister, Simon J. Hollis, and Jeremy Bennett. Identifying compiler options to minimize energy consumption for embedded platforms. The Computer Journal, 58(1):95–109, January 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/95.

Pennycook:2012:AWA

[PHM^{+}12] S. J. Pennycook, S. D.

Hammond, G. R. Mudalige, S. A. Wright, and S. A. Jarvis. On the acceleration of wavefront applications using distributed many-core architectures. The Computer Journal, 55(2):138– 153, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/2/138.full.
pdf+html.

Park:2011:DCM

[PiLCH11] Heewan Park, Hyun il Lim, Seokwoo Choi, and Taisook Han. Detecting common modules in Java packages based on static object trace birthmark. The Computer Journal, 54(1):108– 124, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/108.full.
pdf+html.

Patel:2018:EEA

[PK18] Nileshkumar R. Patel and Shishir Kumar. Energyefficient approach for effective estimation of delimited node position with limited references. The Computer Journal, 61(6):881– 895, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/6/881/4608877.

Prakash:2018:OCT

[PKM18] Chandra Prakash, Rajesh Kumar, and Namita Mittal. Optimized clustering techniques for gait profiling in children with cerebral palsy for rehabilitation. The Computer Journal, 61 (11):1683–1694, November 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/11/1683/4959079.

Perez-Lantero:2016:APC

[PL16] Pablo Pérez-Lantero. Area and perimeter of the convex hull of stochastic points. The Computer Journal, 59(8):1144–1154, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/8/1144.

Phuttharak:2018:LCP

[PL18] Jurairat Phuttharak and Seng W. Loke. LogicCrowd: Crowd-powered logic programming based mobile applications. The Computer Journal, 61(1):32–46, January 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/32/3073684.

Popa:2011:SCG

[Pop11] Ionut. Popa. SE-compression:

a generalization of dictionarybased compression. The Computer Journal, 54(11): 1876–1881, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Pougajendy:2017:DSB

[PP17] Jayashree Pougajendy and Arun Raj Kumar Parthiban. Detection of SIP-based denial of service attack using dual cost formulation of support vector machine. The Computer Journal, 60 (12):1770–1784, December 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1770/3861968.

Park:2011:FCM

[PR11] Eun-Chan Park and Minjoong Rim. Fair coexistence MAC protocol for contention-based heterogeneous networks. The Computer Journal, 54(8):1382– 1397, August 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

                                        org/content/54/8/1382.
                                        full.pdf+html.

Picking:2010:CSU

[PRG^{+}10] Rich Picking, Alexia Robinet, Vic Grout, John McGinn, Armando Roy, Simon Ellis, and Denise Oram. A case study using a methodological approach to developing user interfaces for elderly and disabled people. The Computer Journal, 53(6):842– 859, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/6/842;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/6/842.

Pinto:2011:SEM

[PRJS11] David Pinto, Paolo Rosso, and Héctor Jiménez-Salazar. A self-enriching methodology for clustering narrow domain short texts. The Computer Journal, 54(7): 1148–1165, July 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/7/1148.
                                        full.pdf+html.

Petrenko:2015:GDM

[PS15] Alexandre Petrenko and Adenilso Simao. Generalizing the DS-methods for testing non-deterministic FSMs. The Computer Journal, 58(7):1656–1672, July 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/7/1656.

Perez:2017:UCP

[PS17] Alejandro Pérez and Pablo Sánchez. On the use of C# partial classes for the implementation of software product lines. The Computer Journal, 60(1):86–109, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Perazzo:2015:DRL

[PSD15] Pericle Perazzo, Pavel Skvortsov, and Gianluca Dini. On designing resilient location-privacy obfuscators. The Computer Journal, 58(10):2649–2664, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2649.

Park:2014:EED

[PSP14] Yong Tae Park, Pranesh Sthapit, and Jae-Young Pyun. Energy efficient data fragmentation for ubiquitous computing. The Computer Journal, 57(2):263– 272, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/2/263.full.
                                        pdf+html.

Paterno:2010:AIS

[PSS10] Fabio Paternò, Carmen Santoro, and Antonio Scorcia. Ambient intelligence for supporting task continuity across multiple devices and implementation languages. The Computer Journal, 53(8):1210–1228, October 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/8/1210.
                                        full.pdf+html.

Pacharawongsakda:2013:MLC

[PT13] Eakasit Pacharawongsakda and Thanaruk Theeramunkong. Multi-label classification using dependent and independent dual space reduction. The Computer Journal, 56(9):1113–1135, September 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/9/1113.
                                        full.pdf+html.

Pinna:2018:PNM

[PTOM18] Andrea Pinna, Roberto Tonelli, Matteo Orrú, and Michele Marchesi. A Petri nets model for blockchain analysis. The Computer Journal, 61(9):1374–1388, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1374/4824744.

Petrou:2010:NCI

[PTP10] Maria Petrou, Marco Elio Tabacchi, and Roberta Piroddi. Networks of concepts and ideas. The Computer Journal, 53(10):1738– 1751, December 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/10/1738.
full.pdf+html.

Poulain:2014:PPA

[PTWB14] David Poulain, Joanna Tomasik, Marc-Antoine Weisser, and Dominique Barth. A packing problem approach to lightpath assignment in an optical ring. The Computer Journal, 57(8):1155–1166, August 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/8/1155.

Pao:2012:MSS

[PW12] Derek Pao and Xing Wang. Multi-stride string searching for high-speed content inspection. The Computer Journal, 55(10):1216–1231, October 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/10/1216.
                                        full.pdf+html.

Palade:2019:RDS

[PW19] Vasile Palade and J. Gerard Wolff. A roadmap for the development of the ‘SP Machine’ for artificial intelligence. The Computer Journal, 62(11):1584–1604, November 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/11/1584/5298771.

Peng:2013:AQA

[PWY^{+}13] Bingyue Peng, Junjie Wu, Hua Yuan, Qingwei Guo, and Dacheng Tao. ANEEC: a quasi-automatic system for massive named entity extraction and categorization. The Computer Journal, 56(11):1328–1346, November 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.

org/content/56/11/1328.
full.pdf+html.

Peng:2017:SMA

[PXG^{+}17] Zhiping Peng, Bo Xu, Antonio Marcel Gates, Delong Cui, and Weiwei Lin. A study of a multi-agent organizational framework with virtual machine clusters as the unit of granularity in cloud computing. The Computer Journal, 60(7):1032– 1043, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/1032/2608048.

Pyle:2019:SLR

[Pyl19] Ian C. Pyle. Software for the Linesman radar data processing system. The Computer Journal, 62(6):806– 819, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/806/5114354.

Pei:2015:SWT

[PYM^{+}15] Qingqi Pei, Dingyu Yan, Lichuan Ma, Zi Li, and Yang Liao. A strong and weak ties feedback-based trust model in multimedia social networks. The Computer Journal, 58(4):627– 643, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/627.

Phuong:2018:CBE

[PYS18] Tran Viet Xuan Phuong, Guomin Yang, and Willy Susilo. Criteria-based encryption. The Computer Journal, 61(4):512– 525, April 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/4/512/4430306.

Pandey:2018:QRD

[PZ18] Saurabh K. Pandey and Mukesh A. Zaveri. Quasi random deployment and localization in layered framework for the Internet of Things. The Computer Journal, 61(2):159–179, February 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/2/159/3788610.

Pandey:2019:DBE

[PZ19] Saurabh K Pandey and Mukesh A Zaveri. DoAbased event localization using uniform concentric circular array in the IoT environment. The Computer Journal, 62(10):1403–1425, October 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/10/1403/5423662.

Podpecan:2012:OES

[PZL12] Vid Podpecan, Monika Zemenova, and Nada Lavrac. Orange4Ws environment for service-oriented data mining. The Computer Journal, 55(1):82–98, January 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/1/82.full.
pdf+html.

Patsakis:2015:PSM

[PZPS15] Constantinos Patsakis, Athanasios Zigomitros, Achilleas Papageorgiou, and Agusti Solanas. Privacy and security for multimedia content shared on OSNs: Issues and countermeasures. The Computer Journal, 58(4):518– 535, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/518.

Peng:2017:FEE

[PZZ^{+}17] Jia Peng, Yanmin Zhu, Qingwen Zhao, Hongzi Zhu, Jian Cao, Guangtao Xue, and Bo Li. Fair energyefficient sensing task allocation in participatory sensing with Smartphones. The Computer Journal, 60(6): 850–865, June 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/850/3051822.

Queiroz:2019:WBF

[QF19] Jordan S. Queiroz and Eduardo L. Feitosa. A Web browser fingerprinting method based on the Web audio API. The Computer Journal, 62(8):1106– 1120, August 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1106/5298776.

Qi:2018:ECP

[QLZ18] Yi Qi, Huan Li, and Zhongzhi Zhang. Extended corona product as an exactly tractable model for weighted heterogeneous networks. The Computer Journal, 61(5):745–760, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/5/745/4602863.

Quiroga:2017:STF

[QO17] Jose Quiroga and Francisco Ortin. SSA transformations to facilitate type inference in dynamically typed code. The Computer Journal, 60 (9):1300–1315, September 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/9/1300/2770526.

Quirchmayr:2015:DCS

[QS15] Thomas Quirchmayr and Mark Strembeck. A discussion of communication schemes for process execution histories to enforce entailment constraints in process-driven SOAs. The Computer Journal, 58(10):2255–2279, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2255.

Qi:2019:TSP

[QYZ19] Yi Qi, Yuhao Yi, and Zhongzhi Zhang. Topological and spectral properties of small-world hierarchical graphs. The Computer Journal, 62(5):769– 784, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/769/5105858.

Qin:2015:LII

[QZXR15] Guangjun Qin, Mingfa Zhu, Limin Xiao, and Li Ruan. Lessen interflow interference using virtual channels partitioning. The Computer Journal, 58(8):1826– 1841, August 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1826.

Qin:2018:BRO

[QZZ18] Baodong Qin, Qinglan Zhao, and Dong Zheng. Bounded revocable and outsourceable ABE for secure data sharing. The Computer Journal, 61(8):1259–1268, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1259/5045945.

Rodriguez-Aguilar:2014:SIO

[RA14] Juan A. Rodriguez-Aguilar. Special issue on optimization in multi-agent systems: Guest Editor’s introduction. The Computer Journal, 57(6):797– 798, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/6/797.full.
                                        pdf+html.

Rezvani:2015:RAC

[RAJ15] Mostafa Rezvani, Mohammad Kazem Akbari, and Bahman Javadi. Resource allocation in cloud computing environments based on integer linear programming. The Computer Journal, 58(2):300–314, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/300.

RahimiZadeh:2017:WAP

[RAKJ17] Keyvan RahimiZadeh, Morteza AnaLoui, Peyman Kabiri, and Bahman Javadi. Workloadaware placement of multitier applications in virtualized datacenters. The Computer Journal, 60(2): 60–??, February 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/2/210/2725497/Workload-
Aware-Placement-of-Multi-
Tier.

Rastegar:2017:LAS

[RASM17] Seyed Hamed Rastegar, Aliazam Abbasfar, and Vahid Shah-Mansouri. Latencyaware sum-rate maximization for 5G softwaredefined radio access networks. The Computer Journal, 60(10):1482–1497, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1482/3769277.

Rincon:2013:NCP

[RATB^{+}13] D. Rincón, A. Agustí-Torra, J. F. Botero, F. Raspall, D. Remondo, X. Hesselbach, M. T. Beck, H. de Meer, F. Niedermeier, and G. Giuliani. A novel collaboration paradigm for reducing energy consumption and carbon dioxide emissions in data centres. The Computer Journal, 56(12):1518–1536, December 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/12/1518.
                                        full.pdf+html.

Radke:2017:SPP

[RB17] Kenneth Radke and Colin Boyd. Security proofs for protocols involving humans. The Computer Journal, 60(4):527–540, March 23, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/4/527/2354604.

Radke:2015:CFA

[RBNB15] Kenneth Radke, Colin Boyd, Juan Gonzalez Nieto, and Harry Bartlett. CHURNs: Freshness assurance for humans. The Computer Journal, 58(10):2404– 2425, October 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/10/2404.

Radivojevic:2016:CBP

[RCS16] Zaharije Radivojević, Miloš Cvetanović, and Saša Stojanović. Comparison of binary procedures: a set of techniques for evading compiler transformations. The Computer Journal, 59(1): 106–118, January 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/1/106.

Rai:2018:SIC

[RCTK18] Sunny Rai, Shampa Chakraverty, Devendra K. Tayal, and Yash Kukreti. A study on impact of context on metaphor detection. The Computer Journal, 61(11): 1667–1682, November 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/11/1667/4958214.

Radi:2014:NIL

[RDB^{+}14a] Marjan Radi, Behnam Dezfouli, Kamalrulnizam Abu Bakar, Shukor Abd Razak, and Malrey Lee. Network initialization in low-power wireless networks: a comprehensive study. The Computer Journal, 57(8):1238– 1261, August 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/8/1238.

Ramadan:2014:OFV

[RDB14b] Wassim Ramadan, Eugen Dedu, and Julien Bourgeois. Oscillation-free video adaptation at application layer on server side and experiments using DCCP. The Computer Journal, 57(8):1195– 1210, August 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/8/1195.

Rivera:2012:SSI

[RDMRM12] Mariano Rivera, Oscar Dalmau, Washington Mio, and Alonso Ramirez-Manzanares. Spatial sampling for image segmentation. The Computer Journal, 55(3):313– 324, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/3/313.full.
                                        pdf+html.

Ren:2016:IBE

[RDZ^{+}16] Yanli Ren, Ning Ding, Xinpeng Zhang, Haining Lu, and Dawu Gu. Identitybased encryption with verifiable outsourced revocation. The Computer Journal, 59(11):1659–1668, November 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1659.

Ramchurn:2010:DCR

[RFMJ10] Sarvapali D. Ramchurn, Alessandro Farinelli, Kathryn S. Macarthur, and Nicholas R. Jennings. Decentralized coordination in RoboCup Rescue. The Computer Journal, 53(9):1447–1461, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1447.
full.pdf+html.

Reeder:2014:GNN

[RG14] John Reeder and Michael Georgiopoulos. Generative neural networks for multitask life-long learning. The Computer Journal, 57(3): 427–450, March 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/3/427.full.
pdf+html.

Rextin:2017:DUP

[RH17] Aimal Rextin and Patrick Healy. Dynamic upward planarity testing of single source embedded digraphs. The Computer Journal, 60 (1):45–59, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Rajmic:2015:SPM

[RHF^{+}15] Pavel Rajmic, Jirí Hošek, Michal Fusek, Sergey Andreev, and Július Stecík. Simplified probabilistic modelling and analysis of enhanced distributed coordination access in IEEE 802.11. The Computer Journal, 58(6):1456–1468, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1456.

Rensfelt:2011:REM

[RHG^{+}11] Olof Rensfelt, Frederik Hermans, Per Gunningberg, Lars-Åke Larzon, and Erik Björnemo. Repeatable experiments with mobile nodes in a relocatable WSN testbed. The Computer Journal, 54(12):1973–1986, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/12/1973.
full.pdf+html.

Raza:2012:NNB

[RHH12] Muhammad Raza, Farookh Khadeer Hussain, and Omar Khadeer Hussain. Neural networkbased approach for predicting trust values based on non-uniform input in mobile applications. The Computer Journal, 55(3):347– 378, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/3/347.full.
pdf+html.

Richards:2013:HBE

[Ric13]  Martin Richards.  How

BCPL evolved from CPL. The Computer Journal, 56 (5):664–670, May 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/5/664.full.
pdf+html;   http://www.
cl.cam.ac.uk/users/mr/
BCPL.

Rivera-illingworth:2010:DNN

[RiCH10] F. Rivera-illingworth, V. Callaghan, and H. Hagras. Detection of normal and novel behaviours in ubiquitous domestic environments. The Computer Journal, 53(2):142–151, February 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/2/142;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        content/full/bxm078/DC1;
                                         http://comjnl.oxfordjournals.
                                        org/cgi/reprint/53/2/142.

Riguzzi:2014:SIP

[Rig14] Fabrizio Riguzzi. Speeding up inference for probabilistic logic programs. The Computer Journal, 57(3): 347–363, March 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/3/347.full.
                                        pdf+html.

Renuga:2018:EPP

[RJ18] S. Renuga and S. S. K. Jagatheeshwari. Efficient privacy-preserving data sanitization over cloud using optimal GSA algorithm. The Computer Journal, 61(10):1577–1588, October 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/10/1577/5053261.

Rademacher:2017:SDW

[RJS^{+}17] Michael Rademacher, Karl Jonas, Florian Siebertz, Adam Rzyska, Moritz Schlebusch, and Markus Kessel. Software-defined wireless mesh networking: Current status and challenges. The Computer Journal, 60 (10):1520–1535, October 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1520/3959606.

Romero:2013:TMB

[RJV13] José Raúl Romero, Juan Ignacio Jaén, and Antonio Vallecillo. A tool for the model-based specification of open distributed systems. The Computer Journal, 56 (7):793–818, July 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/7/793.full.
pdf+html.

Roelleke:2015:HAI

[RKBY15] Thomas Roelleke, Andreas Kaltenbrunner, and Ricardo Baeza-Yates. Harmony assumptions in information retrieval and social networks. The Computer Journal, 58 (11):2982–2999, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Rios:2011:ECA

[RL11] Ruben Rios and Javier Lopez. Exploiting contextawareness to enhance sourcelocation privacy in wireless sensor networks. The Computer Journal, 54(10):1603– 1615, October 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/10/1603.
                                        full.pdf+html.

Rosado:2015:SIS

[RLJ15] David G. Rosado, Nadira Lammari, and Jan Jürjens. Special issue on secure information systems engineering. The Computer Journal, 58(10):2193–2194, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2193.

Ramezani:2017:MOL

[RLTZ17] Fahimeh Ramezani, Jie Lu, Javid Taheri, and Albert Y. Zomaya. A multiobjective load balancing system for cloud environments. The Computer Journal, 60(9):1316–1337, September 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/9/1316/2770527.

Rubio-Largo:2015:PMA

[RLVRGÁ15] Álvaro Rubio-Largo, Miguel A. Vega-Rodríguez, and David L. González-Álvarez. A parallel multiobjective approach based on honey bees for traffic grooming in optical networks. The Computer Journal, 58(9):2171–2191, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/2171.

Rokach:2008:DMD

[RM08] Lior Rokach and Oded Z.

Maimon. Data mining with decision trees: theory and applications, volume 69 of Series in machine perception and artificial intelligence. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2008. ISBN 981-277-171-9. xviii + 244 pp. LCCN QA76.9.D343 R654 2008.

Rezaeibagha:2019:EMC

[RM19] Fatemeh Rezaeibagha and Yi Mu. Efficient micropayment of cryptocurrency from blockchains. The Computer Journal, 62(4):507– 517, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/507/5155318.

Rosas:2011:CBC

[RMB11] Erika Rosas, Olivier Marin, and Xavier Bonnaire. Corps:

Building a community of reputable peers in distributed hash tables. The Computer Journal, 54(10): 1721–1735, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/54/10/1721.
                                        full.pdf+html.

Rodriguez:2015:MSC

[RMB15] Ricardo J. Rodríguez, José Merseguer, and Simona Bernardi. Modelling security of critical infrastructures: a survivability assessment. The Computer Journal, 58(10):2313–2327, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2313.

Rebollo:2015:ISG

[RMFM15] Oscar Rebollo, Daniel Mellado, and Eduardo Fernandez-Medina. ISGcloud: a security governance framework for cloud computing. The Computer Journal, 58(10):2233–2254, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2233.

Rodriguez:2011:API

[RMGT11] Gabriel Rodríguez, María J.

Martín, Patricia González, and Juan Touriño. Analysis of performance-impacting factors on checkpointing frameworks: the CPPC case study. The Computer Journal, 54(11):1821–1837, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Rahaman:2010:STB

[RMP10] H. Rahaman, J. Mathew, and D. K. Pradhan. Secure testable S-box architecture for cryptographic hardware implementation. The Computer Journal, 53(5):581– 591, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/581;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/5/581.

Rajasingh:2016:TBN

[RMP^{+}16] Indra Rajasingh, Paul Manuel, N. Parthiban, D. Azubha Jemilet, and R. Sundara Rajan. Transmission in butterfly networks. The Computer Journal, 59(8):1174–1179, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/8/1174.

Rajan:2015:LBD

[RMR^{+}15a] R. Sundara Rajan, Paul Manuel, Indra Rajasingh, N. Parthiban, and Mirka Miller. A lower bound for dilation of an embedding. The Computer Journal, 58(12):3271–3278, December 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/12/3271.

Ruiz:2015:ISS

[RMR15b] Jose Fran. Ruiz, Antonio Maña, and Carsten Rudolph. An integrated security and systems engineering process and modelling framework. The Computer Journal, 58(10):2328–2350, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2328.

Rocchi:2012:WIB

[Roc12] Paolo Rocchi. What is information: Beyond the jungle of information theories. The Computer Journal, 55 (7):856–860, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.

org/content/55/7/856.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Rogers:2011:ATS

[Rog11] A. Rogers. Agent technologies for sensor networks. The Computer Journal, 54(3):307–308, March 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/307.full.
pdf+html.

Rosenbloom:2012:CC

[Ros12a] Paul S. Rosenbloom. Computing and computation. The Computer Journal, 55 (7):820–824, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/820.full.

pdf+html. Special Focus on the Centenary of Alan Turing. See erratum [Ros12b].

Rosenbloom:2012:ECC

[Ros12b] Paul S. Rosenbloom. Erratum: Computing and computation. The Computer Journal, 55(10):1266, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

org/content/55/10/1266.
full.pdf+html. See [Ros12a].

Rosenberg:2014:RMF

[Ros14] Arnold L. Rosenberg. Region management by finitestate robots. The Computer Journal, 57(1):59–72, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/59.full.
pdf+html.

Reddy:2016:DND

[RR16] K. Hemant Kumar Reddy and Diptendu Sinha Roy. DPPACS: a novel data partitioning and placement aware computation scheduling scheme for dataintensive cloud applications. The Computer Journal, 59(1):64–82, January 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/1/64.

Rodriguez-Rodriguez:2015:WAR

[RRCC^{+}15] R. Rodríguez-Rodríguez, F. Castro, D. Chaver, R. Gonzalez-Alberquilla, L. Piñuel, and F. Tirado. Write-aware replacement policies for PCM-based systems. The Computer Journal, 58(9):2000–2025, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/2000.

Rodriguez-Rodriguez:2018:RDI

[RRDC^{+}18] R. Rodríguez-Rodríguez, J. Díaz, F. Castro, P. Ibáñez, D. Chaver, V. Viñals, J. C. Saez, M. Prieto-Matias, L. Piñuel, T. Monreal, and J. M. Llabería. Reuse detector: Improving the management of STT-RAM SLLCs. The Computer Journal, 61(6):856– 880, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/856/4568418.

Rastegari:2019:ECS

[RSD19] Parvin Rastegari, Willy Susilo, and Mohammad Dakhlalian. Efficient certificateless signcryption in the standard model: Revisiting Luo and Wan’s scheme from wireless personal communications (2018). The Computer Journal, 62(8):1178– 1193, August 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1178/5485598.

Reniers:2014:REB

[RSW14] Michel A. Reniers, Rob Schoren, and Tim A. C.

Willemse. Results on embeddings between statebased and event-based systems. The Computer Journal, 57(1):73–92, January 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/73.full.
pdf+html.

Ruijters:2012:GPA

[RT12] Daniel Ruijters and Philippe Thévenaz. GPU prefilter for accurate cubic B-spline interpolation. The Computer Journal, 55(1):15–20, January 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/1/15.full.
pdf+html.

Ramos:2013:DSJ

[RTE^{+}13] Sabela Ramos, Guillermo L. Taboada, Roberto R. Expósito, Juan Touriño, and Ramón Doallo. Design of scalable Java communication middleware for multi-core systems. The Computer Journal, 56(2):214–228, February 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/2/214.full.
pdf+html.

Soysal:2011:JUA

[SA11] Medeni Soysal and A. Aydin Alatan. Joint utilization of appearance, geometry and chance for scene logo retrieval. The Computer Journal, 54(7):1221– 1231, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1221.
full.pdf+html.

Sabah:2011:NLU

[Sab11] Gérard Sabah. Natural language understanding, where are we going? where could we go? The Computer Journal, 54(9):1505–1513, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1505.
full.pdf+html.

Sakellari:2010:CPN

[Sak10]  Georgia Sakellari.  The

cognitive packet network: a survey. The Computer Journal, 53(3):268– 279, March 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/3/268;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/3/268.

Shao:2016:LDC

[SAK16] Zhendong Shao, Igor Averbakh, and Sandi Klavžar. Labeling dot-Cartesian and dot-lexicographic product graphs with a condition at distance two. The Computer Journal, 59(1):151– 158, January 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/1/151.

Sarbazi-Azad:2011:TPG

[SAKOK11] H. Sarbazi-Azad, A. Khonsari, and M. Ould-Khaoua. On the topological properties of grid-based interconnection networks: Surface area and volume of radial spheres. The Computer Journal, 54(5):726– 737, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/5/726.full.
                                        pdf+html.

Sampaio:2019:CAS

[SAPS19] Sandra Sampaio, Mashael Aljubairah, Hapsoro Adi Permana, and Pedro Sampaio. A conceptual approach for supporting traffic data wrangling tasks. The Computer Journal, 62(3):461– 480, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/3/461/5158248.

Serhani:2012:SFB

[SBBB12] Mohamed Adel Serhani, Abdelghani Benharref, Elarbi Badidi, and Salah Bouktif. Scalable federated broker management for selection of Web services. The Computer Journal, 55(12):1420– 1439, December 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/12/1420.
full.pdf+html.

Shankar:2019:FBV

[SBV19] N. Shankar, S. Sathish Babu, and C. Viswanathan. Femur bone volumetric estimation for osteoporosis classification using optimizationbased deep belief network in X-ray images. The Computer Journal, 62(11):1656– 1670, November 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/11/1656/5380592.

Shih:2010:SCP

[SC10] Yu-Ying Shih and Daniel Chao. Sequence of control in S^{3}PMR. The Computer Journal, 53(10):1691–1703, December 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/10/1691.
                                        full.pdf+html.

Sevinc:2011:EGA

[SC11] Ender Sevinç and Ahmet Coşar. An evolutionary genetic algorithm for optimization of distributed database queries. The Computer Journal, 54(5):717– 725, May 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/5/717.full.
                                        pdf+html.

Sepehri:2015:PPQ

[SCD15] Maryam Sepehri, Stelvio Cimato, and Ernesto Damiani. Privacy-preserving query processing by multiparty computation. The Computer Journal, 58(10): 2195–2212, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/58/10/2195.

Sarker:2018:ITS

[SCKH18] Iqbal H. Sarker, Alan Colman, Muhammad Ashad Kabir, and Jun Han. Individualized time-series segmentation for mining mobile phone user behavior. The Computer Journal, 61(3):349–368, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/3/349/4107199.

Shao:2018:TPQ

[SCT18a] Zhou Shao, Muhammad Aamir Cheema, and David Taniar. Trip planning queries in indoor venues. The Computer Journal, 61(3):409– 426, March 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/3/409/4656254.

Sohail:2018:SAS

[SCT18b] Ammar Sohail, Muhammad Aamir Cheema, and David Taniar. Social-aware spatial top-k and skyline queries. The Computer Journal, 61(11):1620–1638, November 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/11/1620/4931622.

Saadatfar:2015:ISE

[SDN15] Hamid Saadatfar, Hossein Deldari, and Mahmoud Naghibzadeh. Improving the scheduler’s energy saving capability by noting both job and resource characteristics. The Computer Journal, 58(6):1482–1493, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1482.

Sadeq:2013:DDT

[SDW13] Muhammad Jafar Sadeq, Matt Duckham, and Michael F. Worboys. Decentralized detection of topological events in evolving spatial regions. The Computer Journal, 56(12):1417–1431, December 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/12/1417.
                                        full.pdf+html.

Seradji:2017:BGW

[SF17] Shabnam Seradji and Mehran S. Fallah. A Bayesian game of whitewashing in reputation systems. The Computer Journal, 60(8):1223–1237, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/8/1223/3064965.

Sun:2013:ITE

[SGG^{+}13] Yi Sun, Yang Guo, Yuming Ge, Shan Lu, Jihua Zhou, and Eryk Dutkiewicz. Improving the transmission efficiency by considering noncooperation in ad hoc networks. The Computer Journal, 56(8):1034–1042, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/8/1034.
full.pdf+html.

Sun:2015:FSW

[SGH15] Shi-Feng Sun, Dawu Gu, and Zhengan Huang. Fully secure wicked identity-based encryption against key leakage attacks. The Computer Journal, 58(10):2520–2536, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2520.

Seda:2010:GDF

[SH10] Anthony Karel Seda and Pascal Hitzler. Generalized distance functions in the theory of computation. The Computer Journal, 53 (4):443–464, May 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/4/443;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/4/443.

Song:2015:ADT

[SH15] WeiTao Song and Bin Hu.

Approach to detecting typeflaw attacks based on extended strand spaces. The Computer Journal, 58(4): 572–587, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/4/572.

Simko:2015:FVA

[SHH^{+}15] Viliam Simko, David Hauzar, Petr Hnetynka, Tomas Bures, and Frantisek Plasil. Formal verification of annotated textual use-cases. The Computer Journal, 58(7): 1495–1529, July 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/7/1495.

Shih:2008:DWS

[Shi08] Frank Y. Shih. Digital Watermarking and Steganography: Fundamentals and Techniques. Taylor and Francis, Boca Raton, FL, USA, 2008. ISBN 1-4200-4757-4. 180 pp. LCCN QA76.9.A25 S467

                                        2008. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy0745/2007034224-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0725/2007034224.
                                        html.

Su:2015:EEE

[SHL^{+}15] Sen Su, Qingjia Huang, Jian Li, Xiang Cheng, Peng Xu, and Kai Shuang. Enhanced energy-efficient scheduling for parallel tasks using partial optimal slacking. The Computer Journal, 58(2):246–257, February 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/2/246.

Stedmon:2011:MER

[SHR^{+}11] Alex W. Stedmon, Benjamin Hasseldine, David Rice, Mark Young, Steve Markham, Michael Hancox, Edward Brickell, and Joanna Noble. ‘MotorcycleSim’: an evaluation of rider interaction with an innovative motorcycle simulator. The Computer Journal, 54(7):1010–1025, July 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1010.
full.pdf+html.

Sharifi:2014:STM

[SIK14] Beaux P. Sharifi, David I.

Inouye, and Jugal K. Kalita. Summarization of Twitter microblogs. The Computer Journal, 57(3):378– 402, March 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/3/378.full.
                                        pdf+html.

Sinanoglu:2012:FMI

[Sin12] Ozgur Sinanoglu. Fault model independent, maximal compaction of test responses in the presence of unknown response bits. The Computer Journal, 55(12):1525–1537, December 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/12/1525.
                                        full.pdf+html.

Saginbekov:2014:TES

[SJ14] Sain Saginbekov and Arshad Jhumka. Towards efficient stabilizing code dissemination in wireless sensor networks. The Computer Journal, 57(12):1790–1816, December 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/12/1790.

Shah:2018:PPV

[SJ18a] Kaushal A. Shah and Devesh C. Jinwala. Privacy preserving, verifiable and resilient data aggregation in grid-based networks. The Computer Journal, 61(4): 614–628, April 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/614/4855686.

Stephan:2018:PSO

[SJ18b] Thompson Stephan and K. Suresh Joseph. Particle swarm optimization-based energy efficient channel assignment technique for clustered cognitive radio sensor networks. The Computer Journal, 61(6):926– 936, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/926/4796922.

Saifan:2017:OCS

[SJA17] Ramzi Saifan, Iyad Jafar, and Ghazi Al Sukkar. Optimized cooperative spectrum sensing algorithms in cognitive radio networks. The Computer Journal, 60(6): 835–849, June 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/835/3051821.

Sun:2012:HAR

[SJS12] Yan Sun, Qiangfeng Jiang, and Mukesh Singhal. A hillarea-restricted geographic routing protocol for mobile ad hoc and sensor networks. The Computer Journal, 55(8):932–949, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/932.full.
                                        pdf+html.

Sonowal:2018:SAS

[SK18a] Gunikhan Sonowal and K. S. Kuppusamy. SmiDCA: an anti-smishing model with machine learning approach. The Computer Journal, 61(8):1143–1157, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1143/4985552.

Srivastava:2018:CBI

[SK18b] Prashant Srivastava and Ashish Khare. Contentbased image retrieval using local binary curvelet co-occurrence pattern — a multiresolution technique. The Computer Journal, 61(3):369–385, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/3/369/4191409.

Son:2012:FSB

[SKK^{+}12] Hyejin Son, Taeyoon Kang, Hwangnam Kim, Jong-Bae Park, and Jae Hyung Roh. A fair and secure bandwidth allocation for Ami Mesh Network in Smart Grid. The Computer Journal, 55(10):1232–1243, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1232.
full.pdf+html.

Subramanian:2018:GTB

[SKK18] Chitra M. Subramanian, Aneesh Krishna, and Arshinder Kaur. Game theory-based requirements analysis in the i^{∗} framework. The Computer Journal, 61(3):427–446, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/3/427/4671076.

Sievi-Korte:2015:TGS

[SKKM15] Outi Sievi-Korte, Kai Koskimies, and Erkki Mäkinen. Techniques for genetic software architecture design. The Computer Journal, 58(11): 3141–3170, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Sumesh:2019:GTB

[SKS19] Sreenithya Sumesh, Aneesh Krishna, and Chitra M. Subramanian. Game theorybased reasoning of opposing non-functional requirements using inter-actor dependencies. The Computer Journal, 62(11):1557–1583, November 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/11/1557/5288329.

Sloman:2010:EPB

[SL10a] Morris Sloman and Emil Lupu. Engineering policybased ubiquitous systems. The Computer Journal, 53(7):1113–1127, September 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/7/1113;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/7/1113.

Smyrnakis:2010:DOM

[SL10b] Michalis Smyrnakis and David S. Leslie. Dynamic opponent modelling in fictitious play. The Computer Journal, 53(9):1344–1359, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/9/1344.
                                        full.pdf+html.

Sallam:2014:MOV

[SL14] Ahmed Sallam and Kenli Li. A multi-objective virtual machine migration policy in cloud systems. The Computer Journal, 57(2):195– 204, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/195.full.
pdf+html.

Song:2015:NPA

[SLL15] Yinglei Song, Chunmei Liu, and Yongzhong Li. A new parameterized algorithm for predicting the secondary structure of RNA sequences including pseudoknots. The Computer Journal, 58(11): 3114–3125, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Smith:2011:MIF

[SLP11] Connie U. Smith, Catalina M.

Lladó, and Ramon Puigjaner. Model interchange format specifications for experiments, output and results. The Computer Journal, 54(5):674–690, May 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/5/674.full.
pdf+html.

Sensoy:2011:RDA

[ŞLV^{+}11] Murat Şensoy, Thao Le, Wamberto W. Vasconcelos, Timothy J. Norman, and Alun D. Preece. Resource determination and allocation in sensor networks: a hybrid approach. The Computer Journal, 54(3):356– 372, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/356.full.
pdf+html.

Song:2017:ABF

[SLW^{+}17] Tao Song, Yuchen Liu, Yiding Wang, Ruhui Ma, Alei Liang, Zhengwei Qi, and Haibing Guan. Ashman: a bandwidth fragmentationbased dynamic flow scheduling for data center networks. The Computer Journal, 60(10):1498–1509, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/10/1498/3867578.

Sun:2016:RSP

[SLY^{+}16] Shi-Feng Sun, Joseph K. Liu, Yu Yu, Baodong Qin, and Dawu Gu. RKA-secure public key encryptions against efficiently invertible functions. The Computer Journal, 59(11):1637–1658, November 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1637.

Sun:2014:RTS

[SLZ14] Linjia Sun, Xiaohui Liang, and Qinping Zhao. Recursive templates segmentation and exemplars matching for human parsing. The Computer Journal, 57(3):364– 377, March 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/3/364.full.
pdf+html.

Salles:2012:SMR

[SM12] Ronaldo M. Salles and Donato A. Marino, Jr. Strategies and metric for resilience in computer networks. The Computer Journal, 55(6): 728–739, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/6/728.full.
pdf+html.

Sivagami:2016:CBM

[SM16] L. Sivagami and J. Martin Leo Manickam. Clusterbased MAC protocol for collision avoidance and TDMA scheduling in underwater wireless sensor networks. The Computer Journal, 59(10):1527–1535, October 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/10/1527.

Sun:2014:LCC

[SMLM14] Jianshan Sun, Jian Ma, Zhiying Liu, and Yajun Miao. Leveraging content and connections for scientific article recommendation in social computing contexts. The Computer Journal, 57(9):1331–1342, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/9/1331.

Scriney:2019:ADM

[SMM^{+}19] Michael Scriney, Suzanne McCarthy, Andrew McCarren, Paolo Cappellari, and Mark Roantree. Automating data mart construction from semi-structured data sources. The Computer Journal, 62(3):394– 413, March 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/3/394/5038272.

Sycara:2010:ASP

[SNG^{+}10] Katia Sycara, Timothy J. Norman, Joseph A. Giampapa, Martin J. Kollingbaum, Chris Burnett, Daniele Masato, Mairi McCallum, and Michael H. Strub. Agent support for policydriven collaborative mission planning. The Computer Journal, 53(5):528– 540, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/528;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/5/528.

Simao:2010:FCD

[SP10] Adenilso Simão and Alexandre Petrenko. Fault coverage-driven incremental test generation. The Computer Journal, 53(9):1508– 1522, November 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/53/9/1508.
full.pdf+html.

Shiaeles:2015:FII

[SP15] Stavros N. Shiaeles and Maria Papadaki. FHSD: an improved IP spoof detection method for Web DDoS attacks. The Computer Journal, 58(4):892– 903, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/892.

Saez:2018:IBT

[SPdGPM18] J. C. Saez, A. Pousa, A. E. de Giusti, and M. Prieto-Matias. On the interplay between throughput, fairness and energy efficiency on asymmetric multicore processors. The Computer Journal, 61(1):74–94, January 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/1/74/3769278.

Suksomboon:2011:PNQ

[SPJA11] Kalika Suksomboon, Panita Pongpaibool, Yusheng Ji, and Chaodit Aswakul. PC-Nash: QoS provisioning framework with pathclassification scheme under Nash equilibrium. The Computer Journal, 54(6): 931–943, June 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/6/931.full.
                                        pdf+html.

Saez:2017:PDP

[SPRR^{+}17] J. C. Saez, A. Pousa, R. Rodriíguez-Rodriíguez, F. Castro, and M. Prieto-Matias. PMCTrack: Delivering performance monitoring counter support to the OS scheduler. The Computer Journal, 60(1):60–85, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Sarma:2018:TAS

[SPS^{+}18] Himangshu Sarma, Robert Porzel, Jan D. Smeddinck, Rainer Malaka, and Arun Baran Samaddar. A text to animation system for physical exercises. The Computer Journal, 61 (11):1589–1604, November 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/11/1589/4883866.

Sajeev:2010:MEO

[SR10] A. S. M. Sajeev and Sakgasit Ramingwong. Mum effect as an offshore outsourcing risk: a study of differences in perceptions. The Computer Journal, 53(1): 120–126, January 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/120;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/120.

Skandhakumar:2012:AFU

[SRD^{+}12] Nimalaprakasan Skandhakumar, Jason Reid, Ed Dawson, Robin Drogemuller, and Farzad Salim. An authorization framework using building information models. The Computer Journal, 55(10):1244–1264, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1244.
full.pdf+html.

Sadri:2010:SIA

[SS10a] Fariba Sadri and Kostas Stathis. Special issue on artificial societies for ambient intelligence editorial introduction. The Computer Journal, 53(8):1136–1137, October 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1136.
full.pdf+html.

Sukthankar:2010:ATD

[SS10b] Gita Sukthankar and Katia Sycara. Analyzing team decision-making in tactical scenarios. The Computer Journal, 53(5):503– 512, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/503;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/5/503.

Seo:2012:ACI

[SSK12] Hyungjung Seo, Jaewon Seo, and Taewhan Kim. Algorithms for combined interand intra-task dynamic voltage scaling. The Computer Journal, 55(11):1367–1382, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1367.
full.pdf+html.

Subbarayalu:2019:HNI

[SSK19] Venkatraman Subbarayalu, B Surendiran, and P Arun Raj Kumar. Hybrid network intrusion detection system for smart environments based on Internet of Things. The Computer Journal, 62(12):1822–1839, December 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/12/1822/5618851.

Spanakis:2012:EWK

[SSS12a] Gerasimos Spanakis, Georgios Siolas, and Andreas Stafylopatis. Exploiting Wikipedia knowledge for conceptual hierarchical clustering of documents. The Computer Journal, 55(3): 299–312, March 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/3/299.full.
                                        pdf+html.

Starka:2012:ACS

[SSS^{+}12b] Jakub Stárka, Martin Svoboda, Jan Sochna, Jirí Schejbal, Irena Mlýnková, and David Bednárek. Analyzer: a complex system for data analysis. The Computer Journal, 55(5):590– 615, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/5/590.full.
                                        pdf+html.

Saikkonen:2016:CSM

[SSS16] Riku Saikkonen and Eljas Soisalon-Soininen. Cachesensitive memory layout for dynamic binary trees. The Computer Journal, 59(5): 630–649, May 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/5/630.

Sheng:2015:ACA

[SSY15] Quan Z. Sheng, Elhadi M. Shakshuki, and Jian Yu. Ambient and context-aware services for the future Web. The Computer Journal, 58(8):1687–1688, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/8/1687.

Shen:2016:RMM

[ST16] Wuqiang Shen and Shaohua Tang. RGB, a mixed multivariate signature scheme. The Computer Journal, 59(4):439– 451, April 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/439.

Shoaran:2017:EZK

[ST17] Maryam Shoaran and Alex Thomo. Editorial: Zeroknowledge-private counting of group triangles in social networks. The Computer Journal, 60(1):126– 134, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Stakhov:2018:MCS

[Sta18] Alexey Stakhov. Missioncritical systems, paradox of Hamming code, row hammer effect, ‘Trojan Horse’ of the binary system and numeral systems with irrational bases. The Computer Journal, 61(7):1038– 1063, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/1038/4430323.

Szczerbak:2014:KNS

[STBB14] Michał K. Szczerbak, François Toutain, Ahmed Bouabdallah, and Jean-Marie Bonnin. KRAMER: New social medium based on collaborative recognition of important situations. The Computer Journal, 57(9):1296– 1317, September 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/9/1296.

Sandes:2018:FBP

[STW^{+}18] Edans F. O. Sandes, George L. M. Teodoro, Maria Emilia M. T. Walter, Xavier Martorell, Eduard Ayguade, and Alba C. M. A. Melo. Formalization of block pruning: Reducing the number of cells computed in exact biological sequence comparison algorithms. The Computer Journal, 61(5):687– 713, May 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/5/687/4539903.

Sarathambekai:2018:MOO

[SU18] S. Sarathambekai and K. Umamaheswari. Multi-objective optimization techniques for task scheduling problem in distributed systems. The Computer Journal, 61(2):248–263, February 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/2/248/3867581.

Suzuki:2013:SID

[Suz13]  Einoshin Suzuki.  Spe-

cial issue on discovery science: Guest Editor’s introduction. The Computer Journal, 56(3):271– 273, March 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/3/271.full.
pdf+html.

Sabri:2015:TRG

[SV15] Ahmad Sabri and Vincent Vajnovszki. Two reflected Gray code-based orders on some restricted growth sequences. The Computer Journal, 58(5):1099–1111, May 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/5/1099.

Steiner:2015:CMI

[SVG^{+}15] Thomas Steiner, Ruben Verborgh, Joaquim Gabarro, Erik Mannens, and Rik Van de Walle. Clustering media items stemming from multiple social networks. The Computer Journal, 58(9):1861–1875, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/1861.

Serral:2013:CAC

[SVP13] Estefanía Serral, Pedro Valderas, and Vicente Pelechano. Context-adaptive coordination of pervasive services by interpreting models during runtime. The Computer Journal, 56(1):87–114, January 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/1/87.full.
                                        pdf+html.

Stoilos:2015:FEO

[SVS15] Giorgos Stoilos, Tassos Venetis, and Giorgos Stamou. A fuzzy extension to the OWL 2 RL ontology language. The Computer Journal, 58(11):2956–2971, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Subramani:2014:CIZ

[SW14] K. Subramani and James Worthington. On certifying instances of zero-clairvoyant scheduling. The Computer Journal, 57(1):129– 137, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/129.full.
pdf+html.

Swade:2011:IUE

[Swa11] Doron Swade. Inventing the user: EDSAC in context. The Computer Journal, 54(1):143–147, January 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/1/143.full.
pdf+html.

Sun:2013:ISP

[SWG13] Wei Sun, Tao Wen, and Quan Guo. Improving the start-up performance of the TFRC protocol. The Computer Journal, 56(11):1269– 1278, November 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/11/1269.
full.pdf+html.

Sun:2012:SPR

[SWLZ12] Xiaoxun Sun, Hua Wang, Jiuyong Li, and Yanchun Zhang. Satisfying privacy requirements before data anonymization. The Computer Journal, 55(4):422– 437, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/4/422.full.
                                        pdf+html.

Shiah:2013:CCD

[SY13] Chwan-Yi Shiah and Yun-Sheng Yen. Compression of Chinese document images by complex shape matching. The Computer Journal, 56(11):1292–1304, November 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/11/1292.
                                        full.pdf+html.

Savas:2015:GMA

[SY15] Erkay Savaş and Cemal Yılmaz. A generic method for the analysis of a class of cache attacks: a case study for AES. The Computer Journal, 58(10):2716–2737, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2716.

Suenaga:2011:GMM

[SYH11] Shunichiro Suenaga, Nobukazu Yoshioka, and Shinichi Honiden. Group migration by mobile agents in wireless sensor networks. The Computer Journal, 54(3):345– 355, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/345.full.
pdf+html.

Su:2015:DTA

[SZB15] Xing Su, Minjie Zhang, and Quan Bai. Dynamic task allocation for heterogeneous agents in disaster environments under time, space and communication constraints. The Computer Journal, 58(8):1776–1791, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/8/1776.

Su:2019:TMP

[SZB19] Xing Su, Minjie Zhang, and Quan Bai. Two mathematical programming-based approaches for wireless mobile robot deployment in disaster environments. The Computer Journal, 62(6):905– 918, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/905/5363889.

Sun:2015:ECF

[SZL15] Chang-Ai Sun, Yimeng Zai, and Huai Liu. Evaluating and comparing fault-based testing strategies for general Boolean specifications: a series of experiments. The Computer Journal, 58(5): 1199–1213, May 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/5/1199.

Song:2016:CDB

[SZL16] Sulin Song, Shuming Zhou, and Xiaoyan Li. Conditional diagnosability of burnt pancake networks under the PMC model. The Computer Journal, 59(1):91–105, January 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/1/91.

Shen:2014:ERC

[SZS14] Limin Shen, Futai Zhang, and Yinxia Sun. Efficient revocable certificateless encryption secure in the standard model. The Computer Journal, 57(4):592– 601, April 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/4/592.full.
pdf+html.

Shen:2018:EMB

[SZW^{+}18] Yanping Shen, Kangfeng Zheng, Chunhua Wu, Mingwu Zhang, Xinxin Niu, and Yixian Yang. An ensemble method based on selection using bat algorithm for intrusion detection. The Computer Journal, 61(4):526– 538, April 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/526/4582946.

Thenmozhi:2016:ACB

[TA16a] D. Thenmozhi and Chandrabose Aravindan. An automatic and clause-based approach to learn relations for ontologies. The Computer Journal, 59(6):889– 907, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/6/889.

Thenmozhi:2016:PIU

[TA16b] D. Thenmozhi and Chandrabose Aravindan. Paraphrase identification by using clause-based similarity features and machine translation metrics. The Computer Journal, 59(9):1289– 1302, September 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/9/1289.

Toumi:2018:FTM

[TAC^{+}18] Khalifa Toumi, Mohamed Aouadi, Ana R. Cavalli, Wissam Mallouli, Jordi Puiggal Allepuz, and Pol Valletb Montfort. A framework for testing and monitoring security policies: Application to an electronic voting system. The Computer Journal, 61(8):1109–1122, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1109/4925398.

Taherkhani:2011:UDT

[Tah11] Ahmad Taherkhani. Using decision tree classifiers in source code analysis to recognize algorithms: an experiment with sorting algorithms. The Computer Journal, 54(11):1845–1860, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Tamba:2018:HRN

[Tam18] Tua A. Tamba. On Handelman’s representation of network utility maximization. The Computer Journal, 61(6):798–807, June 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/798/4160674.

Tan:2011:LEC

[Tan11]  Zuowen Tan.  An off-

line electronic cash scheme based on proxy blind signature. The Computer Journal, 54(4):505–512, April 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/4/505.full.
pdf+html.

Tang:2015:ETE

[Tan15] Qiang Tang. From ephemerizer to timed-ephemerizer: Achieve assured lifecycle enforcement for sensitive data. The Computer Journal, 58(4):1003–1020, April 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/1003.

Tate:2010:SPT

[TB10] Jonathan Tate and Iain Bate. Sensornet protocol tuning using principled engineering methods. The Computer Journal, 53(7):991– 1019, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.

                                        org/cgi/content/abstract/
                                        53/7/991;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/7/991.

Turgut:2011:HAT

[TB11] Damla Turgut and Ladislau Bölöni. Heuristic approaches for transmission scheduling in sensor networks with multiple mobile sinks. The Computer Journal, 54(3):332– 344, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/3/332.full.
                                        pdf+html.

Taghavi:2018:NIT

[TBBH18] Mona Taghavi, Jamal Bentahar, Kaveh Bakhtiyari, and Chihab Hanachi. New insights towards developing recommender systems. The Computer Journal, 61(3):319–348, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/3/319/3893562.

Tang:2015:CER

[TCL15] Ying-Kai Tang, Sherman S. M. Chow, and Joseph K. Liu. Comments on ‘Efficient Revocable Certificateless Encryption Secure in the Standard Model’. The Computer Journal, 58(4): 779–781, April 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/4/779.

Thomas:2012:IDS

[TD12] Manoj Thomas and Gurpreet Dhillon. Interpreting deep structures of information systems security. The Computer Journal, 55(10):1148–1156, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1148.
full.pdf+html.

Tsakalozos:2016:DUA

[TEP^{+}16] Konstantinos Tsakalozos, Spiros Evangelatos, Fotis Psallidas, Marcos R. Vieira, Vassilis J. Tsotras, and Alex Delis. DiVA: Using application-specific policies to ‘dive’ into vector approximations. The Computer Journal, 59(9):1363–1382, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/9/1363.

Thimbleby:2011:SD

[Thi11] Harold Thimbleby. Signposting in documents. The Computer Journal, 54(7): 1119–1135, July 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/7/1119.
                                        full.pdf+html.

Tian:2011:SHK

[THP^{+}11] Biming Tian, Song Han, Sazia Parvin, Jiankun Hu, and Sajal Das. Selfhealing key distribution schemes for wireless networks: a survey. The Computer Journal, 54(4):549– 569, April 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/4/549.full.
                                        pdf+html.

Tian:2012:ESH

[THP^{+}12] Biming Tian, Song Han, Sazia Parvin, Jiankun Hu, and Sajal Das. Erratum: Self-healing key distribution schemes for wireless networks: a survey. The Computer Journal, 55(4): 438, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/4/438.full.
                                        pdf+html.

Tang:2018:PIH

[THY^{+}18] Zhenjun Tang, Ziqing Huang, Heng Yao, Xianquan Zhang, Lv Chen, and Chunqiang Yu. Perceptual image hashing with weighted DWT features for reduced-reference image quality assessment. The Computer Journal, 61 (11):1695–1709, November 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/11/1695/4993055.

Timotheou:2010:RNN

[Tim10] Stelios Timotheou. The random neural network: a survey. The Computer Journal, 53(3):251– 267, March 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/3/251;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/3/251.

Timotheou:2011:ATA

[Tim11] Stelios Timotheou. Assettask assignment algorithms in the presence of execution uncertainty. The Computer Journal, 54(9):1514–1525, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1514.
full.pdf+html.

Tian:2012:TOE

[TJZF12] Hui Tian, Hong Jiang, Ke Zhou, and Dan Feng. Transparency-orientated encoding strategies for Voiceover-IP steganography. The Computer Journal, 55(6): 702–716, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/6/702.full.
                                        pdf+html.

Takahashi:2015:ROC

[TK15] Takeshi Takahashi and Youki Kadobayashi. Reference ontology for cybersecurity operational information. The Computer Journal, 58(10):2297–2312, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2297.

Turkay:2011:IIT

[TKB11] Cagatay Turkay, Emre Koc, and Selim Balcisoy. Integrating information theory in agent-based crowd simulation behavior models. The Computer Journal, 54(11): 1810–1820, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Tweneboah-Koduah:2018:SRA

[TKB18] Samuel Tweneboah-Koduah and William J. Buchanan. Security risk assessment of critical infrastructure systems: a comparative study. The Computer Journal, 61 (9):1389–1406, September 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1389/4833881.

Taherkhani:2011:RAU

[TKM11] Ahmad Taherkhani, Ari Korhonen, and Lauri Malmi. Recognizing algorithms using language constructs, software metrics and roles of variables: an experiment with sorting algorithms. The Computer Journal, 54(7):1049–1066, July 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1049.
full.pdf+html.

Thayammal:2019:UBO

[TL19] Manju C. Thayammal and M. Mary Linda. Utilitybased optimal resource allocation in LTE-A networks by hybrid ACO-TS with MFA scheme. The Computer Journal, 62(6):931– 942, June 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/6/931/5377388.

Taherkordi:2011:GCB

[TLRE11] Amirhosein Taherkordi, Frédéric Loiret, Romain Rouvoy, and Frank Eliassen. A generic component-based approach for programming, composing and tuning sensor software. The Computer Journal, 54(8):1248–1266, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/8/1248.
                                        full.pdf+html.

Tang:2015:ECP

[TMC15] Qiang Tang, Hua Ma, and Xiaofeng Chen. Extend the concept of public key encryption with delegated search. The Computer Journal, 58(4):724– 734, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/724.

Tynan:2011:CIP

[TMOO11] Richard Tynan, Conor Muldoon, Gregory O’Hare, and Michael O’Grady. Coordinated intelligent power management and the heterogeneous sensing coverage problem. The Computer Journal, 54(3):490– 502, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/3/490.full.
pdf+html.

Tsai:2014:GTT

[TNWT14] Jason Tsai, Thanh H.

Nguyen, Nicholas Weller, and Milind Tambe. Gametheoretic target selection in contagion-based domains. The Computer Journal, 57(6):893–905, June 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/6/893.full.
pdf+html.

Tang:2015:GBR

[TPG^{+}15] Xiaolan Tang, Juhua Pu, Yang Gao, Yu Xie, and Zhang Xiong. GPS-based replica deletion scheme with anti-packet distribution for vehicular networks. The Computer Journal, 58(6): 1399–1415, June 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/6/1399.

Tripathi:2018:EEV

[TPV18] Atul Tripathi, Isha Pathak, and Deo Prakash Vidyarthi. Energy efficient VM placement for effective resource utilization using modified binary PSO. The Computer Journal, 61(6):832– 846, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/832/4643566.

Traub:2012:WRC

[Tra12] Joseph Traub. What is the right computational model for continuous scientific problems? The Computer Journal, 55(7):836– 837, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/7/836.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Trcek:2010:SMF

[Trc10] Denis Trcek. Security metrics foundations for computer security. The Computer Journal, 53(7):1106– 1112, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/1106;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1106.

Trujillo-Rasua:2016:CMA

[TRY16] Rolando Trujillo-Rasua and Ismael G. Yero. Characterizing 1-metric antidimensional trees and unicyclic graphs. The Computer Journal, 59(8):1264– 1273, August 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1264.

Thayammal:2017:EPM

[TS17] S. Thayammal and D. Selvathi. Edge preserved multispectral image compression using extended shearlet transform. The Computer Journal, 60(7):986– 994, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/986/2608041.

Taherizadeh:2019:DML

[TS19] Salman Taherizadeh and Vlado Stankovski. Dynamic multi-level autoscaling rules for containerized applications. The Computer Journal, 62(2):174– 197, February 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/174/4993728.

Tsalapati:2017:QRU

[TSC^{+}17] Eleni Tsalapati, Giorgos Stoilos, Alexandros Chortaras, Giorgos Stamou, and George Koletsos. Query rewriting under ontology change. The Computer Journal, 60(3):60–??, March 2017. CODEN CMPJA6. ISSN ???? URL https://

academic.oup.com/comjnl/
article/60/3/389/2609373/
Query-Rewriting-Under-
Ontology-Change.

Tillem:2017:NMC

[TSK17] Gamze Tillem, Erkay Savaş, and Kamer Kaya. A new method for computational private information retrieval. The Computer Journal, 60(8):1238–1250, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/8/1238/3077150.

Taniar:2011:SNR

[TST^{+}11] David Taniar, Maytham Safar, Quoc Thai Tran, Wenny Rahayu, and Jong Hyuk Park. Spatial network RNN queries in GIS. The Computer Journal, 54(4):617– 627, April 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/4/617.full.
pdf+html.

Tseng:2012:ERI

[TT12] Yuh-Min Tseng and Tung-Tso Tsai. Efficient revocable ID-based encryption with a public channel. The Computer Journal, 55(4):475– 486, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/4/475.full.
pdf+html.

Tseng:2015:LFI

[TTH15] Yuh-Min Tseng, Tung-Tso Tsai, and Sen-Shan Huang. Leakage-free IDbased signature. The Computer Journal, 58(4):750– 757, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/750.

Tatli:2017:WBC

[TU17] Emi˙

n İslam Tatli and Bedi˙

rhan Urgun. WIVET — benchmarking coverage qualities of Web crawlers. The Computer Journal, 60(4):555–572, March 23, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/555/2608071.

Tupakula:2012:DSB

[TV12] Udaya Kiran Tupakula and Vijay Varadharajan. Dynamic state-based security architecture for detecting security attacks in virtual machines. The Computer Journal, 55(4):397– 409, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/4/397.full.
pdf+html.

Tupakula:2015:TES

[TV15] Udaya Tupakula and Vijay Varadharajan. Trust enhanced security for tenant transactions in the cloud environment. The Computer Journal, 58(10):2388–2403, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2388.

Tian:2019:NWR

[TXJ^{+}19] Yu Tian, Ming Xu, Chong Jiang, Ji-Bo Wang, and Xiao-Yuan Wang. No-wait resource allocation flowshop scheduling with learning effect under limited cost availability. The Computer Journal, 62(1):90–96, January 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/90/4959080.

Tong:2014:ANP

[TY14] Sheau-Ru Tong and Yuan-Tse Yu. Adaptation of nearperfect packet-level FEC interleaving in mobile media streaming. The Computer Journal, 57(11):1711–1722, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1711.

Tong:2018:SAD

[TYL^{+}18] Chao Tong, Xiang Yin, Jun Li, Tongyu Zhu, Renli Lv, Liang Sun, and Joel J. P. C. Rodrigues. A shilling attack detector based on convolutional neural network for collaborative recommender system in social aware network. The Computer Journal, 61(7):949– 958, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/949/4835634.

Thomas:2011:MVA

[TZ11] Nigel Thomas and Yishi Zhao. Mean value analysis for a class of Pepa models. The Computer Journal, 54 (5):643–652, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/5/643.full.
                                        pdf+html.

Ullah:2018:SUB

[UKW^{+}18] Noor Ullah, Xiangjie Kong, Liangtian Wan, Honglong Chen, Zhibo Wang, and Feng Xia. A social utilitybased dissemination scheme for emergency warning messages in vehicular social networks. The Computer Journal, 61(7):971–986, July 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/7/971/4931623.

Ulieru:2011:BRW

[Uli11] Mihaela Ulieru. Book review: Wired for Innovation: How Information Technology is Reshaping the Economy. The Computer Journal, 54(2):304–305, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/2/304.full.

pdf+html. See [BS10b].

Usman:2019:IVS

[URHK19] Muhammad Usman, Aris Cahyadi Risdianto, Jungsu Han, and Jongwon Kim. Interactive visualization of SDNenabled multisite cloud playgrounds leveraging SmartX MultiView visibility framework. The Computer Journal, 62(6):838–854, June 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/6/838/5123533.

Veni:2016:MET

[VB16] T. Veni and S. Mary Saira Bhanu. MDedup++: Exploiting temporal and spatial page-sharing behaviors for memory deduplication enhancement. The Computer Journal, 59(3):353– 370, March 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/3/353.

Valmorbida:2016:ULI

[VBBR16] Willian Valmorbida, Jorge Luis Victória Barbosa, Débora Nice Ferrari Barbosa, and Sandro José Rigo. U-Library: an intelligent model for ubiquitous library support. The Computer Journal, 59(9):1330–1344, September 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/9/1330.

Valova:2010:NPA

[VBMH10] Iren Valova, Derek Beaton, Dan MacLean, and John Hammond. NIPSOM: Parallel architecture and implementation of a growing SOM. The Computer Journal, 53(6):753– 771, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/6/753;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/6/753.

Varghese:2014:CBI

[VBVP14] Abraham Varghese, Kannan Balakrishnan, Reji Rajan Varghese, and Joseph S. Paul. Content-based image retrieval of axial brain slices using a novel LBP with a ternary encoding. The Computer Journal, 57(9):1383– 1394, September 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/9/1383.

vanderAalst:2010:MCA

[vdALM^{+}10] Wil M. P. van der Aalst, Niels Lohmann, Peter Massuthe, Christian Stahl, and Karsten Wolf. Multiparty contracts: Agreeing and implementing interorganizational processes. The Computer Journal, 53(1):90–106, January 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/90;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/90.

vanDiggelen:2010:ESI

[vDBvEW10] Jurriaan van Diggelen, Robbert-Jan Beun, Rogier M. van Eijk, and Peter J. Werkhoven. Efficient semantic information exchange for ambient intelligence. The Computer Journal, 53(8):1138–1151, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1138.
full.pdf+html.

vandenHove:2015:ORP

[vdH15] Gauthier van den Hove. On the origin of recursive procedures. The Computer Journal, 58(11):2892–2899, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Velan:2010:MBS

[Vel10] Kumaara Velan. Modelling bidders in sequential automated auctions. The Computer Journal, 53(2):208– 218, February 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/2/208;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/2/208.

Vergnaud:2017:CAB

[Ver17] Damien Vergnaud. Comment on ‘Attribute-Based Signatures for Supporting Anonymous Certification’ by N. Kaaniche and M. Laurent (ESORICS 2016). The Computer Journal, 60 (12):1801–1808, December 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/12/1801/3861971.

Vernize:2015:MNI

[VGA15] Grazielle Vernize, André Luiz Pires Guedes, and Luiz Carlos Pessoa Albini. Malicious nodes identification for complex network based on local views. The Computer Journal, 58(10):2476–2491, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2476.

Vijayarajan:2019:BKB

[VGA19] R. Vijayarajan, P. Gnanasivam, and R. Avudaiammal.

Bio-key based AES for personalized image cryptography. The Computer Journal, 62(11):1695–1705, November 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/11/1695/5436925.

Vazquez:2011:MIS

[VGF11] Francisco Vázquez, Ester M.

Garzón, and José Jesús Fernández. Matrix implementation of Simultaneous Iterative Reconstruction Technique (Sirt) on GPUs. The Computer Journal, 54 (11):1861–1868, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Vu:2015:NAN

[VKC15] Hai L. Vu, Kenneth K.

Khaw, and Tsong Yueh Chen. A new approach for network vulnerability analysis. The Computer Journal, 58(4):878– 891, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/878.

Verma:2010:UQM

[VKZ^{+}10] Dinesh Verma, Bong Jun Ko, Petros Zerfos, Kang won Lee, Ting He, Matthew Duggan, Kristian Stewart, Ananthram Swami, and Nikoletta Sofra. Understanding the quality of monitoring for network management. The Computer Journal, 53(5):541– 550, June 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/5/541;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/5/541.

Vavpetic:2013:SSD

[VL13] Anže Vavpetič and Nada Lavrač. Semantic subgroup discovery systems and workflows in the SDMtoolkit. The Computer Journal, 56(3):304–320, March 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/3/304.full.
pdf+html.

VallejosC:2014:FTM

[VM14] Reinaldo A. Vallejos C. and José M. Martínez V. A fast transformation of Markov chains and their respective steady-state probability distributions. The Computer Journal, 57(1):1–11, January 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/1/1.full.
pdf+html.

Vinyals:2014:MPA

[VMF^{+}14] Meritxell Vinyals, Kathryn S.

Macarthur, Alessandro Farinelli, Sarvapali D. Ramchurn, and Nicholas R. Jennings. A message-passing approach to decentralized parallel machine scheduling. The Computer Journal, 57(6):856– 874, June 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/6/856.full.
pdf+html.

Veloudis:2016:NPH

[VN16] Simeon Veloudis and Nimal Nissanke. A novel permission hierarchy for RBAC for dealing with SoD in MAC models. The Computer Journal, 59(4):462– 492, April 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/462.

Vigueras:2016:UGA

[VO16] Guillermo Vigueras and Juan M. Orduña. On the use of GPU for accelerating communication-aware mapping techniques. The Computer Journal, 59(6):836– 847, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/6/836.

Vinyals:2011:SSN

[VRAC11] Meritxell Vinyals, Juan A. Rodriguez-Aguilar, and Jesus Cerquides. A survey on sensor networks from a multiagent perspective. The Computer Journal, 54(3): 455–470, March 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/3/455.full.
                                        pdf+html.

Vandierendonck:2010:AMS

[VRD10] Hans Vandierendonck, Sean Rul, and Koen De Bosschere. Accelerating multiple sequence alignment with the Cell BE Processor. The Computer Journal, 53(6): 814–826, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/6/814;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/6/814.

Verbeek:2017:DCT

[VvdAMG17] H. M. W. Verbeek, W. M. P. van der Aalst, and J. Munoz-Gama. Divide and conquer: a tool framework for supporting decomposed discovery in process mining. The Computer Journal, 60 (11):1649–1674, November 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/11/1649/3804254.

Vespa:2011:MDM

[VWR11] Lucas Vespa, Ning Weng, and Ramaswamy Ramaswamy. MS-DFA: Multiple-stride pattern matching for scalable deep packet inspection. The Computer Journal, 54(2):285–303, February 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/2/285.full.
pdf+html.

Wakrime:2017:SBP

[Wak17] Abderrahim Ait Wakrime.

Satisfiability-based privacyaware cloud computing. The Computer Journal, 60 (12):1760–1769, December 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1760/3778288.

Wang:2014:IIA

[Wan14] Huaqun Wang. Insecurity of ‘Improved Anonymous Multi-Receiver Identity-Based Encryption’. The Computer Journal, 57(4): 636–638, April 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/4/636.full.

pdf+html. See [Chi12].

Wang:2016:SDM

[WB16] You Wang and Jun Bi. Software-defined mobility support in IP networks. The Computer Journal, 59(2):159–177, February 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/2/159.

Wu:2015:CPM

[WBS15] Xiuchao Wu, Kenneth N. Brown, and Cormac J. Sreenan. Contact probing mechanisms for opportunistic sensor data collection. The Computer Journal, 58(8):1792–1810, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1792.

Wu:2013:RUA

[WCCL13] Ro-Yu Wu, Jou-Ming Chang, An-Hang Chen, and Chun-Liang Liu. Ranking and unranking t-ary trees in a Gray-code order. The Computer Journal, 56(11):1388– 1395, November 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/11/1388.
full.pdf+html.

Wang:2017:EIS

[WCCL17] Ming-Hung Wang, Alex Chuan-Hsien Chang, Kuan-Ta Chen, and Chin-Laung Lei. Estimating ideological scores of Facebook pages: an empirical study in Taiwan. The Computer Journal, 60(11):1675–1686, November 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/11/1675/3836918.

Wang:2019:IFT

[WCD19] Gaoli Wang, Zhenfu Cao, and Xiaolei Dong. Improved fault-tolerant aggregate signatures. The Computer Journal, 62(4):481– 489, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/481/5139676.

Weng:2010:BBP

[WCKH10] Chuan-Chi Weng, Ching-Wen Chen, Chang-Jung Ku, and Shiow-Fen Hwang. A bandwidth-based poweraware routing protocol with low route discovery overhead in mobile ad hoc networks. The Computer Journal, 53(7):969–990, September 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/969;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/969.

Wei:2011:SRS

[WCL^{+}11] Hongxing Wei, Youdong Chen, Miao Liu, Yingpeng Cai, and Tianmiao Wang. Swarm robots: from self-assembly to locomotion. The Computer Journal, 54(9):1465–1474, September 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/9/1465.
full.pdf+html.

Wang:2015:EMR

[WCL15] Shenlu Wang, Muhammad Aamir Cheema, and Xuemin Lin. Efficiently monitoring reverse k-nearest neighbors in spatial networks. The Computer Journal, 58(1):40–56, January 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/1/40.

Wu:2010:LGN

[WCW10] Ro-Yu Wu, Jou-Ming Chang, and Yue-Li Wang. Loopless generation of non-regular trees with a prescribed branching sequence. The Computer Journal, 53(6): 661–666, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/661;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/661.

Wu:2014:DGC

[WCW^{+}14] Zhiang Wu, Jie Cao, Junjie Wu, Youquan Wang, and Chunyang Liu. Detecting genuine communities from large-scale social networks: a pattern-based method. The Computer Journal, 57(9):1343–1357, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1343.

Wu:2018:ARP

[WCW^{+}18] Jigang Wu, Long Chen, Xu Wang, Guiyuan Jiang, Siew kei Lam, and Thambipillai Srikanthan. Algorithms for replica placement and update in tree network. The Computer Journal, 61(2):273–287, February 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/2/273/4055637.

Wu:2017:HMA

[WCXZ17] Ying Wu, Jinyong Chang, Rui Xue, and Rui Zhang. Homomorphic MAC from algebraic one-way functions for network coding with small key size. The Computer Journal, 60 (12):1785–1800, December 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1785/3885827.

Wang:2018:SEA

[WDCL18] Haijiang Wang, Xiaolei Dong, Zhenfu Cao, and Dongmei Li. Secure and efficient attribute-based encryption with keyword search. The Computer Journal, 61(8):1133–1142, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/8/1133/4975828.

Weidlich:2012:BEC

[WDW12] Matthias Weidlich, Remco Dijkman, and Mathias Weske. Behaviour equivalence and compatibility of business process models with complex correspondences. The Computer Journal, 55(11):1398–1418, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/11/1398.
full.pdf+html.

Wang:2015:HPD

[WEFJ15] Xi Wang, Alejandro Erickson, Jianxi Fan, and Xiaohua Jia. Hamiltonian properties of DCell networks. The Computer Journal, 58 (11):2944–2955, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wegner:2012:EC

[Weg12] Peter Wegner. The evolution of computation. The Computer Journal, 55(7): 811–813, July 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/7/811.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Wetherfield:2010:PRP

[Wet10] Michael Wetherfield. Personal recollections of programming Deuce in the late 1950s. The Computer Journal, 53(9):1535–1549, November 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/9/1535.
full.pdf+html.

Wu:2010:MRS

[WF10] Jian Wu and Steve Furber. A multicast routing scheme for a universal spiking neural network architecture. The Computer Journal, 53(3):280–288, March 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/3/280;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/3/280.

Wang:2018:SMC

[WGL^{+}18] Ji-Bo Wang, Xin-Na Geng, Lu Liu, Jian-Jun Wang, and Yuan-Yuan Lu. Single machine CON/SLK due date assignment scheduling with controllable processing time and job-dependent learning effects. The Computer Journal, 61(9):1329–1337, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1329/4781602.

Weiler:2017:ESE

[WGS17] Andreas Weiler, Michael Grossniklaus, and Marc H. Scholl. Editorial: Survey and experimental analysis of event detection techniques for Twitter. The Computer Journal, 60(3): 60–??, March 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/3/329/2608058/Editorial-
Survey-and-Experimental-
Analysis-of.

Wei:2014:NEA

[WGZW14] Qingting Wei, Jihong Guan, Shuigeng Zhou, and Xin Wang. A new and effective approach to GML documents compression. The Computer Journal, 57(11):1723–1740, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1723.

Whittle:2012:NCC

[Whi12a] P. Whittle. A natural channel coding for the finite and infinite time axes. The Computer Journal, 55(7):788– 798, July 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/7/788.full.

pdf+html. Special Focus on the Centenary of Alan Turing.

Whittle:2012:NMO

[Whi12b]  Peter Whittle.  Neural

memories and oscillatory operation. The Computer Journal, 55(5):515– 521, May 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/5/515.full.
                                        pdf+html.

Wei:2016:PAB

[WHLH16] Jianghong Wei, Xinyi Huang, Wenfen Liu, and Xuexian Hu. Practical attributebased signature: Traceability and revocability. The Computer Journal, 59(11):1714–1734, November 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/11/1714.

Wright:2013:PFS

[WHP^{+}13] S. A. Wright, S. D. Hammond, S. J. Pennycook, R. F. Bird, J. A. Herdman, I. Miller, A. Vadgama, A. Bhalerao, and S. A. Jarvis. Parallel file system analysis through application I/O tracing. The Computer Journal, 56(2):141– 155, February 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/2/141.full.
pdf+html.

Wang:2016:DRS

[WHS^{+}16] Zhuo Wang, Tingting Hou, Dawei Song, Zhun Li, and Tianqi Kong. Detecting review spammer groups via bipartite graph projection. The Computer Journal, 59(6):861– 874, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/6/861.

Wu:2015:MQA

[WHSW15] Lin Wu, Xiaodi Huang, John Shepherd, and Yang Wang. Multi-query augmentationbased Web landmark photo retrieval. The Computer Journal, 58(9):2120–2134, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/2120.

Wang:2012:GAS

[WHYH12] Dan Wang, Xiaowei Han, Hong Yang, and Binbin Hao. On global adaptive synchronization in weighted complex networks. The Computer Journal, 55(12):1486–1491, December 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/12/1486.
                                        full.pdf+html.

Winskel:2011:ECS

[Win11] Glynn Winskel. Events, causality and symmetry. The Computer Journal, 54 (1):42–57, January 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/54/1/42.full.
                                        pdf+html.

Wang:2016:TAR

[WJ16] Weifeng Wang and Li Jiao. Trace abstraction refinement for solving Horn clauses. The Computer Journal, 59(8):1236–1251, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/8/1236.

Wang:2019:MBN

[WJ19] Qian Wang and Chenhui Jin. A method to bound the number of active Sboxes for a kind of AESlike structure. The Computer Journal, 62(8):1121– 1131, August 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1121/5319150.

Wang:2013:EVT

[WL13] Fasheng Wang and Mingyu Lu. Efficient visual tracking via Hamiltonian Monte Carlo Markov chain. The Computer Journal, 56(9): 1102–1112, September 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.
org/content/56/9/1102.
full.pdf+html.

Wang:2018:MSD

[WL18] Ji-Bo Wang and Lin Li. Machine scheduling with deteriorating jobs and modifying maintenance activities. The Computer Journal, 61(1): 47–53, January 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/47/3867592.

Wang:2019:SFT

[WLC^{+}19] Guijuan Wang, Cheng-Kuan Lin, Baolei Cheng, Jianxi Fan, and Weibei Fan. Structure fault-tolerance of the generalized hypercube. The Computer Journal, 62(10):1463–1476, October 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/10/1463/5480919.

Wang:2015:RGD

[WLH15a] Yang Wang, Hua Li, and Menglan Hu. Reusing garbage data for efficient workflow computation. The Computer Journal, 58(1): 110–125, January 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/1/110.

Wei:2015:FST

[WLH15b] Jianghong Wei, Wenfen Liu, and Xuexian Hu. Forward-secure threshold attribute-based signature scheme. The Computer Journal, 58(10):2492–2506, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2492.

Wen:2018:CRF

[WLHH18] Yunhua Wen, Shengli Liu, Ziyuan Hu, and Shuai Han. Computational robust fuzzy extractor. The Computer Journal, 61(12):1794–1805, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/12/1794/5060317.

Widanapathirana:2014:ASS

[WLI^{+}14] Chathuranga Widanapathirana, Jonathan C. Li, Milosh V. Ivanovich, Paul G. Fitzpatrick, and Y. Ahmet Şekercioğlu. Adaptive statistical signatures of network soft-failures in user devices. The Computer Journal, 57(8):1262–1278, August 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/8/1262.

Wang:2018:SSA

[WLW^{+}18] Jiaqi Wang, Yunyao Lu, Xiaojie Wang, Jing Dong, and Xiping Hu. SAR: a socialaware route recommendation system for intelligent transportation. The Computer Journal, 61(7):987– 997, July 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/7/987/4987043.

Wang:2018:MMS

[WLWL18] Ji-Bo Wang, Lu Liu, Jian-Jun Wang, and Lin Li. Makespan minimization scheduling with ready times, group technology and shortening job processing times. The Computer Journal, 61 (9):1422–1428, September 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/9/1422/4835633.

Wu:2015:TRM

[WLY^{+}15] Guowei Wu, Zuosong Liu, Lin Yao, Jing Deng, and Jie Wang. A trust routing for multimedia social networks. The Computer Journal, 58(4):688– 699, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/688.

Wang:2015:POL

[WLZ^{+}15] Xiang Wang, Ying Lu, Yi Zhang, Zexi Zhao, Tongsheng Xia, and Limin Xiao. Power optimization in logic synthesis for mixed polarity Reed–Muller logic circuits. The Computer Journal, 58(6):1306–1313, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1306.

Wu:2018:MMO

[WLZ^{+}18] Chin-Chia Wu, Shang-Chia Liu, Chuanli Zhao, Sheng-Zhi Wang, and Win-Chin Lin. A multi-machine order scheduling with learning using the genetic algorithm and particle swarm optimization. The Computer Journal, 61(1):14–31, January 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/1/14/3073680.

Wang:2019:OSS

[WM19] Jinyong Wang and Xiaoping Mi. Open source software reliability model with the decreasing trend of fault detection rate. The Computer Journal, 62(9):1301–1312, September 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/9/1301/5146169.

Wu:2012:PSC

[WMS^{+}12] Wei Wu, Yi Mu, Willy Susilo, Xinyi Huang, and Li Xu. A provably secure construction of certificatebased encryption from certificateless encryption. The Computer Journal, 55(10):1157–1168, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1157.
full.pdf+html.

Waldock:2011:FCC

[WN11] Antony Waldock and David Nicholson. A framework for cooperative control applied to a distributed sensor network. The Computer Journal, 54(3):471– 481, March 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/3/471.full.
                                        pdf+html.

Wang:2017:MVC

[WNNZ17] Shuling Wang, Flemming Nielson, Hanne Riis Nielson, and Naijun Zhan. Modelling and verifying communication failure of hybrid systems in HCSP. The Computer Journal, 60(8):1111– 1130, August 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/8/1111/2354609.

Wang:2015:HAD

[WOLP15] Xing Wang, Nga Lam Or, Ziyan Lu, and Derek Pao. Hardware accelerator to detect multi-segment virus patterns. The Computer Journal, 58(10):2443–2460, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2443.

Wang:2010:APF

[WOV^{+}10] You-Chiun Wang, Tomoaki Ohtsuki, Athanasios (Thanos) Vasilakos, Ashutosh Sabharwal, Yuh-Shyan Chen, and Yu-Chee Tseng. Algorithms, protocols and future applications of wireless sensor networks. The Computer Journal, 53(10):1551–1552, December 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/10/1551.
full.pdf+html.

Wang:2017:PPK

[WP17] Yujue Wang and Hwee-Hwa Pang. Probabilistic public key encryption for controlled equijoin in relational databases. The Computer Journal, 60(4):600– 612, March 23, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/600/2354608.

Wood:2012:ERV

[WRSV12] Timothy Wood, K. K.

Ramakrishnan, Prashant Shenoy, and Jacobus Van der Merwe. Enterpriseready virtual cloud pools: Vision, opportunities and challenges. The Computer Journal, 55(8):995– 1004, August 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/995.full.
                                        pdf+html.

Wu:2010:ABE

[WS10] Zheng-Yu Wu and Han-Tao Song. Ant-based energy-aware disjoint multipath routing algorithm for MANETs. The Computer Journal, 53(2):166– 176, February 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/2/166;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/2/166.

Wang:2015:DBS

[WS15] Yang Wang and Wei Shi. Dataflow-based scheduling for scientific workflows in HPC with storage constraints. The Computer Journal, 58(7):1628–1644, July 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/7/1628.

Wang:2015:RSA

[WSA15] Yang Wang, Willy Susilo, and Man Ho Au. Revisiting security against the arbitrator in optimistic fair exchange. The Computer Journal, 58(10):2665–2676, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/10/2665.

Wu:2011:FEO

[WSR11] Yanbo Wu, Quan Z. Sheng, and Damith C. Ranasinghe. Facilitating efficient object tracking in large-scale traceability networks. The Computer Journal, 54(12):2053– 2071, November 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/12/2053.
full.pdf+html.

Wen:2019:PLU

[WSY19] Tangliu Wen, Lan Song, and Zhen You. Proving linearizability using reduction. The Computer Journal, 62(9):1342–1364, September 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/9/1342/5187914.

Wu:2010:IBM

[WT10] Tsu-Yang Wu and Yuh-Min Tseng. An ID-based mutual authentication and key exchange protocol for low-power mobile devices.

The Computer Journal, 53(7):1062–1070, September 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/1062;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/1062.

Wu:2012:PCH

[WT12] Xingfu Wu and Valerie Taylor. Performance characteristics of hybrid MPI/ OpenMP implementations of NAS Parallel Benchmarks SP and BT on large-scale multicore clusters. The Computer Journal, 55(2):154–167, February 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/2/154.full.
pdf+html.

Wang:2018:EHI

[WT18] Husen Wang and Qiang Tang. Efficient homomorphic integer polynomial evaluation based on GSW FHE. The Computer Journal, 61(4):575–585, April 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/575/4791881.

Wu:2011:SPF

[WVGP11] Yuqing Wu, Dirk Van Gucht, Marc Gyssens, and Jan Paredaens. A study of a positive fragment of path queries: Expressiveness, normal form and minimization. The Computer Journal, 54(7):1091–1118, July 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1091.
full.pdf+html.

Wang:2017:IST

[WW17] Shuo-I Wang and Fu-Hsing Wang. Independent spanning trees in RTCCpyramids. The Computer Journal, 60(1):13–26, January 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wang:2019:SCB

[WW19] Shiying Wang and Mujiangshan Wang. The strong connectivity of bubble-sort star graphs. The Computer Journal, 62(5):715– 729, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/715/5074542.

Wiese:2017:RSM

[WWB17] Lena Wiese, Tim Waage, and Ferdinand Bollwein. A replication scheme for multiple fragmentations with overlapping fragments. The Computer Journal, 60(3): 60–??, March 2017. CO-DEN CMPJA6. ISSN ????

URL https://academic.
oup.com/comjnl/article/
60/3/308/2608047/A-Replication-
Scheme-for-Multiple-Fragmentations.

Wang:2011:MMW

[WWC^{+}11] Yini Wang, Sheng Wen, Silvio Cesare, Wanlei Zhou, and Yang Xiang. The microcosmic model of worm propagation. The Computer Journal, 54(10):1700–1720, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1700.
full.pdf+html.

Wang:2012:HAR

[WWHL12] Yang Wang, Lin Wu, Xiaodi Huang, and Xuemin Lin. Human action recognition from video sequences by enforcing tri-view constraints. The Computer Journal, 55(9):1030–1040, September 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

org/content/55/9/1030.
full.pdf+html.

Wu:2018:STP

[WWJ18] Haiqin Wu, Liangmin Wang, and Shunrong Jiang. Secure top-k preference query for location-based services in crowd-outsourcing environments. The Computer Journal, 61(4):496–511, April 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/496/4539928.

Wang:2016:ECE

[WWW16] Shiying Wang, Zhenhua Wang, and Mujiangshan Wang. The 2-extra connectivity and 2-extra diagnosability of bubble-sort star graph networks. The Computer Journal, 59(12):1839– 1856, December 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/12/1839.

Wang:2018:MFA

[WWXH18] Biao Wang, Xueqing Wang, Rui Xue, and Xinyi Huang. Matrix FHE and its application in optimizing bootstrapping. The Computer Journal, 61(12):1845–1861, December 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/12/1845/5086682.

Wu:2017:DNR

[WWZ^{+}17] Wei Wu, Jiguang Wan, Ling Zhan, Jibin Wang, and Changsheng Xie. DROP: a new RAID architecture for enhancing shared RAID performance. The Computer Journal, 60(6):777– 790, June 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/6/777/2953329.

Wang:2018:SMT

[WXLL18] Zhenyou Wang, Cuntao Xiao, Xianwei Lin, and Yuan-Yuan Lu. Single machine total absolute differences penalties minimization scheduling with a deteriorating and resourcedependent maintenance activity. The Computer Journal, 61(1):105–110, January 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/1/105/3867579.

Wu:2010:SMS

[WXP^{+}10] Zhao Wu, Naixue Xiong, Jong Hyuk Park, Tai-Hoon Kim, and Lei Yuan. A simulation model supporting time and non-time metrics for Web service composition. The Computer Journal, 53(2):219–233, February 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/2/219;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/2/219.

Wu:2012:ICA

[WXZ^{+}12] ZongDa Wu, GuanDong Xu, YanChun Zhang, Peter Dolog, and ChengLang Lu. An improved contextual advertising matching approach based on Wikipedia knowledge. The Computer Journal, 55(3):277– 292, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/3/277.full.
pdf+html.

Wan:2013:CPS

[WYL^{+}13] Jiafu Wan, Hehua Yan, Di Li, Keliang Zhou, and Lu Zeng. Cyber-physical systems for optimal energy management scheme of autonomous electric vehicle. The Computer Journal, 56(8):947–956, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/8/947.full.
                                        pdf+html.

Wei:2016:APS

[WYML16] Jiannan Wei, Guomin Yang, Yi Mu, and Kaitai Liang. Anonymous proxy signature with hierarchical traceability. The Computer Journal, 59(4):559–569, April 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/4/559.

Wan:2017:CHB

[WZ17] Shaohua Wan and Yin Zhang. Coverage hole bypassing in wireless sensor networks. The Computer Journal, 60(10):1536–1544, October 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/10/1536/4055589.

Wang:2018:AMB

[WZCC18] Rong Wang, Yan Zhu, Tung-Shou Chen, and Chin-Chen Chang. An authentication method based on the turtle shell algorithm for privacy-preserving data mining. The Computer Journal, 61(8):1123–1132, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/8/1123/4925401.

Wang:2018:EFT

[WZF18] Shiying Wang, Guozhen Zhang, and Kai Feng. Edge fault tolerance of Cartesian product graphs on super restricted edge connectivity. The Computer Journal, 61(5):761–772, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/5/761/4643583.

Zhang:2015:LPS

[wZfG15] Yi wen Zhang and Rui feng Guo. Low-power scheduling algorithms for sporadic task with shared resources in hard real-time systems. The Computer Journal, 58(7): 1585–1597, July 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/7/1585.

Wang:2012:NIS

[WZXL12] Xiaofeng Wang, Nanning Zheng, Jianru Xue, and Zhenli Liu. A novel image signature method for content authentication. The Computer Journal, 55(6): 686–701, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/55/6/686.full.
                                        pdf+html.

Xu:2014:AHA

[XGLM14] Chang Xu, Hua Guo, Zhoujun Li, and Yi Mu. Affiliation-hiding authenticated asymmetric group key agreement based on short signature. The Computer Journal, 57(10):1580–1590, October 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/10/1580.

Xie:2015:NNM

[XHC^{+}15] Ping Xie, Jianzhong Huang, Qiang Cao, Xiao Qin, and Changsheng Xie. A new non-MDS RAID-6 code to support fast reconstruction and balanced I/Os. The Computer Journal, 58(8):1811–1825, August 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/8/1811.

Xie:2018:SFR

[XHQX18] Ping Xie, Jianzhong Huang, Xiao Qin, and Changsheng Xie. SmartRec: Fast recovery from single failures in heterogeneous RAID-coded storage systems. The Computer Journal, 61(6):896– 911, June 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/6/896/4626972.

Xie:2013:FBA

[XHTH13] Yi Xie, J. Hu, S. Tang, and X. Huang. A forward– backward algorithm for nested hidden semi-Markov model and application to network traffic. The Computer Journal, 56(2):229– 238, February 2013. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/56/2/229.full.
pdf+html.

Xie:2011:SRM

[Xie11] Yulai Xie. Some results on minimum discrete bending energy path in simple polygon. The Computer Journal, 54(7):1205– 1210, July 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/7/1205.
full.pdf+html.

Xu:2019:TES

[XLC19] Zhiwu Xu, Ping Lu, and Haiming Chen. Towards an effective syntax and a generator for deterministic standard regular expressions. The Computer Journal, 62(9):1322–1341, September 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/9/1322/5165111.

Xu:2012:AHA

[XLM^{+}12] Chang Xu, Zhoujun Li, Yi Mu, Hua Guo, and Tao Guo. Affiliationhiding authenticated asymmetric group key agreement. The Computer Journal, 55(10):1180–1191, October 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/10/1180.
full.pdf+html.

Xie:2014:MLU

[XLM^{+}14] Haoran Xie, Qing Li, Xudong Mao, Xiaodong Li, Yi Cai, and Qianru Zheng. Mining latent user community for tag-based and content-based search in social media. The Computer Journal, 57(9):1415–1430, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1415.

Xia:2017:DLA

[XLX17] Qiufen Xia, Weifa Liang, and Zichuan Xu. Data locality-aware big data query evaluation in distributed clouds. The Computer Journal, 60(6):791– 809, June 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/791/3058779.

Xia:2017:DCB

[XLXZ17] Xiaofang Xia, Wei Liang, Yang Xiao, and Meng Zheng. Difference-comparison-based malicious meter inspection in neighborhood area networks in Smart Grid. The Computer Journal, 60(12):1852–1870, December 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/12/1852/4107829.

Xiang:2011:MAB

[XS11] Yonghong Xiang and Iain A.

Stewart. A multipath analysis of biswapped networks. The Computer Journal, 54(6):920–930, June 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/920.full.
pdf+html.

Xi:2011:CKV

[XTH11] Kai Xi, Yan Tang, and Jiankun Hu. Correlation keystroke verification scheme for user access control in cloud computing environment. The Computer Journal, 54(10):1632–1644, October 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/10/1632.
                                        full.pdf+html.

Xu:2011:EOK

[XXW11] Chenfeng Xu, Hongsheng Xi, and Fengguang Wu. Evaluation and optimization of kernel file readaheads based on Markov decision models. The Computer Journal, 54(11):1741–1755, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Xiao:2018:KRL

[XY18] Dianyan Xiao and Yang Yu. Klepto for ring-LWE encryption. The Computer Journal, 61(8):1228–1239, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1228/5035449.

Xu:2011:ABA

[XYL^{+}11] Shuo Xu, Fan Yu, Zhe Luo, Ze Ji, Duc Truong Pham, and Renxi Qiu. Adaptive bees algorithm — bioinspiration from honeybee foraging to optimize fuel economy of a semi-track aircushion vehicle. The Computer Journal, 54(9):1416– 1426, September 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/54/9/1416.
full.pdf+html.

Xu:2014:SIS

[XZA14] Guandong Xu, Aoying Zhou, and Nitin Agarwal. Special issue on social computing and its applications. The Computer Journal, 57(9):1279–1280, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1279.

Xu:2017:RCC

[XZL17] Xiang Xu, Shuming Zhou, and Jinqiang Li. Reliability of complete cubic networks under the condition of g-good-neighbor. The Computer Journal, 60(5):625– 635, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/5/625/2608077.

Xu:2018:KCC

[XZLL18] Liqiong Xu, Shuming Zhou, Guanqin Lian, and Zuwen Luo. A kind of conditional connectivity of Cayley graphs generated by 2trees. The Computer Journal, 61(5):714–721, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/5/714/4565507.

Xu:2015:ORA

[XZLW15] Chang Xu, Liehuang Zhu, Zhoujun Li, and Feng Wang. One-round affiliation-hiding authenticated asymmetric group key agreement with semi-trusted group authority. The Computer Journal, 58(10):2509–2519, October 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2509.

Xu:2017:MSB

[XZW^{+}17] Hai-Xia Xu, Wei Zhou, Yao-Nan Wang, Wei Wang, and Yan Mo. Matrix separation based on LMaFit-Seed. The Computer Journal, 60 (11):1609–1618, November 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/11/1609/2990263.

Xia:2010:ITA

[XZY^{+}10] Liang Xia, Yongxin Zhu, Jun Yang, Jingwei Ye, and Zonghua Gu. Implementing a thermal-aware scheduler in Linux kernel on a multicore processor. The Computer Journal, 53(7):895– 903, September 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/895;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/7/895.

Yu:2015:SDS

[YAM^{+}15] Yong Yu, Man Ho Au, Yi Mu, Willy Susilo, and Huai Wu. Secure delegation of signing power from factorization. The Computer Journal, 58(4):867– 877, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/867.

Yang:2019:CMS

[Yan19] Yuxing Yang. Characterizations of minimum structureand substructure-cuts of hypercubes. The Computer Journal, 62(9):1313–1321, September 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/9/1313/5158245.

Yu:2012:VID

[YAQ12] Hongfang Yu, Vishal Anand, and Chunming Qiao. Virtual infrastructure design for surviving physical link failures. The Computer Journal, 55(8):965–978, August 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/8/965.full.
                                        pdf+html.

Yasunaga:2019:ECS

[Yas19]  Kenji Yasunaga.  Er-

ror correction by structural simplicity: Correcting samplable additive errors. The Computer Journal, 62(9):1265–1276, September 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/9/1265/5116109.

Yu:2016:SFA

[YB16] Misun Yu and Doo-Hwan Bae. SimpleLock^{+}: Fast and accurate hybrid data race detection. The Computer Journal, 59(6):793– 809, June 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/6/793.

Yang:2011:HPB

[YC11] Kai-Ting Yang and Ge-Ming Chiu. A hybrid pull-based with piggybacked push protocol for cache sharing. The Computer Journal, 54(12):2017–2032, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/12/2017.
full.pdf+html.

Yang:2014:EIQ

[YC14a] Heejung Yang and Chin-Wan Chung. Efficient Iceberg query processing in sensor networks. The Computer Journal, 57(12):1834– 1851, December 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/12/1834.

Yang:2014:OIS

[YC14b] Jinn-Shyong Yang and Jou-Ming Chang. Optimal independent spanning trees on Cartesian product of hybrid graphs. The Computer Journal, 57(1):93–99, January 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/1/93.full.
                                        pdf+html.

Yang:2019:APG

[YC19] Dong-Yuh Yang and Yi-Chun Cho. Analysis of the N-policy GI/M/1/K queueing systems with working breakdowns and repairs. The Computer Journal, 62(1):130–143, January 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/1/130/5005378.

Yang:2015:ISC

[YCL15] Shiyu Yang, Muhammad Aamir Cheema, and Xuemin Lin. Impact set: Computing influence using query logs. The Computer Journal, 58 (11):2928–2943, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Yi:2017:ZCL

[YCL17] Wentan Yi, Shaozhen Chen, and Yuchen Li. Zerocorrelation linear cryptanalysis of SAFER block cipher family using the undisturbed bits. The Computer Journal, 60(4):613– 624, March 23, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/613/2354610.

Yu:2016:DNF

[YCR16] Jiangshan Yu, Vincent Cheval, and Mark Ryan. DTKI: a new formalized PKI with verifiable trusted parties. The Computer Journal, 59(11):1695–1713, November 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1695.

Yilmaz:2011:IAD

[YDE11] Onur Yilmaz, Orhan Dagdeviren, and Kayhan Erciyes. Interference-aware dynamic algorithms for energy-efficient topology control in wireless ad hoc and sensor networks. The Computer Journal, 54(8):1398–1411, August 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/8/1398.
full.pdf+html.

Yin:2018:TCE

[YDHW18] Liangze Yin, Wei Dong, Fei He, and Ji Wang. A trueconcurrency encoding for BMC of compositional systems. The Computer Journal, 61(5):658–671, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/5/658/4055636.

Yevtushenko:2015:DCD

[YEFVJ15] Nina Yevtushenko, Khaled El-Fakih, Tiziano Villa, and Jie-Hong R. Jiang. Deriving compositionally deadlockfree components over synchronous automata compositions. The Computer Journal, 58(11):2793–2803, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Yang:2015:SHI

[YGFL15] Zhen Yang, Kaiming Gao, Kefeng Fan, and Yingxu Lai. Sensational headline identification by normalized cross entropy-based metric. The Computer Journal, 58(4):644–655, April 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/644.

Yang:2014:IMP

[YGH^{+}14] Xu Yang, Deyuan Guo, Hu He, Haijing Tang, and Yanjun Zhang. An implementation of Message-Passing Interface over Vx-Works for real-time embedded multi-core systems. The Computer Journal, 57(11):1756–1764, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1756.

Yuan:2015:SSS

[YGLW15] Weiwei Yuan, Donghai Guan, Sungyoung Lee, and Jin Wang. Skeleton searching strategy for recommender searching mechanism of trust-aware recommender systems. The Computer Journal, 58(9):1876– 1883, September 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/9/1876.

Ying:2017:PFS

[YHGL17] Chen Ying, Hao Huang, Ajay Gopinathan, and Zongpeng Li. A priorfree spectrum auction for approximate revenue maximization. The Computer Journal, 60(6):898– 910, June 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/898/3073681.

Hsieh:2012:SSP

[yHRT^{+}12] Ming yu Hsieh, Rolf Riesen, Kevin Thompson, William Song, and Arun Rodrigues. SST: a scalable parallel framework for architecturelevel performance, power, area and thermal simulation. The Computer Journal, 55(2):181–191, February 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/2/181.full.
pdf+html.

Yu:2017:PNB

[YHS^{+}17] Jian Yu, Jun Han, Jean-Guy Schneider, Cameron Hine, and Steve Versteeg. A Petri-net-based virtual deployment testing environment for enterprise software systems. The Computer Journal, 60(1):27–44, January 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Yildiz:2012:UDT

[Yil12] Olcay Taner Yildiz. Univariate decision tree induction using maximum margin classification. The Computer Journal, 55(3):293– 298, March 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/55/3/293.full.
pdf+html.

Yasin:2014:OMS

[YIUH14] Waheed Yasin, Hamidah Ibrahim, Nur Izura Udzir, and Nor Asilah Wati Abdul Hamid. An overview of media streams caching in peerto-peer systems. The Computer Journal, 57(8):1167– 1177, August 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/8/1167.

Yang:2018:AIW

[YKK18] Zhi-Fang Yang, Chih-Ting Kuo, and Te-Hsi Kuo. Authorization identification by watermarking in logpolar coordinate system. The Computer Journal, 61 (11):1710–1723, November 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/11/1710/4993056.

Yi:2017:ICM

[YL17] Haibo Yi and Weijian Li. On the importance of checking multivariate public key cryptography for sidechannel attacks: The case of enTTS scheme. The Computer Journal, 60(8):1197– 1209, August 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/8/1197/2996413.

Yuen:2013:ELT

[YLA^{+}13] Tsz Hon Yuen, Joseph K. Liu, Man Ho Au, Willy Susilo, and Jianying Zhou. Efficient linkable and/or threshold ring signature without random oracles. The Computer Journal, 56 (4):407–421, April 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/56/4/407.full.
                                        pdf+html.

Yang:2015:PLB

[YLC15] Jinn-Shyong Yang, Sih-Syuan Luo, and Jou-Ming Chang. Pruning longer branches of independent spanning trees on folded hyper-stars. The Computer Journal, 58(11):2972–2981, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Yang:2012:WSI

[YLL^{+}12] Chunfang Yang, Fenlin Liu, Shiguo Lian, Xiangyang Luo, and Daoshun Wang. Weighted stego-image steganalysis of messages hidden into each bit plane. The Computer Journal, 55(6): 717–727, June 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/6/717.full.
pdf+html.

Yang:2017:SGC

[YLL^{+}17] Zheng Yang, Junyu Lai, Chao Liu, Wanping Liu, and Shuangqing Li. Simpler generic constructions for strongly secure one-round key exchange from weaker assumptions. The Computer Journal, 60(8):1145–1160, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/8/1145/2725496.

Yu:2016:DLR

[YLLS16] Hairong Yu, Guohui Li, Jianjun Li, and Lihchyun Shu. DO_{cyclical}: a latencyresistant cyclic multi-threading approach for automatic program parallelization. The Computer Journal, 59(8):1155–1173, August 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/8/1155.

Yan:2019:ETF

[YLSL19] Xuezhi Yan, Ruige Li, Xiaoying Sun, and Guohong Liu. Effects of touch force profiles and waveforms of electrostatic tactile feedback on touchscreen pan operation. The Computer Journal, 62(7):1016–1035, July 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/7/1016/5165118.

Yang:2017:AES

[YLW^{+}17] Yu Yang, Hongbo Liu, Hua Wang, Ansheng Deng, and Colton Magnant. On algorithms for enumerating subtrees of hexagonal and phenylene chains. The Computer Journal, 60(5):690– 710, April 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

                                        academic.oup.com/comjnl/
                                        article/60/5/690/2632622.

Yu:2011:IRS

[YLX^{+}11] Jie Yu, Zhoujun Li, Peng Xiao, Chengfang Fang, Jia Xu, and Ee-Chien Chang. ID repetition in structured P2P networks. The Computer Journal, 54(6):962– 975, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/6/962.full.
                                        pdf+html.

Yang:2015:MIM

[YMS^{+}15] Chen Yang, Jian Ma, Thushari Silva, Xiaoyan Liu, and Zhongsheng Hua. A multilevel information mining approach for expert recommendation in online scientific communities. The Computer Journal, 58(9):1921–1936, September 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/1921.

Yu:2011:CLE

[YMWS11] Yong Yu, Yi Mu, Guilin Wang, and Ying Sun. Cryptanalysis of an offline electronic cash scheme based on proxy blind signature. The Computer Journal, 54(10):1645–1651, October 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/10/1645.
full.pdf+html.

Yuksel:2011:SKE

[YNN11] Ender Yüksel, Hanne Riis Nielson, and Flemming Nielson. A secure key establishment protocol for Zig-Bee wireless sensor networks. The Computer Journal, 54(4):589–601, April 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.

                                        org/content/54/4/589.full.
                                        pdf+html.

Yen:2015:MSM

[YNP15] Neil Y. Yen, Uyen Trang Nguyen, and Jong Hyuk Park. Mining social media for knowledge discovery. The Computer Journal, 58(9):1859–1860, September 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/9/1859.

Yumbul:2015:EEP

[YS15] Kazim Yumbul and Erkay Savaş. Enhancing an embedded processor core for efficient and isolated execution of cryptographic algorithms. The Computer Journal, 58(10):2368–2387, October 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/10/2368.

Yang:2015:EPS

[YSC^{+}15] Bin Yang, Xingming Sun, Xianyi Chen, Jianjun Zhang, and Xu Li. Exposing photographic splicing by detecting the inconsistencies in shadows. The Computer Journal, 58(4):588– 600, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/588.

Younes:2011:SMI

[YT11] Osama Younes and Nigel Thomas. An SRN model of the IEEE 802.11 DCF MAC protocol in multihop ad hoc networks with hidden nodes. The Computer Journal, 54(6):875– 893, June 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/54/6/875.full.
pdf+html.

Yangui:2016:OCM

[YT16a] Sami Yangui and Samir Tata. An OCCI compliant model for PaaS resources description and provisioning. The Computer Journal, 59(3):308– 324, March 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/3/308.

Yi:2016:VSF

[YT16b] Haibo Yi and Shaohua Tang. Very small FPGA processor for multivariate signatures. The Computer Journal, 59(7):1091–1101, July 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/7/1091.

Yi:2016:FIS

[YTV16] Haibo Yi, Shaohua Tang, and Ranga Vemuri. Fast inversions in small finite fields by using binary trees. The Computer Journal, 59(7): 1102–1112, July 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/7/1102.

Yu:2012:ECR

[YWDW12] Kui Yu, Xindong Wu, Wei Ding, and Hao Wang. Exploring causal relationships with streaming features. The Computer Journal, 55(9):1103–1117, September 2012. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/9/1103.
                                        full.pdf+html.

Yuan:2018:NSB

[YWFQ18] Baoguo Yuan, Junfeng Wang, Zhiyang Fang, and Li Qi. A new software birthmark based on weight sequences of dynamic control flow graph for plagiarism detection. The Computer Journal, 61(8):1202–1215, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/8/1202/5026598.

Yao:2014:GFT

[YWR^{+}14] Lin Yao, Guowei Wu, Jiankang Ren, Yanwei Zhu, and Ying Li. Guaranteeing fault-tolerant requirement load balancing scheme based on VM migration. The Computer Journal, 57(2):225–232, February 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/2/225.full.
pdf+html.

Yi:2010:CGS

[YWSH10] Chih-Wei Yi, Peng-Jun Wan, Chao-Min Su, and Chen-Wei Huang. The critical Grid size and transmission radius for localminimum-free Grid routing in wireless ad hoc and sensor networks. The Computer Journal, 53(10):1621–1631, December 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/10/1621.
full.pdf+html.

Yang:2010:SEP

[YWY10] Xuejun Yang, Ji Wang, and Xiaodong Yi. Slicing execution with partial weakest precondition for model abstraction of C programs. The Computer Journal, 53(1):37–49, January 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/cgi/content/abstract/
53/1/37;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/1/37.

Yu:2017:PFS

[YY17] Huifang Yu and Bo Yang. Pairing-free and secure certificateless signcryption scheme. The Computer Journal, 60(8):1187–1196, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/8/1187/2979229.

Yildiz:2017:BLF

[YYK^{+}17] Muhammet Yildiz, Berrin Yanikoğlu, Alisher Kholmatov, Alper Kanak, Umut Uludağ, and Hakan Erdoğan. Biometric layering with fingerprints: Template security and privacy through multi-biometric template fusion. The Computer Journal, 60(4):573–587, March 23, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/573/2354606.

Yesilyurt:2015:RWM

[YYO15] Murat Yesilyurt, Yildiray Yalman, and A. Turan Ozcerit. A robust watermarking method for MPEG-4 based on kurtosis. The Computer Journal, 58(7): 1645–1655, July 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/58/7/1645.

Zhang:2018:NSG

[yZdZhZ18] Bei yang Zhang, Xie dong Zhang, and Zhi hua Zhang. Nested-stacking genetic algorithm for the optimal placement of sensors in bridge. The Computer Journal, 61(9):1269–1283, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1269/4608881.

Yu:2012:AAW

[YZJH12] Shui Yu, Wanlei Zhou, Weijia Jia, and Jiankun Hu. Attacking anonymous Web browsing at local area networks through browsing dynamics. The Computer Journal, 55(4):410– 421, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/55/4/410.full.
                                        pdf+html.

Yi:2015:SWT

[YZLC15] Yuhao Yi, Zhongzhi Zhang, Yuan Lin, and Guanrong Chen. Small-world topology can significantly improve the performance of noisy consensus in a complex network. The Computer Journal, 58(12):3242– 3254, December 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/12/3242.

Zamazal:2019:MLS

[Zam19]  Ondřej Zamazal.  Ma-

chine learning support for EU funding project categorization. The Computer Journal, 62(11):1684–1694, November 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/11/1684/5389530.

Zhang:2010:LAW

[ZBY^{+}10] Yuan Zhang, Lichun Bao, Shih-Hsien Yang, Max Welling, and Di Wu. Localization algorithms for wireless sensor retrieval. The Computer Journal, 53(10):1594–1605, December 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/10/1594.
full.pdf+html.

Zafar:2010:ROM

[ZC10] Huzaifa Zafar and Daniel D.

Corkill. Reducing online model-development time by agents using constraints between shared observations. The Computer Journal, 53(8):1302–1314, October 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/53/8/1302.
full.pdf+html.

Zhang:2012:SBF

[ZCL^{+}12] Zhang Zhang, Jun Cheng, Jun Li, Wei Bian, and Dacheng Tao. Segmentbased features for time series classification. The Computer Journal, 55(9):1088– 1102, September 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/55/9/1088.
full.pdf+html.

Zhang:2013:KEH

[ZCL13] Yinghui Zhang, Xiaofeng Chen, and Hui Li. Keyevolving hierarchical IDbased signcryption. The Computer Journal, 56(10): 1228–1248, October 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        http://comjnl.oxfordjournals.
                                        org/content/56/10/1228.
                                        full.pdf+html.

Zhao:2016:FSM

[ZCX^{+}16] Xiang Zhao, Yifan Chen, Chuan Xiao, Yoshiharu Ishikawa, and Jiuyang Tang. Frequent subgraph mining based on Pregel. The Computer Journal, 59(8):1113– 1128, August 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/8/1113.

Zhou:2018:RAR

[ZDCZ18] Mosong Zhou, Xiaoshe Dong, Heng Chen, and Xingjun Zhang. A runtime available resource capacity evaluation model based on the concept of similar tasks. The Computer Journal, 61(5):722–744, May 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/5/722/4568417.

Zhou:2017:SOD

[ZDL^{+}17] Junwei Zhou, Hui Duan, Kaitai Liang, Qiao Yan, Fei Chen, F. Richard Yu, Jieming Wu, and Jianyong Chen. Securing outsourced data in the multi-authority cloud with fine-grained access control and efficient attribute revocation. The Computer Journal, 60(8):1210–1222, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/8/1210/3051823.

Zhan:2015:VSA

[ZDM^{+}15] Yongzhao Zhan, Shan Dai, Qirong Mao, Lu Liu, and Wei Sheng. A video semantic analysis method based on kernel discriminative sparse representation and weighted KNN. The Computer Journal, 58(6):1360– 1372, June 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1360.

Zheng:2015:RAC

[ZDZ^{+}15a] Hao Zheng, Xiaoshe Dong, Zhengdong Zhu, Baoke Chen, Yizhi Zhang, and Xingjun Zhang. Research on algorithms to capture drivers’ write operations. The Computer Journal, 58 (11):3035–3056, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Zuo:2015:CRE

[ZDZ^{+}15b] Liyun Zuo, Shoubin Dong, Chunsheng Zhu, Lei Shu, and Guangjie Han. A cloud resource evaluation model based on entropy optimization and ant colony clustering. The Computer Journal, 58(6):1254–1266, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1254.

Zhang:2018:FAA

[ZFL18] Xiaojuan Zhang, Xiutao Feng, and Dongdai Lin. Fault attack on ACORN v3. The Computer Journal, 61(8):1166–1179, August 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/61/8/1166/4993730.

Zhu:2012:PPP

[ZFZ12] Lina Zhu, Li Feng, and Zuochang Zhang. Predicting the propagation path of random worm by subnet infection situation using fuzzy reasoning. The Computer Journal, 55(4):487– 496, April 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.

org/content/55/4/487.full.
pdf+html.

Zhou:2016:HFD

[ZGC16] Peng Zhou, Xiaojing Gu, and Rocky K. C. Chang. Harvesting file download exploits in the Web: a hacker’s view. The Computer Journal, 59(4):522– 540, April 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/4/522.

Zhou:2014:TDP

[ZH14] Huiyu Zhou and Kotaro Hirasawa. Traffic density prediction with timerelated data mining using genetic network programming. The Computer Journal, 57(9):1395–1414, September 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/9/1395.

Zadeh:2015:ASP

[ZH15] Abdulah Abdulah Zadeh and Howard M. Heys. Application of simple power analysis to stream ciphers constructed using feedback shift registers. The Computer Journal, 58(4):961– 972, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/961.

Zhao:2019:GTC

[ZH19] Shu-Li Zhao and Rong-Xia Hao. The generalized threeconnectivity of two kinds of Cayley graphs. The Computer Journal, 62(1):144– 149, January 1, 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/1/144/5005389.

Zhang:2015:STR

[Zha15] Zhiyong Zhang. Security, trust and risk in multimedia social networks. The Computer Journal, 58(4):515– 517, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/515.

Zhu:2015:IDM

[ZHL15] Hui Zhu, Cheng Huang, and Hui Li. Information diffusion model based on privacy setting in online social networking services. The Computer Journal, 58(4):536– 548, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/536.

Zhao:2017:ISM

[ZHL^{+}17] Chuanli Zhao, Chou-Jung Hsu, Win-Chin Lin, Wen-Hsiang Wu, and Chin-Chia Wu. An investigation of single-machine duewindow assignment with time-dependent processing times and a controllable rate-modifying activity. The Computer Journal, 60(9):1353–1362, September 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/9/1353/2990264.

Zhao:2019:GCK

[ZHW19] Shu-Li Zhao, Rong-Xia Hao, and Lidong Wu. The generalized connectivity of (n, k)-bubble-sort graphs. The Computer Journal, 62(9):1277–1283, September 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/9/1277/5139680.

Zhou:2014:PAL

[ZHY^{+}14] Jiufeng Zhou, Lixin Han, Yuan Yao, Xiaoqin Zeng, and Feng Xu. A parallel approach to link sign prediction in large-scale online social networks. The Computer Journal, 57(7):1092– 1104, July 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/7/1092.
                                        full.pdf+html.

Zeng:2014:NFC

[ZJ14] Shengke Zeng and Shaoquan Jiang. A new framework for conditionally anonymous ring signature. The Computer Journal, 57(4):567– 578, April 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/57/4/567.full.
                                        pdf+html.

Zhou:2015:DAJ

[ZJH^{+}15] Wei Zhou, Tao Jing, Yan Huo, Jin Qian, and Zhen Li. Double auction for joint channel and power allocation in cognitive radio networks. The Computer Journal, 58(12):3295–3305, December 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/12/3295.

Zhang:2017:TOE

[ZJHJ17] Fan Zhang, Tao Jing, Yan Huo, and Kaiwei Jiang. Throughput optimization for energy harvesting cognitive radio networks with save-then-transmit protocol. The Computer Journal, 60(6):911–924, June 1, 2017. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/6/911/3073682.

Zhang:2019:JOS

[ZJHJ19] Fan Zhang, Tao Jing, Yan Huo, and Kaiwei Jiang. Joint optimization of spectrum sensing and transmit power in energy harvesting cognitive radio sensor networks. The Computer Journal, 62(2):215–230, February 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/2/215/5050814.

Zhang:2016:LRR

[ZJLC16] Tao Zhang, He Jiang, Xiapu Luo, and Alvin T. S. Chan. A literature review of research in bug resolution: Tasks, challenges and future directions. The Computer Journal, 59(5):741– 773, May 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/5/741.

Zhou:2015:DCD

[ZL15] Lihua Zhou and Kevin Lü.

Detecting communities with different sizes for social network analysis. The Computer Journal, 58(9):1894– 1908, September 2015. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/58/9/1894.

Zhong:2019:ECC

[ZL19] Jianghua Zhong and Dongdai Lin. On equivalence of cascade connections of two nonlinear feedback shift registers. The Computer Journal, 62(12):1793–1804, December 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/12/1793/5554328.

Zeyda:2014:CMS

[ZLCW14] Frank Zeyda, Lalkhumsanga Lalkhumsanga, Ana Cavalcanti, and Andy Wellings. Circus models for safetycritical Java programs. The Computer Journal, 57(7): 1046–1091, July 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/7/1046.
                                        full.pdf+html.

Zupancic:2015:MAA

[ZLG15] Domen Zupančič, Mitja Luštrek, and Matjaž Gams. Multi-agent architecture for control of heating and cooling in a residential space. The Computer Journal, 58(6):1314–1329, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1314.

Zhou:2014:EIM

[ZLL^{+}14] Rui Zhou, Chengfei Liu, Jianxin Li, Junhu Wang, and Jeffrey Xu Yu. Evaluating irredundant maximal contained rewritings for XPath queries on views. The Computer Journal, 57(11):1674–1692, November 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/57/11/1674.

Zhang:2015:CSA

[ZLX^{+}15] Jianhui Zhang, Zhi Li, Feng Xia, Shaojie Tang, Xingfa Shen, and Bei Zhao. Cooperative scheduling for adaptive duty cycling in asynchronous sensor networks. The Computer Journal, 58(6):1267–1279, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/6/1267.

Zhai:2019:CDG

[ZLX^{+}19] Yafei Zhai, Limei Lin, Li Xu, Xinxin Zhang, and Yanze Huang. The conditional diagnosability with ggood-neighbor of exchanged hypercubes. The Computer Journal, 62(5):747– 756, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/5/747/5095651.

Zheng:2010:CLO

[ZLYX10] Meng Zheng, Wei Liang, Haibin Yu, and Yang Xiao. Cross layer optimization for energy-constrained wireless sensor networks: Joint rate control and routing. The Computer Journal, 53(10):1632–1642, December 2010. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/53/10/1632.
                                        full.pdf+html.

Zhang:2016:TLT

[ZM16] Mingwu Zhang and Yi Mu. Token-leakage tolerant and vector obfuscated IPE and application in privacypreserving two-party point/ polynomial evaluations. The Computer Journal, 59(4): 493–507, April 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/59/4/493.

Zhang:2018:SPF

[ZM18] Jianhong Zhang and Jian Mao. On the security of a pairing-free certificateless signcryption scheme. The Computer Journal, 61(4): 469–471, April 1, 2018. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/469/4107198.

Zhao:2019:AFT

[ZM19] Shuang Zhao and Jixiang Meng. Arc fault tolerance of maximally arcconnected networks modeled by digraphs. The Computer Journal, 62(5):706– 714, May 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/5/706/5068282.

Zhao:2013:LLF

[ZMSM13] Wenbing Zhao, P. M.

Melliar-Smith, and L. E. Moser. Low latency fault tolerance system. The Computer Journal, 56(6):716– 740, June 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/6/716.full.
pdf+html.

Zhang:2016:CAH

[ZMW16] Leyou Zhang, Yi Mu, and Qing Wu. Compact anonymous hierarchical identitybased encryption with constant size private keys. The Computer Journal, 59(4): 452–461, April 2016. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/59/4/452.

Zhou:2015:EGE

[ZNQR15] Zhangbing Zhou, Huansheng Ning, Meikang Qiu, and Habib F. Rashvand. Editorial: Green energy management and smart grid. The Computer Journal, 58(6):1225–1226, June 2015. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/6/1225.

Zhang:2013:LIF

[ZQ13] Xiaoling Zhang and Jianguo Qian. L(p, q)-labeling and integer flow on planar graphs. The Computer Journal, 56(6):785– 792, June 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/6/785.full.
                                        pdf+html.

Zhu:2010:WVB

[ZSJ10] Xiaoyan Zhu, Qinbao Song, and Zihan Jia. A weighted voting-based associative classification algorithm. The Computer Journal, 53(6): 786–801, July 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/cgi/content/abstract/
53/6/786;   http://comjnl.
oxfordjournals.org/cgi/
reprint/53/6/786.

Zhu:2019:DFS

[ZSL19] Bing Zhu, Kenneth W.

Shum, and Hui Li. On the duality and file size hierarchy of fractional repetition codes. The Computer Journal, 62(1):150–160, January 1, 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/1/150/5095654.

Zhang:2010:PDS

[ZSX10] Baopeng Zhang, Yuanchun Shi, and Xin Xiao. A policy-driven service composition method for adaptation in pervasive computing environment. The Computer Journal, 53(2):152– 165, February 2010. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.

                                        org/cgi/content/abstract/
                                        53/2/152;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/53/2/152.

Zareei:2011:EME

[ZTBW11] Mahdi Zareei, Alireza Taghizadeh, Rahmat Budiarto, and Tat-Chee Wan. EMS-MAC: Energy efficient contentionbased medium access control protocol for mobile sensor networks. The Computer Journal, 54(12):1963–1972, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/12/1963.
                                        full.pdf+html.

Zhu:2015:PPD

[ZTL15] Hong Zhu, Shengli Tian, and Kevin Lü. Privacypreserving data publication with features of independent ℓ-diversity. The Computer Journal, 58(4):549– 571, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/549.

Zulkufli:2018:WCC

[ZTTM18] Nurul Liyana Mohamad Zulkufli, Sherzod Turaev, Mohd Izzuddin Mohd Tamrin, and Azeddine Messikh. Watson–Crick context-free grammars: Grammar simplifications and a parsing algorithm. The Computer Journal, 61(9):1361–1373, September 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/9/1361/4796924.

Zhang:2015:PSE

[ZV15] Xiaowang Zhang and Jan Van den Bussche. On the power of SPARQL in expressing navigational queries. The Computer Journal, 58(11):2841–2851, November 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Zhou:2016:SRB

[ZVG16] Lan Zhou, Vijay Varadharajan, and K. Gopinath. A secure role-based cloud storage system for encrypted patient-centric health records. The Computer Journal, 59(11):1593–1611, November 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/11/1593.

Zhou:2011:ERB

[ZVH11] Lan Zhou, Vijay Varadharajan, and Michael Hitchens. Enforcing role-based access control for secure data storage in the cloud. The Computer Journal, 54(10):1675– 1687, October 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/54/10/1675.
                                        full.pdf+html.

Zhang:2015:FAA

[ZW15] Zhiyong Zhang and Kanliang Wang. A formal analytic approach to credible potential path and mining algorithms for multimedia social networks. The Computer Journal, 58(4):668– 678, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/58/4/668.

Zhang:2019:CTO

[ZWC^{+}19] Zikai Zhang, Jigang Wu, Long Chen, Guiyuan Jiang, and Siew-Kei Lam. Collaborative task offloading with computation result reusing for mobile edge computing. The Computer Journal, 62(10):1450–1462, October 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        academic.oup.com/comjnl/
                                        article/62/10/1450/5462515.

Zhu:2015:MLS

[ZWFW15] Yanmin Zhu, Yin Wang, George Forman, and Hong Wei. Mining large-scale GPS streams for connectivity refinement of road maps. The Computer Journal, 58(9):2109–2119, September 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/9/2109.

Zhai:2014:NLB

[ZWJ^{+}14] Haibin Zhai, Albert K.

Wong, Hai Jiang, Yi Sun, Jun Li, and Zhongcheng Li. A node-link-based P2P cache deployment algorithm in ISP networks. The Computer Journal, 57(2):183– 194, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.
org/content/57/2/183.full.
pdf+html.

Zhang:2015:FER

[ZWTM15] Mingwu Zhang, Chunzhi Wang, Tsuyoshi Takagi, and Yi Mu. Functional encryption resilient to hard-to-invert leakage. The Computer Journal, 58(4):735– 749, April 2015. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/58/4/735.

Zhao:2016:RFB

[ZX16] Hongzhi Zhao and Yuan Xue. RSD fault block model for highly efficient fault-tolerant Manhattan routing algorithms in 2D mesh. The Computer Journal, 59(10):1511–1526, October 2016. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/59/10/1511.

Zhang:2011:SIR

[ZXZ^{+}11] Jun Zhang, Yang Xiang, Wanlei Zhou, Lei Ye, and Yi Mu. Secure image retrieval based on visual content and watermarking protocol. The Computer Journal, 54(10):1661–1674, October 2011. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/54/10/1661.
                                        full.pdf+html.

Zhou:2017:CLR

[ZY17] Yanwei Zhou and Bo Yang. Continuous leakage-resilient public-key encryption scheme with CCA security. The Computer Journal, 60(8): 1161–1172, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

                                        https://academic.oup.com/

comjnl/article/60/8/1161/
2882687.

Zhou:2017:EAT

[ZYF17] Cheng Zhou, Zhonghai Yin, and Youqian Feng. Events algebra of triggers. The Computer Journal, 60(5): 751–760, April 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/5/751/2937753.

Zhou:2019:CLR

[ZYH^{+}19] Yanwei Zhou, Bo Yang, Hongxia Hou, Lina Zhang, Tao Wang, and Mingxiao Hu. Continuous leakageresilient identity-based encryption with tight security. The Computer Journal, 62(8):1092–1105, August 2019. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/8/1092/5288324.

Zhou:2018:CLR

[ZYM18] Yanwei Zhou, Bo Yang, and Yi Mu. Continuous leakage-resilient identitybased encryption without random oracles. The Computer Journal, 61(4):586– 600, April 1, 2018. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/4/586/4824746.

Zhang:2013:EER

[ZYR^{+}13] Daqiang Zhang, Zhijun Yang, Vaskar Raychoudhury, Zhe Chen, and Jaime Lloret. An energyefficient routing protocol using movement trends in vehicular ad hoc networks. The Computer Journal, 56(8):938–946, August 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/8/938.full.
                                        pdf+html.

Zhang:2013:BLR

[ZYT13] Mingwu Zhang, Bo Yang, and Tsuyoshi Takagi. Bounded leakage-resilient functional encryption with hidden vector predicate. The Computer Journal, 56(4):464– 477, April 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.
                                        org/content/56/4/464.full.
                                        pdf+html.

Zhang:2013:MDP

[ZYWW13] Bin Zhang, Jiahai Yang, Jianping Wu, and Ziyu Wang. MBST: Detecting packet-level traffic anomalies by feature stability. The Computer Journal, 56(10):1176–1188, October 2013. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/10/1176.
full.pdf+html.

Zhang:2013:CEE

[ZYY^{+}13] Jing Zhang, Xi Yang, Qi Yao, Xiaohu Ge, Minho Jo, and Guoqiang Mao. Cooperative energy efficiency modeling and performance analysis in cochannel interference cellular networks. The Computer Journal, 56(8):1010– 1019, August 2013. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/56/8/1010.
full.pdf+html.

Zhao:2019:LRC

[ZYY19] Yi Zhao, Yong Yu, and Bo Yang. Leakage resilient CCA security in stronger model: Branch hidden ABO-LTFs and their applications. The Computer Journal, 62(4):631– 640, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/631/5272751.

Zhang:2017:MDS

[ZZ17] Hailong Zhang and Yongbin Zhou. Mahalanobis distance similarity measure based higher order optimal distinguisher. The Computer Journal, 60(8):1131–1144, August 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/8/1131/2967015.

Zhang:2018:VEG

[ZZLL18] Meng-Jia Zhang, Kang Zhang, Jie Li, and Yi-Na Li. Visual exploration of 3D geospatial networks in a virtual reality environment. The Computer Journal, 61(3):447–458, March 1, 2018. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/61/3/447/4725106.

Zhang:2017:NLR

[ZZM17a] Leyou Zhang, Jingxia Zhang, and Yi Mu. Novel leakage-resilient attributebased encryption from hash proof system. The Computer Journal, 60(4):541– 554, March 23, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/4/541/2608069.

Zheng:2017:AGB

[ZZM17b] Lin Zheng, Fuxi Zhu, and Alshahrani Mohammed. Attribute and global boosting: a rating prediction method in context-aware recommendation. The Computer Journal, 60(7):957– 968, July 1, 2017. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL https://

academic.oup.com/comjnl/
article/60/7/957/2609377.

Zeng:2019:SAE

[ZZQ^{+}19] Ming Zeng, Kai Zhang, Haifeng Qian, Xiaofeng Chen, and Jie Chen. A searchable asymmetric encryption scheme with support for Boolean queries for cloud applications. The Computer Journal, 62(4): 563–578, April 2019. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

academic.oup.com/comjnl/
article/62/4/563/5253754.

Zhao:2016:CGD

[ZZX16] Jie Zhao, Rongcai Zhao, and Jinchen Xu. Code generation for distributed-memory architectures. The Computer Journal, 59(1):119– 132, January 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.
org/content/59/1/119.

Zhu:2014:OML

[ZZZ14] Chao Zhu, Qiang Zhu, and Calisto Zuzarte. Optimization of monotonic linear progressive queries based on dynamic materialized views. The Computer Journal, 57 (5):708–730, May 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/5/708.full.
                                        pdf+html.